### v1

# Heights and periods of algebraic cycles in families

Abstract.
We consider the Beilinson–Bloch heights and Abel–Jacobian periods of homologically trivial Chow cycles in families. For the Beilinson–Bloch heights, we show that for any $g\geq 3$, there is a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$, the coarse moduli of curves of genus $g$ over $\overline{\mathbb{Q}}$, such that the heights of Ceresa cycles and Gross–Schoen cycles over $U$ satisfy the Northcott property. For the Abel–Jacobi periods, we provide an algebraic criterion for the existence of a Zariski open dense subset of any family such that all cycles not defined over $\overline{\mathbb{Q}}$ are non-torsion and verify that this criterion holds true for Ceresa cycles and Gross–Schoen cycles.
1. Introduction
This paper aims to initiate a project to study Beilinson–Bloch heights and Abel–Jacobi periods for homologically trivial cycles in families. We have achieved two goals: the Northcott property for the heights of the Gross–Schoen cycles and Cerasa cycles and an algebraic criterion for the non-degeneracy of Abel–Jacobi periods for the general family of cycles.
1.1. Gross–Schoen and Ceresa cycles
For a smooth projective irreducible curve $C$ defined over a filed $k$ of $g\geq 3$ and a class $\xi\in\mathrm{Pic}^{1}(C)$ such that $(2g-2)\xi=\omega_{C}$, we have two homologically trivial 1-cycles:
  1. (1)
    the Gross–Schoen cycle ${\mathrm{GS}}(C):=\Delta_{\xi}(C)\in{\mathrm{Ch}}_{1}(C^{3})$ obtained by modified the diagonal cycle in $C^{3}$ using base class $\xi$ ([GS95, Zha10]);
  2. (2)
    the Ceresa cycle by ${\mathrm{Ce}}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in{\mathrm{Ch}}_{1}({\mathrm{Jac}}(C))$ defined by embedding $i_{\xi}\colon C{\longrightarrow}{\mathrm{Jac}}(C)$ via $\xi$.
Up to torsions, the definition of these cycles does not depend on the choice of $\xi$.
Let ${\mathscr{M}}_{g}$ be the moduli space of smooth projective curves of genus $g$. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathscr{C}}_{s}$ the curve parametrized by $s$.
Theorem 1.1.
For each $g\geq 3$, there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$, and positive numbers $\epsilon$ and $c$, such that for any $s\in U(\overline{\mathbb{Q}})$,
$\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{\mathrm{Fal}}(s)-c,$
$\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{\mathrm{Fal}}(s)-c,$
where $h_{\mathrm{Fal}}$ is the Faltings height on ${\mathscr{M}}_{g}(\overline{\mathbb{Q}})$.
Theorem 1.1 yields the following Northcott property immediately.
Corollary 1.2 (Northcott property).
For each $g\geq 3$, there exists a Zariski open dense subset of the ${\mathscr{M}}_{g}$ defined over $\overline{\mathbb{Q}}$ such that for any $H,D\in{\mathbb{R}}$,
#$\#\{s\in U(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(s):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}<H\}<\infty,$
#$\#\{s\in U(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(s):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}<H\}<\infty.$
In the course of the proof, we also establish the following geometric result.
Theorem 1.3.
For each $g\geq 3$, there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ such that the followings hold true:
  1. (i)
    ${\mathrm{GS}}({\mathscr{C}}_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s})$ are both non-torsion in the Chow groups for all $s\in U({\mathbb{C}})\setminus U(\overline{\mathbb{Q}})$.
  2. (ii)
    there exist at most countably many $s\in U({\mathbb{C}})$ such that ${\mathrm{AJ}}({\mathrm{GS}}({\mathscr{C}}_{s}))$ or ${\mathrm{AJ}}({\mathrm{Ce}}({\mathscr{C}}_{s}))$ is torsion in the intermediate Jacobians.
Hain [Ha24] also proved Theorem 1.3 with $U$ a non-empty analytic open subset of ${\mathscr{M}}_{g}^{\mathrm{an}}$ with a different method. When $g=3$, Hain’s result is completely explicit, while ours is not.
The proofs of Theorem 1.1, Corollary 1.2 and Theorem 1.3 have two parts: the arithmetic part and the geometric part. Both parts are needed to prove Theorem 1.1 and Corollary 1.2, while Theorem 1.3 only relies on the geometric part. We will explain each part in more detail in §$\S$1.3.
Before moving on, let us mention that the geometric part of our paper works for any family of homologically trivial cycles and, more generally, for any admissible normal function. More specifically, for any smooth projective morphism $f\colon X\rightarrow S$ of algebraic varieties with irreducible fibers and any family of homologically trivial cycles $Z$, all defined over $\overline{\mathbb{Q}}$, we show that there exists a Zariski closed subset $S_{{\mathscr{F}}}(1)$ of $S$ such that: (i) $[Z_{s}]$ is non-torsion in ${\mathrm{Ch}}^{*}(X_{s})$ for every transcendental point $s$ of $S\setminus S_{{\mathscr{F}}}(1)$; (ii) there are at most countably many $s\in S({\mathbb{C}})$ outside $S_{{\mathscr{F}}}(1)$ such that ${\mathrm{AJ}}(Z_{s})$ is torsion in the intermediate Jacobian. We then prove a checkable criterion for $S_{{\mathscr{F}}}(1)\not=S$. A simple case will be presented in Corollary 1.9 and Remark 1.10. We do these by studying the normal function associated with $Z$, which is defined in §$\S$§C.5, and by studying its Betti rank.
1.2. Normal functions
Now, we turn to the geometric part of the framework and explain our results.
Let $S$ be a quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ be a VHS of weight $-1$ over $S$, and consider the intermediate Jacobian (write ${\mathscr{V}}$ for the holomorphic vector associated with ${\mathbb{V}}_{{\mathbb{Z}}}$)
$\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})={\mathbb{V}}_{\mathbb{Z}}\backslash{\mathscr{V}}/{\mathscr{F}}^{0}{\mathscr{V}}{\longrightarrow}S.$
Let $\nu\colon S{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function; then $\nu$ defines an admissible variation of mixed Hodge structures ${\mathbb{E}}_{\nu}$ on $S$ which is an extension of ${\mathbb{Z}}(0)_{S}$ by ${\mathbb{V}}_{{\mathbb{Z}}}$.
The fiberwise isomorphism ${\mathbb{V}}_{{\mathbb{R}},s}{\overset{\sim}{{\longrightarrow}}}{\mathbb{V}}_{{\mathbb{C}},s}/{\mathscr{F}}^{0}_{s}={\mathscr{V}}_{s}/{\mathscr{F}}^{0}_{s}$ makes ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ into a local system of real tori
${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\xrightarrow{\sim}{\mathbb{V}}_{\mathbb{R}}/{\mathbb{V}}_{\mathbb{Z}}.$
Let ${\mathscr{F}}_{\mathrm{Betti}}$ denote the induced foliation, which we call the Betti foliation. More precisely, for any point $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, there is a local section $\sigma\colon U{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ from a neighborhood $\Delta$ of $\pi(x)$ in $S^{\mathrm{an}}$, with $x\in\sigma(U)$, represented by a flat section of ${\mathbb{V}}_{\mathbb{R}}$. The manifolds $\sigma(U)$ gluing together to a foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. See §$\S$§C.2 for more details. In particular, all torsion multi-sections are leaves of ${\mathscr{F}}_{\mathrm{Betti}}$.
1.2.1. Betti strata
The Betti foliation defines a strata on $S$ as follows. For each integer $t\geq 0$, set
(1.1) $S_{\mathscr{F}}(t):=\left\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\right\}$
where by abuse of notation $\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}}$ means the intersection with the leaves. This subset is, by definition, real-analytic. We then have the following Betti strata on $S$
(1.2) $\emptyset=S_{{\mathscr{F}}}(\dim S+1)\subseteq S_{{\mathscr{F}}}(\dim S)\subseteq\cdots\subseteq S_{{\mathscr{F}}}(1)\subseteq S_{{\mathscr{F}}}(0)=S.$
A main theorem of the geometric part is that the Betti strata are algebraic:
Theorem 1.4 (Theorem 3.2).
For each $t\geq 0$, $S_{\mathscr{F}}(t)$ is Zariski closed in $S$.
Remark 1.5.
In the Betti strata, $S_{{\mathscr{F}}}(1)$ plays a particularly important role. For example, by definition of the Betti foliation, $S_{{\mathscr{F}}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion. In particular, there are at most countably many $s\in S({\mathbb{C}})$ outside $S_{{\mathscr{F}}}(1)$ such that $\nu(s)$ is torsion.
Finally, let us point out that $S_{{\mathscr{F}}}(t)$ is closely related to the degeneracy loci defined by the first-named author. Indeed, when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is an abelian scheme, then $S_{{\mathscr{F}}}(t)$ is the $\nu(S)^{\mathrm{deg}}(-t+1)$ in [Gao20a, Defn. 1.6] for each $t\geq 1$.
1.2.2. Betti rank
Now assume that $S$ is smooth. Then the Betti foliation induces a decomposition $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$ for each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. Thus for each $s\in S({\mathbb{C}})$ we have a linear map
(1.3) $\nu_{\mathrm{Betti},s}\colon T_{s}S\xrightarrow{\mathrm{d}\nu}T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}.$
The Betti rank of $\nu$ is defined to be:
(1.4) $r(\nu):=\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S).$
A trivial upper bound for $r(\nu)$ is $r(\nu)\leq\min\{\dim S,\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}\}$ for any $s\in S({\mathbb{C}})$. One can also easily improve this trivial upper bound as will be explained below.
The Betti rank is easily seen to be related to the Betti strata in the following way: $\dim S-r(\nu)$ is the maximum of $t\geq 0$ such that $S_{{\mathscr{F}}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. Thus Theorem 1.4 implies that $r(\nu)=\dim S-\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}$. However, this equality is, in general, not applicable to compute $r(\nu)$. But at this stage, we have
(1.5) $S_{{\mathscr{F}}}(1)\not=S\Leftrightarrow r(\nu)=\dim S.$
The second main theorem for the geometric part is the following formula for $r(\nu)$, which is often computable in practice. The VMHS ${\mathbb{E}}_{\nu}$ on $S$ induces a period map $\varphi=\varphi_{\nu}\colon S{\longrightarrow}\Gamma\backslash{\mathscr{D}}$, with ${\mathscr{D}}$ a mixed Mumford–Tate domain with Mumford–Tate group $\mathbf{G}$ (so ${\mathscr{D}}$ is a $\mathbf{G}({\mathbb{R}})^{+}$-orbit for $\mathbf{G}$ the generic Mumford–Tate group of the VMHS ${\mathbb{E}}_{\nu}$); we refer to §$\S$2.12.2 for the construction of $\varphi$ and ${\mathscr{D}}$.
Denote by $V$ the unipotent radical of $\mathbf{G}$. It equals ${\mathbb{V}}^{\prime}_{{\mathbb{Q}},s}$ for any $s\in S({\mathbb{C}})$, where ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ is the largest sub-VHS of ${\mathbb{V}}_{{\mathbb{Z}}}$ such that $\nu$ becomes torsion under the projection ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}/{\mathbb{V}}^{\prime}_{{\mathbb{Z}}})$; see Remark 2.2. Now the trivial upper bound on $r(\nu)$ can be easily improved to be $r(\nu)\leq\min\{\dim\varphi(S),\frac{1}{2}\dim V\}$.
Here is our formula to compute $r(\nu)$.
Theorem 1.6 (Theorem 3.1).
The Betti rank is given by
(1.6) $r(\nu)=\min_{N}\left(\dim\varphi_{/N}(S)+\frac{1}{2}\dim_{\mathbb{Q}}(V\cap N)\right),$
where $N$ runs through the set of normal subgroup of $\mathbf{G}$, and $\varphi_{/N}$ is the induced period map
[…]
with $[p_{N}]$ the quotient by $N$ (see §$\S$§B.3).
Notice that the trivial upper bound on $r(\nu)$ above is recovered by taking $N=\{1\}$ and $N=\mathbf{G}$.
Here are two applications of Theorem 1.6, on two cases where the trivial upper bound on $r(\nu)$ is attained.
Corollary 1.7 (Theorem 5.1).
Assume: (i) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is irreducible, i.e. the only sub-VHSs are trivial or itself; (ii) $\nu(S)$ is not a torsion section. Then $r(\nu)=\min\{\dim\varphi(S),\frac{1}{2}\dim_{{\mathbb{Q}}}{\mathbb{V}}_{{\mathbb{Q}},s}\}$ for one (and hence all) $s\in S({\mathbb{C}})$.
Notice that in the situation of Corollary 1.7, we have $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
Corollary 1.8 (Theorem 5.2).
Assume: (i) the connected algebraic monodromy group $H$ of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple; (ii) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ has no isotrivial sub-VHS, i.e. locally constant VHS. Then $r(\nu)=\min\{\dim\varphi(S),\frac{1}{2}\dim_{{\mathbb{Q}}}V\}$.
An immediate corollary on the torsion locus, combining (1.5) and Remark 1.5 and Theorem 1.4, is:
Corollary 1.9.
Under the assumptions of either Corollary 1.7 or Corollary 1.8. Assume that $\varphi$ is generically finite and that $\dim S\leq\frac{1}{2}\dim_{{\mathbb{Q}}}V$. Then there exists a Zariski open dense subset $U$ of $S$ such that $\nu(s)$ is torsion for at most countably many $s\in U({\mathbb{C}})$. Indeed, one can take $U=S\setminus S_{{\mathscr{F}}}(1)$.
Remark 1.10.
If $\varphi$ is not necessarily generically finite, then the conclusion becomes: there exists a Zariski open dense subset $U$ of $S$ such that $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is contained in at most countably many fibers of $\varphi$. Indeed, $\varphi(S)$ is an algebraic variety by [BBT23]. Then we can conclude by applying Corollary 1.9 to $\varphi(S)$ and the induced normal function $\nu^{\prime}$ (see §$\S$3.1 for the construction of $\nu^{\prime}$).
Applying either Corollary 1.7 or Corollary 1.8 to the Gross–Schoen and the Ceresa normal functions, we obtain the following corollary will be used to prove Theorem 1.1, Corollary 1.2 and Theorem 1.3.
Corollary 1.11 (Corollary 5.3).
If $g\geq 3$, then the Betti rank of the Gross-Schoen (resp. of the Ceresa) normal function is $33g-$ 3.
Hain [Ha24] gave a different proof of Corollary 1.11. When $g=3$, he even gave an explicit description on the locus where $\dim\nu_{\mathrm{Betti},s}(T_{s}S)<3g-3$.
1.3. Plan and Ingredients of proofs
Our plan to prove results like Theorem 1.1, Corollary 1.2 and Theorem 1.3 has two parts: the arithmetic part and the geometric part. Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers and let $Z$ be a family of homologically trivial cycles, all defined over $\overline{\mathbb{Q}}$
The arithmetic part requires the construction of a suitable adelic line bundle $\overline{{\mathscr{L}}}$ over the base, a theory initiated by the second-named author and developed in further joint work with Yuan [YZ21], such that the height function $h_{\overline{{\mathscr{L}}}}$ is the Beilinsin–Bloch height. Then, the desired height inequalities will follow from suitable bigness properties of $\overline{{\mathscr{L}}}$. In the context of Theorem 1.1 and Corollary 1.2, by the second-named author’s [Zha10, Thm. 2.5.5] the desired $\overline{{\mathscr{L}}}$ can be obtained from a suitable Deligne pairing. The required bigness property is the bigness of the generic fiber $\tilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$. We furthermore show in Proposition 6.3 that the bigness of $\tilde{{\mathscr{L}}}$ will follow from the non-vanishing of $c_{1}(\overline{{\mathscr{L}}})^{\wedge\dim S}$, with $c_{1}(\overline{{\mathscr{L}}})$ the curvature form. The arithmetic part of this paper is confined to the Gross-Shoen and the Ceresa cycles.
The geometric part studies the (admissible) normal function $\nu$ associated with $Z$. The main results are the Zariski closedness of the Betti strata and the formula for the Betti rank $r(\nu)$. This part has been explained in §$\S$1.2. The proofs of Theorem 1.4 and Theorem 1.6 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the generic rank of the Betti map for abelian schemes. A core of our proof is Ax–Schanuel for VMHS independently proved by Chiu [Chi21] and Gao–Klingler [GK24], which will be used multiple times.
Hain’s works on the Hodge-theoretic computation of the Archimedean local height pairing and the Betti form are the key ingredients to bridging the geometric and arithmetic parts. Indeed, Hain in [Hai90] proved that the archimedean local height pairing could be computed using the metricized biextension line bundle on $S$, and in [HR04] computed (joint with Reed), the curvature form of the metrized biextension line bundle. In our situation, we work with the height pairing of $Z_{s}$ with itself for any $s\in S(\overline{\mathbb{Q}})$, and the metrized biextension line bundle in question is the pullback of the metrized tautological line bundle under $\nu$. Hain in [Hai13] proved that its curvature form $\beta_{\nu}$, called the Betti form, is a semi-positive $(1,1)$-form. We show in Corollary D.7 that $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$ if and only if $r(\nu)=\dim S$. Finally, back to the situation of Theorem 1.1, Corollary 1.2, and Theorem 1.3. Hain’s formula [Hai90, Prop. 3.3.12] and the second-named author’s [Zha10, Thm. 2.5.5] together imply that the curvature $c_{1}(\overline{{\mathscr{L}}})$ is precisely the Betti form. This finishes the bridge between the geometric part and the arithmetic part. Again, the bridge between the geometric part and the arithmetic part works in much larger generality, provided that the arithmetic part is solved.
Acknowledgements
We would like to thank Richard Hain for numerous inspiring discussions and for sharing his notes and Xinyi Yuan for his help during the preparation of this project. Part of this work was done when ZG visited SZ at Princeton University in January, February, and May 2024. ZG would like to thank Princeton University for its hospitality. Finally, we would like to thank Richard Hain for mentioning our work and method at the conference “The Ceresa Cycle in Arithmetic and Geometry” at ICERM in May, 2024.
Ziyang Gao received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n 945714). Shouwu Zhang is supported by an NSF award, DMS-2101787.
2. Period map associated with normal functions
Let $S$ be a smooth irreducible quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a polarized VHS on $S$ of weight $-1$. Let $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function.
The first goal of this section is to define the period map $\varphi_{\nu}$ associated with $\nu$ (when $\nu$ is clearly in context, we simply denote it by $\varphi$). This map fits into the diagram in the category of complex varieties
(2.1) ${{\mathscr{D}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${\Gamma\backslash{\mathscr{D}}}$
where
  1. (i)
[…]
  2. (ii)
    $\Gamma$ is a suitable arithmetic subgroup of $\mathbf{G}({\mathbb{Q}})$.
We emphasize that $V$ is, in general, not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$, and its geometric meaning will be given in Remark 2.2.
[…]
2.1. Universal period map to the classifying space
Recall that $\nu$ induces a VMHS $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ on $S$ of weight $-1$ and 00 which is graded-polarized (better, admissible), fitting into the short exact sequence in the category of graded-polarized VMHS $0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{E}}_{\nu}\rightarrow{\mathbb{Z}}(0)_{S}\rightarrow 0$ with the canonical polarization on ${\mathbb{Z}}(0)_{S}$.
Let $u_{S}\colon\tilde{S}\rightarrow S^{\mathrm{an}}$ be the universal covering map. Then the pullback $u_{S}^{*}{\mathbb{E}}_{\nu}$ (resp. $u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}$) is canonically trivialized as a local system $u_{S}^{*}{\mathbb{E}}_{\nu}\cong\tilde{S}\times E_{\nu}$ (resp. $u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}\cong\tilde{S}\times M_{-1,{\mathbb{Z}}}$), with $E_{\nu}=H^{0}(\tilde{S},u_{S}^{*}{\mathbb{E}}_{\nu})$ (resp. with $M_{-1,{\mathbb{Z}}}=H^{0}(\tilde{S},u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}})$). Then for each $s\in S({\mathbb{C}})$, the fiber ${\mathbb{E}}_{\nu,s}$ can be canonically identified with $E_{\nu}$ and ${\mathbb{V}}_{{\mathbb{Z}},s}$ can be canonically identified with $M_{-1,{\mathbb{Z}}}$. The pullback under $u_{S}$ of the short exact sequence of VMHS defined by $\nu$ becomes split after $\otimes{\mathbb{Q}}$ since $\tilde{S}$ is simply connected.
Each $\tilde{s}\in\tilde{S}$ gives rise to a polarized mixed Hodge structure $(E_{\nu,{\mathbb{Q}}},(W_{\bullet})_{\tilde{s}},({\mathscr{F}}_{{\mathbb{E}}}^{\bullet})_{\tilde{s}})$ on $E_{\nu,{\mathbb{Q}}}$, And this induces the universal period map
$\tilde{\varphi}=\tilde{\varphi}_{\nu}\colon\tilde{S}\longrightarrow{\mathscr{M}}$
[…]
(2.2) ${{\mathscr{M}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${\tilde{S}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varphi}}$$\scriptstyle{\tilde{\varphi}_{0}}$${{\mathscr{M}}_{0}.}$
2.2. Construction of the Mumford–Tate domain
In practice, we need to refine this period map and replace the classifying space ${\mathscr{M}}$ by the smallest Mumford–Tate domain ${\mathscr{D}}$, which contains $\tilde{\varphi}(\tilde{S})$. It is constructed as follows.
[…]
Fix $\tilde{s}\in\tilde{S}$. Define
(2.3) ${\mathscr{D}}:=\mathbf{G}({\mathbb{R}})^{+}\cdot\tilde{\varphi}(\tilde{s})\subseteq{\mathscr{M}}.$
[…]
Finally, since $S$ is a quasi-projective variety, there exists an arithmetic subgroup $\Gamma$ of $\mathbf{G}({\mathbb{Q}})$ such that $\tilde{\varphi}$ descends to a morphism $\varphi\colon S^{\mathrm{an}}\rightarrow\Gamma\backslash{\mathscr{D}}$ fitting into the diagram (2.1).
2.3. Setup for o-minimality
Let $\mathfrak{F}\subseteq{\mathscr{D}}$ be a fundamental set for the quotient $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$, i.e. $u|_{\mathfrak{F}}$ is surjective and $(u|_{\mathfrak{F}})^{-1}(\bar{x})$ is finite for each $\bar{x}\in\Gamma\backslash{\mathscr{D}}$. If $\mathfrak{F}$ is a semi-algebraic subset of ${\mathscr{D}}$, then we have a semi-algebraic structure on $\Gamma\backslash{\mathscr{D}}$ induced by $u|_{\mathfrak{F}}$.
By the main result of [BBKT24], there exists a semi-algebraic fundamental set $\mathfrak{F}\subseteq{\mathscr{D}}$ for the quotient $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$ with the following properties: $\varphi$ is ${\mathbb{R}}_{\mathrm{an},\exp}$-definable for the semi-algebraic structure on $\Gamma\backslash{\mathscr{D}}$ defined by $\mathfrak{F}$.
2.4. Relating the Betti foliation and the period map
Recall the projection to the pure part $p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}\subseteq{\mathscr{M}}_{0}$ from (2.2) and the semi-algebraic structure ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ as in Proposition B.3.(i). The diagram (2.1) can be complete into:
[…]
and $\varphi_{0}=p\circ\varphi$ is the period map for the VHS ${\mathbb{V}}_{{\mathbb{Z}}}\rightarrow S$.
For each $\tilde{s}\in\tilde{S}$, write $\tilde{s}_{V}$ for the image $\tilde{\varphi}(\tilde{s})$ under the natural projection ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}\rightarrow V({\mathbb{R}})$. This gives a map $\tilde{\varphi}_{V}\colon\tilde{S}\rightarrow V({\mathbb{R}})$ sending $\tilde{s}\mapsto\tilde{s}_{V}$. Notice that
(2.5) $\tilde{\varphi}=(\tilde{\varphi}_{V},\tilde{\varphi}_{0})\colon\tilde{S}\longrightarrow V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}.$
Recall $S_{{\mathscr{F}}}(t)=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ and define $S_{\mathrm{rk}}(t)=\{s\in S({\mathbb{C}}):\dim\nu_{\mathrm{Betti},s}(T_{s}S)\leq t\}$ for $\nu_{\mathrm{Betti},s}\colon T_{s}S\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ defined in (1.3).
Lemma 2.1.
For each $t\geq 0$, we have
$u_{S}^{-1}\left(S_{{\mathscr{F}}}(t)\right)$ =$=\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}\tilde{\varphi}^{-1}(\{\tilde{s}_{V}\}\times{\mathscr{D}}_{0})\geq t\}$
(2.6) $=\bigcup\nolimits_{r\geq 0}\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}\tilde{\varphi}^{-1}(\tilde{\varphi}(\tilde{s}))=r,~{}\{\tilde{s}_{V}\}\times\tilde{C}\subseteq\tilde{\varphi}(\tilde{S})\text{ for }$
[…]
and
(2.7) $u_{S}^{-1}\left(S_{\mathrm{rk}}(t)\right)=\{\tilde{s}\in\tilde{S}:\mathrm{rank}(\mathrm{d}\tilde{\varphi}_{V})_{\tilde{s}}\leq t\}.$
Notice that in the union in (2.1), the second condition $\{\tilde{s}_{V}\}\times\tilde{C}\subseteq\tilde{\varphi}(\tilde{S})$ always holds true for $r>t$.
Proof.
Consider $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$. Each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lies in ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ for $s=\pi(x)$, which is canonically isomorphic to $\mathrm{Ext}_{\mathrm{MHS}}({\mathbb{Z}}(0),{\mathbb{V}}_{{\mathbb{Z}},s})$ by Carlson [Car85]. Hence each $x$ gives rise to a ${\mathbb{Z}}$-mixed Hodge structure of weight $-1$ and 00, and for the universal covering map $u_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ we obtain a period map $\tilde{\varphi}_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}\rightarrow{\mathscr{M}}$. The map $\nu\circ u_{S}\colon\tilde{S}\rightarrow S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lifts to $\tilde{\nu}\colon\tilde{S}\rightarrow\tilde{{\mathscr{J}}}$, and $\tilde{\varphi}=\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu}$.
[…]
The zero section of $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ gives rise to a Levi decomposition of $\mathbf{G}^{{\mathscr{M}}}$ and hence an identification ${\mathscr{M}}=M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}$ as in Proposition B.3.(i) applied to ${\mathscr{M}}$, and $\tilde{\varphi}_{{\mathscr{J}}}$ becomes
(2.8) $\tilde{\varphi}_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}=M_{-1}({\mathbb{R}})\times\tilde{S}\xrightarrow{(1,\tilde{\varphi}_{0})}M_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}.$
By the last paragraph, the leaves of ${\mathscr{F}}_{\mathrm{Betti}}$ are precisely $u_{{\mathscr{J}}}\left(\tilde{\varphi}_{{\mathscr{J}}}^{-1}(\{a\}\times{\mathscr{M}}_{0})\right)$ for all $a\in M_{-1}({\mathbb{R}})$.
Thus $u_{{\mathscr{J}}}^{-1}\left(\nu(S_{{\mathscr{F}}}(t))\right)=\{\tilde{s}^{\prime}\in\tilde{\nu}(\tilde{S}):\dim_{\tilde{s}^{\prime}}\tilde{\varphi}_{{\mathscr{J}}}^{-1}(\{\tilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}$, where $\tilde{s}^{\prime}_{V}$ is the image of $\tilde{s}\in\tilde{{\mathscr{J}}}\xrightarrow{\tilde{\varphi}_{{\mathscr{J}}}}M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}\rightarrow M_{-1}({\mathbb{R}})$ with the last map being the natural projection. Applying $\tilde{\nu}^{-1}$ (whose fibers are of dimension 00 because $\nu$ is injective) to the set above and noticing that $u_{{\mathscr{J}}}\circ\tilde{\nu}=\nu\circ u_{S}$, we have
$u_{S}^{-1}\left(S_{{\mathscr{F}}}(t)\right)=\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}(\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu})^{-1}(\{\tilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}.$
Hence the first equality in (2.1) holds true because $\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu}=\tilde{\varphi}$ and by Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$. Similarly, we get (2.7).
The second equality in (2.1) clearly holds. Hence, we are done. ∎
Remark 2.2.
The proof of Lemma 2.1 (with Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$ and the identification $\tilde{{\mathscr{J}}}=M_{-1}({\mathbb{R}})\times\tilde{S}$) also yields the following assertion: ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})\subseteq{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is the intermediate Jacobian of a sub-VHS translated by a torsion multisection; it contains $\nu(S)$ and is the minimal one containing $\nu(S)$ with respect to inclusion. The relative dimension $\dim{\mathscr{J}}_{\nu}-\dim S$ equals $\frac{1}{2}\dim V$.
Similarly, $V$ equals ${\mathbb{V}}^{\prime}_{{\mathbb{Q}},s}$ for any $s\in S({\mathbb{C}})$, where ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ is the largest sub-VHS of ${\mathbb{V}}_{{\mathbb{Z}}}$ such that $\nu$ is torsion under the projection ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}/{\mathbb{V}}^{\prime}_{{\mathbb{Z}}})$.
Moreover for the natural projection $\varphi_{{\mathscr{J}}_{\nu}}\colon{\mathscr{J}}_{\nu}\rightarrow\Gamma\backslash{\mathscr{D}}$, we have $\varphi(S)=\varphi_{{\mathscr{J}}_{\nu}}(\nu(S))$.
Since $\tilde{\varphi}_{V}$ factors through $\tilde{\varphi}$ and has target $V({\mathbb{R}})$, (2.7) immediately yields the following trivial upper bound
(2.9) $\nu_{\mathrm{Betti},s}(T_{s}S)\leq\min\left\{\dim\varphi(S),\frac{1}{2}\dim V\right\}\quad\text{ for all }s\in S({\mathbb{C}}).$
3. The Betti rank and Zariski closedness of the rank-strata
The goal of this section is to prove one of the main theorems of this paper. We will prove a formula to compute the Betti rank $r(\nu)$, and in the process, we show that the Betti foliation on the intermediate Jacobian defines Zariski closed strata.
Let $S$ be a smooth irreducible quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a polarized VHS on $S$ of weight $-1$. Let $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function.
[…]
$r(\nu)=\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S).$
Retain the notation from (2.1). In particular, we have the period map $\varphi=\varphi_{\nu}\colon S\rightarrow\Gamma\backslash{\mathscr{D}}$ for the (mixed) Mumford–Tate domain ${\mathscr{D}}$, the ${\mathbb{Q}}$-group $\mathbf{G}$ and its unipotent radical $V$ which is a vector group. We emphasize that $V$ is, in general, not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$.
The main result of this section is the following formula for $r(\nu)$. The advantage is that it is often computable in practice.
Theorem 3.1.
The Betti rank is given by the following formula:
(3.1) $r(\nu)=\min_{N}\left\{\dim\varphi_{/N}(S)+\frac{1}{2}\dim_{{\mathbb{Q}}}(V\cap N)\right\},$
where $N$ runs through the set of normal subgroups of $\mathbf{G}$, and $\varphi_{/N}$ is the induced map
[…]
Taking $N=\{1\}$ and $N=\mathbf{G}$, we recover the trivial upper bound (2.9) of $r(\nu)$.
We also show that the Betti foliation defines Zariski’s closed strata on $S$.
Theorem 3.2.
For each $t\geq 0$, the set $S_{{\mathscr{F}}}(t):=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ is Zariski closed in $S$. In particular, $r(\nu)=\dim S-\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}$.
The proofs of Theorem 3.1 and Theorem 3.2 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the case when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is polarizable. A key ingredient for our proof is the mixed Ax-Schanuel theorem, which is used multiple times. In [Gao20a], the version for universal abelian varieties [Gao20b, Thm. 1.1] was used. In the current paper, we need the version for admissible VMHS independently proved by Chiu [Chi21] and Gao–Klingler [GK24]. We also invoke [BBT23] on the algebraicity of $\varphi(S)$, whose proof builds up on o-minimal GAGA, for two reasons. First, it eases the notation for the proof. Second and more importantly, [BBT23] is necessary to prove the full version of Theorem 1.6. Still, we point out that [BBT23] is not needed in the case which we care the most about in this paper, i.e. $r(\nu)=\dim S$ if and only if the RHS of (1.6) equals $\dim S$ because we can easily reduce the proof to the case where every fiber of $\varphi$ has dimension 00.
Another crucial input for the proof is the o-minimal structure associated with the period map [BBKT24], which we will recall in §$\S$2.3. This allows us to apply (o-minimal) definable Chow.
3.1. Replacing $S$ by $\varphi(S)$
We shall replace $S$ by $\varphi(S)$ in the proof using [BBT23]. This largely eases the notation.
[…]
${S^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Gamma\backslash{\mathscr{D}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[p]}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi_{0}}$$\scriptstyle{\varphi}$${\Gamma_{0}\backslash{\mathscr{D}}_{0}}$
with the dotted arrow being the restriction $[p]|_{S^{\prime}}$.
Recall from Remark 2.2 that ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})$ is an intermediate Jacobian over $S$. Set ${\mathscr{J}}^{\prime}:=S^{\prime}\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})$; it is an intermediate Jacobian over $S^{\prime}$. Then the inclusion $S^{\prime}\subseteq\Gamma\backslash{\mathscr{D}}$ yields a section $\nu^{\prime}$ of ${\mathscr{J}}^{\prime}\rightarrow S^{\prime}$, and thus we can define $S^{\prime}_{{\mathscr{F}}}(t)$ for each $t\geq 0$ with respect to $\nu^{\prime}$. We have the following commutative diagram, with $\varphi_{{\mathscr{J}}}$ induced by $\varphi$, such that $\varphi_{{\mathscr{J}}}\circ\nu=\nu^{\prime}\circ\varphi$:
[…]
For each $r\geq 0$, denote by
(3.2) $S_{\geq r}:=\{s\in S({\mathbb{C}}):\dim_{s}\varphi^{-1}(\varphi(s))\geq r\}.$
It is a closed algebraic subset of $S$ by upper semi-continuity.
By (the proof of) Lemma 2.1, more precisely the second equality of (2.1), we have
(3.3) $S_{{\mathscr{F}}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime}_{{\mathscr{F}}}(t-r)\right).$
This equality allows us to replace $S$ by $\varphi(S)$ to study the Betti rank and the Betti strata.
3.2. Bi-algebraic system and Ax–Schanuel
From now on, in the whole section, we replace $S$ by $\varphi(S)$ and view $S$ as an algebraic subvariety of the complex analytic space $\Gamma\backslash{\mathscr{D}}$, unless otherwise stated.
[…]
Proposition 3.3.
Let ${\mathscr{D}}_{N}$ be a weak Mumford–Tate domain. Then $u({\mathscr{D}}_{N})\cap S$ is a closed algebraic subset of $S$.
Definition 3.4.
  1. (i)
    Let $\tilde{Y}\subseteq{\mathscr{D}}$ be a complex analytic irreducible subset. The weakly special closure of $\tilde{Y}$, denoted by $\tilde{Y}^{\mathrm{ws}}$, is the smallest weak Mumford–Tate domain in ${\mathscr{D}}$ which contains $\tilde{Y}$.
  2. (ii)
    Let $Y\subseteq S$ be an irreducible subvariety. The weakly special closure of $Y$, denoted by $Y^{\mathrm{ws}}$, is $u(\widetilde{Y}^{\mathrm{ws}})$ for one (hence any) complex analytic irreducible component $\widetilde{Y}$ of $u^{-1}(Y)$.
The following Ax–Schanuel theorem for VMHS was independently proved by Chiu [Chi21] and Gao–Klingler [GK24]. We refer to Definition B.2 for the algebraic structure on ${\mathscr{D}}$.
Theorem 3.5 (weak Ax–Schanuel for VMHS).
Let $\tilde{Z}\subseteq u^{-1}(S)$ be a complex analytic irreducible subset. Then
(3.4) $\dim\tilde{Z}^{\mathrm{Zar}}+\dim u(\tilde{Z})^{\mathrm{Zar}}\geq\dim\tilde{Z}^{\mathrm{ws}}+\dim\tilde{Z},$
where $\tilde{Z}^{\mathrm{ws}}$ is the smallest weak Mumford–Tate domain which contains $\tilde{Z}$.
Proof.
Let $Y:=u(\tilde{Z})^{\mathrm{Zar}}$. Let $\mathscr{Z}:=\{(z,y)\in\tilde{Z}\times Y({\mathbb{C}}):u(z)=y\}$, then $\mathscr{Z}$ is a complex analytic irreducible subset of ${\mathscr{D}}\times_{\Gamma\backslash{\mathscr{D}}}Y^{\prime}$. The Zariski closure of $\mathscr{Z}$ in ${\mathscr{D}}\times Y$ is contained in $\tilde{Z}^{\mathrm{Zar}}\times Y$, and $\dim\mathscr{Z}=\dim\tilde{Z}$. Then (3.4) is a direct consequence of the mixed Ax–Schanuel theorem [GK24, Thm. 1.1] applied to $\mathscr{Z}$. ∎
We close this introductory subsection with the following definition. In practice, we often need to work with algebraic subvarieties $Y\subseteq S$, which are not weak Mumford–Tate domains, and the following number measures how far it is from being one.
(3.5) $\delta_{\mathrm{ws}}(Y):=\dim Y^{\mathrm{ws}}-\dim Y.$
If we do not replace $S$ by $\varphi(S)$, then each $Y$ on the right-hand side should be replaced by $\varphi(Y)$.
Definition 3.6.
An irreducible algebraic subvariety $Y$ of $S$ is called weakly optimal if the following holds true: $Y\subsetneq Y^{\prime}\subseteq S\Rightarrow\delta_{\mathrm{ws}}(Y)<\delta_{\mathrm{ws}}(Y^{\prime})$, for any $Y^{\prime}\subseteq S$ irreducible.
3.3. Applications of Ax–Schanuel
Retain the notation in (2.4). We start with the following application of mixed Ax–Schanuel.
Proposition 3.7.
For each $t\geq 0$, $S_{{\mathscr{F}}}(t)$ is contained in the union of weakly optimal subvarieties $Y\subseteq S$ satisfying
(3.6) $\dim Y\geq\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})+t.$
Proof.
It suffices to prove two things:
  1. (i)
    $S_{{\mathscr{F}}}(t)$ is covered by the union of irreducible subvarieties $Y\subseteq S$ satisfying (3.6) (without requiring $Y$ to be weakly optimal);
  2. (ii)
    If $Y\subseteq S$ is an irreducible subvariety satisfying (3.6) and is maximal for this property with respect to inclusions, then $Y$ is weakly optimal.
Let us prove (i). By Lemma 2.1, more precisely the first equality of (2.1), $S_{{\mathscr{F}}}(t)$ is covered by irreducible subvarieties $Y\subseteq S$ such that
$Y:=\overline{u(\{a\}\times\tilde{C})}^{\mathrm{Zar}},$
for some complex analytic irreducible $\tilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\tilde{C}=t$ and some $a\in V({\mathbb{R}})$.
Apply mixed Ax–Schanuel in this context (Theorem 3.5 to $\{a\}\times\tilde{C}$). Then we get
$\dim\overline{\{a\}\times\tilde{C}}^{\mathrm{Zar}}+\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}+t.$
By Lemma B.4, $\overline{\{a\}\times\tilde{C}}^{\mathrm{Zar}}=\{a\}\times\overline{\tilde{C}}^{\mathrm{Zar}}$. Hence
$\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\overline{\tilde{C}}^{\mathrm{Zar}}+t\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\tilde{C}^{\mathrm{ws}}+t.$
The last inequality holds because $\overline{\tilde{C}}^{\mathrm{Zar}}\subseteq\tilde{C}^{\mathrm{ws}}$. So
$\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\tilde{C}^{\mathrm{ws}}+t.$
Now, to prove (i), it suffices to prove $\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$ and $\dim\tilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$.
Let us prove $u((\{a\}\times\tilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$; the upshot is $\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$. By definition of $Y$, we have $u(\{a\}\times\tilde{C})\subseteq Y$. Hence $u((\{a\}\times\tilde{C})^{\mathrm{ws}})\subseteq Y^{\mathrm{ws}}$. On the other hand, $u((\{a\}\times\tilde{C})^{\mathrm{ws}})$ is closed algebraic by Proposition 3.3, so $Y\subseteq u((\{a\}\times\tilde{C})^{\mathrm{ws}})$. So $Y^{\mathrm{ws}}\subseteq u((\{a\}\times\tilde{C})^{\mathrm{ws}})$. Now we have established $u((\{a\}\times\tilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$.
Similarly, we have $\dim\tilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$. Hence, we are done for (i).
For (ii), let $Y\subseteq Y^{\prime}\subseteq X$. Assume $\delta_{\mathrm{ws}}(Y)\geq\delta_{\mathrm{ws}}(Y^{\prime})$, i.e.
$\dim Y^{\mathrm{ws}}-\dim Y\geq\dim Y^{\prime,\mathrm{ws}}-\dim Y^{\prime}.$
The assumption on $Y$ implies $\dim Y^{\mathrm{ws}}-\dim Y\leq\dim[p](Y^{\mathrm{ws}})-t$. Combined with the inequality above, we obtain $\dim Y^{\prime,\mathrm{ws}}-\dim Y^{\prime}\leq\dim[p](Y^{\prime,\mathrm{ws}})-t$ because $Y\subseteq Y^{\prime}$. Therefore $Y=Y^{\prime}$ by maximality of $Y$. Hence, (ii) is established. ∎
Next, we state a finiteness proposition for weakly optimal subvarieties in $S$ la Ullmo, which we give a proof using twice mixed Ax–Schanuel in the next section. When ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is polarizable, i.e. in the mixed Shimura case, this is [Gao20b, Thm. 1.4].
Proposition 3.8.
There exist finitely many pairs $({\mathscr{D}}^{\prime}_{1},N_{1}),\ldots,({\mathscr{D}}^{\prime}_{k},N_{k})$, with each ${\mathscr{D}}^{\prime}_{j}$ a Mumford–Tate domain contained in ${\mathscr{D}}$ and $N_{j}$ a normal subgroup of $\mathrm{MT}({\mathscr{D}}^{\prime}_{j})$, such that the following holds true. For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}$ is the image of an $N_{j}({\mathbb{R}})^{+}$-orbit contained in ${\mathscr{D}}^{\prime}_{j}$ under $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$ for some $j\in\{1,\ldots,k\}$.
The same statement (with $Y^{\mathrm{ws}}$ replaced by $\varphi(Y)^{\mathrm{ws}}$) still holds true without replacing $S$ by $\varphi(S)$.
Denote by $\Gamma_{j}=\Gamma\cap\mathrm{MT}({\mathscr{D}}^{\prime}_{j})({\mathbb{Q}})$ and $\Gamma_{j,/N_{j}}=\Gamma_{j}/(\Gamma_{j}\cap N_{j}({\mathbb{Q}}))$. Then equivalently, each such $Y^{\mathrm{ws}}$ is a fiber of the quotient $[p_{N_{j}}]\colon u({\mathscr{D}}^{\prime}_{j})=\Gamma_{j}\backslash{\mathscr{D}}^{\prime}_{j}\rightarrow\Gamma_{j,/N_{j}}\backslash({\mathscr{D}}^{\prime}_{j}/N_{j})$.
[…]
3.4. Proof of Theorem 3.1
For each $j\in\{1,\ldots,k\}$, Proposition 3.3 says that $u({\mathscr{D}}^{\prime}_{j})\cap S$ is a closed algebraic subset of $S$. The restriction
$[p_{N_{j}}]|_{S}\colon u({\mathscr{D}}^{\prime}_{j})\cap S\rightarrow\Gamma_{j,/N_{j}}\backslash({\mathscr{D}}^{\prime}_{j}/N_{j})$
[…]
For each $t\geq 0$, the subset
(3.7) $E_{j}(t):=\left\{s\in u({\mathscr{D}}^{\prime}_{j})\cap S:\dim_{s}[p_{N_{j}}]|_{S}^{-1}([p_{N_{j}}](s))\geq\frac{1}{2}\dim(V\cap N_{j})+t\right\}$
is both definable and complex analytic in $u({\mathscr{D}}^{\prime}_{j})\cap S$. Hence, $E_{j}(t)$ is algebraic by definable Chow. Moreover, it is closed in $u({\mathscr{D}}^{\prime}_{j})\cap S$ by the upper semi-continuity of fiber dimensions. So $E_{j}(t)$ is a closed algebraic subset of $S$.
Proposition 3.9.
For each $t\geq 0$, we have
$S_{{\mathscr{F}}}(t)\subseteq\bigcup\nolimits_{j=1}^{k}E_{j}(t).$
Proof.
Let $t\geq 0$. By Proposition 3.7, $S_{{\mathscr{F}}}(t)$ is covered by weakly optimal $Y\subseteq S$ such that $\dim Y\geq\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})+t$. Then by Proposition 3.8, $Y^{\mathrm{ws}}$ is a fiber $[p_{N_{j}}]$ for some $j\in\{1,\ldots,k\}$, and hence $\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})=\frac{1}{2}\dim(V\cap N_{j})$. So $\dim Y\geq\frac{1}{2}\dim(V\cap N_{j})+t$. So $Y\subseteq E_{j}(t)$ because $[p_{N_{j}}](Y)$ is a point. ∎
Now we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
[…]
Let us prove $\leq$”. By (2.7) we have $r(\nu)=\max_{\tilde{s}\in\tilde{S}}(\mathrm{d}\tilde{\varphi}_{V})_{\tilde{s}}$. Hence $r(\nu)\leq\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)$ for any normal subgroup $N\lhd\mathbf{G}$.
Let us prove $\geq$”. Let $t=\dim S-r(\nu)$. Then $S_{{\mathscr{F}}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. By (3.3) and Proposition 3.9 (which should be applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
$S_{{\mathscr{F}}}(t)\subseteq S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
Each $S_{\geq r}$ is Zariski closed in $S$, and each $E_{j}(t-r)$ is Zariski closed in $S^{\prime}=\varphi(S)$. Hence, each member in the union on the right-hand side is Zariski closed in $S$. Taking the Zariski closure of both sides, we then have $S$ equal to a member on the right-hand side.
If $S=S_{\geq t}$, then $\geq$” holds true already for $N=\{1\}$.
Assume $S=S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right)$ for some $10\leq r\leq t-$ 1 and some $j$. Then $S=S_{\geq r}=\varphi^{-1}\left(E_{j}(t-r)\right)$, So each fiber of $\varphi$ has dimension $\geq r$, and $S^{\prime}=E_{j}(t-r)$. Moreover, $\mathrm{MT}({\mathscr{D}}^{\prime}_{j})=\mathbf{G}$ because $S^{\prime}=\varphi(S)$ is Hodge generic in $\Gamma\backslash{\mathscr{D}}$. Set $N=N_{j}$. Each fiber of the map
[…]
has ${\mathbb{C}}$-dimension $\geq r+\left(\frac{1}{2}\dim(V\cap N)+(t-r)\right)=\frac{1}{2}\dim(V\cap N)+t$ by definition of $E_{j}(t-r)$. So
$r(\nu)=\dim S-t\geq\left(\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)+t\right)-t=\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N).$
So, $\geq$” is established. ∎
3.5. Zariski closedness of the degeneracy loci
We start with the following lemma, which is the converse of Proposition 3.9.
Lemma 3.10.
For each $t\geq 0$ and each $j\in\{1,\ldots,k\}$, we have $E_{j}(t)\subseteq S_{{\mathscr{F}}}(t)$.
Proof.
Fix $j$. Denote by $\mathbf{H}_{j}=\mathrm{MT}({\mathscr{D}}^{\prime}_{j})$, $V_{j}:=V\cap\mathbf{H}_{j}$, and $\mathbf{H}_{j,0}:=\mathbf{H}_{j}/V_{j}$. Under the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), we have ${\mathscr{D}}^{\prime}_{j}=(V_{j}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime}_{j})$ by Lemma B.5 (applied to ${\mathscr{D}}^{\prime}_{j}\subseteq{\mathscr{D}}$).
Because $N_{j}\lhd\mathbf{H}_{j}$, we have: (i) $V\cap N_{j}=V_{j}\cap N_{j}$ is a $\mathbf{H}_{j,0}$-module; (ii) the action of $p(N_{j})\lhd\mathbf{H}_{j,0}$ on $V_{j}/(V_{j}\cap N_{j})$ is trivial. Let $x\in{\mathscr{D}}^{\prime}_{j}$. Under ${\mathscr{D}}^{\prime}_{j}=(V_{j}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime}_{j})$, write $x=(v,x_{0})$. Then $N_{j}({\mathbb{R}})^{+}x$ becomes $((V\cap N_{j})({\mathbb{R}})+v)\times p(N_{j})({\mathbb{R}})^{+}x_{0}$. Notice that this $v\in V({\mathbb{R}})$ is fixed.
For each $s\in E_{j}(t)$, by definition there exist an irreducible $\tilde{Y}\subseteq u^{-1}(S)\cap{\mathscr{D}}^{\prime}_{j}$ such that $s\in u(\tilde{Y})$, $\dim\tilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, and that $\tilde{Y}$ is contained in a fiber of the quotient ${\mathscr{D}}^{\prime}_{j}\rightarrow{\mathscr{D}}^{\prime}_{j}/N_{j}$. The last condition implies that $\tilde{Y}\subseteq N_{j}({\mathbb{R}})^{+}x$ for some $x\in{\mathscr{D}}^{\prime}_{j}$. Hence by the discussion above, $\tilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$ for a fixed $v\in V({\mathbb{R}})$. Now that $\dim_{{\mathbb{C}}}\tilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, the following property holds true: For each $(a,x_{0})\in\tilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$, there exists a complex analytic subset $\tilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\tilde{C}\geq t$ such that $\{a\}\times\tilde{C}\subseteq\tilde{Y}$. Hence $s\in S_{{\mathscr{F}}}(t)$ by Lemma 2.1 (more precisely, the first equality in (2.1)). Now, the conclusion of the lemma holds as $s$ runs over $E_{j}(t)$. ∎
Now we are ready to prove Theorem 3.2.
Proof of Theorem 3.2.
[…]
By (3.3), Proposition 3.9 and Lemma 3.10 (both applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
(3.8) $S_{{\mathscr{F}}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
So $S_{{\mathscr{F}}}(t)$ is Zariski closed in $S$ because each member in the union on the right-hand side is. The “In particular” part is easy to check once we have established the Zariski closedness of $S_{{\mathscr{F}}}(t)$. ∎
4. Proof of the finiteness result à la Ullmo
The goal of this section is to prove Proposition 3.8, the finiteness result regarding weakly optimal subvarieties for admissible VMHS of weight $-1$ and 00.
Before moving on to the proof, let us take a step back and look at this proposition from a historical point of view. In studying the André–Oort conjecture, Ullmo [Ull14, Thm. 4.1] proved this finiteness result for maximal weakly special subvarieties (a particular kind of weakly optimal subvarieties) in the case of pure Shimura varieties as an application of the pure Ax–Lindemann theorem (a special case of Ax–Schanuel). The finiteness is ultimately obtained by the following fact: any countable set which is definable in an o-minimal structure is finite.
[…]
Recently, Baldi–Urbanik [BU24, Thm. 7.1] proved Proposition 3.8 for general admissible VMHS in a different way (but also uses Ax–Schanuel as a core). Weakly optimal subvarieties are called monodromically atypical maximal, and Proposition 3.8 is called Geometric Zilber–Pink as in [BKU24]. Their proof does not use o-minimality, and gives some effective results.
Retain the notation in (2.4). In this section, we also replace $S$ by $\varphi(S)$ to ease notation. The proof also works without doing so, except that when we apply Ax–Schanuel for Theorem 4.2, we need to combine it with [BBT23].
4.1. Zariski optimal subsets and an application of mixed Ax–Schanuel
Definition 4.1.
A complex analytic irreducible subset $\tilde{Y}$ of $u^{-1}(S)$ is called Zariski optimal if the following holds true: $\tilde{Y}\subsetneq\tilde{Y}^{\prime}\subseteq u^{-1}(S)\Rightarrow\delta_{\mathrm{Zar}}(\tilde{Y})<\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$, with $\tilde{Y}^{\prime}$ complex analytic irreducible.
The following theorem is, in fact, an equivalent statement to the weak mixed Ax–Schanuel theorem (Theorem 3.5). In this paper, we only need one deduction.
Theorem 4.2.
Assume $\tilde{Y}\subseteq u^{-1}(S)$ is Zariski optimal. Then $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ and $\tilde{Y}$ is a complex analytic irreducible component of $\tilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$.
Proof.
Let $\tilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal. Then $\tilde{Y}$ is a complex analytic irreducible component of $\tilde{Y}^{\mathrm{Zar}}\cap u^{-1}(S)$. Hence it suffices to prove $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$.
Let $\tilde{Y}^{\prime}$ be a complex analytic irreducible component of $\tilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$ which contains $\tilde{Y}$. Then $u(\tilde{Y}^{\prime})$ is closed algebraic in $S$ because $u(\tilde{Y}^{\mathrm{ws}})\cap S$ is a closed algebraic subset of $S$ by Proposition 3.3.
Assume $\tilde{Y}\not=\tilde{Y}^{\prime}$. Then $\delta_{\mathrm{Zar}}(\tilde{Y})<\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$ by the Zariski optimality of $\tilde{Y}$. So
$\dim\tilde{Y}^{\mathrm{Zar}}-\dim\tilde{Y}<\dim\tilde{Y}^{\prime\mathrm{Zar}}-\dim\tilde{Y}^{\prime}\leq\dim\tilde{Y}^{\mathrm{ws}}-\dim\tilde{Y}^{\prime}.$
Now that $u(\tilde{Y})\subseteq u(\tilde{Y}^{\prime})$, we have $u(\tilde{Y})^{\mathrm{Zar}}\subseteq u(\tilde{Y}^{\prime})$ because the right hand side is closed algebraic in $S$. So
$\dim u(\tilde{Y})^{\mathrm{Zar}}\leq\dim u(\tilde{Y}^{\prime})=\dim\tilde{Y}^{\prime}.$
These two inequalities together yield.
$\dim\tilde{Y}^{\mathrm{Zar}}+\dim u(\tilde{Y})^{\mathrm{Zar}}<\dim\tilde{Y}^{\mathrm{ws}}+\dim\tilde{Y}.$
This contradicts the weak mixed Ax–Schanuel theorem (Theorem 3.5) for $\tilde{Y}$.
Hence we must have $\tilde{Y}=\tilde{Y}^{\prime}$. In particular, $u(\tilde{Y})$ is algebraic. So $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$. ∎
For $Y\subseteq S$ irreducible, let $\tilde{Y}$ be a complex analytic irreducible component of $u^{-1}(Y)$.
Corollary 4.3.
If $Y\subseteq X$ is weakly optimal, then $\tilde{Y}$ is Zariski optimal in $u^{-1}(S)$.
Proof.
Let $\tilde{Y}^{\prime}\supseteq\tilde{Y}$ be a complex analytic irreducible subset of $u^{-1}(S)$ such that $\delta_{\mathrm{Zar}}(\tilde{Y})\geq\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$. We may and do assume that $\tilde{Y}^{\prime}$ is Zariski optimal.
Set $Y^{\prime\prime}:=u(\tilde{Y}^{\prime})^{\mathrm{Zar}}$. Then $Y^{\prime\prime}\subseteq u(\tilde{Y}^{\prime,\mathrm{ws}})\cap S$ because $u(\tilde{Y}^{\prime,\mathrm{ws}})\cap S$ is closed and algebraic. Thus
$\delta_{\mathrm{ws}}(Y^{\prime\prime})=\dim Y^{\prime\prime,\mathrm{ws}}-\dim Y^{\prime\prime}\leq\dim\tilde{Y}^{\prime,\mathrm{ws}}-\dim\tilde{Y}^{\prime}.$
Since $\tilde{Y}^{\prime}$ is Zariski optimal, we have $\tilde{Y}^{\prime,\mathrm{Zar}}=\tilde{Y}^{\prime,\mathrm{ws}}$ by Theorem 4.2. So, the inequality above further implies
$\delta_{\mathrm{ws}}(Y^{\prime\prime})\leq\dim\tilde{Y}^{\prime,\mathrm{Zar}}-\dim\tilde{Y}^{\prime}=\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})\leq\delta_{\mathrm{Zar}}(\tilde{Y})\leq\dim\tilde{Y}^{\mathrm{ws}}-\dim\tilde{Y}=\delta_{\mathrm{ws}}(Y).$
So $Y^{\prime\prime}=Y$ because $Y$ is weakly optimal. Thus $\tilde{Y}=\tilde{Y}^{\prime}$ is Zariski optimal. ∎
4.2. A parametrization of Zariski optimal subsets
The goal of this subsection is to construct a space ${\mathscr{N}}^{0}$ which parametrizes Zariski optimal subsets of $u^{-1}(S)$.
Fix a Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$. Then for the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), the action of $\mathbf{G}({\mathbb{R}})$ on ${\mathscr{D}}$ is given by $(v,g_{0})\cdot(x_{V},x_{0})=(v+g_{0}x_{V},g_{0}x_{0})$.
[…]
Consider the set ${\mathscr{N}}$ of pairs $(x,H)$ consisting of $x\in(u|_{\mathfrak{F}})^{-1}(S)$ and $H$ a connected subgroup of $\mathbf{G}_{{\mathbb{R}}}$ with the following properties: $H_{0}:=H/(V_{{\mathbb{R}}}\cap H)$ is a semi-simple subgroup of $\mathbf{G}_{0,{\mathbb{R}}}$, and $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}_{{\mathbb{R}}}}(H)$.
For each $(x,H)\in{\mathscr{N}}$, denote by $V_{H}:=V_{{\mathbb{R}}}\cap H$. Then $V_{H}$ is the unipotent radical of $H$. It can be easily shown that each weak Mumford–Tate domain is $N({\mathbb{R}})x$ for some $(x,N_{{\mathbb{R}}})\in{\mathscr{N}}$ (with $N_{{\mathbb{R}}}$ defined over ${\mathbb{Q}}$).
Lemma 4.4.
Let $(x,H)\in{\mathscr{N}}$. Then $H({\mathbb{R}})x$ is:
  1. (i)
    an algebraic subset of ${\mathscr{D}}$;
  2. (ii)
    $(v^{\prime}+V_{H}({\mathbb{R}}))\times H_{0}({\mathbb{R}})x_{0}\subseteq V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}$ for some $v^{\prime}\in V({\mathbb{R}})$ and $x_{0}=p(x)$.
Proof.
$H({\mathbb{R}})x$ is by definition semi-algebraic, and is complex analytic since $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}_{{\mathbb{R}}}}(H)$. So (i) holds.
Any two Levi decompositions differ from the conjugation by an element of the unipotent radical. Hence there exists a $v\in V({\mathbb{R}})$ such that
$H=(v,1)(V_{H}\rtimes H_{0})(-v,1).$
Thus the condition $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}_{{\mathbb{R}}}}(H)$ implies $\mathrm{Int}(-v)(h_{x}({\mathbb{S}}))\subseteq\mathrm{N}_{\mathbf{G}_{{\mathbb{R}}}}(V_{H}\rtimes H_{0})=V^{\prime}\rtimes G_{0}^{\prime}$ for some $V^{\prime}<V_{{\mathbb{R}}}$ and $G_{0}^{\prime}<\mathbf{G}_{0,{\mathbb{R}}}$. Then $H_{0}$ acts trivially on $V^{\prime}/V_{H}$ by the normality condition.
Write $x=(x_{V},x_{0})\in V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}$, then $\mathrm{Int}(-v)(h_{x}({\mathbb{S}}))=h_{(x_{V}-v,x_{0})}({\mathbb{S}})$ by Proposition B.3.(i). So by the last paragraph $x_{V}-v\in V_{H}({\mathbb{R}})+v^{\prime\prime}$ for some $v^{\prime\prime}\in V({\mathbb{R}})$ with $H_{0}({\mathbb{R}})\cdot v^{\prime\prime}=v^{\prime\prime}$. Thus $(v,1)(V_{H}\rtimes H_{0})({\mathbb{R}})(-v,1)\cdot x=(V_{H}({\mathbb{R}})+v^{\prime\prime}+v)\times H_{0}({\mathbb{R}})x_{0}$. So (ii) holds. ∎
Define the following two functions on ${\mathscr{N}}$:
(4.1) $d\colon{\mathscr{N}}\rightarrow{\mathbb{R}},$ $\quad(x,H)\mapsto\dim_{x}\left(u^{-1}(S)\bigcap H({\mathbb{R}})x\right),$
$\delta\colon{\mathscr{N}}\rightarrow{\mathbb{R}},$ $\quad(x,H)\mapsto\dim_{x}H({\mathbb{R}})x-\dim_{x}\left(u^{-1}(S)\bigcap H({\mathbb{R}})x\right).$
Finally, we are ready to define
[…]
$H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x\Rightarrow d(x,H)>d(x,H^{\prime}),$
$H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x\Rightarrow\delta(x,H)<\delta(x,H^{\prime})\big{\}}.$
The proof of the following proposition uses Theorem 4.2 (twice) and hence mixed Ax–Schanuel.
Proposition 4.5.
The following two sets are equal:
    the set of orbits $\{H({\mathbb{R}})x:(x,H)\in{\mathscr{N}}^{0}\}$;
    $\{\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}:\tilde{Y}\subseteq u^{-1}(S)\text{ Zariski optimal, with }\tilde{Y}\cap\mathfrak{F}\not=\emptyset\}$.
Moreover, $H({\mathbb{R}})x$ is a weak Mumford–Tate domain for each $(x,H)\in{\mathscr{N}}^{0}$.
Proof.
Take $(x,H)\in{\mathscr{N}}^{0}$. We wish to prove that $H({\mathbb{R}})x$ equals $\tilde{Y}^{\mathrm{Zar}}$ for some Zariski optimal $\tilde{Y}\subseteq u^{-1}(S)$ which passes through $x\in\mathfrak{F}$.
Let $\tilde{Y}^{\prime}$ be a complex analytic irreducible component of $u^{-1}(S)\cap H({\mathbb{R}})x$ which passes through $x$ with $\dim\tilde{Y}^{\prime}=d(x,H)$.
Take $\tilde{Y}\supseteq\tilde{Y}^{\prime}$ with $\tilde{Y}\subseteq u^{-1}(S)$ complex analytic irreducible and $\delta_{\mathrm{Zar}}(\tilde{Y})\leq\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$. We may and do assume that $\tilde{Y}$ is Zariski optimal. Then $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain by Theorem 4.2, and hence $\tilde{Y}^{\mathrm{Zar}}=H^{\prime}({\mathbb{R}})x$ for some $(x,H^{\prime})\in{\mathscr{N}}$. Thus $\delta_{\mathrm{Zar}}(\tilde{Y})=\delta(x,H^{\prime})$.
Thus $\tilde{Y}^{\prime}$ is contained in some irreducible component of $H({\mathbb{R}})x\cap H^{\prime}({\mathbb{R}})x$, which by Lemma 4.4.(ii) is $H^{\prime\prime}({\mathbb{R}})x$ for some $(x,H^{\prime\prime})\in{\mathscr{N}}$. Hence $d(x,H)=d(x,H^{\prime\prime})$, so by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime\prime}({\mathbb{R}})x$. So $H({\mathbb{R}})x\subseteq H^{\prime}({\mathbb{R}})x$. But $\delta(x,H)=\delta(x,H^{\prime})$. So by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime}({\mathbb{R}})x$, and hence $H({\mathbb{R}})x=\tilde{Y}^{\mathrm{Zar}}$. Now we have proved that the first set is contained in the second set and the “Moreover” part of the proposition.
Conversely let $\tilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal with $\tilde{Y}\cap\mathfrak{F}\not=\emptyset$. By Theorem 4.2, $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain, and hence equals $H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}$. We may furthermore choose this $x$ to be a smooth point of $H({\mathbb{R}})x\cap u^{-1}(S)\cap\mathfrak{F}$; then $\dim\tilde{Y}=d(x,H)$.
We wish to prove $(x,H)\in{\mathscr{N}}^{0}$. Let us check the two properties which define ${\mathscr{N}}^{0}$ by contradiction.
Assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x$ and $d(x,H)=d(x,H^{\prime})$. Then $\dim\tilde{Y}=d(x,H^{\prime})$ and hence $\tilde{Y}\subseteq H^{\prime}({\mathbb{R}})x$ by analytic continuation. But then $\tilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$ by Lemma 4.4.(i). So $\tilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x=\tilde{Y}^{\mathrm{Zar}}$, which is impossible. So for any $(x,H^{\prime})\in{\mathscr{N}}$, we have $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x\Rightarrow d(x,H)>d(x,H^{\prime})$.
Next, assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x$ and $\delta(x,H)\geq\delta(x,H^{\prime})$. Take a complex analytic irreducible component $\tilde{Y}^{\prime}$ of $u^{-1}(S)\bigcap H^{\prime}({\mathbb{R}})x$ such that $\dim\tilde{Y}^{\prime}=d(x,H^{\prime})$. Then $\tilde{Y}\subseteq\tilde{Y}^{\prime}$ by analytic continuation. By Lemma 4.4.(i), we have $\tilde{Y}^{\prime,\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$. So $\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})\leq\delta(x,H^{\prime})\leq\delta(x,H)=\delta_{\mathrm{Zar}}(\tilde{Y})$. Hence $\tilde{Y}^{\prime}=\tilde{Y}$ because $\tilde{Y}$ is Zariski optimal. So $\delta(x,H^{\prime})=\delta(x,H)$ and $d(x,H^{\prime})=d(x,H)$, and so
$\dim H({\mathbb{R}})x=d(x,H)+\delta(x,H)=d(x,H^{\prime})+\delta(x,H^{\prime})=\dim H^{\prime}({\mathbb{R}})x.$
This contradicts $H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x$. So for any $(x,H^{\prime})\in{\mathscr{N}}$, we have $H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x\Rightarrow\delta(x,H)<\delta(x,H^{\prime})$.
This finishes the proof of $(x,H)\in{\mathscr{N}}^{0}$. So, the second set is contained in the first set. Now we are done. ∎
4.3. A finiteness result for ${\mathscr{N}}^{0}$
Let ${\mathscr{N}}^{0}$ be as defined in (4.2). In this subsection, we prove that the following finiteness result. The proof relies on o-minimality.
Proposition 4.6.
There are only finitely many subgroups $H$ of $\mathbf{G}_{{\mathbb{R}}}$ such that $(x,H)\in{\mathscr{N}}^{0}$ for some $x\in(u|_{\mathfrak{F}})^{-1}(S)$.
We start with the following classical result.
Lemma 4.7.
  1. (i)
    There exists a finite set $\Omega_{V}=\{V_{1},\ldots,V_{n}\}$ of subspaces of $V_{{\mathbb{R}}}$ with the following property: Each subspace of $V_{{\mathbb{R}}}$ equals $g_{V}V_{j}$ for some $g_{V}\in{\mathrm{GL}}(V_{{\mathbb{R}}})$ and some $j\in\{1,\ldots,n\}$.
  2. (ii)
    There exists a finite set $\Omega_{0}=\{G_{1},\ldots,G_{n}\}$ of semi-simple subgroups of $\mathbf{G}_{0,{\mathbb{R}}}$, with no compact factors, such that the following holds: Each semi-simple subgroups of $\mathbf{G}_{0,{\mathbb{R}}}$ equals $g_{0}G_{j}g_{0}^{-1}$ for some $g_{0}\in\mathbf{G}_{0}({\mathbb{R}})$ and some $j\in\{1,\ldots,n\}$.
Consider the set $\Upsilon$ consisting of elements $(x,g_{V},g_{0},V_{j},G_{j},v)\in(u|_{\mathfrak{F}})^{-1}(S)\times{\mathrm{GL}}(V_{{\mathbb{R}}})\times\mathbf{G}_{0}({\mathbb{R}})\times\Omega_{V}\times\Omega_{0}\times V({\mathbb{R}})$ satisfying the following properties:
  1. (a)
    $g_{0}G_{j}g_{0}^{-1}$ stabilizes $g_{V}V_{j}$; hence $g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1}$ is a subgroup of $\mathbf{G}_{{\mathbb{R}}}=(V\rtimes\mathbf{G}_{0})_{{\mathbb{R}}}$;
  2. (b)
    $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}({\mathbb{R}})}\left((v,1)(g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1})(-v,1)\right)$.
[…]
For each $(x,g_{V},g_{0},V_{j},G_{j},v)\in\Upsilon$, denote by $H^{(j)}_{v}:=(v,1)(g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1})(-v,1)$. Then we obtain a map that is surjective by Lemma 4.7
(4.3) $\psi\colon\Upsilon\rightarrow{\mathscr{N}},\quad(x,g_{V},g_{0},V_{j},G_{j},v)\mapsto\left(x,H^{(j)}_{v}\right).$
Composing this map with the functions $d\colon{\mathscr{N}}\rightarrow{\mathbb{R}}$ and $\delta\colon{\mathscr{N}}\rightarrow{\mathbb{R}}$ from (4.1), we obtain two functions on $\Upsilon$
$d_{\Upsilon}\colon\Upsilon\rightarrow{\mathbb{R}},$ $\quad(x,g_{V},g_{0},V_{j},G_{j},v)\mapsto\dim_{x}\left((u|_{\mathfrak{F}})^{-1}(S)\bigcap H^{(j)}_{v}({\mathbb{R}})x\right),$
$\delta_{\Upsilon}\colon\Upsilon\rightarrow{\mathbb{R}},$ $\quad(x,g_{V},g_{0},V_{j},G_{j},v)\mapsto\dim_{x}H^{(j)}_{v}({\mathbb{R}})x-\dim_{x}\left((u|_{\mathfrak{F}})^{-1}(S)\bigcap H^{(j)}_{v}({\mathbb{R}})x\right).$
The general theory of o-minimal geometry says that $x\mapsto\dim_{x}X$ is a definable function for any definable set $X$. Hence $d_{\Upsilon}$ and $\delta_{\Upsilon}$ are definable functions on $\Upsilon$.
For the subset ${\mathscr{N}}^{0}\subseteq{\mathscr{N}}$, the inverse $\psi^{-1}({\mathscr{N}}^{0})\subseteq\Upsilon$ is
$\Xi=\big{\{}\normalsize(x,g_{V}$ $,g_{0},V_{j},G_{j},v)\in\Upsilon:\text{ for any }(x,g^{\prime}_{V},g^{\prime}_{0},V_{j^{\prime}},G_{j^{\prime}},v^{\prime})\in\Upsilon,$
$H^{(j^{\prime})}_{v^{\prime}}({\mathbb{R}})x\subsetneq H^{(j)}_{v}({\mathbb{R}})x\Rightarrow d_{\Upsilon}(x,g_{V},g_{0},V_{j},G_{j},v)>d_{\Upsilon}(x,g^{\prime}_{V},g^{\prime}_{0},V_{j^{\prime}},G_{j^{\prime}},v^{\prime}),$
$H^{(j)}_{v}({\mathbb{R}})x\subsetneq H^{(j^{\prime})}_{v^{\prime}}({\mathbb{R}})x\Rightarrow\delta_{\Upsilon}(x,g_{V},g_{0},V_{j},G_{j},v)<\delta_{\Upsilon}(x,g^{\prime}_{V},g^{\prime}_{0},V_{j^{\prime}},G_{j^{\prime}},v^{\prime})\big{\}},$
which is a definable subset of $\Upsilon$ since both $d_{\Upsilon}$ and $\delta_{\Upsilon}$ are definable functions.
With these preparations, we are ready to prove Proposition 4.6.
Proof of Proposition 4.6.
Consider the map
$\begin{array}[]{rcl}\rho\colon\Xi&\rightarrow&\bigcup_{i=1}^{n}\left({\mathrm{GL}}(V_{{\mathbb{R}}})/\mathrm{Stab}_{{\mathrm{GL}}(V_{{\mathbb{R}}})}(V_{j})\right)\times\left(\mathbf{G}_{0}({\mathbb{R}})/N_{\mathbf{G}_{0}({\mathbb{R}})}(G_{j})\right)\times V({\mathbb{R}})\\ (x,g_{V},g_{0},V_{j},G_{j},v)&\mapsto&\qquad\qquad(g_{V}V_{j},g_{0}G_{j}g_{0}^{-1},v).\end{array}$
By the surjectivity of $\psi$, for any $(x,H)\in{\mathscr{N}}^{0}$, the group $H$ equals $H^{(j)}_{v}=(v,1)(g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1})(-v,1)$ for some $(g_{V}V_{j},g_{0}G_{j}g_{0}^{-1},v)\in\rho(\Xi)$.
So, it suffices to prove that $\rho(\Xi)$ is finite.
First, the map $\rho$ is clearly definable. Hence $\rho(\Xi)$ is definable.
Next, Proposition 4.5 says that $H^{(j)}_{v}({\mathbb{R}})x$ is a weak Mumford–Tate domain. Hence there exists a ${\mathbb{Q}}$-subgroup $N$ of $\mathbf{G}$ such that $N_{{\mathbb{R}}}=(v,1)(g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1})(-v,1)$. This implies that $g_{V}V_{j}=(V\cap N)_{{\mathbb{R}}}$ and $g_{0}G_{j}g_{0}^{-1}=N_{0,{\mathbb{R}}}$. Moreover, $(v,1)\left((V\cap N)\rtimes N_{0}\right)(-v,1)=N$ and hence $v\in V({\mathbb{Q}})$. So $\rho(\Xi)$ is countable since ${\mathbb{Q}}$ is countable.
Therefore, $\rho(\Xi)$ is finite because it is definable and countable. ∎
4.4. Proof of Proposition 3.8
Let $Y$ be a weakly optimal subvariety of $S$. Take a complex analytic irreducible component $\tilde{Y}$ of $u^{-1}(Y)$ such that $\tilde{Y}\cap\mathfrak{F}\not=\emptyset$. Then $\tilde{Y}$ is Zariski optimal in $u^{-1}(S)$ by Corollary 4.3. Then by Proposition 4.5, we have $\tilde{Y}^{\mathrm{ws}}=H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}^{0}$.
Hence by the finiteness result Proposition 4.6, there exist finitely many ${\mathbb{Q}}$-groups $N_{1},\ldots,N_{k}<\mathbf{G}$ satisfying the following property: For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}=u(\tilde{Y}^{\mathrm{ws}})$ equals $u(N_{j}({\mathbb{R}})^{+}x)$ for some $j\in\{1,\ldots,k\}$ and some $x\in{\mathscr{D}}$ with $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})$. Then ${\mathscr{D}}^{\prime}_{j}:=\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})^{+}x$ is a Mumford–Tate domain and is independent of the choice of such $x$. We are done. ∎
5. Application to non-degeneracy in case of irreducible VHSs
In this section, we give two applications of Theorem 3.1: when the VHS is irreducible and when it has a simple algebraic monodromy group. Both cases apply to the Gross-Shoen and the Ceresa normal functions.
Let $S$ be a smooth irreducible quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a polarized VHS on $S$ of weight $-1$. Let $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function. Then we have the associated Betti form $\beta_{\nu}$ defined in Definition D.5; it is a semi-positive $(1,1)$-form on $S$.
We wish to check whether $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$, which by Corollary D.7 becomes $r(\nu)=\dim S$ for the Betti rank $r(\nu)$ defined in (1.4). Now Theorem 3.1 gives a checkable criterion. Recall the period map $\varphi=\varphi_{\nu}\colon S\rightarrow\Gamma\backslash{\mathscr{D}}$, the ${\mathbb{Q}}$-group $\mathbf{G}$ and its unipotent radical $V$ which is a vector group from (2.1). We emphasize that $V$ is in general not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$ and its geometric meaning is given by Remark 2.2: $\frac{1}{2}\dim V$ is the relative dimension of ${\mathscr{J}}_{\nu}$, the smallest intermediate Jacobian of sub-VHS of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ translated by a torsion multisection which contains $\nu(S)$.
Notice that the trivial upper bound (2.9) yields the following necessary condition for $r(\nu)=\dim S$:
${\tt(Hyp):}\qquad\dim\varphi(S)=\dim S\leq\frac{1}{2}\dim_{{\mathbb{Q}}}V.$
Theorem 5.1.
Assume: (i) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is irreducible, i.e. the only sub-VHSs are trivial or itself; (ii) $\nu(S)$ is not a torsion section. Then
$r(\nu)=\min\left\{\dim\varphi(S),\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}\right\}$
for one (and hence for all) $s\in S({\mathbb{C}})$.
In particular, if futhermore $\dim\varphi(S)=\dim S$ and $\dim{\mathbb{V}}_{{\mathbb{Q}},s}\geq 2\dim S$, then we have $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$.
Proof.
Set $\mathbf{G}_{0}=\mathbf{G}/V$. Then $V$ is a $\mathbf{G}_{0}$-submodule of ${\mathbb{V}}_{{\mathbb{Q}},s}$ for one (and hence all) $s\in S({\mathbb{C}})$.
By (i), ${\mathbb{V}}_{{\mathbb{Q}},s}$ is irreducible as a $\mathbf{G}_{0}$-module. By (ii), $V\not=\{0\}$. Hence $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
Let $N$ be a normal subgroup of $\mathbf{G}$. Then $V\cap N$ is a $\mathbf{G}_{0}$-submodule of $V$, and hence gives rise to a sub-VHS of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$. Hence by (i), either $V\cap N=\{0\}$ or $V\cap N=V$.
Assume $V\cap N=\{0\}$. Then $N\lhd\mathbf{G}_{0}=\mathbf{G}/V$ is reductive and it acts trivially on $V=V/(V\cap N)$. Hence, $N({\mathbb{R}})^{+}x_{0}$ is a point for any $x_{0}\in{\mathscr{D}}_{0}$, and therefore $N$ is contained in the center of $\mathbf{G}_{0}$. Using again the fact that $N$ acts trivially on $V$, we see that $N$ is contained in the center of $\mathbf{G}$. So $\dim\varphi_{/N}(S)=\dim\varphi(S)$ in this case.
On the other hand, it is clearly true that $\min_{N,~{}V\cap N=V}\left\{\dim\varphi_{/N}(S)\right\}+\frac{1}{2}\dim V$ is attained at $N=\mathbf{G}$ and hence equals $\frac{1}{2}\dim V$. Now we are done. ∎
Theorem 5.2.
Assume: (i) the connected algebraic monodromy group $H$ of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple; (ii) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ has no isotrivial sub-VHS, i.e. locally constant VHS. Then
$r(\nu)=\min\left\{\dim\varphi(S),\frac{1}{2}\dim V\right\}.$
Proof.
Set $\mathbf{G}_{0}=\mathbf{G}/V$. By Deligne, $H\lhd\mathbf{G}^{\mathrm{der}}_{0}$.
Let $N$ be a normal subgroup of $\mathbf{G}$. We may and do assume $N\lhd\mathbf{G}^{\mathrm{der}}$. The reductive part $\mathbf{G}_{N}:=N/(V\cap N)$ is a normal subgroup of $\mathbf{G}$. Now that $\mathbf{G}_{N}\cap H$ is a normal subgroup of $H$, we have either $\mathbf{G}_{N}\cap H=\{1\}$ or $H<\mathbf{G}_{N}$ by (i).
Assume $H<\mathbf{G}_{N}$. Since $\mathbf{G}_{0}$ is reductive, we can decompose $V=(V\cap N)\bigoplus(V\cap N)^{\perp}$ as $\mathbf{G}_{0}$-modules. Now $(V\cap N)^{\perp}$ gives rise to a sub-VHS of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$. But $H$ acts trivially on $(V\cap N)^{\perp}$ because $\mathbf{G}_{N}$ acts trivially on $V/(V\cap N)$. So $(V\cap N)^{\perp}=\{0\}$ by (ii). Hence $V\cap N=V$, and hence $V<N$. Thus $\varphi_{/N}(S)$ is a point. So $\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)=\frac{1}{2}\dim V$.
If $\mathbf{G}_{N}\cap H=\{1\}$, then $\dim\varphi_{/N}(S)=\dim\varphi(S)$. Now we are done. ∎
Now, we apply these two theorems to the Gross-Shoen and the Ceresa normal functions. Let $\nu_{{\mathrm{GS}}}$ (resp. $\nu_{{\mathrm{Ce}}}$) be the Gross–Schoen (resp. Ceresa) normal function ${\mathscr{M}}_{g}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defined in (C.5), with ${\mathbb{V}}_{{\mathbb{Z}}}$ from Proposition C.10. Let $\beta_{{\mathrm{GS}}}$ and $\beta_{{\mathrm{Ce}}}$ be the associated Betti forms.
Corollary 5.3.
Assume $g\geq 3$. The Betti ranks $r(\nu_{{\mathrm{GS}}})=r(\nu_{{\mathrm{Ce}}})=3g-3$. Equivalently, $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\not\equiv 0$ and $\beta_{{\mathrm{Ce}}}^{\wedge(3g-3)}\not\equiv 0$.
Proof.
We shall apply Theorem 5.1 and let us check the assumptions. Assumption (i) holds by Proposition C.10. (iii). Assumption (ii) holds by Proposition C.10. (ii). Moreover the VHS ${\mathbb{V}}_{{\mathbb{Z}}}$ on $S$ has maximal moduli, so $\dim\varphi_{{\mathrm{GS}}}({\mathscr{M}}_{g})=\dim\varphi_{{\mathrm{Ce}}}({\mathscr{M}}_{g})=\dim{\mathscr{M}}_{g}=3g-3$ where $\varphi_{{\mathrm{GS}}}$ and $\varphi_{{\mathrm{Ce}}}$ are the period maps. Finally, $\dim{\mathbb{V}}_{{\mathbb{Q}},s}=\frac{2g(2g-1)(2g-2)}{6}-2g$, which is $>2(3g-3)$ when $g\geq 3$. Hence, we can conclude by Theorem 5.1. ∎
We are ready to prove Theorem 1.3.
Proof of Theorem 1.3.
Denote for simplicity by $S={\mathscr{M}}_{g}$. Let $\nu$ be either $\nu_{{\mathrm{GS}}}$ or $\nu_{{\mathrm{Ce}}}$.
Set $U:=S\setminus S_{{\mathscr{F}}}(1)$, which is Zariski open by Theorem 3.2.
By Corollary 5.3 and the “In particular” part of Theorem 3.2, we have that $\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}=0$. Thus $S_{{\mathscr{F}}}(1)\not=S$. Now, part (i) follows from Proposition C.7.
For (ii), notice that by definition $S_{{\mathscr{F}}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion. Thus, the set $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is discrete, and hence at most countable. ∎
6. Northcott property for Gross–Schoen and Ceresa cycles
In this section, we turn to the arithmetic applications.
6.1. Construction of the adelic line bundle
Let ${\mathscr{M}}_{g}$ be the moduli scheme of smooth curves of genus $g$ over ${\mathbb{Z}}$, and let ${\mathscr{C}}_{g}\rightarrow{\mathscr{M}}_{g}$ be the universal curve.
Denote by ${\mathscr{J}}_{g}=\mathrm{Jac}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ the relative Jacobian. Identify ${\mathscr{J}}_{g}$ with its dual via the principal polarization given by a suitable theta divisor.
The Poincaré line bundle ${\mathscr{P}}$ on ${\mathscr{J}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{J}}_{g}$ extends to an integrable adelic line bundle $\overline{{\mathscr{P}}}$ as follows. Define ${\mathscr{P}}^{\Delta}:=\Delta^{*}{\mathscr{P}}$ for the diagonal $\Delta\colon{\mathscr{J}}_{g}\rightarrow{\mathscr{J}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{J}}_{g}$. Then ${\mathscr{P}}^{\Delta}$ is relatively ample on ${\mathscr{J}}_{g}\rightarrow{\mathscr{M}}_{g}$, and $[2]^{*}{\mathscr{P}}^{\Delta}=({\mathscr{P}}^{\Delta})^{\otimes 4}$. So $({\mathscr{J}}_{g},[2],{\mathscr{P}}^{\Delta})$ is a polarized dynamical system over ${\mathscr{M}}_{g}$ in the sense of [YZ21, §$\S$2.6.1]. Thus, Tate’s limit process gives a nef adelic line bundle $\overline{{\mathscr{P}}}^{\Delta}$ on ${\mathscr{J}}_{g}$, as executed by [YZ21, Thm. 6.1.1]. Now we obtain the desired $\overline{{\mathscr{P}}}\in\widehat{\mathrm{Pic}}({\mathscr{J}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}$ by letting
$2\overline{{\mathscr{P}}}:=m^{*}\overline{{\mathscr{P}}}^{\Delta}-p_{1}^{*}\overline{{\mathscr{P}}}^{\Delta}-p_{2}^{*}\overline{{\mathscr{P}}}^{\Delta},$
where $m,p_{1},p_{2}\colon{\mathscr{J}}_{g}\rightarrow{\mathscr{J}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{J}}_{g}$ with $m$ being the addition and $p_{1}$ (resp. $p_{2}$) being the projection to the first (resp. second) factor.
Take $\mathbf{\xi}\in\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ such that $(2g-2)\mathbf{\xi}=\omega_{{\mathscr{C}}_{g}/{\mathscr{M}}_{g}}$. Let
$i_{\mathbf{\xi}}\colon{\mathscr{C}}_{g}\longrightarrow{\mathscr{J}}_{g}$
be the Abel–Jacobi map based at $\mathbf{\xi}$. Then we have an ${\mathscr{M}}_{g}$-morphism $(i_{\mathbf{\xi}},i_{\mathbf{\xi}})\colon{\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}\rightarrow{\mathscr{J}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{J}}_{g}$, and hence get an integrable adelic line bundle on ${\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}$
$\overline{{\mathscr{Q}}}:=(i_{\mathbf{\xi}},i_{\mathbf{\xi}})^{*}\overline{{\mathscr{P}}}\in\widehat{\mathrm{Pic}}({\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}},$
and we can compute
(6.1) ${\mathscr{Q}}={\mathscr{O}}(\Delta)-p_{1}^{*}\xi-p_{2}^{*}\xi.$
with $\Delta$ the diagonal of ${\mathscr{C}}_{g}\rightarrow{\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}$, and $p_{1},p_{2}$ the projections ${\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}\rightarrow{\mathscr{C}}_{g}$.
Finally the Deligne pairing gives an adelic line bundle on ${\mathscr{M}}_{g}$ by [YZ21, Thm. 4.1.3]
(6.2) ¯$\overline{{\mathscr{L}}}:=\langle\overline{{\mathscr{Q}}},\overline{{\mathscr{Q}}},\overline{{\mathscr{Q}}}\rangle\in\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}.$
The following theorem is a reformulation of the second-named author’s [Zha10, Thm. 2.3.5].
Theorem 6.1.
For any $s\in{\mathscr{M}}_{g}(\overline{\mathbb{Q}})$, we have
$\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}=h_{\overline{{\mathscr{L}}}}(s).$
Proof.
In some neighborhood $U$ of $s$ in ${\mathscr{M}}_{g}$, $\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})=\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is an element of $Z^{2}({\mathscr{C}}_{U}^{3})$, i.e. it is a $2$-cocycle of ${\mathscr{C}}_{U}^{3}$.
Let $\ell$ be a rational section of ${\mathscr{Q}}$ over ${\mathscr{C}}_{U}^{2}$. By (6.1), the divisor $\mathrm{div}(\ell_{s^{\prime}})$ at each $s^{\prime}\in U({\mathbb{C}})$ is not the pullback of a divisor under the two natural projections ${\mathscr{C}}_{U}^{2}\rightarrow{\mathscr{C}}_{U}$. So $\mathrm{div}(\ell)$ can be seen as a correspondence of ${\mathscr{C}}_{U}$. Up to shrinking $U$ we can take rational sections $\ell_{1},\ell_{2},\ell_{3}$ of ${\mathscr{Q}}$ over ${\mathscr{C}}_{U}^{2}$ with the following property: For their divisors $t_{1},t_{2},t_{3}\in\mathrm{Div}({\mathscr{C}}_{U}^{2})$, we have
$|t_{1}|\cap|t_{2}|\cap|t_{3}|=\emptyset,\qquad|\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})|\cap|(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})|=\emptyset,$
where $t_{1}\otimes t_{2}\otimes t_{3}$ is seen as a correspondence of ${\mathscr{C}}_{U}^{3}$. Notice that we have a rational section $\langle\ell_{1},\ell_{2},\ell_{3}\rangle$ of ${\mathscr{L}}$ on $U$.
We shall apply [Zha10, Thm. 2.3.5] to ${\mathscr{C}}_{s}$. We have seen that $t_{1},t_{2},t_{3}$ restricted to the fiber over $s\in U(\overline{\mathbb{Q}})$ satisfies the assumption of [Zha10, Thm. 2.3.5], so
(6.3) $\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),(t_{1,s}\otimes t_{2,s}\otimes t_{3,s})^{*}\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}=\hat{t}_{1,s}\cdot\hat{t}_{2,s}\cdot\hat{t}_{3,s}$
where $\hat{t}_{1,s}$, $\hat{t}_{2,s}$, $\hat{t}_{3,s}$ are suitable arithmetic divisors on some model of ${\mathscr{C}}_{s}^{2}$ extending $t_{1,s}$, $t_{2,s}$, $t_{3,s}$. The right hand side is precisely $\frac{1}{[{\mathbb{Q}}(s):{\mathbb{Q}}]}\widehat{\deg}\overline{{\mathscr{L}}}|_{s}=h_{\overline{{\mathscr{L}}}}(s)$. The left hand side is $\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}$ because, by (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$. Hence, we are done. ∎
6.2. Relating with the Gross–Schoen normal function
Let $\nu_{{\mathrm{GS}}}$ be the Gross–Schoen normal function ${\mathscr{M}}_{g}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defined in (C.5) (with ${\mathbb{V}}_{{\mathbb{Z}}}$ from Proposition C.10).
By (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ Hence, they define the same normal functions $\nu_{{\mathrm{GS}}}$ on $U$. Set
$\overline{{\mathscr{P}}^{{\mathrm{GS}}}}:=(\nu_{{\mathrm{GS}}},\nu_{{\mathrm{GS}}})^{*}\overline{{\mathscr{P}}}=\nu_{{\mathrm{GS}}}^{*}\overline{{\mathscr{P}}^{\Delta}}$
for the metrized Poincaré bundle $\overline{{\mathscr{P}}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{{\mathscr{M}}_{g}}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defined in Definition D.1 (or the metrized tautological bundle $\overline{{\mathscr{P}}^{\Delta}}$ from Definition D.3). Set
$\beta_{{\mathrm{GS}}}:=c_{1}(\overline{{\mathscr{P}}^{{\mathrm{GS}}}})$
to be the Betti form, i.e. the curvature of $\overline{{\mathscr{P}}^{{\mathrm{GS}}}}$. It is a semi-positive $(1,1)$-form. The following Proposition can also be deduced from R. de Jong’s work [dJ16, (9.3)]:
Proposition 6.2.
We have the following identity of $(1,1)$-forms on ${\mathscr{M}}_{g}$:
$c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})=\beta_{{\mathrm{GS}}}.$
In particular, $c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})$ is semi-positive.
Proof.
We use the notation in the proof of Theorem 6.1. For any $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, we obtain a neighborhood $U$ such that (6.3) holds for any $s^{\prime}\in U({\mathbb{C}})$. So we get
$\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U}),(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})\rangle_{\infty}=-\log\|\langle\ell_{1},\ell_{2},\ell_{3}\rangle\|_{\infty}\quad\text{ on }U.$
By Proposition D.2, we have a section $\beta\in H^{0}(U,{\mathscr{P}}^{{\mathrm{GS}}})$ such that
$\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U}),(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})\rangle_{\infty}=-\log\|\beta\|\quad\text{ on }U.$
So $\log\|\beta\|=\log\|\langle\ell_{1},\ell_{2},\ell_{3}\rangle\|_{\infty}$ on $U$. Taking ¯$\frac{\partial\bar{\partial}}{\pi i}$, we get that $c_{1}(\overline{{\mathscr{P}}^{{\mathrm{GS}}}})=c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})$ on $U$. Now we can conclude by letting $s$ run over ${\mathscr{M}}_{g}({\mathbb{C}})$. ∎
6.3. Bigness of the generic fiber of the adelic line bundle
By [YZ21, Thm. 5.3.5], lower bounds of the height function $h_{\overline{{\mathscr{L}}}}$ correspond to bigness properties of $\overline{{\mathscr{L}}}$. In this paper, we prove the bigness of the generic fiber of $\overline{{\mathscr{L}}}$ and deduce the desired height comparison from it.
There is a natural base change of adelic line bundles $\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}\rightarrow\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Q}})_{{\mathbb{Q}}}$. Denote by $\tilde{{\mathscr{L}}}$ the image of $\overline{{\mathscr{L}}}$ under this base change map; $\tilde{{\mathscr{L}}}$ is called the generic fiber of $\overline{{\mathscr{L}}}$.
We start by stating the following formula for $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})$. Notice that it does not follow from arithmetic Hilbert–Samuel because we have not proved the nefness of $\tilde{{\mathscr{L}}}$.
Proposition 6.3.
We have
$\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})^{3g-3}.$
The proof of this proposition shall be postponed to §$\S$6.4.
Theorem 6.4.
Assume $g\geq 3$. Then the adelic line bundle $\tilde{{\mathscr{L}}}$ is big, i.e. $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})>0$.
Proof.
By Proposition 6.3 and Proposition 6.2, we have $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}$. Now $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\geq 0$ since $\beta_{{\mathrm{GS}}}$ is semi-positive. By Corollary 5.3, $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\not\equiv 0$. Hence $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})>0$. ∎
Proof of Theorem 1.1.
The claim for $\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}$ is a direct consequence of Theorem 6.4 and [YZ21, Thm. 5.3.5.(iii)]. Notice that one can either take the adelic line bundle on ${\mathscr{M}}_{g}$, which defines the Faltings height (it exists by [YZ21, §$\S$2.6.2]), or one can take any ample line bundle on a suitable compactification of ${\mathscr{M}}_{g}$ and then use the comparison of the logarithmic Weil height and the Faltings height.
The claim for $\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}$ follows because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple. ∎
6.4. Proof of Proposition 6.3
Denote for simplicity by ${\mathscr{S}}={\mathscr{M}}_{g}$.
By the flatness of extension ${\mathbb{Q}}\subseteq{\mathbb{C}}$, we have $\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{\mathbb{Q}})=\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{\mathbb{C}})$. By [YZ21, Theorem 5.2.1], it suffices to construct a sequence of model Hermitian line bundles $({\mathscr{S}}_{i},\overline{\mathscr{L}}_{i})$ with limit $({\mathscr{S}},\overline{\mathscr{L}})$ so that
$\lim_{i{\longrightarrow}\infty}\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{i})=\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})^{3g-3}.$
Let $\overline{\mathscr{N}}$ be an ample Hermitian line bundle on some integral model ${\mathscr{S}}$ of ${\mathscr{S}}$. Following proof of [YZ21, Theorem 6.1.1], we have a sequence of model line bundles $({\mathscr{J}}_{i},\overline{\mathscr{P}}^{\Delta}_{i},\ell_{i})$ with the limit $({\mathscr{J}},\overline{\mathscr{P}}^{\Delta})$, so that $\overline{\mathscr{P}}^{\Delta}_{i}+4^{-i}\pi^{*}\overline{\mathscr{N}}$ is nef.
For any $n\in{\mathbb{N}}$, we have an action of ${\mathscr{J}}[n]$ on ${\mathscr{J}}$ by translating torsion points:
$m_{n}\colon{\mathscr{J}}[n]\times{\mathscr{J}}\overset{(m,p)}{{\longrightarrow}}{\mathscr{J}}\times{\mathscr{J}},\qquad(t,x)\mapsto(x+t,x)$
By the Theorem of square, ${\mathrm{N}}_{p}m_{n}^{*}{\mathscr{P}}^{\Delta}=({\mathscr{P}}^{\Delta})^{n^{2g}}$. Thus, we obtain an adelic metrized line bundle
$\overline{\mathscr{P}}^{\Delta}_{i,n}:=n^{-2g}{\mathrm{N}}_{p}m_{n}^{*}\overline{\mathscr{P}}^{\Delta}_{i}.$
This bundle is again realized on some model ${\mathscr{J}}_{i,n}$ of ${\mathscr{J}}$ with connection morphism $\ell_{i,n}:{\mathscr{P}}^{\Delta}{\longrightarrow}{\mathscr{P}}^{\Delta}_{i,n,{\mathscr{J}}}$. Moreover ${\mathrm{div}}(\ell_{i,n})$ is in fact bounded by ${\mathrm{div}}(\ell_{i})$. Over ${\mathbb{C}}$, the curvature form $c_{1}(\overline{\mathscr{P}}^{\Delta}_{i,n,{\mathbb{C}}})$ is obtained from the curvature form $c_{1}(\overline{\mathscr{P}}^{\Delta}_{i,{\mathbb{C}}})$ by taking over average over $n$-torsion points. It follows that these forms converge to $c_{1}(\overline{\mathscr{P}}^{\Delta})$ uniformly in any compact subset of ${\mathscr{J}}({\mathbb{C}})$. These bundles also induce a double sequence of model line bundles $({\mathscr{S}}_{i},\overline{\mathscr{L}}_{i,n})$ of $({\mathscr{S}},\overline{\mathscr{L}})$ so that they convergent to $({\mathscr{S}},\overline{\mathscr{L}})$ as $i{\longrightarrow}\infty$, and that the metric $c_{1}(\overline{\mathscr{L}}_{i,n,{\mathbb{C}}})$ uniformly convergent to $c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})$ as $n{\longrightarrow}\infty$.
More precisely, we let $\Omega_{i}$ be an increasing sequence of relatively compact open subsets of ${\mathscr{S}}({\mathbb{C}})$ so that ${\mathscr{S}}({\mathbb{C}})=\cup\Omega_{i}$ and $\epsilon_{i}$ be a decreasing sequence of positive numbers convergent to 00 so that on $\Omega_{i}$,
$c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})\leq\epsilon_{i}^{-1}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
Then for each $i$, we choose $n_{i}$ so that
(6.4) $-\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\leq c_{1}(\overline{\mathscr{L}}_{i,n_{i},{\mathbb{C}}})-c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})\leq\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})$
as hermitian forms on the tangent bundle on $\Omega_{i}$. We simply write ¯$\overline{\mathscr{L}}_{i}=\overline{\mathscr{L}}_{i,n_{i}}$. Using the reference curvature $c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})$, we may talk about eigenvalues of $c_{1}(\overline{\mathscr{L}}_{{\mathbb{C}}})$ and $c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})$.
Now we apply Demailly’s Morse inequality [Dem91] for the bundle ¯$\overline{\mathscr{L}}_{i}$ on ${\mathscr{S}}_{i}({\mathbb{C}})$. For each $q\in{\mathbb{N}}$, let ${\mathscr{S}}_{i,q}$ denote the subset of ${\mathscr{S}}_{i}({\mathbb{C}})$ of points where $c_{1}(\overline{\mathscr{L}}_{i})$ has $q$-negative eigenvalues and $n-q$ positive eigenvalues. Then by [Dem91, (1.3), (1.5)], we have the following estimate as $k{\longrightarrow}\infty$:
$h^{q}(k{\mathscr{L}}_{i})\leq\frac{k^{d}}{d!}\left|\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}\right|+o(k^{d}),\qquad d:=3g-3$
$\sum_{q}(-1)^{q}h^{q}(k{\mathscr{L}}_{i})=\frac{k^{d}}{d!}\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}+o(k^{d}).$
It follows that
$\left|\widehat{\mathrm{vol}}({\mathscr{L}}_{i})-\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}\right|\leq\sum_{q>0}\left|\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}\right|.$
By [YZ21, Theorem 5.4.4],
$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}=\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}})^{d}.$
Thus, it remains to the following estimate for each $q>0$:
$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}=0.$
We want to prove this on $\Omega_{i}\cap{\mathscr{S}}_{i,q}$ and on its complement ${\mathscr{S}}_{i,q}\setminus\Omega_{i}$ respectively.
By (6.4), we have
$-\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\leq c_{1}(\overline{\mathscr{L}}_{i})\leq(\epsilon^{-1}+\epsilon^{d})c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
Thus on $\Omega_{i}\cap{\mathscr{S}}_{i,q}$, $c_{1}(\overline{\mathscr{L}}_{i})$ has all eigenvalue $\leq\epsilon^{-1}+\epsilon^{d}$ and one negative eigenvalue with absolute value bounded by $\leq\epsilon^{d}$. It follows that $|c_{1}({\mathscr{L}}_{i})^{d}|$ is bounded by
$\epsilon^{d}(\epsilon^{-1}+\epsilon^{d})^{d-1}c_{1}(\overline{\mathscr{N}})^{d}=\epsilon_{i}(1+\epsilon_{i}^{d+1})^{d-1}c_{1}(\overline{\mathscr{N}})^{d}.$
It follows that
$\int_{{\mathscr{S}}_{i,q}\cap\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{d}|=O(\epsilon_{i}).$
It remains to treat the integral over ${\mathscr{S}}_{i,q}\setminus\Omega_{i}$. Using decomposition
${\mathscr{P}}=\frac{1}{2}(m^{*}{\mathscr{P}}^{\Delta}-p_{1}^{*}{\mathscr{P}}^{\Delta}-p_{2}^{*}{\mathscr{P}}^{\Delta}),$
we may write ¯$\overline{\mathscr{L}}_{i}=\overline{\mathscr{E}}_{i}-\overline{\mathscr{F}}_{i}$, where ¯$\overline{\mathscr{E}}_{i}$ and ¯$\overline{\mathscr{F}}_{i}$ are two sequences of new bundles convergent to $\overline{\mathscr{E}}$ and $\overline{\mathscr{F}}$ with smooth metrics respectively. Then we have
$\int_{{\mathscr{S}}_{i,q}\setminus\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{d}|\leq\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}.$
Now let $\Omega_{i}^{\prime}\subset\Omega_{i}$ be another increasing sequence of relatively compact open subsets so that $\cup\Omega_{i}^{\prime}={\mathscr{S}}({\mathbb{C}})$. Then we can construct an increasing sequence of continuous functions $f_{i}$ so that $f_{i}(x)=1$ on ${\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}$ and $f_{i}=0$ on $\Omega_{i}^{\prime}$. Then for any $i\geq j$ we have
$\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{i}\cdot c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}.$
Fix $j$ and take $i{\longrightarrow}\infty$, we get
$\limsup_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{j}^{\prime}}c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{d}.$
Let $j{\longrightarrow}\infty$. We get
$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}=0.$
We are done. ∎
Appendix A Variation of mixed Hodge structures
A.1. Definitions
Let $R$ be a subring of ${\mathbb{R}}$.
Definition A.1.
Let $M$ be a free $R$-module of finite rank.
  1. (i)
    An $R$-pure Hodge structure on $M$ of weight $n$ is a decreasing filtration $F^{\bullet}$ on $M_{{\mathbb{C}}}$ (the Hodge filtration) such that $M_{{\mathbb{C}}}=F^{p}M_{{\mathbb{C}}}\oplus\overline{F^{n+1-p}M_{{\mathbb{C}}}}$ for all $p\in{\mathbb{Z}}$.
  2. (ii)
    An $R$-mixed Hodge structure on $M$ is a triple $(M,W_{\bullet},F^{\bullet})$ consisting of two filtrations, an increasing filtration $W_{\bullet}$ on $M_{{\mathbb{Q}}}$ (the weight filtration) and a decreasing filtration $F^{\bullet}$ on $M_{{\mathbb{C}}}$ (the Hodge filtration), such that for each $k\in{\mathbb{Z}}$, $\mathrm{Gr}_{k}^{W}M_{{\mathbb{Q}}}=W_{k}/W_{k-1}$ is a ${\mathbb{Q}}$-pure Hodge structure of weight $k$ for the filtration on $\mathrm{Gr}_{k}^{W}(M_{\mathbb{C}})$ deduced from $F^{\bullet}$.
Pure Hodge structures of weight $n$ can be defined in terms of bigradings. Indeed, set $M^{p,n-p}:=F^{p}M_{{\mathbb{C}}}\cap\overline{F^{p+1}M_{{\mathbb{C}}}}$, then $M_{{\mathbb{C}}}=\oplus_{p}M^{p,n-p}$ (the Hodge decomposition) and $\overline{M^{n-p,p}}=M^{p,n-p}$. We have $F^{p}=\oplus_{p^{\prime}\geq p}M^{p^{\prime},n-p^{\prime}}$.
For a mixed Hodge structure $(M,W_{\bullet},F^{\bullet})$, the numbers $k\in{\mathbb{Z}}$ such that $\mathrm{Gr}_{k}^{W}M_{{\mathbb{Q}}}\not=0$ are called its weights, and the numbers $h^{p,q}(M)=\dim_{{\mathbb{C}}}F^{p}\mathrm{Gr}^{W}_{p+q}(M_{{\mathbb{C}}})/F^{p+1}\mathrm{Gr}^{W}_{p+q}(M_{{\mathbb{C}}})$ are called its Hodge numbers .
For each $n\in{\mathbb{Z}}$, define $R(n)$ to be the pure Hodge structure on $R$ of weight $-2n$ such that $R(n)^{-n,-n}={\mathbb{C}}$ and $R(n)^{p,q}=0$ for all $(p,q)\not=(-n,-n)$.
A polarization on a pure Hodge structure $V$ of weight $n$ is a morphism of Hodge structures
$Q\colon V_{{\mathbb{Q}}}\otimes V_{{\mathbb{Q}}}\longrightarrow{\mathbb{Q}}(-n)$
such that the Hermitian form on $V_{{\mathbb{C}}}$ given by $Q(Cu,\bar{v})$ is positive-definite where $C$ is the Weil operator ($C|_{H^{p,q}}=i^{p-q}$ for all $p,q$).
A.2. Mumford–Tate group
Now, let us turn to a more group theoretical point of view on mixed Hodge structures. Let ${\mathbb{S}}=\mathrm{Res}_{{\mathbb{C}}/{\mathbb{R}}}{\mathbb{G}}_{\mathrm{m},{\mathbb{C}}}$ be the Deligne torus, i.e. the real algebraic group such that ${\mathbb{S}}({\mathbb{R}})={\mathbb{C}}^{*}$ and ${\mathbb{S}}({\mathbb{C}})={\mathbb{C}}^{*}\times{\mathbb{C}}^{*}$, and that the complex conjugation on ${\mathbb{S}}({\mathbb{C}})$ sends $(z_{1},z_{2})\mapsto(\bar{z}_{2},\bar{z}_{1})$.
As for pure Hodge structures, mixed Hodge structures can also be equivalently defined in terms of bigradings by Deligne [Del71, 1.2.8]. Given a ${\mathbb{Q}}$-vector space $M$ of finite dimension, a bigrading $M_{{\mathbb{C}}}=\oplus_{p,q\in{\mathbb{Z}}}I^{p,q}$ is equivalent to a homomorphism $h\colon{\mathbb{S}}_{{\mathbb{C}}}\rightarrow\mathrm{GL}(M_{{\mathbb{C}}})$. In particular, any mixed Hodge structure on $M$ defines a unique homomorphism $h\colon{\mathbb{S}}_{{\mathbb{C}}}\rightarrow\mathrm{GL}(M_{{\mathbb{C}}})$, and we use $(M,h)$ to denote this mixed Hodge structure.
Definition A.2.
For any mixed Hodge structure $(M,h)$, its Mumford–Tate group is the smallest ${\mathbb{Q}}$-subgroup $\mathbf{G}$ of $\mathrm{GL}(M_{{\mathbb{Q}}})$ such that $h({\mathbb{S}}_{{\mathbb{C}}})\subseteq\mathbf{G}({\mathbb{C}})$.
Now we assume that $M$ has weight 00 and $-1$ with $M_{-1}:={\mathrm{Gr}}^{W}_{-1}M$ and $M_{0}:={\mathrm{Gr}}^{W}_{0}M$. Then, we have an exact sequence
$00{\longrightarrow}M_{-1}{\longrightarrow}M{\longrightarrow}M_{0}{\longrightarrow}$ 0 .
It is clear that $h({\mathbb{S}})$ stabilizes this exact sequence and induced identity on $N$. Let $\mathbf{G}_{0}$ denote the Mumford–Tate group of $M_{-1}$. Then, we have an exact sequence of reductive groups:
$0{\longrightarrow}V{\longrightarrow}\mathbf{G}{\longrightarrow}\mathbf{G}_{0}{\longrightarrow}0,$
where $V$ is a vector group included into ${\mathrm{Hom}}(M_{0},M_{-1})$. If $M_{-1}$ is polarized, then $\mathbf{G}_{0}$ is reductive by Deligne. Thus, $V$ is the unipotent radical of $\mathbf{G}$.
The base change to ${\mathbb{R}}$ has a natural splitting $M_{\mathbb{R}}=M_{-1,{\mathbb{R}}}\oplus M_{0,{\mathbb{R}}}$ of Hodge structures given by the inverse of the isomorphism $M_{\mathbb{R}}\cap F^{0}M_{\mathbb{C}}{\overset{\sim}{{\longrightarrow}}}M_{0,{\mathbb{R}}}$. This induces splittings
$\mathbf{G}_{\mathbb{R}}=V_{\mathbb{R}}\rtimes\mathbf{G}_{0,{\mathbb{R}}}.$
Moreover $h({\mathbb{S}})$ is included into $G_{0,{\mathbb{R}}}$. Thus, we have proved the following:
Lemma A.3.
Assume $(M,h)$ has weight $-1$ and 00. The following holds.
  1. (i)
    The $h$ is defined over ${\mathbb{R}}$.
  2. (ii)
    The unipotent radical $V$ of $\mathbf{G}$ is a vector group.
A.3. Variation of mixed Hodge structures and admissibility
Definition A.4.
Let $S$ be a connected complex manifold. A variation of mixed Hodge structures (VMHS) on $S$ is a triple $({\mathbb{M}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$ consisting of
    a local system ${\mathbb{M}}_{{\mathbb{Z}}}$ of free ${\mathbb{Z}}$-modules of finite rank on $S$ ,
    a finite increasing filtration (weight filtration) $W_{\bullet}$ of the local system ${\mathbb{M}}:={\mathbb{M}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}_{S}}{\mathbb{Q}}_{S}$ by local subsystems,
    a finite decreasing filtration (Hodge filtration) ${\mathscr{F}}^{\bullet}$ of the holomorphic vector bundle ${\mathscr{M}}:={\mathbb{M}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}_{S}}{\mathscr{O}}_{S}$ by holomorphic subbundles
satisfying the following properties
  1. (i)
    for each $s\in S$, the triple $({\mathbb{M}}_{s},W_{\bullet},{\mathscr{F}}^{\bullet}_{s})$ defines a mixed Hodge structure on ${\mathbb{M}}_{s}$,
  2. (ii)
    the connection $\nabla\colon{\mathscr{M}}\rightarrow{\mathscr{M}}\otimes_{{\mathscr{O}}_{S}}\Omega_{S}^{1}$ whose sheaf of horizontal sections is ${\mathbb{M}}_{{\mathbb{C}}}:={\mathbb{M}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}_{S}}{\mathbb{C}}_{S}$ satisfies the Griffiths’ transversality condition
    $\nabla({\mathscr{F}}^{p})\subseteq{\mathscr{F}}^{p-1}\otimes\Omega_{S}^{1}.$
The weights and Hodge numbers of $({\mathbb{M}}_{s},W_{\bullet},{\mathscr{F}}^{\bullet}_{s})$ are the same for all $s\in S$. We call them the weights and the Hodge numbers of the VMHS $({\mathbb{M}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$.
If there is only one $n\in{\mathbb{Z}}$ such that $\mathrm{Gr}^{W}_{n}{\mathbb{M}}\not=0$, then each fiber of this VMHS is a pure Hodge structure of weight $n$. In this case, the VMHS is said to be pure. More precisely, we have the following definition.
Definition A.5.
A variation of Hodge structures (VHS) of weight $n$ on $S$ is a pair $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ such that $({\mathbb{V}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$ is a VMHS, where $W_{\bullet}$ is the increasing filtration on ${\mathbb{V}}:={\mathbb{V}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}_{S}}{\mathbb{Q}}_{S}$ defined by $W_{n-1}=0$ and $W_{n}={\mathbb{V}}$.
To each VMHS $({\mathbb{M}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$ on $S$, we can associate variations of pure Hodge structures obtained from $\mathrm{Gr}_{k}^{W}{\mathbb{M}}$.
We shall use the following convention: For each $n\in{\mathbb{Z}}$ and any VHS $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$, define ${\mathbb{V}}_{{\mathbb{Z}}}(n)$ to be the VHS $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet-n})$. In particular ${\mathbb{Z}}(n)_{S}$ be the VHS on $S$ of weight $-2n$ such that $({\mathbb{Z}}(n)_{S})_{s}={\mathbb{Z}}(n)$ for each $s\in S$.
Definition A.6.
A polarization of VHS $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ on $S$ of weight $n$ is a morphism of VHS ${\mathbb{V}}\otimes{\mathbb{V}}\rightarrow{\mathbb{Q}}(-n)_{S}$ inducing on each fiber a polarization.
We say that a VMHS $({\mathbb{M}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$ is graded-polarizable if $\mathrm{Gr}_{k}^{W}{\mathbb{M}}$ has a polarization for each $k\in{\mathbb{Z}}$.
Example A.7.
Let $f\colon X\rightarrow S$ be a projective smooth morphism of algebraic varieties over ${\mathbb{C}}$ with irreducible fibers of dimension $d$. For each $s\in S({\mathbb{C}})$, the cohomology group $H^{n}(X_{s},{\mathbb{Z}})$ is endowed with a natural Hodge structure of weight $n$ by the de Rham–Betti comparison, and this makes $R^{n}f_{*}{\mathbb{Z}}_{X}$ into a VHS on $S^{\mathrm{an}}$ of weight $n$. It is polarizable by the hard Lefschetz theorem and the Lefschetz decomposition.
We close this subsection with a discussion on the admissibility of VMHS. Let $S$ be a smooth, complex quasi-projective variety. Given a VMHS on $S^{\mathrm{an}}$, we often need to extend it to suitable compactifications of $S^{\mathrm{an}}$ and hence study the asymptotic behavior of the VMHS near the boundary. This leads to the definition of admissible VMHS, which are the graded-polarized VMHSs with good asymptotic properties. This concept was introduced by Steenbrick–Zucker [SZ85, Prop. 3.13] on a curve and Kashiwara [Kas86, 1.8 and 1.9] in general, and the property for a VMHS on $S^{\mathrm{an}}$ to be admissible does not depend on the choice of the compactification. We shall not recall the precise definition here (see, for example, [PS08, Defn. 14.49]), but point out the following:
  1. (1)
    Any VMHS arising from geometry is admissible [EZ86] (see also [BZ14]).
  2. (2)
    Any VHS is admissible.
Appendix B Classfying space and Mumford–Tate domain
We relate the Betti foliation on intermediate Jacobians to the fibered structure of certain Mumford–Tate domain parametrizing ${\mathbb{Q}}$-mixed Hodge structure of weight $-1$ and 00. In this appendix, we recall and prove some results about such classifying spaces and Mumford–Tate domains. The main result is Proposition B.3, which explains and compares the semi-algebraic structure and the complex structure of such Mumford–Tate domains. This comparison is used to study the Betti foliation of intermediate Jacobians.
B.1. Classifying space
Let $M_{-1}$ be a finite dimensional ${\mathbb{Q}}$-vector space. Let $M=M_{-1}\oplus{\mathbb{Q}}$.
B.1.1. Pure Hodge structures
Consider the polarized Hodge data on $M_{-1}$: a non-degenerate skew pairing $Q_{-1}\colon M_{-1}\otimes M_{-1}\rightarrow{\mathbb{Q}}(1)$, and a partition $\{h^{p,q}_{M_{-1}}\}_{p,q\in{\mathbb{Z}}}$ of $\dim M_{-1,{\mathbb{C}}}$ into non-negative integers with $p+q=-1$ such that $h^{p,q}_{M_{-1}}=h^{q,p}_{M_{-1}}$. Then there exists a classifying space ${\mathscr{M}}_{0}$ parametrizing ${\mathbb{Q}}$-Hodge structures on $M_{-1}$ of weight $-1$ with a polarization by $Q_{-1}$ such that the $(p,q)$-constituent of $M_{-1,{\mathbb{C}}}$ has complex dimension $h^{p,q}$. Moreover, the ${\mathbb{Q}}$-group $\mathbf{G}_{0}^{{\mathscr{M}}}:=\mathrm{Aut}(M_{-1},Q_{-1})$, the associated real Lie group $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ acts transitively on ${\mathscr{M}}_{0}$, i.e.
${\mathscr{M}}_{0}=\mathbf{G}_{0}({\mathbb{R}})^{+}x_{0}$
for any point $x_{0}\in{\mathscr{M}}_{0}$. This makes ${\mathscr{M}}_{0}$ into a semi-algebraic open subset of a flag variety ${\mathscr{M}}_{0}^{\vee}$, which is a suitable $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{C}})$-orbit, and hence endows ${\mathscr{M}}_{0}$ with a complex structure.
We can be more explicit on the action of $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ on ${\mathscr{M}}_{0}$. For each $x_{0}\in{\mathscr{M}}_{0}$, we have a Hodge decomposition and a Hodge filtration $F^{\bullet}_{x_{0}}$
(B.1) $M_{-1,{\mathbb{C}}}=\bigoplus\nolimits_{p+q=-1}(M_{-1,x_{0}})^{p,q},\qquad F^{p}_{x_{0}}M_{-1,{\mathbb{C}}}=\bigoplus\nolimits_{p^{\prime}\geq p}(M_{-1,x_{0}})^{p^{\prime},q^{\prime}}$
with $(M_{-1,x_{0}})^{q,p}=\overline{(M_{-1,x_{0}})^{p,q}}$. The inclusion ${\mathscr{M}}_{0}\subseteq{\mathscr{M}}_{0}^{\vee}$ is given by $x_{0}\mapsto F^{\bullet}_{x_{0}}$.
For the Deligne torus ${\mathbb{S}}=\mathrm{Res}_{{\mathbb{C}}/{\mathbb{R}}}{\mathbb{G}}_{\mathrm{m},{\mathbb{C}}}$, the bi-grading decomposition above defines a morphism $h_{x_{0}}\colon{\mathbb{S}}\rightarrow\mathrm{GL}(M_{-1,{\mathbb{R}}})$, with $(M_{-1,x_{0}})^{p,q}$ the eigenspace of the character $z\mapsto z^{-p}\bar{z}^{-q}$ of ${\mathbb{S}}$. It is known that $h_{x_{0}}({\mathbb{S}})<\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})$ for all $x_{0}\in{\mathscr{M}}_{0}$. Hence we have a $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map, which is known to be injective,
(B.2) ${\mathscr{M}}_{0}\rightarrow{\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}}),\quad x_{0}\mapsto h_{x_{0}}$
with the action of $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ on ${\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}})$ given by conjugation. So we will view ${\mathscr{M}}_{0}$ as a subset of ${\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}})$.
We close this subsubsection with the following remark on the Mumford–Tate group $\mathrm{MT}_{x_{0}}$ of the pure Hodge structure on $M_{-1}$ determined by $x_{0}$. We have that $\mathrm{MT}_{x_{0}}$ is a subgroup of $\mathbf{G}_{0}^{{\mathscr{M}}}$ for all $x_{0}\in{\mathscr{M}}_{0}$ and equals $\mathbf{G}_{0}^{{\mathscr{M}}}$ for some $x_{0}\in{\mathscr{M}}_{0}$.
B.1.2. Weight $-1$ and 00
Next, we turn to mixed Hodge structures of weight of $-1$ and 00.
Fix the following data on $M=M_{-1}\oplus{\mathbb{Q}}$: the weight filtration $W_{\bullet}:=(0=W_{-2}M\subseteq W_{-1}M=M_{-1}\subseteq W_{0}M=M)$; the partition $\{h^{p,q}\}_{p,q\in{\mathbb{Z}}}$ of $\dim M_{{\mathbb{C}}}$ into non-negative integers, with $h^{p,q}=h^{p,q}_{M_{-1}}$ for $p+q=-1$ and $h^{0,0}=1$ and $h^{p,q}=0$ otherwise.
There exists the classifying space ${\mathscr{M}}$ parametrizing ${\mathbb{Q}}$-mixed Hodge structures $(M,W_{\bullet},F^{\bullet})$ of weight $-1$ and 00 such that:
  1. (a)
    the $(p,q)$-constituent $\mathrm{Gr}^{p}_{F}\mathrm{Gr}^{W}_{p+q}M_{\mathbb{C}}$ has complex dimension $h^{p,q}$;
  2. (b)
    $\mathrm{Gr}^{W}_{-1}M=M_{-1}$ is polarized by $Q_{-1}$.
See for example [Pea00, below (3.7) to the Remark below Lem. 3.9]. Notice that $\mathrm{Gr}^{W}_{0}M={\mathbb{Q}}$ is polarized by $Q_{0}\colon{\mathbb{Q}}\otimes{\mathbb{Q}}\rightarrow{\mathbb{Q}}$ is $a\otimes b\mapsto ab$. So $M$ is graded-polarized.
In our case, we need a better understanding of the structure of ${\mathscr{M}}$ than [Pea00]. The map $x\mapsto F_{x}^{\bullet}$ realizes ${\mathscr{M}}$ as a semi-algebraic open subset of a suitable flag variety ${\mathscr{M}}^{\vee}$, which is easily seen to be an orbit under $\mathbf{G}({\mathbb{C}})$ for the ${\mathbb{Q}}$-group
(B.3) $\mathbf{G}^{{\mathscr{M}}}:=M_{-1}\rtimes\mathrm{Aut}(M_{-1},Q_{-1})=M_{-1}\rtimes\mathbf{G}_{0}^{{\mathscr{M}}}.$
[…]
(B.4) ${\mathscr{M}}=\mathbf{G}^{{\mathscr{M}}}({\mathbb{R}})^{+}x$
and a $\mathbf{G}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map, which is known to be injective
(B.5) ${\mathscr{M}}\longrightarrow{\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{{\mathbb{R}}}),\quad x\mapsto h_{x}.$
As in the pure case, for the Mumford–Tate group $\mathrm{MT}_{x}$ for the mixed Hodge structure on $M$ determined by $x$, we have that $\mathrm{MT}_{x}<\mathbf{G}^{{\mathscr{M}}}$ for all $x\in{\mathscr{M}}$, and $\mathrm{MT}_{x}=\mathbf{G}^{{\mathscr{M}}}$ for some $x\in{\mathscr{M}}$.
B.1.3.
For the quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}=\mathbf{G}^{{\mathscr{M}}}/M_{-1}$, consider the following surjective $\mathbf{G}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map
(B.6) $p\colon{\mathscr{M}}\rightarrow{\mathscr{M}}_{0},\quad g\cdot x\mapsto p(g)\cdot\bar{h}_{x}$
with $\bar{h}_{x}=p\circ h_{x}\colon{\mathbb{S}}\rightarrow\mathbf{G}^{{\mathscr{M}}}_{{\mathbb{R}}}\rightarrow\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}}$. This map sends the mixed Hodge structure $(M,W_{\bullet}M,F^{\bullet}M_{{\mathbb{C}}})$ to the associated pure Hodge structure on $M_{-1}=\mathrm{Gr}_{W}^{-1}M$.
We will study the fibered structure given by $p$ more carefully. We will do this in the finer setting of Mumford–Tate domains.
B.2. Mumford–Tate domains
Definition B.1.
A subset ${\mathscr{D}}$ of the classifying space ${\mathscr{M}}$ is called a (mixed) Mumford–Tate domain if there exists an element $x\in{\mathscr{D}}$ such that ${\mathscr{D}}=\mathbf{G}({\mathbb{R}})^{+}x$, where $\mathbf{G}=\mathrm{MT}(h_{x})$.
The group $\mathbf{G}$ in the definition above is called the generic Mumford–Tate group of ${\mathscr{D}}$ and is denoted by $\mathrm{MT}({\mathscr{D}})$. It is known that $\mathbf{G}<\mathbf{G}^{{\mathscr{M}}}$.
[…]
Recall that ${\mathscr{M}}$ is a semi-algebraic open subset in some algebraic variety ${\mathscr{M}}^{\vee}$ over ${\mathbb{C}}$. Hence, ${\mathscr{D}}$ is a semi-algebraic open subset in some algebraic variety ${\mathscr{D}}^{\vee}$ over ${\mathbb{C}}$. This endows ${\mathscr{D}}$ with a semi-algebraic structure and a complex structure.
Definition B.2.
A subset of ${\mathscr{D}}$ is said to be irreducible algebraic if it is both complex analytic irreducible and semi-algebraic.
In view of [KUY16, Lem. B.1 and its proof], a subset of ${\mathscr{D}}$ is irreducible algebraic if and only if it is a component of $U\cap{\mathscr{D}}$ with $U$ an algebraic subvariety of ${\mathscr{D}}^{\vee}$.
Now, let us take a closer look at the semi-algebraic structure and a complex structure on ${\mathscr{D}}$.
The unipotent radical $V$ of $\mathbf{G}=\mathrm{MT}({\mathscr{D}})$ equals $M_{-1}\cap\mathbf{G}$ by reason of weight. Let $\mathbf{G}_{0}:=\mathbf{G}/V$ be the reductive part. Set ${\mathscr{D}}_{0}:=p({\mathscr{D}})\subseteq{\mathscr{M}}_{0}$ for the map $p$ defined in (B.6). Then ${\mathscr{D}}_{0}$ is a $\mathbf{G}_{0}({\mathbb{R}})^{+}$-orbit and is in fact a (pure) Mumford–Tate domain in the classifying space ${\mathscr{M}}_{0}$, and $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$ for all $x_{0}\in{\mathscr{D}}_{0}$.
By abuse of notation, we also use $p$ to denote the natural projections
(B.7) $p\colon\mathbf{G}\rightarrow\mathbf{G}_{0}=\mathbf{G}/V\quad\text{ and }\quad p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}.$
Fix a Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$. Identify $\mathbf{G}_{0}$ with $\{0\}\times\mathbf{G}_{0}$.
Recall that each $x_{0}\in{\mathscr{M}}_{0}$ endows $M_{-1}$ with a Hodge structure of weight $-1$. For $x_{0}\in{\mathscr{D}}_{0}$, $V$ is a sub-Hodge structure because $V$ is a $\mathbf{G}_{0}$-submodule of $M_{-1}$ and that $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$.
Finally, consider the constant bundle $V({\mathbb{C}})\times{\mathscr{D}}_{0}\rightarrow{\mathscr{D}}_{0}$. Define the holomorphic subbundle $F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})$ to be such that the fiber over each $x_{0}\in{\mathscr{D}}_{0}$ is $F^{0}_{x_{0}}V_{{\mathbb{C}}}$.
Proposition B.3.
The following is true:
  1. (i)
    The semi-algebraic structure on ${\mathscr{D}}\subseteq{\mathrm{Hom}}({\mathbb{S}},\mathbf{G}_{{\mathbb{R}}})$ is given by
    $V({\mathbb{R}})\times{\mathscr{D}}_{0}\xrightarrow{\sim}{\mathscr{D}},\quad(v,x_{0})\mapsto\mathrm{Int}(v)\circ h_{x_{0}}.$
  2. (ii)
    The complex structure on ${\mathscr{D}}$ is given by ${\mathscr{D}}=(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})$.
  3. (iii)
    These two structures are related by the natural bijection
    (B.8) $V({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq V({\mathbb{C}})\times{\mathscr{D}}_{0}\longrightarrow(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0}).$
Proof.
Fix $x\in{\mathscr{D}}$ and let $h_{x}\colon{\mathbb{S}}\rightarrow\mathbf{G}_{{\mathbb{R}}}$ be the corresponding homomorphism.
The Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ induces a bijection $\mathbf{G}({\mathbb{R}})^{+}x\cong V({\mathbb{R}})\times{\mathscr{D}}_{0}$, which is semi-algebraic. This establishes (i).
Set ${\mathscr{X}}:=\mathbf{G}({\mathbb{R}})^{+}V({\mathbb{C}})\cdot h_{x}\subseteq{\mathrm{Hom}}({\mathbb{S}}_{{\mathbb{C}}},\mathbf{G}_{{\mathbb{C}}})$, where the action is given by conjugation. The quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}$ induces a natural surjective map ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and by [Pin89, 1.8(a)] each fiber of this map is a $V({\mathbb{C}})$-torsor.
The Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ induces a global section of ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and hence an isomorphism ${\mathscr{X}}\cong V({\mathbb{C}})\times{\mathscr{D}}_{0}$ over ${\mathscr{D}}_{0}$.
Consider the following surjective equivariant map
(B.9) $\varphi\colon{\mathscr{X}}\rightarrow{\mathscr{D}},\qquad gh_{x}g^{-1}\mapsto g\cdot x.$
By [Pin89, 1.8(b)], for each $x\in{\mathscr{D}}$, the fiber $\varphi^{-1}(x)$ is a principle homogeneous space under $F^{0}_{x_{0}}V_{{\mathbb{C}}}$ with $x_{0}=p(x)$. Hence $\varphi$ gives a bijection $(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})\rightarrow{\mathscr{D}}$ over ${\mathscr{D}}_{0}$. Moreover, the complex structure of ${\mathscr{D}}$ is precisely given by this bijection; see [GK24, Proof of Prop. 2.6 in Appendix A] which is a consequence of [Pea00, Thm. 3.13]. This establishes (ii).
[…]
Corollary B.4.
Let $\tilde{Z}_{0}\subseteq{\mathscr{D}}_{0}$ be an irreducible algebraic subset. Then for any $a\in V({\mathbb{R}})$, the subset $\{a\}\times\tilde{Z}_{0}\subseteq V({\mathbb{R}})\times{\mathscr{D}}_{0}\cong{\mathscr{D}}$ is irreducible algebraic.
Proof.
$\{a\}\times\tilde{Z}_{0}$ is clearly semi-algebraic in ${\mathscr{D}}$. In view of (B.8), $\{a\}\times\tilde{Z}_{0}$ is also complex analytic in ${\mathscr{D}}$. Hence, we are done. ∎
Lemma B.5.
Let ${\mathscr{D}}^{\prime}$ be a sub-Mumford–Tate domain of ${\mathscr{D}}$. Then under the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), we have ${\mathscr{D}}^{\prime}=(V^{\prime}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime})$ for some $v_{0}\in V({\mathbb{Q}})$.
Proof.
Denote by $\mathbf{G}^{\prime}=\mathrm{MT}({\mathscr{D}}^{\prime})$, $V^{\prime}:=V\cap\mathbf{G}^{\prime}$, and $\mathbf{G}^{\prime}_{0}:=\mathbf{G}^{\prime}/V^{\prime}$. Then, since ${\mathscr{D}}^{\prime}$ is a sub-Mumford–Tate domain of ${\mathscr{D}}$ and by reason of weight, $V^{\prime}$ is the unipotent radical of $\mathbf{G}^{\prime}$.
Set $\mathbf{G}^{\prime\prime}:=V\mathbf{G}^{\prime}$. Then $\mathbf{G}^{\prime\prime}$ is a subgroup of $\mathbf{G}$, with unipotent radical $V$ and reductive part $\mathbf{G}^{\prime}_{0}$. We can construct two Levi decompositions of $\mathbf{G}^{\prime\prime}$ as follows.
First, $\mathbf{G}^{\prime}_{0}=\mathbf{G}^{\prime}/V^{\prime}=\mathbf{G}^{\prime}/(V\cap\mathbf{G}^{\prime})<\mathbf{G}/V=\mathbf{G}_{0}$. Under the Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ fixed above Proposition B.3, we have $\{0\}\times\mathbf{G}^{\prime}_{0}<\mathbf{G}^{\prime\prime}$. The composite $\{0\}\times\mathbf{G}^{\prime}_{0}<\mathbf{G}^{\prime\prime}\rightarrow\mathbf{G}^{\prime\prime}/V=\mathbf{G}^{\prime}_{0}$ is the natural isomorphism (the projection to the second factor). Thus, we get a Levi decomposition $\mathbf{G}^{\prime\prime}=V\rtimes\mathbf{G}^{\prime}_{0}$, which is compatible with our fixed Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$.
Next, fix a Levi decomposition $\mathbf{G}^{\prime}=V^{\prime}\rtimes^{\prime}\mathbf{G}^{\prime}_{0}$. It is uniquely determined by a section of $\mathbf{G}^{\prime}\rightarrow\mathbf{G}^{\prime}_{0}$, i.e. an injective morphism $s^{\prime}\colon\mathbf{G}^{\prime}_{0}\rightarrow\mathbf{G}^{\prime}$ such that $s^{\prime}\circ p_{\mathbf{G}^{\prime}}=1_{\mathbf{G}^{\prime}_{0}}$. Now abuse of notation denote by $s^{\prime}$ the composite $\mathbf{G}^{\prime}_{0}\xrightarrow{s^{\prime}}\mathbf{G}^{\prime}<\mathbf{G}^{\prime\prime}$. Then $s^{\prime}$ defines a Levi decomposition $\mathbf{G}^{\prime\prime}=V\rtimes^{\prime}\mathbf{G}_{0}^{\prime}$.
By the general theory of algebraic groups, any two Levi decompositions of $\mathbf{G}^{\prime\prime}$ differ from the conjugation by some $v_{0}\in V({\mathbb{Q}})$. Now, the conclusion follows from Proposition B.3.(i).∎
B.3. Quotient by a normal subgroup and weak Mumford–Tate domains
Let ${\mathscr{D}}$ be a Mumford–Tate domain in ${\mathscr{M}}$ with $\mathrm{MT}({\mathscr{D}})=\mathbf{G}$. Let $N\lhd\mathrm{MT}({\mathscr{D}})$. By [GK24, Prop. 5.1], we have a quotient in the category of complex varieties
$p_{N}\colon{\mathscr{D}}\rightarrow{\mathscr{D}}/N$
with the following properties: (i) ${\mathscr{D}}/N$ is a Mumford–Tate domain in some classifying space of mixed Hodge structures (which must be of weight $-1$ and 00) and $\mathrm{MT}({\mathscr{D}}/N)=\mathbf{G}/N$; (ii) each fiber of $p_{N}$ is an $N({\mathbb{R}})^{+}$-orbit. It is clearly true that $p_{N}$ is semi-algebraic. More precisely, let $M_{-1}/N$ be the maximal quotient of $M_{-1}$ on which $N$ acts trivially, and let $M/N=M_{-1}/N\oplus{\mathbb{Q}}$. For each $x\in{\mathscr{D}}$, the composition
$\bar{h}_{x}:{\mathbb{S}}\overset{h_{x}}{{\longrightarrow}}{\mathrm{GL}}(M_{{\mathbb{R}}}){\longrightarrow}{\mathrm{GL}}((M/N)_{{\mathbb{R}}})$
defines some Hodge structure on $M/N$ with weight $-1$ on $M_{-1}/N$ and weight 00 on the quotient ${\mathbb{Q}}$. The Hodge numbers on $M/N$ defined by $h_{x}$ do not depend on the choice of $x$. Let ${\mathscr{M}}^{\prime}$ denote the corresponding classification space. Then $\bar{h}_{x}$ defined a point $\bar{x}$ in ${\mathscr{M}}^{\prime}$. The images of such $\bar{x}$ form a Mumford–Tate domain denoted by ${\mathscr{D}}/N$. Thus, we have the desired quotient map $p_{N}:{\mathscr{D}}{\longrightarrow}{\mathscr{D}}/N$ which is both complex analytic and real semi-algebraic.
Assume $\Gamma$ is an arithmetic subgroup of $\mathbf{G}({\mathbb{Q}})$. Then the quotient $\Gamma\backslash{\mathscr{D}}$ is an orbifold. Denote by $\Gamma_{/N}$ the image of $\Gamma$ under the quotient $\mathbf{G}\rightarrow\mathbf{G}/N$. Then the quotient $p_{N}$ induces
$[p_{N}]\colon\Gamma\backslash{\mathscr{D}}\rightarrow\Gamma_{/N}\backslash({\mathscr{D}}/N).$
We will pay special attention to the fibers of $p_{N}$ (and $[p_{N}]$). More generally, we define:
Definition B.6.
A subset ${\mathscr{D}}_{N}$ of ${\mathscr{D}}$ is called a weak Mumford–Tate domain if there exist $x\in{\mathscr{D}}$ and a normal subgroup $N$ of $\mathrm{MT}(x)$ such that ${\mathscr{D}}_{N}=N({\mathbb{R}})^{+}x$.
In this definition, if $x$ is taken to be a Hodge generic point, i.e. $\mathrm{MT}(x)=\mathbf{G}$, then the weakly Mumford–Tate domain thus obtained is a fiber of $p_{N}$.
Appendix C Intermediate Jacobians and normal functions
[…]
Write ${\mathscr{V}}={\mathbb{V}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}}{\mathscr{O}}_{S}$ for the holomorphic vector bundle obtained from ${\mathbb{V}}_{{\mathbb{Z}}}$.
C.1. Definition and basic property of Intermediate Jacobians
Definition C.1.
The quotient ${\mathscr{V}}/({\mathscr{F}}^{n}+{\mathbb{V}}_{{\mathbb{Z}}})$ is called the (relative) intermediate Jacobian of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$, and is denoted by ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ (or ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ when the Hodge filtration is clear in the context).
Lemma C.2.
$\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is torus fibration, i.e. a holomorphic family of compact complex torus.
Proof.
For each $s\in S$, we have ${\mathbb{V}}_{{\mathbb{C}},s}={\mathscr{F}}^{n}_{s}\oplus\overline{{\mathscr{F}}^{n}_{s}}$ because ${\mathbb{V}}_{{\mathbb{Z}},s}$ has weight $12n-$ 1. In particular, $\dim{\mathscr{F}}^{n}_{s}=\frac{1}{2}\dim{\mathscr{V}}_{s}$ and ${\mathbb{V}}_{{\mathbb{Z}},s}\cap{\mathscr{F}}^{n}_{s}=\{0\}$. Hence, each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{V}}_{s}/({\mathscr{F}}^{n}_{s}+{\mathbb{V}}_{{\mathbb{Z}},s})$ is a compact complex torus, and this yields the claim. ∎
Example C.3.
Let us look at the following example from geometry. Let $f\colon X\rightarrow S$ be a smooth projective morphism of relative dimension $d$ over a complex quasi-projective variety such that each fiber is irreducible. For each $n\geq 1$, the relative intermediate Jacobian of the (polarizable) VHS $R^{2n-1}f_{*}{\mathbb{Z}}(n)_{X}$ is called the $n$-th intermediate Jacobian of $X\rightarrow S$ and is denoted by ${\mathscr{J}}^{n}(X/S)$. If $S$ is a point and $X=X$, then we simply write ${\mathscr{J}}^{n}(X)$.
For each $s\in S$ we have $H_{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{Z}})={\mathbb{V}}_{{\mathbb{Z}},s}$ as ${\mathbb{Z}}$-modules. So the local system $(R^{1}\pi_{*}{\mathbb{Z}}_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})})^{\vee}$ is ${\mathbb{V}}_{{\mathbb{Z}}}$. In particular, ${\mathscr{V}}_{s}=H_{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}},{\mathbb{C}}))$.
The dual of $H^{1,0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\subseteq H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$ gives $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})^{\vee}\rightarrow\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$. Thus the de Rham–Betti comparison $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\cong H^{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{C}})$ gives rise to a Hodge filtration ${\mathscr{F}}_{{\mathscr{J}}}^{\bullet}$ on ${\mathscr{V}}$ defined by ${\mathscr{F}}_{{\mathscr{J}}}^{1}=0$, ${\mathscr{F}}_{{\mathscr{J}}}^{-1}={\mathscr{V}}$ and
$({\mathscr{F}}_{{\mathscr{J}}}^{0})_{s}=\ker({\mathscr{V}}_{s}=H_{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{C}})\rightarrow\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}))\quad\text{for all }s\in S.$
Lemma C.4.
${\mathscr{F}}_{{\mathscr{J}}}^{0}={\mathscr{F}}^{n}$.
Proof.
We computed in the proof of Lemma C.2 that ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{V}}_{s}/({\mathscr{F}}^{n}_{s}+{\mathbb{V}}_{{\mathbb{Z}},s})$. Hence $\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})={\mathscr{V}}_{s}/{\mathscr{F}}^{n}_{s}$. So $({\mathscr{F}}_{{\mathscr{J}}}^{0})_{s}={\mathscr{F}}^{n}_{s}$ by defintion. We are done. ∎
By definition, ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}(n))$ and ${\mathbb{V}}_{{\mathbb{Z}}}(n)$ is of weight $-1$. Hence in the rest of the paper, without loss of generality, we assume $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ to have weight $-1$ for the discussion on intermediate Jacobians.
C.2. Betti foliation
Next, we discuss the Betti foliation on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. Write $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ for the natural projection and $d$ for the relative dimension.
The holomorphic vector bundle ${\mathscr{V}}$ is endowed with a connection $\nabla$ whose sheaf of the horizontal sections is ${\mathbb{V}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}}{\mathbb{C}}$.
For each open subset $\Delta\subseteq S$, We have the following exact sequence.
$0\longrightarrow{\mathscr{F}}_{{\mathscr{J}}}^{0}{\mathscr{V}}_{\Delta}\longrightarrow{\mathscr{V}}_{\Delta}\longrightarrow\mathrm{Lie}(\pi^{-1}(\Delta)/\Delta)\rightarrow 0$
Locally on $S$, the local system ${\mathbb{V}}_{{\mathbb{Z}}}|_{\Delta}$ is trivial and so is $({\mathscr{V}},\nabla)|_{\Delta}$. Thus ${\mathbb{V}}_{{\mathbb{Z}}}|_{\Delta}\subseteq{\mathscr{V}}_{\Delta}$ becomes ${\mathbb{Z}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta$, which extends to ${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta$. Notice that $({\mathbb{R}}^{2d}\times\Delta)\cap{\mathscr{F}}_{{\mathscr{J}}}^{0}{\mathscr{V}}_{\Delta}$ is 00 on each fiber by weight reasons. Hence ${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta\rightarrow\mathrm{Lie}(\pi^{-1}(\Delta)/\Delta)$ is a real analytic diffeomorphism, and over each $s\in\Delta$ it becomes a group homomorphism. Moreover, the image of $\{r\}\times\Delta$ is easily seen to be complex analytic for any $r\in{\mathbb{R}}^{2d}$.
Taking the quotient of ${\mathbb{V}}_{{\mathbb{Z}}}$ on both sides, we obtain a real analytic diffeomorphism $({\mathbb{R}}^{2d}/{\mathbb{Z}}^{2d})\times\Delta\xrightarrow{\sim}\pi^{-1}(\Delta)$. Let $b_{\Delta}\colon\pi^{-1}(\Delta)\rightarrow{\mathbb{R}}^{2d}/{\mathbb{Z}}^{2d}$ be the composite of its inverse with the projection to the first factor $({\mathbb{R}}^{2d}/{\mathbb{Z}}^{2d})\times\Delta\rightarrow{\mathbb{R}}^{2d}/{\mathbb{Z}}^{2d}$. Then, each fiber of $b_{\Delta}$ is complex analytic by the last sentence of the last paragraph.
The construction above patches to a real analytic homeomorphism
(C.1) ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\xrightarrow{\sim}{\mathbb{V}}_{{\mathbb{R}}}/{\mathbb{V}}_{{\mathbb{Z}}}.$
This gives a foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ which we call the Betti foliation. More concretely ${\mathscr{F}}_{\mathrm{Betti}}$ is defined as follows: for each $x\in\pi^{-1}(\Delta)$, the local leaf through $x$ is the fiber $b_{\Delta}^{-1}(b_{\Delta}(x))$. Each leaf is holomorphic to the discussion above. In fact, the Betti foliation is the unique foliation on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ which is everywhere transverse to the fibers of $\pi$ and whose set of leaves contains all torsion multisections.
For each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, the Betti foliation induces a decomposition $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$, and $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ induces a natural isomorphism $T_{x}{\mathscr{F}}_{\mathrm{Betti}}\cong T_{\pi(x)}S$. The translation on the torus ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$ yields a canonical isomorphism $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})=T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})$. Hence, we have a linear map
(C.2) $q_{x}\colon T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})\rightarrow T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})=T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)}),$
whose kernel is $T_{x}{\mathscr{F}}_{\mathrm{Betti}}$.
We close this subsection with the following discussion. Take a holomorphic section $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ of the intermediate Jacobian. Then we have a linear map, at each $s\in S$,
(C.3) $\nu_{\mathrm{Betti},s}\colon T_{s}S\xrightarrow{\mathrm{d}\nu}T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\xrightarrow{q_{\nu(s)}}T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s}).$
Notice that, since $\pi$ induces canonically $T_{\nu(s)}{\mathscr{F}}_{\mathrm{Betti}}=T_{s}S$, the map $\mathrm{d}\nu$ is exactly $(1,\nu_{\mathrm{Betti},s})\colon T_{s}S\rightarrow T_{s}S\oplus T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})=T_{s}S\oplus T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})=T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$.
C.3. Sections of intermediate Jacobians
In this subsection, we explain how any holomorphic section $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defines a family of mixed Hodge structures on $S$ which varies holomorphically (Definition A.4 without the Griffiths’ transversality) of weight $-1$ and 00.
First we define a local system ${\mathbb{E}}_{\nu}$ on $S$ associated with $\nu$. Write ${\mathscr{J}}$ for the sheaf of holomorphic sections of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$. Then we have the following exact sequence
$0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathscr{V}}/{\mathscr{F}}^{0}{\mathscr{V}}\rightarrow{\mathscr{J}}\rightarrow 0$
Taking cohomology yields the boundary map.
$c\colon H^{0}(S,{\mathscr{J}})\rightarrow H^{1}(S,{\mathbb{V}}_{{\mathbb{Z}}}).$
On the other hand, it is known that $H^{1}(S,{\mathbb{V}}_{{\mathbb{Z}}})$ can be canonically identified with $\mathrm{Ext}_{\mathrm{loc.sys}}({\mathbb{Z}}_{S},{\mathbb{V}}_{{\mathbb{Z}}})$, the isomorphism classes of the extensions of local systems (${\mathbb{Z}}_{S}$ by ${\mathbb{V}}_{{\mathbb{Z}}}$) on $S$, so $c(\nu)\in H^{1}(S,{\mathbb{V}}_{{\mathbb{Z}}})$ defines a local system ${\mathbb{E}}_{\nu}$ on $S$ fitting into the short exact sequence $0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{E}}_{\nu}\rightarrow{\mathbb{Z}}_{S}\rightarrow 0$. Notice that this defines a weight filtration $W_{\bullet}$ on ${\mathbb{E}}_{\nu}$ of weight $-1$ and 00, by letting $W_{-2}{\mathbb{E}}_{\nu}=0$ and $W_{-1}{\mathbb{E}}_{\nu}={\mathbb{V}}_{{\mathbb{Z}}}$ and $W_{0}{\mathbb{E}}_{\nu}={\mathbb{E}}_{\nu}$. Notice that in the category of local systems, this exact sequence is split after $\otimes{\mathbb{Q}}$ if $S$ is simply connected.
Carlson [Car85] proved that ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ is canonically isomorphic to $\mathrm{Ext}_{\mathrm{MHS}}({\mathbb{Z}}(0),{\mathbb{V}}_{{\mathbb{Z}},s})$, the set of congruence classes of extensions of ${\mathbb{Z}}(0)$ by ${\mathbb{V}}_{{\mathbb{Z}},s}$ in the category of mixed Hodge structures; see for example [BZ14, Thm. 8.4.2]. In our context, this says the following. For each $s\in S({\mathbb{C}})$, $\nu(s)$ defines a mixed Hodge structure on ${\mathbb{E}}_{\nu,s}$, and hence a Hodge filtration ${\mathscr{F}}_{{\mathbb{E}}}^{\bullet}$ on the fiber ${\mathbb{E}}_{\nu,s}\otimes{\mathbb{C}}$. Since $\nu$ is holomorphic, the fiberwise Hodge filtrations give rise to a Hodge filtration on ${\mathbb{E}}_{\nu}$. So $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ is a family of mixed Hodge structures on $S$ which varies holomorphically and has weight $-1$ and 00. We are done.
Moreover, if ${\mathbb{V}}_{{\mathbb{Z}}}$ is polarized, then ${\mathbb{E}}_{\nu}$ is graded-polarized with this polarization on ${\mathbb{V}}_{{\mathbb{Z}}}$ and the canonical polarization on ${\mathbb{Z}}(0)$ given by ${\mathbb{Q}}\otimes{\mathbb{Q}}\rightarrow{\mathbb{Q}}$, $a\otimes b\mapsto ab$.
C.4. Normal functions
[…]
Definition C.5.
A holomorphic section $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is called an admissible normal function if $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ is an admissible VMHS.
C.5. Normal functions arising from families of algebraic cycles
Let $X$ be a smooth irreducible projective variety over ${\mathbb{C}}$ with $\dim X=d$. For each $n$, the Chow group of $n$-cocycles $\mathrm{CH}^{n}(X)$ is the group of algebraic cycles of codimension $n$ on $X$ modulo rational equivalence. Denote by $\mathrm{CH}^{n}(X)_{\hom}$ the kernel of the cycle map $\mathrm{CH}^{n}(X)\rightarrow H^{2n}(X,{\mathbb{Z}})$. An $n$-cocycle $Z$ is said to be homologically trivial if its Chow class lies in $\mathrm{CH}^{n}(X)_{\hom}$. By abuse of notation, we will denote the Chow class of a (co)cycle $Z$ also by $Z$.
Let ${\mathscr{J}}^{n}(X)$ be the intermediate Jacobian corresponding to $H^{2n-1}(X,{\mathbb{Z}})$ from Example C.3.
The Abel–Jacobi map
(C.4) ${\mathrm{AJ}}\colon\mathrm{CH}^{n}(X)_{\hom}\longrightarrow{\mathscr{J}}^{n}(X)$
[…]
Now, we turn to the family version and define the corresponding normal function. Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers of dimension $d$. Let $Z$ be a family of homologically trivial $n$-cocycle in $X/S$, i.e. a formal sum of integral subschemes of $X$ which are flat and dominant over $S$ such that each fiber $Z_{s}$ is a homologically trivial $n$-cocycle of $X_{s}$.
Theorem C.6 ([EZ86]).
The holomorphic section
$\nu_{Z}\colon S\rightarrow{\mathscr{J}}^{n}(X/S),\qquad s\mapsto{\mathrm{AJ}}(Z_{s})$
It is an admissible normal function.
The following proposition is a simple application of the Abel–Jacobi map and the Betti foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on intermediate Jacobians introduced in §$\S$§C.2.
Proposition C.7.
Set $S^{\circ}:=S\setminus S_{{\mathscr{F}}}(1)$.
Assume $X/S$ and $Z$ are defined over $\overline{\mathbb{Q}}$. Then $[Z_{s}]\in\mathrm{CH}^{n}(X_{s})$ is non-torsion for any $s\in S^{\circ}({\mathbb{C}})\setminus S^{\circ}(\overline{\mathbb{Q}})$.
Proof.
Take $s\in S({\mathbb{C}})$, and set $\bar{s}$ to be the $\overline{\mathbb{Q}}$-Zariski closure of $s$ in $S$.
Assume $[Z_{s}]$ is a torsion point of $\mathrm{CH}^{n}(X_{s})$. Then $[Z_{t}]$ is a torsion point of $\mathrm{CH}^{n}(X_{t})$ for each $t\in\bar{s}(\overline{\mathbb{Q}})$, and $t\mapsto{\mathrm{AJ}}(Z_{t})$ is a torsion section of ${\mathscr{J}}^{n}(X/S)\times_{S}\bar{s}\rightarrow\bar{s}$. So $\nu_{Z}(\bar{s})\subseteq{\mathscr{F}}_{\mathrm{Betti}}$. If $s\not\in S(\overline{\mathbb{Q}})$, then $\dim\bar{s}\geq 1$, and hence $\bar{s}\subseteq S_{{\mathscr{F}}}(1)$ by definition. Now we are done. ∎
C.6. Normal functions associated with Gross–Schoen cycles and Ceresa cycles
Let $C$ be an irreducible smooth projective curve defined over a field $k$ of genus $g\geq 3$. Let $\xi\in C(k)$. We have:
    For each subset $T\subseteq\{1,2,3\}$, the modified diagonal $\Delta_{T}(C)=\{(x_{1},x_{2},x_{3}):x_{i}=\xi\text{ for }i\not\in T,~{}x_{j}=x_{j^{\prime}}\text{ for all }j,j^{\prime}\in T\}\in\mathrm{CH}^{2}(C^{3})$.
    The (classical) Abel–Jacobi map $i_{\xi}\colon C\rightarrow\mathrm{Jac}(C)$ sending $x\mapsto[x-\xi]$.
Both $\Delta_{T}(C)$ and $i_{\xi}$ can be extended to any $\xi=\sum_{i}n_{i}e_{i}\in\mathrm{Div}^{1}(C)(k)=\mathrm{Pic}^{1}(C)(k)$. This is classical and direct for $i_{\xi}$, and for $\Delta_{T}(C)$ one can define $\Delta_{123}(C)=\{(c,c,c):c\in C\}$, $\Delta_{12}(C)=\sum n_{i}\{(c,c,e_{i}):c\in C\}$, $\Delta_{1}(C)=\sum_{i,j}n_{i}n_{j}\{(c,e_{i},e_{j}):c\in C\}$, etc.
For any $\xi\in\mathrm{Pic}^{1}(C)$, define the Gross–Schoen cycle based at $\xi$ (resp. the Ceresa cycle based at $\xi$) to be:
  • (Gross–Schoen cycle based at $\xi$) $\Delta_{{\mathrm{GS}},\xi}(C):=\Delta_{123}(C)-\Delta_{12}(C)-\Delta_{13}(C)-\Delta_{23}(C)+\Delta_{1}(C)+\Delta_{2}(C)+\Delta_{3}(C)\in\mathrm{CH}^{2}(C^{3})$.
  • (Ceresa cycle based at $\xi$) ${\mathrm{Ce}}_{\xi}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in\mathrm{CH}^{g-1}(\mathrm{Jac}(C))$.
${\mathrm{Ce}}_{\xi}(C)$ is clearly homologically trivial because $[-1]$ acts trivially on even-degree cohomology groups. It is not hard to check that $\Delta_{{\mathrm{GS}},\xi}(C)$ is also homologically trivial; this follows for example from [GS95, Prop. 3.1] and because the map $\mathrm{Pic}^{1}(C)\rightarrow\mathrm{CH}^{2}(C^{3})$, $\xi\mapsto\Delta_{{\mathrm{GS}},\xi}(C)$ is a group homomorphism.
Let $\omega_{C}$ be the canonical divisor on $C$.
Definition C.8.
Let $\xi\in\mathrm{Pic}^{1}(C)(k)$ such that $(2g-2)\xi=\omega_{C}$. Define the (canonical) Gross–Schoen cycle and the (canonical) Ceresa cycle of $C$ to be:
  • (Gross–Schoen cycle) $\Delta_{{\mathrm{GS}}}(C):=\Delta_{{\mathrm{GS}},\xi}(C)\in\mathrm{CH}^{2}(C^{3})_{\hom}$.
  • (Ceresa cycle) ${\mathrm{Ce}}(C):={\mathrm{Ce}}_{\xi}(C)\in\mathrm{CH}^{g-1}(\mathrm{Jac}(C))_{\hom}$.
There are finitely many $\xi$’s with $(2g-2)\xi=\omega_{C}$ and each two differ from a $(2g-2)$-torsion. Hence, the Gross–Schoen cycle and the Ceresa cycle of $C$ are well-defined at $(2g-2)$-torsions.
We will associate normal functions with the Schoen cycles and Ceresa cycles. First we have ${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))\in{\mathscr{J}}^{2}(C^{3})={\mathscr{J}}(H^{3}(C^{3},{\mathbb{Z}})(2))$ and ${\mathrm{AJ}}({\mathrm{Ce}}(C))\in{\mathscr{J}}^{g-1}(\mathrm{Jac}(C))={\mathscr{J}}(H^{2g-3}(\mathrm{Jac}(C),{\mathbb{Z}})(g-1))$. Using the Poincaré duality on $\mathrm{Jac}(C)$ and on $C$, we get $H^{2g-3}(\mathrm{Jac}(C),{\mathbb{Z}})(g-1)=H_{3}(\mathrm{Jac}(C),{\mathbb{Z}})(-1)=\bigwedge^{3}H_{1}(C,{\mathbb{Z}})(-1)=\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2)$. So ${\mathrm{AJ}}({\mathrm{Ce}}(C))\in{\mathscr{J}}(\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2))$.
We have $\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2)\subseteq H^{3}(C^{3},{\mathbb{Z}})(2)$ in the following way. First, the Künneth formula gives a decomposition $H^{3}(C^{3},{\mathbb{Z}})=H^{1}(C,{\mathbb{Z}})^{\otimes 3}\bigoplus H^{1}(C,{\mathbb{Z}})(-1)^{\oplus 6}$. Next, $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$, as a subspace of $H^{3}(C^{3},{\mathbb{Z}})$, has a basis consisting of $\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3}$, with $\alpha_{j}$ the pullback of an element in $H^{1}(C,{\mathbb{Z}})$ under the $j$-th projection $C^{3}\rightarrow C$. The symmetric group $S_{3}$ acts naturally on $C^{3}$, and this induces an action of $S_{3}$ on $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$ with $\sigma(\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3})=\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$ for each $\sigma\in S_{3}$. Then $(H^{1}(C,{\mathbb{Z}})^{\otimes 3})^{S_{3}}=\bigwedge^{3}H^{1}(C,{\mathbb{Z}})$ for this action, with each member having a basis consisting of $\sum_{\sigma\in S_{3}}\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$.
It is easy to check that the pushforward of $\Delta_{{\mathrm{GS}}}(C)$ to any two factors of $C^{3}$ is trivial. Hence ${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))\in{\mathscr{J}}(H^{1}(C,{\mathbb{Z}})^{\otimes 3}(2))$. Moreover, the modified diagonal is easily seen to be invariant under the action of $S_{3}$ on $C^{3}$. Thus ${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))\in{\mathscr{J}}(\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2))$.
Lemma C.9.
${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))=3{\mathrm{AJ}}({\mathrm{Ce}}(C))$.
Proof.
Consider $\iota_{3}\colon C^{3}\rightarrow\mathrm{Jac}(C)$, $(c_{1},c_{2},c_{3})\mapsto i_{\xi}(c_{1})+i_{\xi}(c_{2})+i_{\xi}(c_{3})$. The difference $(\iota_{3})_{*}\Delta_{{\mathrm{GS}}}(C)-3{\mathrm{Ce}}(C)$ was computed by the second-named author [Zha10, Thm. 1.5.5]. More precisely, the Fourier–Mukai transformation yields a spectrum decomposition $C=\sum_{j=0}^{g-1}C_{j}$ in $\mathrm{CH}^{g-1}(\mathrm{Jac}(C))$, with $[n]_{*}C_{j}=n^{2+j}C_{j}$ for all $n\in{\mathbb{Z}}$ and $j\in\{0,\ldots,g-1\}$, and [Zha10, Thm. 1.5.5 and its proof] implies that
$(\iota_{3})_{*}\Delta_{{\mathrm{GS}}}(C)-3{\mathrm{Ce}}(C)=\sum\nolimits_{j\geq 2}a_{j}C_{j}$
For appropriate numbers $a_{j}$. The multiplication $[n]\colon\mathrm{Jac}(C)\rightarrow\mathrm{Jac}(C)$ induces $\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})\rightarrow\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})$, $x\mapsto n^{3}x$. Hence $n^{2+j}{\mathrm{AJ}}(C_{j})=n^{3}{\mathrm{AJ}}(C_{j})$ for all $n\in{\mathbb{Z}}$, and so ${\mathrm{AJ}}(C_{j})=0$ for all $j\geq 2$.
Hence the conclusion follows because the induced map of $\iota_{3}$ is the natural projection.∎
Now, we turn to defining and studying the normal functions. Let ${\mathscr{M}}_{g}$ be the moduli space of curves of genus $g$, and let $f\colon{\mathscr{C}}_{g}\rightarrow{\mathscr{M}}_{g}$ be the universal curve. We have a VHS $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}\rightarrow{\mathscr{M}}_{g}$. The fiberwise polarization $q_{H}\colon H^{1}(C,{\mathbb{Z}})\otimes H^{1}(C,{\mathbb{Z}})\xrightarrow{\cup}H^{2}(C,{\mathbb{Z}})\cong{\mathbb{Z}}(-1)$, with $\cup$ the cup product (the dual of the intersection pairing) on $H^{1}(C,{\mathbb{Z}})$, induces a polarization on $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}$.
Now we have two normal functions
(C.5) $\nu_{{\mathrm{GS}}}\colon\mathbb{M}_{g}\longrightarrow{\mathscr{J}}(\bigwedge\nolimits^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}),$ $\qquad s\mapsto{\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})),$
$\nu_{{\mathrm{Ce}}}\colon\mathbb{M}_{g}\longrightarrow{\mathscr{J}}(\bigwedge\nolimits^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}),$ $\qquad s\mapsto{\mathrm{AJ}}({\mathrm{Ce}}({\mathscr{C}}_{s})).$
We can do better.
Proposition C.10.
Let ${\mathbb{V}}_{{\mathbb{Z}}}$ be the kernel of the morphism of VHS (called the contractor)
(C.6) $c\colon\bigwedge\nolimits^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}(2)\longrightarrow R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}(1)$
fiberwise defined by $x\wedge y\wedge z\mapsto q_{H}(y,z)x+q_{H}(z,x)y+q_{H}(x,y)z$. Then
  1. (i)
    both $\nu_{{\mathrm{GS}}}$ and $\nu_{{\mathrm{Ce}}}$ have images in ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$;
  2. (ii)
    neither $\nu_{{\mathrm{GS}}}$ nor $\nu_{{\mathrm{Ce}}}$ is a torsion section;
  3. (iii)
    ${\mathbb{V}}_{{\mathbb{Z}}}$ is an irreducible VHS on ${\mathscr{M}}_{g}$, i.e. the only sub-VHSs of ${\mathbb{V}}_{{\mathbb{Z}}}$ are 00 and itself.
Proof.
For each $s\in S({\mathbb{C}})$, $\nu_{{\mathrm{GS}}}(s)\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ by [Zha10, Lem. 5.1.5], and hence $\nu_{{\mathrm{Ce}}}(s)\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ by Lemma C.9. So (i) holds.
For (ii), it suffices to find a curve $C$ of genus $g$ such that ${\mathrm{AJ}}({\mathrm{Ce}}(C))$ is not torsion. There are many examples of such curves in existing literature. Alternatively, the result for $\nu_{{\mathrm{Ce}}}$ can be already deduced from Ceresa’s original argument in [Cer83].
For (iii), it suffices to prove that $V:={\mathbb{V}}_{{\mathbb{Z}},s}\otimes{\mathbb{Q}}$ is a simple $\mathrm{Sp}_{2g}$-module for one (and hence all) $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, or equivalently $V$ is an irreducible representation of $\mathrm{Sp}_{2g}$. This is a standard result of the representation theory for $\mathfrak{sp}_{2g}$, see for example [FH13, Thm. 17.5]. ∎
Appendix D Metrized Poincaré bundle and local height pairing
Let $S$ be a smooth irreducible quasi-projective variety.
D.1. Height pairing at Archimedean places
Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers of dimension $d$. Let $p,q$ be non-negative integers such that $p+q=d+1$.
Let $Z$, resp. $W$, be a family of homologically trivial $p$-cocycle in $X/S$, resp. a family of homologically trivial $q$-cocycle in $X/S$. Assume $Z$ and $W$ have disjoint supports over the generic fiber, then up to replacing $S$ by a Zariski open dense subset, we may assume that $Z_{s}$ and $W_{s}$ have disjoint supports for all $s\in S({\mathbb{C}})$.
In Arakelov’s theory, the local height pairing at archimedean places in this context is defined by
(D.1) $\langle Z_{s},W_{s}\rangle_{\infty}=\int_{Z_{s}}\eta_{W_{s}}.$
where $\eta_{W_{s}}$ is a Green’s current for $W_{s}$. This pairing is known to be symmetric.
The two families of homologically trivial cocycles give rise to two admissible normal functions.
$\nu_{Z}\colon S\rightarrow{\mathscr{J}}^{p}(X/S)\quad\text{ and }\quad\nu_{W}\colon S\rightarrow{\mathscr{J}}^{q}(X/S).$
A particularly important case is when $p=q=\frac{d+1}{2}$ and that $Z$ and $W$ are rationally equivalent. In this case, we obtain the height pairings $\langle Z_{s},Z_{s}\rangle_{\infty}$, or simply the height of $Z_{s}$. More precisely, let $Z$ be a family of homologically trivial $p$-cocycle in $X/S$ with $p=\frac{d+1}{2}$, and let $\eta$ be the generic point of $S$. By Moving Lemma, there exists a $p$-cocycle $W_{\eta}$ of $X_{\eta}$ whose class in $\mathrm{CH}^{p}(X_{\eta})$ is the same as the class of $Z_{\eta}$ such that $W_{\eta}$ and $Z_{\eta}$ have disjoint supports. Then $W_{\eta}$ extends to a family of homologically trivial $p$-cocycle $W$ in $X/S$, and $W_{s}$ and $Z_{s}$ have disjoint supports for $s$ in a Zariski open dense subset of $S$. Then we set $\langle Z_{s},Z_{s}\rangle_{\infty}=\langle Z_{s},W_{s}\rangle_{\infty}$. Notice that in this case $\nu_{Z}=\nu_{W}$.
Hain [Hai90, §$\S$3.3] related the local height pairing at archimedean places (D.1) to the metrized Poincaré bundle. We will review this in the next subsection.
D.2. Metrized Poincaré bundle
Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a polarized VHS on $S$ of weight $-1$. The intermediate Jacobian ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is a torus fibration by Lemma C.2. The dual torus fibration $\mathrm{Pic}^{0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}))$ can be described as follows. Set ${\mathbb{V}}_{{\mathbb{Z}}}^{\vee}:={\mathrm{Hom}}_{\mathrm{VHS}}({\mathbb{V}}_{{\mathbb{Z}}},{\mathbb{Z}}(1)_{S})$ with the natural Hodge filtration. Then ${\mathbb{V}}_{{\mathbb{Z}}}^{\vee}$ is a VHS on $S$ of weight $-1$, and there is a canonical isomorphism ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})=\mathrm{Pic}^{0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}))$.
The general theory of biextension says that there exists a unique line bundle ${\mathscr{P}}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})$ satisfying the following properties:
  1. (i)
    Over each $s\in S({\mathbb{C}})$, we have ${\mathscr{P}}|_{\{0\}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}\cong{\mathscr{O}}_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}$,
  2. (ii)
    over each $s\in S({\mathbb{C}})$, ${\mathscr{P}}|_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}\times\{\lambda\}}$ represents $\lambda\in\mathrm{Pic}^{0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}_{s}$,
  3. (iii)
    $\epsilon^{*}{\mathscr{P}}\cong{\mathscr{O}}_{S}$ for the zero section $\epsilon$ of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}\rightarrow S$.
Moreover, ${\mathscr{P}}$ can be endowed a canonical Hermitian metric $\|\cdot\|_{\mathrm{can}}$ uniquely determined by the following properties: (i) the curvature of $\overline{{\mathscr{P}}}$ is translation invariant on each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}$; (ii) $\epsilon^{*}\overline{{\mathscr{P}}}\cong({\mathscr{O}}_{S},\|\cdot\|_{\mathrm{triv}})$ for the trivial metric on ${\mathscr{O}}_{S}$.
Definition D.1.
The metrized line bundle $\overline{{\mathscr{P}}}:=({\mathscr{P}},\|\cdot\|_{\mathrm{can}})$ is called the metrized Poincaré bundle.
[…]
Going from ${\mathbb{Z}}$ to ${\mathbb{R}}$-coefficients, one has the set ${\mathscr{B}}({\mathbb{V}}_{{\mathbb{R}},s})$ of mixed ${\mathbb{R}}$-Hodge structures of weight $0,-1,-2$ whose weight graded pieces are ${\mathbb{R}}(0)$, ${\mathbb{V}}_{{\mathbb{R}},s}$ and ${\mathbb{R}}(1)$. One can check that ${\mathscr{B}}({\mathbb{V}}_{{\mathbb{R}},s})$ is canonically isomorphic to ${\mathbb{R}}$. Then on ${\mathscr{P}}^{*}_{s}={\mathscr{B}}({\mathbb{V}}_{{\mathbb{Z}},s})$,
$\|p\|_{\mathrm{can}}=e^{\mathrm{f}_{s}(p)}$
where $\mathrm{f}_{s}$ is the forgetful map ${\mathscr{B}}({\mathbb{V}}_{{\mathbb{Z}},s})\rightarrow{\mathscr{B}}({\mathbb{V}}_{{\mathbb{R}},s})={\mathbb{R}}$.
Now, let us go back to the setting of §$\S$§D.1. Recall the section $(\nu_{Z},\nu_{W})\colon S\rightarrow{\mathscr{J}}^{p}(X/S)\times_{S}{\mathscr{J}}^{q}(X/S)$ obtained from the two families of homologically trivial cocycles $Z$ and $W$.
[…]
Hain [Hai90, Prop. 3.3.2] constructed a section $\beta_{Z,W}$ of the line bundle $(\nu_{Z},\nu_{W})^{*}{\mathscr{P}}\rightarrow S$ in view of the Hodge-theoretic construction of ${\mathscr{P}}$ explained below Definition D.1, and proved [Hai90, Prop. 3.3.12] $\log\|\beta_{Z,W}(s)\|=-\int_{Z_{s}}\eta_{W_{s}}$ for all $s\in S({\mathbb{C}})$. To summarize, we have
Proposition D.2 (Hain).
We have $-\log\|\beta_{Z,W}(s)\|=\langle Z_{s},W_{s}\rangle_{\infty}$ for all $s\in S({\mathbb{C}})$.
D.3. Metrized tautological bundle
As explained at the end of §$\S$§D.1, we are often more interested in the case where ${\mathscr{J}}^{p}(X/S)$ is self-dual and $\nu_{Z}=\nu_{W}$. In this case, $(\nu_{Z},\nu_{W})^{*}\overline{{\mathscr{P}}}=\nu_{Z}^{*}\Delta^{*}\overline{{\mathscr{P}}}$ for the diagonal $\Delta\colon{\mathscr{J}}^{p}(X/S)\rightarrow{\mathscr{J}}^{p}(X/S)\times_{S}{\mathscr{J}}^{p}(X/S)$. We discuss this case in this and the next subsections in a more general setting.
Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a VHS on $S$ of weight $-1$, with a polarization ${\mathscr{Q}}\colon{\mathbb{V}}_{{\mathbb{Z}}}\otimes{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{Z}}(1)_{S}$. Then ${\mathscr{Q}}$ induces a morphism of VHS ${\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}^{\vee}$, and hence a morphism between the intermediate Jacobians $i_{{\mathscr{Q}}}\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\longrightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}$. We thus have a morphism
$\Delta_{{\mathscr{Q}}}=(1,i_{{\mathscr{Q}}})\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}.$
This is the case, for example, for Poincaré duality, where the polarization is given by the cup product.
Definition D.3.
The metrized line bundle $\overline{{\mathscr{P}}^{\Delta_{\mathscr{Q}}}}:=\Delta_{{\mathscr{Q}}}^{*}\overline{{\mathscr{P}}}$ is called the metrized tautological bundle on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. When the polarization ${\mathscr{Q}}$ is clear in context, we simply denote it by $\overline{{\mathscr{P}}^{\Delta}}$.
By [HR04, Prop. 7.1 and 7.3], the curvature form $c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ is a closed $2$-form, uniquely determined by the following properties.
Proposition D.4.
[…]
  1. (i)
    Restricted to each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}$, $c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ is the unique translation invariant $2$-form $\omega_{s}$ given by $2{\mathscr{Q}}_{s}$,
  2. (ii)
    $c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ vanishes along each leaf of the Betti foliation.
Here is a more explicit formula for $c_{1}(\overline{{\mathscr{P}}^{\Delta}})$. Write $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ for the natural projection. Recall the linear map (C.2)
$q_{x}\colon T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)}).$
with kernel $T_{x}{\mathscr{F}}_{\mathrm{Betti}}$. Now for any $v_{1},v_{2}\in T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, we have
(D.2) $c_{1}(\overline{{\mathscr{P}}^{\Delta}})(v_{1},v_{2})=2{\mathscr{Q}}_{\pi(x)}(q_{x}(v_{1}),q_{x}(v_{2})).$
D.4. Pullback by admissible normal functions
Retain the setup in §$\S$§D.3. Now we turn to admissible normal functions $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$.
Hain proved [Hai13, Thm. 13.1] that $\nu^{*}c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ is a semi-positive $(1,1)$-form.
Definition D.5.
The semi-positive $(1,1)$-form $\nu^{*}c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ is called the Betti form associated with $\nu$, which we denote by $\beta_{\nu}$.
The Betti foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ induces a linear map (C.3) at each $s\in S({\mathbb{C}})$
$\nu_{\mathrm{Betti},s}\colon T_{s}S\xrightarrow{\mathrm{d}\nu}T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\xrightarrow{q_{\nu(s)}}T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s}).$
By (D.2), $\beta_{\nu}(u,\bar{u})=2{\mathscr{Q}}_{s}(\nu_{\mathrm{Betti},s}(u),\overline{\nu_{\mathrm{Betti},s}(u)})$ for any $u\in T_{s}S$. By Griffiths’ transversality, $\nu_{\mathrm{Betti},s}(u)\in{\mathbb{V}}_{s}^{-1,0}$. So
(D.3) $\beta_{\nu}(u,\bar{u})\geq 0\text{ for all }u\in T_{s}S,\text{ with equality if and only if }\nu_{\mathrm{Betti},s}(u)=0.$
Notice that this also explains the semi-positivity of $\beta_{\nu}$. Now we use (D.3) to prove the following proposition.
Proposition D.6.
For any $s\in S({\mathbb{C}})$, the following are equivalent:
  1. (i)
    $(\beta_{\nu}^{\wedge\dim S})_{s}\not=0$;
  2. (ii)
    $\dim\nu_{\mathrm{Betti},s}(T_{s}S)=\dim S$.
This proposition has the following immediate corollary.
Corollary D.7.
$\beta_{\nu}^{\wedge\dim S}\not\equiv 0$ if and only if $\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S)=\dim S$.
We will call $\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S)$ the Betti rank of $\nu$ and denote it by $r(\nu)$.
Proof of Proposition D.6.
Assume (i) is false, i.e. $(\beta_{\nu}^{\wedge\dim S})_{s}=0$. By (D.3) there exists $0\not=u\in T_{s}S$ with $\nu_{\mathrm{Betti},s}(u)=0$. Thus $\ker\nu_{\mathrm{Betti},s}\not=0$, and therefore $\dim\nu_{\mathrm{Betti},s}(T_{s}S)<\dim S$. So (ii) is also false.
Assume (ii) is false. Then there exists $0\not=u\in\ker\nu_{\mathrm{Betti},s}$. By (D.3), $\beta_{\nu}(u,\bar{u})=0$. Thus $u$ is an eigenvector of the Hermitian matrix defining $\beta_{\nu}$ with eigenvalue 00. Hence the determinant of this matrix is 00, so $\beta_{\nu}^{\wedge\dim S}=0$ at $s$. So (i) is also false. ∎

### v2

@@
 We consider the Beilinson–Bloch heights and Abel–Jacobian periods of homologically trivial Chow cycles in families.
-For the Beilinson–Bloch heights, we show that for any $g\geq 3$, there is a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$, the coarse moduli of curves of genus $g$ over $\overline{\mathbb{Q}}$, such that the heights of Ceresa cycles and Gross–Schoen cycles over $U$ satisfy the Northcott property.
-For the Abel–Jacobi periods, we provide an algebraic criterion for the existence of a Zariski open dense subset of any family such that all cycles not defined over $\overline{\mathbb{Q}}$ are non-torsion and verify that this criterion holds true for Ceresa cycles and Gross–Schoen cycles.
+For the Beilinson–Bloch heights, we show that for any $g\geq 3$, we can construct a Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of ${\mathscr{M}}_{g}$, the coarse moduli of curves of genus $g$ over ${\mathbb{Q}}$, such that the heights of Ceresa cycles and Gross–Schoen cycles over ${\mathscr{M}}_{g}^{\mathrm{amp}}$ have a lower bound and satisfy the Northcott property.
+For the Abel–Jacobi periods, we prove that the Ceresa and Gross–Schoen cycle over every $\overline{\mathbb{Q}}$ point in ${\mathscr{M}}_{g}^{\mathrm{amp}}$ is non-torsion.
+We also prove results and make conjectures for any family $X\rightarrow S$ of smooth projective varieties and homologically trivial cocycles $Z$ defined over a number field.
+Using the Betti strata, we construct a Zariski open subset $S^{\mathrm{amp}}$ of $S_{{\mathbb{C}}}$ over which $Z_{s}$ is non-torsion for all transcendental points $s$ and prove a criterion for $S^{\mathrm{amp}}$ to be dense.
+We conjecture the rationality of $S^{\mathrm{amp}}$ and also conjecture that $S^{\mathrm{amp}}$ is precisely the open locus on which the Beilinson–Bloch height of $Z_{s}$ has a lower bound and satisfies the Northcott property.
+Dedicated to Gerd Faltings’s 70th birthday
 1. Introduction
-This paper aims to initiate a project to study Beilinson–Bloch heights and Abel–Jacobi periods for homologically trivial cycles in families.
-We have achieved two goals: the Northcott property for the heights of the Gross–Schoen cycles and Cerasa cycles and an algebraic criterion for the non-degeneracy of Abel–Jacobi periods for the general family of cycles.
-1.1. Gross–Schoen and Ceresa cycles
-For a smooth projective irreducible curve $C$ defined over a filed $k$ of $g\geq 3$ and a class $\xi\in\mathrm{Pic}^{1}(C)$ such that $(2g-2)\xi=\omega_{C}$, we have two homologically trivial 1-cycles:
+For a curve over ${\mathbb{C}}$, the classical Abel–Jacobi map is a bijection
+${\mathrm{AJ}}\colon\mathrm{Pic}^{0}(C){\overset{\sim}{{\longrightarrow}}}{\mathrm{Jac}}(C).$
+When $C$ is defined over a number field $k$, one can define the Néron–Tate height on ${\mathrm{Jac}}(C)(\bar{k})$ using the canonical polarization.
+This height has a Northcott property, which gives positivity of the height pairing, which in turn is an important ingredient for the proof of the Mordell–Weil theorem and the formulation of the Birch and Swinnerton-Dyer conjecture.
+Notice that by the Hodge index theorem of Faltings and Hriljac, these heights can also be defined as arithmetic intersections on integral models of $C$.
+In the 1950s-1960s, Weil and Griffiths extended the Abel–Jacobi map to homologically trivial cycles (say of codimension $n$) on smooth and projective varieties $X$ over ${\mathbb{C}}$
+${\mathrm{AJ}}^{n}\colon{\mathrm{Ch}}^{n}(X)_{\mathrm{hom}}{\longrightarrow}{\mathrm{Jac}}^{n}(X)$
+where ${\mathrm{Jac}}^{n}(X)$ are called intermediate Jacobians and are only complex tori in general.
+In the 1980s, Beilinson [Blo84] and Bloch [Blo84] independently proposed a conditional definition of heights for these cycles using integral models.
+They conjectured the positivity of their heights, an extension of the Mordell–Weil theorem, and an extension of the Birch and Swinnerton-Dyer conjecture.
+However, the positivity cannot be proved via a Northcott property because of the lack of algebraicity of ${\mathrm{Jac}}^{n}(X)$.
+This paper aims to initiate a project to study Beilinson–Bloch heights and Abel–Jacobi periods for homologically trivial cycles in a family of smooth and projective varieties $f\colon X\rightarrow S$ parametrized by a quasi-projective variety $S$.
+We have achieved two goals:
   1. (1)
-    the Gross–Schoen cycle ${\mathrm{GS}}(C):=\Delta_{\xi}(C)\in{\mathrm{Ch}}_{1}(C^{3})$ obtained by modified the diagonal cycle in $C^{3}$ using base class $\xi$ ([GS95, Zha10]);
+    The generic positivity and Northcott property for the Beilinson–Bloch heights of the Gross–Schoen cycles and Cerasa cycles parametrized by moduli of curves.
+These are the simplest situations where the Beilinson–Bloch heights are unconditionally defined.
   2. (2)
-    the Ceresa cycle by ${\mathrm{Ce}}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in{\mathrm{Ch}}_{1}({\mathrm{Jac}}(C))$ defined by embedding $i_{\xi}\colon C{\longrightarrow}{\mathrm{Jac}}(C)$ via $\xi$.
-Up to torsions, the definition of these cycles does not depend on the choice of $\xi$.
-Let ${\mathscr{M}}_{g}$ be the moduli space of smooth projective curves of genus $g$. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathscr{C}}_{s}$ the curve parametrized by $s$.
+    An algebraic criterion for the non-degeneracy of Abel–Jacobi periods for general families of cycles.
+Our method of proof, roughly speaking, is to utilize the intermediate Jacobians to get some geometric positivity and then extend it to some arithmetic positivity.
+In the following, we describe the main results and the methods of proof in detail.
+1.1. Main results on Gross–Schoen and Ceresa cycles
+For a smooth projective irreducible curve $C$ defined over a field $k$ and a class $\xi\in\mathrm{Pic}^{1}(C)$, we have two homologically trivial 1-cycles:
+  • the Gross–Schoen cycle based at $\xi$ defined by ${\mathrm{GS}}_{\xi}(C):=\Delta_{{\mathrm{GS}},\xi}(C)\in{\mathrm{Ch}}_{1}(C^{3})$ obtained by modified the diagonal cycle in $C^{3}$ using base class $\xi$ ([GS95, Zha10]);
+  • the Ceresa cycle based at $\xi$ defined by ${\mathrm{Ce}}_{\xi}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in{\mathrm{Ch}}_{1}({\mathrm{Jac}}(C))$ defined by embedding $i_{\xi}\colon C{\longrightarrow}{\mathrm{Jac}}(C)$ via $\xi$ ([Cer83]).
+Let ${\mathscr{M}}_{g}$ be the moduli space of smooth projective curves of genus $g$, let ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$ be the universal curve, and let $\pi\colon\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g}){\longrightarrow}{\mathscr{M}}_{g}$ be the relative Picard group of degree $1$ on ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$.
+For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathscr{C}}_{s}$ the curve parametrized by $s$.
+For each $\xi\in\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})({\mathbb{C}})$, to ease notation denote by
+${\mathrm{GS}}(\xi):={\mathrm{GS}}_{\xi}({\mathscr{C}}_{s})\quad\text{ and }\quad{\mathrm{Ce}}(\xi):={\mathrm{Ce}}_{\xi}({\mathscr{C}}_{s})$
+for $s=\pi(\xi)\in{\mathscr{M}}_{g}({\mathbb{C}})$.
+For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, a particularly important element $\xi_{s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ is such that the divisor class of $(2g-2)\xi_{s}$ is the canonical divisor class on ${\mathscr{C}}_{s}$; it is well-defined up to $(2g-2)$-torsions.
+Here are our main results about the Beilinson–Bloch heights for these cycles unconditionally defined by Gross–Schoen [GS95] and Künnemann [Kün01] respectively.
 Theorem 1.1.
-For each $g\geq 3$, there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$, and positive numbers $\epsilon$ and $c$, such that for any $s\in U(\overline{\mathbb{Q}})$,
-$\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{\mathrm{Fal}}(s)-c,$
-$\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{\mathrm{Fal}}(s)-c,$
-where $h_{\mathrm{Fal}}$ is the Faltings height on ${\mathscr{M}}_{g}(\overline{\mathbb{Q}})$.
+For each $g\geq 3$, there exist a Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$, such that for any $s\in{\mathscr{M}}_{g}^{\mathrm{amp}}(\overline{\mathbb{Q}})$ and any $\xi\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$, we have
+(1.1) $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{s}))-c,$
+$\langle{{\mathrm{Ce}}(\xi),{\mathrm{Ce}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{s}))-c,$
+where $h_{\mathrm{Fal}}$ is the Faltings height on ${\mathscr{M}}_{g}(\overline{\mathbb{Q}})$ and $h_{\mathrm{NT}}$ is the canonical Néron–Tate height on ${\mathrm{Jac}}({\mathscr{C}}_{s})(\overline{\mathbb{Q}})$.
+The proof will be executed in §$\S$6.3, in which we give a construction of ${\mathscr{M}}_{g}^{\mathrm{amp}}$.
+Indeed, we will take ${\mathscr{M}}_{g}^{\mathrm{amp}}$ to be ${\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, where ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is the Betti degeneracy locus (1.3) with the normal function $\nu$ being the Gross–Schoen normal function (C.6).
+A priori ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is only a real analytic set.
+But we show that ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is proper Zariski closed in ${\mathscr{M}}_{g,{\mathbb{C}}}$ by applying Ax-Schanuel multiple times; see Theorem 1.4 and Corollary 1.8.
+Finally, we show that it is defined over ${\mathbb{Q}}$ by volume calculation for adelic line bundles; see Corollary 6.4.
+Because of the importance of this Zariski open dense set ${\mathscr{M}}_{g}^{\mathrm{amp}}$ on which (1.1) holds, we will make the following definition: We call
+  1. (i)
+    ${\mathscr{M}}_{g}^{\mathrm{amp}}$ the ample locus of the Gross–Schoen and Ceresa cycles;
+  2. (ii)
+    ${\mathscr{M}}_{g}^{\mathrm{slim}}:={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{amp}}$ the slim locus of the Gross–Schoen and Ceresa cycles.
+By the explanation above, it is the ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+It certainly contains the hyperelliptic locus and also the locus of Picard curves by Laga–Shnidman [LS24, Thm.
+C].
+We expect ${\mathscr{M}}_{g}^{\mathrm{amp}}$ is the largest open subset of ${\mathscr{M}}_{g}$ so that the height inequality (1.1) in Theorem 1.1 holds; see Conjecture 1.12.
 Theorem 1.1 yields the following Northcott property immediately.
 Corollary 1.2 (Northcott property).
-For each $g\geq 3$, there exists a Zariski open dense subset of the ${\mathscr{M}}_{g}$ defined over $\overline{\mathbb{Q}}$ such that for any $H,D\in{\mathbb{R}}$,
-#$\#\{s\in U(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(s):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}<H\}<\infty,$
-#$\#\{s\in U(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(s):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}<H\}<\infty.$
-In the course of the proof, we also establish the following geometric result.
+For each $g\geq 3$ and the Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of the ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ as above, we have the following Northcott property: For any $H,D\in{\mathbb{R}}$,
+#$\#\{\xi\in\pi^{-1}({\mathscr{M}}_{g}^{\mathrm{amp}})(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(\xi):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}<H\}<\infty,$
+#$\#\{\xi\in\pi^{-1}({\mathscr{M}}_{g}^{\mathrm{amp}})(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(\xi):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{Ce}}(\xi),{\mathrm{Ce}}(\xi)}\rangle_{\mathrm{BB}}<H\}<\infty.$
+In the course of the proof, we also establish the following geometric result on the torsion locus.
+For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s}):={\mathrm{Ce}}(\xi_{s})$.
 Theorem 1.3.
-For each $g\geq 3$, there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ such that the followings hold true:
+For each $g\geq 3$ and the Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of the ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ as above, the followings hold true:
   1. (i)
-    ${\mathrm{GS}}({\mathscr{C}}_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s})$ are both non-torsion in the Chow groups for all $s\in U({\mathbb{C}})\setminus U(\overline{\mathbb{Q}})$.
+    ${\mathrm{GS}}({\mathscr{C}}_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s})$ are both non-torsion in the Chow groups for all $s\in{\mathscr{M}}_{g}^{\mathrm{amp}}({\mathbb{C}})\setminus{\mathscr{M}}_{g}^{\mathrm{amp}}(\overline{\mathbb{Q}})$.
   2. (ii)
-    there exist at most countably many $s\in U({\mathbb{C}})$ such that ${\mathrm{AJ}}({\mathrm{GS}}({\mathscr{C}}_{s}))$ or ${\mathrm{AJ}}({\mathrm{Ce}}({\mathscr{C}}_{s}))$ is torsion in the intermediate Jacobians.
-Hain [Ha24] also proved Theorem 1.3 with $U$ a non-empty analytic open subset of ${\mathscr{M}}_{g}^{\mathrm{an}}$ with a different method.
-When $g=3$, Hain’s result is completely explicit, while ours is not.
-The proofs of Theorem 1.1, Corollary 1.2 and Theorem 1.3 have two parts: the arithmetic part and the geometric part.
-Both parts are needed to prove Theorem 1.1 and Corollary 1.2, while Theorem 1.3 only relies on the geometric part.
-We will explain each part in more detail in §$\S$1.3.
+    there exist at most countably many $s\in{\mathscr{M}}_{g}^{\mathrm{amp}}({\mathbb{C}})$ such that ${\mathrm{AJ}}({\mathrm{GS}}({\mathscr{C}}_{s}))$ or ${\mathrm{AJ}}({\mathrm{Ce}}({\mathscr{C}}_{s}))$ is torsion in the intermediate Jacobians.
+Independently, Hain [Hai24] and Kerr–Tayou [KT24] proved Theorem 1.3 with different methods, with ${\mathscr{M}}_{g}^{\mathrm{amp}}$ replaced by a non-empty open subset $U$ of ${\mathscr{M}}_{g,{\mathbb{C}}}$ which is real analytic and complex algebraic respectively.
+The proofs of Theorem 1.1, Corollary 1.2, and Theorem 1.3 have two parts: the arithmetic part and the geometric part.
+Both parts are needed to prove Theorem 1.1, Corollary 1.2, and Theorem 1.3.
+We will explain each part in more detail in §$\S$1.4.
 Before moving on, let us mention that the geometric part of our paper works for any family of homologically trivial cycles and, more generally, for any admissible normal function.
-More specifically, for any smooth projective morphism $f\colon X\rightarrow S$ of algebraic varieties with irreducible fibers and any family of homologically trivial cycles $Z$, all defined over $\overline{\mathbb{Q}}$, we show that there exists a Zariski closed subset $S_{{\mathscr{F}}}(1)$ of $S$ such that: (i) $[Z_{s}]$ is non-torsion in ${\mathrm{Ch}}^{*}(X_{s})$ for every transcendental point $s$ of $S\setminus S_{{\mathscr{F}}}(1)$; (ii) there are at most countably many $s\in S({\mathbb{C}})$ outside $S_{{\mathscr{F}}}(1)$ such that ${\mathrm{AJ}}(Z_{s})$ is torsion in the intermediate Jacobian.
-We then prove a checkable criterion for $S_{{\mathscr{F}}}(1)\not=S$.
+We prove a formula for its Betti rank and prove the Zariski closedness of the Betti strata.
+A simple application is the following result on the torsion locus: For any smooth projective morphism $f\colon X\rightarrow S$ of algebraic varieties with irreducible fibers and any family of homologically trivial cycles $Z$, all defined over $\overline{\mathbb{Q}}$, we show that there exists a Zariski closed subset $S^{\mathrm{Betti}}(1)$ of $S_{{\mathbb{C}}}$ such that: (i) $[Z_{s}]$ is non-torsion in ${\mathrm{Ch}}^{*}(X_{s})$ for every transcendental point $s$ of $S\setminus S^{\mathrm{Betti}}(1)$; (ii) there are at most countably many $s\in S({\mathbb{C}})$ outside $S^{\mathrm{Betti}}(1)$ such that ${\mathrm{AJ}}(Z_{s})$ is torsion in the intermediate Jacobian.
+We then prove a checkable criterion for $S^{\mathrm{Betti}}(1)\not=S$.
 A simple case will be presented in Corollary 1.9 and Remark 1.10.
-We do these by studying the normal function associated with $Z$, which is defined in §$\S$§C.5, and by studying its Betti rank.
-1.2. Normal functions
+In the case of Gross–Schoen or Ceresa cycles, we will show that ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is defined over ${\mathbb{Q}}$.
+1.2. Normal functions, Betti strata and Betti rank
 Now, we turn to the geometric part of the framework and explain our results.
 Let $S$ be a quasi-projective variety.
-Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ be a VHS of weight $-1$ over $S$, and consider the intermediate Jacobian (write ${\mathscr{V}}$ for the holomorphic vector associated with ${\mathbb{V}}_{{\mathbb{Z}}}$)
+Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ be a VHS of weight $-1$ over $S$,[1][1][1]Any VHS of odd weight can be reduced to weight $-1$ by applying a suitable twist; see §$\S$§C.1.
+and consider the intermediate Jacobian (write ${\mathscr{V}}$ for the holomorphic vector associated with ${\mathbb{V}}_{{\mathbb{Z}}}$)
 $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})={\mathbb{V}}_{\mathbb{Z}}\backslash{\mathscr{V}}/{\mathscr{F}}^{0}{\mathscr{V}}{\longrightarrow}S.$
-Let $\nu\colon S{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function; then $\nu$ defines an admissible variation of mixed Hodge structures ${\mathbb{E}}_{\nu}$ on $S$ which is an extension of ${\mathbb{Z}}(0)_{S}$ by ${\mathbb{V}}_{{\mathbb{Z}}}$.
+Let $\nu\colon S{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function; then $\nu$ defines an admissible variation of mixed Hodge structures (VMHS) ${\mathbb{E}}_{\nu}$ on $S$ which is an extension of ${\mathbb{Z}}(0)_{S}$ by ${\mathbb{V}}_{{\mathbb{Z}}}$.
+To any family of homologically trivial cycles, we can associate such an admissible normal function; see §$\S$§C.5.
 The fiberwise isomorphism ${\mathbb{V}}_{{\mathbb{R}},s}{\overset{\sim}{{\longrightarrow}}}{\mathbb{V}}_{{\mathbb{C}},s}/{\mathscr{F}}^{0}_{s}={\mathscr{V}}_{s}/{\mathscr{F}}^{0}_{s}$ makes ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ into a local system of real tori
@@
 Let ${\mathscr{F}}_{\mathrm{Betti}}$ denote the induced foliation, which we call the Betti foliation.
-More precisely, for any point $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, there is a local section $\sigma\colon U{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ from a neighborhood $\Delta$ of $\pi(x)$ in $S^{\mathrm{an}}$, with $x\in\sigma(U)$, represented by a flat section of ${\mathbb{V}}_{\mathbb{R}}$.
+More precisely, for any point $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, there is a local section $\sigma\colon U{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ from a neighborhood $U$ of $\pi(x)$ in $S^{\mathrm{an}}$, with $x\in\sigma(U)$, represented by a section of the local system ${\mathbb{V}}_{\mathbb{R}}$.
 The manifolds $\sigma(U)$ gluing together to a foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$.
@@
 The Betti foliation defines a strata on $S$ as follows. For each integer $t\geq 0$, set
-(1.1) $S_{\mathscr{F}}(t):=\left\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\right\}$
+(1.2) $S^{\mathrm{Betti}}(t):=\left\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\right\}$
 where by abuse of notation $\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}}$ means the intersection with the leaves.
@@
 We then have the following Betti strata on $S$
-(1.2) $\emptyset=S_{{\mathscr{F}}}(\dim S+1)\subseteq S_{{\mathscr{F}}}(\dim S)\subseteq\cdots\subseteq S_{{\mathscr{F}}}(1)\subseteq S_{{\mathscr{F}}}(0)=S.$
+(1.3) $\emptyset=S^{\mathrm{Betti}}(\dim S+1)\subseteq S^{\mathrm{Betti}}(\dim S)\subseteq\cdots\subseteq S^{\mathrm{Betti}}(1)\subseteq S^{\mathrm{Betti}}(0)=S.$
 A main theorem of the geometric part is that the Betti strata are algebraic:
 Theorem 1.4 (Theorem 3.2).
-For each $t\geq 0$, $S_{\mathscr{F}}(t)$ is Zariski closed in $S$.
+For each $t\geq 0$, $S^{\mathrm{Betti}}(t)$ is Zariski closed in $S$.
 Remark 1.5.
-In the Betti strata, $S_{{\mathscr{F}}}(1)$ plays a particularly important role.
-For example, by definition of the Betti foliation, $S_{{\mathscr{F}}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion.
-In particular, there are at most countably many $s\in S({\mathbb{C}})$ outside $S_{{\mathscr{F}}}(1)$ such that $\nu(s)$ is torsion.
-Finally, let us point out that $S_{{\mathscr{F}}}(t)$ is closely related to the degeneracy loci defined by the first-named author.
-Indeed, when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is an abelian scheme, then $S_{{\mathscr{F}}}(t)$ is the $\nu(S)^{\mathrm{deg}}(-t+1)$ in [Gao20a, Defn.
+In the Betti strata, $S^{\mathrm{Betti}}(1)$ plays a particularly important role.
+For example, $S^{\mathrm{Betti}}(1)$ contains any complex analytic curve $C\subseteq S$ such that $\nu(C)$ is torsion because all torsion multi-sections are leaves of the Betti foliation.
+In particular, there are at most countably many $s\in S({\mathbb{C}})$ outside $S^{\mathrm{Betti}}(1)$ such that $\nu(s)$ is torsion.
+Finally, let us point out that $S^{\mathrm{Betti}}(t)$ is closely related to the degeneracy loci defined by the first-named author.
+Indeed, when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is an abelian scheme, then $S^{\mathrm{Betti}}(t)$ is the $\nu(S)^{\mathrm{deg}}(-t+1)$ in [Gao20a, Defn.
 1.6] for each $t\geq 1$.
@@
 (1.5) $r(\nu):=\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S).$
-A trivial upper bound for $r(\nu)$ is $r(\nu)\leq\min\{\dim S,\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}\}$ for any $s\in S({\mathbb{C}})$.
-One can also easily improve this trivial upper bound as will be explained below.
-The Betti rank is easily seen to be related to the Betti strata in the following way: $\dim S-r(\nu)$ is the maximum of $t\geq 0$ such that $S_{{\mathscr{F}}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$.
-Thus Theorem 1.4 implies that $r(\nu)=\dim S-\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}$.
+A trivial upper bound is $r(\nu)\leq\min\{\dim S,\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}\}$ for any $s\in S({\mathbb{C}})$.
+One can also easily improve this trivial upper bound, as explained below.
+The Betti rank is easily seen to be related to the Betti strata in the following way: $\dim S-r(\nu)$ is the maximum of $t\geq 0$ such that $S^{\mathrm{Betti}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$.
+Thus Theorem 1.4 implies that $r(\nu)=\dim S-\min\{t\geq 0:S^{\mathrm{Betti}}(t)=S\}$.
 However, this equality is, in general, not applicable to compute $r(\nu)$.
 But at this stage, we have
-(1.5) $S_{{\mathscr{F}}}(1)\not=S\Leftrightarrow r(\nu)=\dim S.$
+(1.6) $S^{\mathrm{Betti}}(1)\not=S\Longleftrightarrow r(\nu)=\dim S.$
 The second main theorem for the geometric part is the following formula for $r(\nu)$, which is often computable in practice.
-The VMHS ${\mathbb{E}}_{\nu}$ on $S$ induces a period map $\varphi=\varphi_{\nu}\colon S{\longrightarrow}\Gamma\backslash{\mathscr{D}}$, with ${\mathscr{D}}$ a mixed Mumford–Tate domain with Mumford–Tate group $\mathbf{G}$ (so ${\mathscr{D}}$ is a $\mathbf{G}({\mathbb{R}})^{+}$-orbit for $\mathbf{G}$ the generic Mumford–Tate group of the VMHS ${\mathbb{E}}_{\nu}$); we refer to §$\S$2.12.2 for the construction of $\varphi$ and ${\mathscr{D}}$.
-Denote by $V$ the unipotent radical of $\mathbf{G}$.
-It equals ${\mathbb{V}}^{\prime}_{{\mathbb{Q}},s}$ for any $s\in S({\mathbb{C}})$, where ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ is the largest sub-VHS of ${\mathbb{V}}_{{\mathbb{Z}}}$ such that $\nu$ becomes torsion under the projection ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}/{\mathbb{V}}^{\prime}_{{\mathbb{Z}}})$; see Remark 2.2.
-Now the trivial upper bound on $r(\nu)$ can be easily improved to be $r(\nu)\leq\min\{\dim\varphi(S),\frac{1}{2}\dim V\}$.
-Here is our formula to compute $r(\nu)$.
+The VMHS ${\mathbb{E}}_{\nu}$ on $S$ induces a period map $\varphi=\varphi_{\nu}\colon S{\longrightarrow}\Gamma\backslash{\mathscr{D}}$, with ${\mathscr{D}}$ a mixed Mumford–Tate domain whose Mumford–Tate group we denote by $\mathbf{G}$ (so ${\mathscr{D}}$ is a $\mathbf{G}({\mathbb{R}})^{+}$-orbit and $\mathbf{G}$ is the generic Mumford–Tate group of the VMHS ${\mathbb{E}}_{\nu}$); we refer to §$\S$2.12.2 for the construction of $\varphi$ and ${\mathscr{D}}$.
+Denote by $V$ the unipotent radical of $\mathbf{G}$. It equals ${\mathbb{V}}^{\prime}_{{\mathbb{Q}},s}$ for any $s\in S({\mathbb{C}})$, where
+[…]
+See Remark 2.2. Now the trivial upper bound on $r(\nu)$ can be easily improved to be $r(\nu)\leq\min\{\dim\varphi(S),\frac{1}{2}\dim V\}$.
+Here is our formula to compute $r(\nu)$. The trivial upper bound on $r(\nu)$ above is recovered by taking $N=\{1\}$ and $N=\mathbf{G}$.
 Theorem 1.6 (Theorem 3.1).
@@
 with $[p_{N}]$ the quotient by $N$ (see §$\S$§B.3).
-Notice that the trivial upper bound on $r(\nu)$ above is recovered by taking $N=\{1\}$ and $N=\mathbf{G}$.
-Here are two applications of Theorem 1.6, on two cases where the trivial upper bound on $r(\nu)$ is attained.
+We will give two applications of Theorem 1.6 on two cases where the trivial upper bound on $r(\nu)$ is attained: Theorem 5.1 when $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is irreducible, and Theorem 5.2 when the algebraic monodromy group of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple (or when the Mumford–Tate group of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple).
+To highlight these applications, we state Theorem 5.1 in the introduction. We also point out that Theorem 5.2 could be more applicable in geometric situations.
 Corollary 1.7 (Theorem 5.1).
@@
 Then $r(\nu)=\min\{\dim\varphi(S),\frac{1}{2}\dim_{{\mathbb{Q}}}{\mathbb{V}}_{{\mathbb{Q}},s}\}$ for one (and hence all) $s\in S({\mathbb{C}})$.
-Notice that in the situation of Corollary 1.7, we have $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
-Corollary 1.8 (Theorem 5.2).
-Assume: (i) the connected algebraic monodromy group $H$ of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple; (ii) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ has no isotrivial sub-VHS, i.e.
-locally constant VHS.
-Then $r(\nu)=\min\{\dim\varphi(S),\frac{1}{2}\dim_{{\mathbb{Q}}}V\}$.
+It is easy to see that (i) and (ii) of Corollary 1.7 imply ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}={\mathbb{V}}_{{\mathbb{Z}}}$, and hence $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
+Moreover, if $r(\nu)=\frac{1}{2}\dim_{{\mathbb{Q}}}{\mathbb{V}}_{{\mathbb{Q}},s}$, then ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ must be an abelian scheme by Griffiths’ transversality; see the comment below (2.8).
+Applying Corollary 1.7 to the Gross–Schoen and the Ceresa normal functions, we obtain the following corollary will be used to prove Theorem 1.1, Corollary 1.2 and Theorem 1.3.
+Corollary 1.8 (Corollary 5.3).
+If $g\geq 3$, then the Betti rank of the Gross-Schoen (resp. of the Ceresa) normal function over ${\mathscr{M}}_{g}$ is $33g-$ 3.
+Hain [Hai24] gave a different proof of Corollary 1.8. Our Theorem 1.6 can also be applied to any subvariety $S\subseteq{\mathscr{M}}_{g}$.
+1.2.3. Consequence on torsion loci
 An immediate corollary on the torsion locus, combining (1.6) and Remark 1.5 and Theorem 1.4, is:
 Corollary 1.9.
-Under the assumptions of either Corollary 1.7 or Corollary 1.8.
+Under the assumptions of either Corollary 1.7 or Theorem 5.2.
 Assume that $\varphi$ is generically finite and that $\dim S\leq\frac{1}{2}\dim_{{\mathbb{Q}}}V$.
 Then there exists a Zariski open dense subset $U$ of $S$ such that $\nu(s)$ is torsion for at most countably many $s\in U({\mathbb{C}})$.
-Indeed, one can take $U=S\setminus S_{{\mathscr{F}}}(1)$.
+Indeed, one can take $U=S\setminus S^{\mathrm{Betti}}(1)$.
 Remark 1.10.
-If $\varphi$ is not necessarily generically finite, then the conclusion becomes: there exists a Zariski open dense subset $U$ of $S$ such that $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is contained in at most countably many fibers of $\varphi$.
+In general, $\varphi$ is not necessarily generically finite.
+Then the conclusion becomes: there exists a Zariski open dense subset $U$ of $S$ such that $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is contained in at most countably many fibers of $\varphi$.
 Indeed, $\varphi(S)$ is an algebraic variety by [BBT23].
 Then we can conclude by applying Corollary 1.9 to $\varphi(S)$ and the induced normal function $\nu^{\prime}$ (see §$\S$3.1 for the construction of $\nu^{\prime}$).
-Applying either Corollary 1.7 or Corollary 1.8 to the Gross–Schoen and the Ceresa normal functions, we obtain the following corollary will be used to prove Theorem 1.1, Corollary 1.2 and Theorem 1.3.
-Corollary 1.11 (Corollary 5.3).
-If $g\geq 3$, then the Betti rank of the Gross-Schoen (resp. of the Ceresa) normal function is $33g-$ 3.
-Hain [Ha24] gave a different proof of Corollary 1.11. When $g=3$, he even gave an explicit description on the locus where $\dim\nu_{\mathrm{Betti},s}(T_{s}S)<3g-3$.
+In Corollary 1.9 and Remark 1.10, the condition “$\dim S\leq\frac{1}{2}\dim_{{\mathbb{Q}}}V$” can be removed if ${\mathscr{J}}({\mathbb{V}}^{\prime}_{{\mathbb{Z}}})\rightarrow S$ is not an abelian scheme, with ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ defined above Theorem 1.6.
+This follows from the comments below Corollary 1.7 (or Theorem 5.2) and is ultimately a consequence of Griffiths’ transversality.
+Kerr–Tayou [KT24] also proved Corollary 1.9 and Remark 1.10, with a different method, under the assumption of Theorem 5.2.
+1.3. Conjectures on the rationality of the Betti strata, the height inequality, and the volume identity.
+Let $S$ be a quasi-projective complex variety, $f\colon X\rightarrow S$ be a smooth family of projective varieties, and $N$ be a relatively ample line bundle on $X$.
+Let $Z$ be a family of homologically trivial $n$-cocycles in $X/S$ with $n\leq\frac{d+1}{2}$, which is homologically trivial and primitive on each geometric fiber $X_{s}$ in the sense that $c_{1}(N_{s})^{d+2-2n}\cdot Z_{s}=0$ in ${\mathrm{Ch}}^{d+2-n}(X_{s})$.
+Then we have a normal function $\nu\colon S\rightarrow{\mathscr{J}}^{n}(X/S)_{\mathrm{prim}}$ as defined by (C.5), where ${\mathscr{J}}^{n}(X/S)_{\mathrm{prim}}$ is the intermediate Jacobian associate with the following VHS:
+${\mathbb{H}}^{n}:=\ker\left(c_{1}(N)^{d+2-2n}\colon R^{2n-1}f_{*}{\mathbb{Z}}(n){\longrightarrow}R^{2d-2n+3}f_{*}{\mathbb{Z}}(d+2-n)\right).$
+We would like to propose the following conjecture on the rationality of the Betti strata.
+Conjecture 1.11.
+Let $k$ be a subfield of ${\mathbb{C}}$ over which $f$ is defined. Then $S^{\mathrm{Betti}}(t)$ is defined over $k$ for each $t\geq 0$.
+In this paper, we prove Conjecture 1.11 for $t=1$ for the Gross–Schoen cycles and Ceresa cycles for any subvariety $S$ of ${\mathscr{M}}_{g}$ by a volume identity (6.4); see Corollary 6.4.
+We expect this volume identity to be true for general family $f\colon X\rightarrow S$.
+We point out that in the particular case where $k$ is algebraically closed (for example $k=\overline{\mathbb{Q}}$), some partial results towards Conjecture 1.11 can be deduced from [KOU23], whose proof is by transcendental method.
+Now, we assume that $f$ and $N$ are defined over a number field $K$; in general, any family $X/S$ can be embedded into one defined over a number field $K$ by the Lefschetz principle.
+Assume that the Beilinson–Bloch pairing is well-defined for the cycle $Z_{s}$ for each $s\in S(\overline{\mathbb{Q}})$.
+This means that $Z_{s}$ can be extended to a cycle ${\mathscr{Z}}_{s}$ on an integral model ${\mathscr{X}}_{s}$ of $X_{s}$ over $O_{K(s)}$ such that the restrictions of ${\mathscr{Z}}_{s}$ on special fibers are homologically trivial.
+Let $g_{s}$ be the Green’s current on $X_{s}({\mathbb{C}})$ of degree $(n-1,n-1)$ so that ¯$\frac{\partial\bar{\partial}}{\pi i}g_{s}=\delta_{Z_{s}({\mathbb{C}})}$.
+Then we have an arithmetic cycle $\overline{\mathscr{Z}}_{s}=({\mathscr{Z}}_{s},g_{s})$.
+The Beilinson–Bloch height of $Z_{s}$ is defined in terms of arithmetic intersection number on ${\mathscr{X}}_{s}$:
+(1.8) $\langle{Z_{s},Z_{s}}\rangle_{\mathrm{BB}}=(-1)^{i}\widehat{c}_{1}(\overline{\mathscr{N}}_{s})^{d+1-2n}\overline{\mathscr{Z}}_{s}\cdot\overline{\mathscr{Z}}_{s},$
+where $\overline{\mathscr{N}}_{s}$ is a Hermitian line bundle on ${\mathscr{X}}_{s}$ extending $N_{s}$.
+They even conjectured that this pairing is non-negative and is zero if and only if $Z_{s}$ is torsion in the Chow group.
+See [Bei87, Blo84, GS91, GS94, Zha22] for details and various connections to the Grothendieck standard conjectures.
+Besides Gross–Schoen [GS95] for the triple product of curves, we know that the height pairing is unconditionally defined on abelian varieties by Künnemann [Kün01], and the product of a curve and a surface by the second-named author [Zha24].
+In the following, we conjecture a height inequality as in our main Theorem 1.1.
+We say that a function $\mathbf{h}\colon S(\overline{\mathbb{Q}})\rightarrow{\mathbb{R}}$ is a dominant height on an open subvariety $U$ of $S$ over $K$, if there exist a immersion $S\subseteq{\mathbb{P}}^{N}$ defined over $\overline{\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$ such that
+(1.9) $\mathbf{h}(s)\geq\epsilon h_{\mathrm{Weil}}(s)-c\qquad\text{ for all }s\in U(\overline{\mathbb{Q}}),$
+where $h_{\mathrm{Weil}}$ is the logarithmic Weil height on ${\mathbb{P}}^{N}(\overline{\mathbb{Q}})$.
+Notice that if this condition is satisfied, then the height inequality (1.9) remains true with $h_{\mathrm{Weil}}$ replaced by $h_{\overline{M}}$ for any integrable adelic line bundle $\overline{M}$ on $S$ (see §$\S$§E.6 for definition) up to changing $\epsilon>0$ and $c>0$.
+Write $S^{\mathrm{amp}}=S\setminus S^{\mathrm{Betti}}(1)$, and $S^{\mathrm{slim}}=S^{\mathrm{Betti}}(1)$ with respect to the normal function $\nu$ constructed from $Z$. We conjecture:
+Conjecture 1.12.
+The Beilinson–Bloch height function $s\mapsto\langle{Z_{s},Z_{s}}\rangle_{\mathrm{BB}}$ is a dominant height on $S^{\mathrm{amp}}$.
+Moreover, $S^{\mathrm{amp}}$ is the largest open subvariety of $S$ on which $s\mapsto\langle{Z_{s},Z_{s}}\rangle_{\mathrm{BB}}$ is a dominant height.
+The second part of Conjecture 1.12 is equivalent to: For any immersion $S\subseteq{\mathbb{P}}^{N}$ defined over $\overline{\mathbb{Q}}$, there exists a Zariski dense sequence $s_{i}\in S^{\mathrm{slim}}(\bar{K})$ such that
+$\lim_{i}h_{\mathrm{Weil}}(s_{i})=\infty,\qquad\lim_{i}\frac{\langle{Z_{s_{i}},Z_{s_{i}}}\rangle_{\mathrm{BB}}}{h_{\mathrm{Weil}}(s_{i})}=0.$
+The proof of Theorem 1.1 executed in §$\S$6.3 in fact proves the first part of Conjecture 1.12 in the case of Gross–Schoen and Ceresa cycles for any subvariety of ${\mathscr{M}}_{g}$.
+A key ingredient is the volume identity (6.4).
+The same argument can be used to show that the following conjectured volume identity (1.10) implies the full Conjecture 1.12 with the help of [YZ21, Thm.
+6.3.6].
+To check whether $S^{\mathrm{amp}}$ is non-empty, one can use (1.6) and Theorem 1.6.
+In [Blo89], Bloch constructed a biextension line bundle $L$ on $S$ defined over $K$.
+At each closed point $s\in S(\overline{\mathbb{Q}})$ where the Beilinson–Bloch height is well-defined, this line bundle has an adelic structure $\overline{L}_{s}$ over $K(s)$ whose degree gives the Beilinson–Bloch height (1.8).
+The archimedean part of this adelic structure has been studied by Hain [Hai90].
+Conjecture 1.13.
+The line bundle $L$ has a natural extension to a nef adelic line bundle $\overline{L}$ on $S$ in the sense of [YZ21], i.e.
+the image of $\overline{L}$ under the composite of (E.6) and (E.5) is $L$, such that its restriction to each $s\in S(\overline{\mathbb{Q}})$ where the Beilinson–Bloch height is well-defined is given by $\overline{L}_{s}$.
+By the arithmetic Hilbert–Samuel formula [YZ21, Thm.
+5.2.2] (which we recall as Theorem E.8.(i)), one consequence of this conjecture is the following volume identity for the generic fiber $\widetilde{L}$ of $\overline{L}$ (see (E.6) for the definition): For any archimedean place $v$ of $K$ and any subvariety $S^{\prime}$ of $S_{v,{\mathbb{C}}}$, we have
+(1.10) $\widehat{\mathrm{vol}}(\widetilde{L}|_{S^{\prime}})=\int_{S^{\prime}({\mathbb{C}})}c_{1}(\bar{L}_{v})^{\dim S}.$
+For the Gross–Schoen and Ceresa cycles for any subvariety of ${\mathscr{M}}_{g}$, we prove the existence of the extension in §$\S$6.1 and the volume identity as Theorem 6.3, but not the nefness of the adelic line bundle in question.
 1.4. Plan and Ingredients of proofs
@@
 Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers and let $Z$ be a family of homologically trivial cycles, all defined over $\overline{\mathbb{Q}}$
-The arithmetic part requires the construction of a suitable adelic line bundle $\overline{{\mathscr{L}}}$ over the base, a theory initiated by the second-named author and developed in further joint work with Yuan [YZ21], such that the height function $h_{\overline{{\mathscr{L}}}}$ is the Beilinsin–Bloch height.
+The arithmetic part requires the construction of a suitable adelic line bundle $\overline{{\mathscr{L}}}$ over the base, a theory initiated by the second-named author [Zha95] and developed in further joint work with Yuan [YZ21], such that the height function $h_{\overline{{\mathscr{L}}}}$ is the Beilinson–Bloch height.
 Then, the desired height inequalities will follow from suitable bigness properties of $\overline{{\mathscr{L}}}$.
@@
 2.5.5] the desired $\overline{{\mathscr{L}}}$ can be obtained from a suitable Deligne pairing.
-The required bigness property is the bigness of the generic fiber $\tilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$.
-We furthermore show in Proposition 6.3 that the bigness of $\tilde{{\mathscr{L}}}$ will follow from the non-vanishing of $c_{1}(\overline{{\mathscr{L}}})^{\wedge\dim S}$, with $c_{1}(\overline{{\mathscr{L}}})$ the curvature form.
+This $\overline{{\mathscr{L}}}$ has been constructed by Yuan [Yua21, §$\S$3.3.4] is slightly different steps.
+The required bigness property is the bigness of the generic fiber $\widetilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$.
+We furthermore prove the volume identity as Theorem 6.3, which implies that the bigness of $\widetilde{{\mathscr{L}}}$ will follow from the non-vanishing of $c_{1}(\overline{{\mathscr{L}}})^{\wedge\dim S}$ with $c_{1}(\overline{{\mathscr{L}}})$ the curvature form.
 The arithmetic part of this paper is confined to the Gross-Shoen and the Ceresa cycles.
@@
 The proofs of Theorem 1.4 and Theorem 1.6 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the generic rank of the Betti map for abelian schemes.
-A core of our proof is Ax–Schanuel for VMHS independently proved by Chiu [Chi21] and Gao–Klingler [GK24], which will be used multiple times.
+A core of our proof is Ax–Schanuel for VMHS independently proved by Chiu [Chi24] and Gao–Klingler [GK24], which will be used multiple times.
 Hain’s works on the Hodge-theoretic computation of the Archimedean local height pairing and the Betti form are the key ingredients to bridging the geometric and arithmetic parts.
-Indeed, Hain in [Hai90] proved that the archimedean local height pairing could be computed using the metricized biextension line bundle on $S$, and in [HR04] computed (joint with Reed), the curvature form of the metrized biextension line bundle.
+Indeed, Hain in [Hai90] proved that the archimedean local height pairing could be computed using the metrized biextension line bundle on $S$, and in [HR04] computed (joint with Reed), the curvature form of the metrized biextension line bundle.
 In our situation, we work with the height pairing of $Z_{s}$ with itself for any $s\in S(\overline{\mathbb{Q}})$, and the metrized biextension line bundle in question is the pullback of the metrized tautological line bundle under $\nu$.
@@
 Again, the bridge between the geometric part and the arithmetic part works in much larger generality, provided that the arithmetic part is solved.
+1.5. Organization of the paper
+We will work on the geometric part of our proof in §$\S$25.
+More precisely, we will define the period map associated with a normal function in §$\S$2 and explain how to see the Betti foliation in this language, prove the formula to compute the Betti rank in §$\S$3 assumption a finiteness result à la Ullmo which will be proved in §$\S$4, and then explain two applications of this formula in geometric situations (including the Gross–Schoen and Ceresa normal functions as particular cases) in §$\S$5.
+Then, in §$\S$6, we will explain the arithmetic part on how to construct the desired adelic line bundle, prove the volume identity, and prove our main theorem by combining the results of both the arithmetic and the geometric part.
+For readers’ convenience, we include several appendices, for which we do not claim originality, to make our paper as self-contained as possible.
 Acknowledgements
-We would like to thank Richard Hain for numerous inspiring discussions and for sharing his notes and Xinyi Yuan for his help during the preparation of this project.
-Part of this work was done when ZG visited SZ at Princeton University in January, February, and May 2024.
+We would like to thank Ari Shnidman for asking us the question about the Northcott property for Ceresa cycles, Richard Hain for numerous inspiring discussions and for sharing his notes, Xinyi Yuan for his help during the preparation of this project, Dick Gross for suggesting we look at $\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ instead of just ${\mathscr{M}}_{g}$, and Peter Sarnak for asking the question about the positivity of the Beilinson–Bloch heights.
+Part of this work was completed when ZG visited SZ at Princeton University in January, February, and May 2024.
 ZG would like to thank Princeton University for its hospitality.
-Finally, we would like to thank Richard Hain for mentioning our work and method at the conference “The Ceresa Cycle in Arithmetic and Geometry” at ICERM in May, 2024.
-Ziyang Gao received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n 945714).
+We also appreciate your comments on an earlier draft.
+Ziyang Gao received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement n 945714).
 Shouwu Zhang is supported by an NSF award, DMS-2101787.
@@
 Let $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function.
-The first goal of this section is to define the period map $\varphi_{\nu}$ associated with $\nu$ (when $\nu$ is clearly in context, we simply denote it by $\varphi$).
-This map fits into the diagram in the category of complex varieties
-(2.1) ${{\mathscr{D}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${\Gamma\backslash{\mathscr{D}}}$
-where
-  1. (i)
 […]
-  2. (ii)
-    $\Gamma$ is a suitable arithmetic subgroup of $\mathbf{G}({\mathbb{Q}})$.
-We emphasize that $V$ is, in general, not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$, and its geometric meaning will be given in Remark 2.2.
 […]
@@
 2.1. Universal period map to the classifying space
-Recall that $\nu$ induces a VMHS $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ on $S$ of weight $-1$ and 00 which is graded-polarized (better, admissible), fitting into the short exact sequence in the category of graded-polarized VMHS $0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{E}}_{\nu}\rightarrow{\mathbb{Z}}(0)_{S}\rightarrow 0$ with the canonical polarization on ${\mathbb{Z}}(0)_{S}$.
-Let $u_{S}\colon\tilde{S}\rightarrow S^{\mathrm{an}}$ be the universal covering map.
+Recall that $\nu$ induces a VMHS $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ on $S$ of weight $-1$ and 00 which is graded-polarized (better, admissible), fitting into the short exact sequence in the category of graded-polarized VMHS
+$0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{E}}_{\nu}\rightarrow{\mathbb{Z}}(0)_{S}\rightarrow 0$
+with the canonical polarization on ${\mathbb{Z}}(0)_{S}$.
+Let $u_{S}\colon\widetilde{S}\rightarrow S^{\mathrm{an}}$ be the universal covering map.
 Then the pullback $u_{S}^{*}{\mathbb{E}}_{\nu}$ (resp.
-$u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}$) is canonically trivialized as a local system $u_{S}^{*}{\mathbb{E}}_{\nu}\cong\tilde{S}\times E_{\nu}$ (resp.
-$u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}\cong\tilde{S}\times M_{-1,{\mathbb{Z}}}$), with $E_{\nu}=H^{0}(\tilde{S},u_{S}^{*}{\mathbb{E}}_{\nu})$ (resp.
-with $M_{-1,{\mathbb{Z}}}=H^{0}(\tilde{S},u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}})$).
-Then for each $s\in S({\mathbb{C}})$, the fiber ${\mathbb{E}}_{\nu,s}$ can be canonically identified with $E_{\nu}$ and ${\mathbb{V}}_{{\mathbb{Z}},s}$ can be canonically identified with $M_{-1,{\mathbb{Z}}}$.
-The pullback under $u_{S}$ of the short exact sequence of VMHS defined by $\nu$ becomes split after $\otimes{\mathbb{Q}}$ since $\tilde{S}$ is simply connected.
-Each $\tilde{s}\in\tilde{S}$ gives rise to a polarized mixed Hodge structure $(E_{\nu,{\mathbb{Q}}},(W_{\bullet})_{\tilde{s}},({\mathscr{F}}_{{\mathbb{E}}}^{\bullet})_{\tilde{s}})$ on $E_{\nu,{\mathbb{Q}}}$, And this induces the universal period map
-$\tilde{\varphi}=\tilde{\varphi}_{\nu}\colon\tilde{S}\longrightarrow{\mathscr{M}}$
+$u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}$) is canonically trivialized as a local system $u_{S}^{*}{\mathbb{E}}_{\nu}\simeq\widetilde{S}\times E_{\nu}$ (resp.
+$u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}\simeq\widetilde{S}\times V_{{\mathbb{Z}}}$), with $E_{\nu}=H^{0}(\widetilde{S},u_{S}^{*}{\mathbb{E}}_{\nu})$ (resp.
+with $V_{{\mathbb{Z}}}=H^{0}(\widetilde{S},u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}})$).
+Then for each $s\in S({\mathbb{C}})$, the fiber ${\mathbb{E}}_{\nu,s}$ can be canonically identified with $E_{\nu}$ and ${\mathbb{V}}_{{\mathbb{Z}},s}$ can be canonically identified with $V_{{\mathbb{Z}}}$.
+The pullback under $u_{S}$ of the short exact sequence of VMHS defined by $\nu$ becomes split after $\otimes{\mathbb{Q}}$ since $\widetilde{S}$ is simply connected.
+Each $\widetilde{s}\in\widetilde{S}$ gives rise to a polarized mixed Hodge structure $(E_{\nu,{\mathbb{Q}}},(W_{\bullet})_{\widetilde{s}},({\mathscr{F}}_{{\mathbb{E}}}^{\bullet})_{\widetilde{s}})$ on $E_{\nu,{\mathbb{Q}}}$, And this induces the universal period map.
+$\widetilde{\varphi}=\widetilde{\varphi}_{\nu}\colon\widetilde{S}\longrightarrow{\mathscr{M}}$
 […]
 […]
-(2.2) ${{\mathscr{M}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${\tilde{S}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varphi}}$$\scriptstyle{\tilde{\varphi}_{0}}$${{\mathscr{M}}_{0}.}$
+(2.1) ${{\mathscr{M}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${\widetilde{S}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{\varphi}}$$\scriptstyle{\widetilde{\varphi}_{0}}$${{\mathscr{M}}_{0}.}$
+In practice, we need to refine this period map $\widetilde{\varphi}$ and replace the classifying space ${\mathscr{M}}$ by the smallest Mumford–Tate domain ${\mathscr{D}}$ which contains $\widetilde{\varphi}(\widetilde{S})$.
+This construction is done in the next subsection.
 2.2. Construction of the Mumford–Tate domain
-In practice, we need to refine this period map and replace the classifying space ${\mathscr{M}}$ by the smallest Mumford–Tate domain ${\mathscr{D}}$, which contains $\tilde{\varphi}(\tilde{S})$.
-It is constructed as follows.
 […]
-Fix $\tilde{s}\in\tilde{S}$. Define
-(2.3) ${\mathscr{D}}:=\mathbf{G}({\mathbb{R}})^{+}\cdot\tilde{\varphi}(\tilde{s})\subseteq{\mathscr{M}}.$
+Fix $\widetilde{s}\in\widetilde{S}$. Define
+(2.2) ${\mathscr{D}}:=\mathbf{G}({\mathbb{R}})^{+}\cdot\widetilde{\varphi}(\widetilde{s})\subseteq{\mathscr{M}}.$
 […]
-Finally, since $S$ is a quasi-projective variety, there exists an arithmetic subgroup $\Gamma$ of $\mathbf{G}({\mathbb{Q}})$ such that $\tilde{\varphi}$ descends to a morphism $\varphi\colon S^{\mathrm{an}}\rightarrow\Gamma\backslash{\mathscr{D}}$ fitting into the diagram (2.1).
+Finally, since $S$ is a quasi-projective variety, there exists an arithmetic subgroup $\Gamma$ of $\mathbf{G}({\mathbb{Q}})$ such that $\widetilde{\varphi}$ descends to a morphism $\varphi\colon S^{\mathrm{an}}\rightarrow\Gamma\backslash{\mathscr{D}}$.
 2.3. Setup for o-minimality
@@
 2.4. Relating the Betti foliation and the period map
-Recall the projection to the pure part $p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}\subseteq{\mathscr{M}}_{0}$ from (2.2) and the semi-algebraic structure ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ as in Proposition B.3.(i).
-The diagram (2.1) can be complete into:
+Denote by ${\mathscr{D}}_{0}:=p({\mathscr{D}})\subseteq{\mathscr{M}}_{0}$ for the projection to the pure part $p$ from (2.1), and by abuse of notation $p=p|_{{\mathscr{D}}}\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}$.
+Recalling the semi-algebraic structure ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ as in Proposition B.3.(i), the period maps $\varphi$ and $\widetilde{\varphi}$ can be complete into:
 […]
 and $\varphi_{0}=p\circ\varphi$ is the period map for the VHS ${\mathbb{V}}_{{\mathbb{Z}}}\rightarrow S$.
-For each $\tilde{s}\in\tilde{S}$, write $\tilde{s}_{V}$ for the image $\tilde{\varphi}(\tilde{s})$ under the natural projection ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}\rightarrow V({\mathbb{R}})$.
-This gives a map $\tilde{\varphi}_{V}\colon\tilde{S}\rightarrow V({\mathbb{R}})$ sending $\tilde{s}\mapsto\tilde{s}_{V}$.
+For each $\widetilde{s}\in\widetilde{S}$, write $\widetilde{s}_{V}$ for the image $\widetilde{\varphi}(\widetilde{s})$ under the natural projection ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}\rightarrow V({\mathbb{R}})$.
+This gives a map $\widetilde{\varphi}_{V}\colon\widetilde{S}\rightarrow V({\mathbb{R}})$ sending $\widetilde{s}\mapsto\widetilde{s}_{V}$.
 Notice that
-(2.5) $\tilde{\varphi}=(\tilde{\varphi}_{V},\tilde{\varphi}_{0})\colon\tilde{S}\longrightarrow V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}.$
-Recall $S_{{\mathscr{F}}}(t)=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ and define $S_{\mathrm{rk}}(t)=\{s\in S({\mathbb{C}}):\dim\nu_{\mathrm{Betti},s}(T_{s}S)\leq t\}$ for $\nu_{\mathrm{Betti},s}\colon T_{s}S\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ defined in (1.3).
+(2.4) $\widetilde{\varphi}=(\widetilde{\varphi}_{V},\widetilde{\varphi}_{0})\colon\widetilde{S}\longrightarrow V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}.$
+Recall $S^{\mathrm{Betti}}(t)=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ and define
+$S_{\mathrm{rk}}(t)=\{s\in S({\mathbb{C}}):\dim\nu_{\mathrm{Betti},s}(T_{s}S)\leq t\}$
+for $\nu_{\mathrm{Betti},s}\colon T_{s}S\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ defined in (1.4).
 Lemma 2.1.
 For each $t\geq 0$, we have
-$u_{S}^{-1}\left(S_{{\mathscr{F}}}(t)\right)$ =$=\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}\tilde{\varphi}^{-1}(\{\tilde{s}_{V}\}\times{\mathscr{D}}_{0})\geq t\}$
-(2.6) $=\bigcup\nolimits_{r\geq 0}\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}\tilde{\varphi}^{-1}(\tilde{\varphi}(\tilde{s}))=r,~{}\{\tilde{s}_{V}\}\times\tilde{C}\subseteq\tilde{\varphi}(\tilde{S})\text{ for }$
+$u_{S}^{-1}\left(S^{\mathrm{Betti}}(t)\right)$ =$=\{\widetilde{s}\in\widetilde{S}:\dim_{\widetilde{s}}\widetilde{\varphi}^{-1}(\{\widetilde{s}_{V}\}\times{\mathscr{D}}_{0})\geq t\}$
+(2.5) $=\bigcup\nolimits_{r\geq 0}\{\widetilde{s}\in\widetilde{S}:\dim_{\widetilde{s}}\widetilde{\varphi}^{-1}(\widetilde{\varphi}(\widetilde{s}))=r,~{}\{\widetilde{s}_{V}\}\times\widetilde{C}\subseteq\widetilde{\varphi}(\widetilde{S})\text{ for }$
 […]
 and
-(2.7) $u_{S}^{-1}\left(S_{\mathrm{rk}}(t)\right)=\{\tilde{s}\in\tilde{S}:\mathrm{rank}(\mathrm{d}\tilde{\varphi}_{V})_{\tilde{s}}\leq t\}.$
-Notice that in the union in (2.1), the second condition $\{\tilde{s}_{V}\}\times\tilde{C}\subseteq\tilde{\varphi}(\tilde{S})$ always holds true for $r>t$.
+(2.6) $u_{S}^{-1}\left(S_{\mathrm{rk}}(t)\right)=\{\widetilde{s}\in\widetilde{S}:\mathrm{rank}(\mathrm{d}\widetilde{\varphi}_{V})_{\widetilde{s}}\leq t\}.$
+Notice that in the union in (2.1), the second condition $\{\widetilde{s}_{V}\}\times\widetilde{C}\subseteq\widetilde{\varphi}(\widetilde{S})$ always holds for $r>t$.
 Proof.
@@
 Each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lies in ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ for $s=\pi(x)$, which is canonically isomorphic to $\mathrm{Ext}_{\mathrm{MHS}}({\mathbb{Z}}(0),{\mathbb{V}}_{{\mathbb{Z}},s})$ by Carlson [Car85].
-Hence each $x$ gives rise to a ${\mathbb{Z}}$-mixed Hodge structure of weight $-1$ and 00, and for the universal covering map $u_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ we obtain a period map $\tilde{\varphi}_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}\rightarrow{\mathscr{M}}$.
-The map $\nu\circ u_{S}\colon\tilde{S}\rightarrow S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lifts to $\tilde{\nu}\colon\tilde{S}\rightarrow\tilde{{\mathscr{J}}}$, and $\tilde{\varphi}=\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu}$.
+Hence each $x$ gives rise to a ${\mathbb{Z}}$-mixed Hodge structure of weight $-1$ and 00, and for the universal covering map $u_{{\mathscr{J}}}\colon\widetilde{{\mathscr{J}}}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ we obtain a period map $\widetilde{\varphi}_{{\mathscr{J}}}\colon\widetilde{{\mathscr{J}}}\rightarrow{\mathscr{M}}$.
+The map $\nu\circ u_{S}\colon\widetilde{S}\rightarrow S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lifts to $\widetilde{\nu}\colon\widetilde{S}\rightarrow\widetilde{{\mathscr{J}}}$, and $\widetilde{\varphi}=\widetilde{\varphi}_{{\mathscr{J}}}\circ\widetilde{\nu}$.
 […]
-The zero section of $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ gives rise to a Levi decomposition of $\mathbf{G}^{{\mathscr{M}}}$ and hence an identification ${\mathscr{M}}=M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}$ as in Proposition B.3.(i) applied to ${\mathscr{M}}$, and $\tilde{\varphi}_{{\mathscr{J}}}$ becomes
-(2.8) $\tilde{\varphi}_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}=M_{-1}({\mathbb{R}})\times\tilde{S}\xrightarrow{(1,\tilde{\varphi}_{0})}M_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}.$
-By the last paragraph, the leaves of ${\mathscr{F}}_{\mathrm{Betti}}$ are precisely $u_{{\mathscr{J}}}\left(\tilde{\varphi}_{{\mathscr{J}}}^{-1}(\{a\}\times{\mathscr{M}}_{0})\right)$ for all $a\in M_{-1}({\mathbb{R}})$.
-Thus $u_{{\mathscr{J}}}^{-1}\left(\nu(S_{{\mathscr{F}}}(t))\right)=\{\tilde{s}^{\prime}\in\tilde{\nu}(\tilde{S}):\dim_{\tilde{s}^{\prime}}\tilde{\varphi}_{{\mathscr{J}}}^{-1}(\{\tilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}$, where $\tilde{s}^{\prime}_{V}$ is the image of $\tilde{s}\in\tilde{{\mathscr{J}}}\xrightarrow{\tilde{\varphi}_{{\mathscr{J}}}}M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}\rightarrow M_{-1}({\mathbb{R}})$ with the last map being the natural projection.
-Applying $\tilde{\nu}^{-1}$ (whose fibers are of dimension 00 because $\nu$ is injective) to the set above and noticing that $u_{{\mathscr{J}}}\circ\tilde{\nu}=\nu\circ u_{S}$, we have
-$u_{S}^{-1}\left(S_{{\mathscr{F}}}(t)\right)=\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}(\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu})^{-1}(\{\tilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}.$
-Hence the first equality in (2.1) holds true because $\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu}=\tilde{\varphi}$ and by Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$.
+The zero section of $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ gives rise to a Levi decomposition of $\mathbf{G}^{{\mathscr{M}}}$ and hence an identification ${\mathscr{M}}=V({\mathbb{R}})\times{\mathscr{M}}_{0}$ as in Proposition B.3.(i) applied to ${\mathscr{M}}$, and $\widetilde{\varphi}_{{\mathscr{J}}}$ becomes
+(2.7) $\widetilde{\varphi}_{{\mathscr{J}}}\colon\widetilde{{\mathscr{J}}}=V({\mathbb{R}})\times\widetilde{S}\xrightarrow{(1,\widetilde{\varphi}_{0})}V({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq V({\mathbb{R}})\times{\mathscr{M}}_{0}.$
+By the last paragraph, the leaves of ${\mathscr{F}}_{\mathrm{Betti}}$ are precisely $u_{{\mathscr{J}}}\left(\widetilde{\varphi}_{{\mathscr{J}}}^{-1}(\{a\}\times{\mathscr{M}}_{0})\right)$ for all $a\in V({\mathbb{R}})$.
+Thus $u_{{\mathscr{J}}}^{-1}\left(\nu(S^{\mathrm{Betti}}(t))\right)=\{\widetilde{s}^{\prime}\in\widetilde{\nu}(\widetilde{S}):\dim_{\widetilde{s}^{\prime}}\widetilde{\varphi}_{{\mathscr{J}}}^{-1}(\{\widetilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}$, where $\widetilde{s}^{\prime}_{V}$ is the image of $\widetilde{s}\in\widetilde{{\mathscr{J}}}\xrightarrow{\widetilde{\varphi}_{{\mathscr{J}}}}V({\mathbb{R}})\times{\mathscr{M}}_{0}\rightarrow V({\mathbb{R}})$ with the last map being the natural projection.
+Applying $\widetilde{\nu}^{-1}$ (whose fibers are of dimension 00 because $\nu$ is injective) to the set above and noticing that $u_{{\mathscr{J}}}\circ\widetilde{\nu}=\nu\circ u_{S}$, we have
+$u_{S}^{-1}\left(S^{\mathrm{Betti}}(t)\right)=\{\widetilde{s}\in\widetilde{S}:\dim_{\widetilde{s}}(\widetilde{\varphi}_{{\mathscr{J}}}\circ\widetilde{\nu})^{-1}(\{\widetilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}.$
+Hence the first equality in (2.1) holds because $\widetilde{\varphi}_{{\mathscr{J}}}\circ\widetilde{\nu}=\widetilde{\varphi}$ and by Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$.
 Similarly, we get (2.6).
@@
 Remark 2.2.
-The proof of Lemma 2.1 (with Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$ and the identification $\tilde{{\mathscr{J}}}=M_{-1}({\mathbb{R}})\times\tilde{S}$) also yields the following assertion: ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})\subseteq{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is the intermediate Jacobian of a sub-VHS translated by a torsion multisection; it contains $\nu(S)$ and is the minimal one containing $\nu(S)$ with respect to inclusion.
+The proof of Lemma 2.1 (with Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$ and the identification $\widetilde{{\mathscr{J}}}=V({\mathbb{R}})\times\widetilde{S}$) also yields the following assertion: ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})\subseteq{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is the intermediate Jacobian of a sub-VHS translated by a torsion multisection; it contains $\nu(S)$ and is the minimal one containing $\nu(S)$ with respect to inclusion.
 The relative dimension $\dim{\mathscr{J}}_{\nu}-\dim S$ equals $\frac{1}{2}\dim V$.
@@
 Moreover for the natural projection $\varphi_{{\mathscr{J}}_{\nu}}\colon{\mathscr{J}}_{\nu}\rightarrow\Gamma\backslash{\mathscr{D}}$, we have $\varphi(S)=\varphi_{{\mathscr{J}}_{\nu}}(\nu(S))$.
-Since $\tilde{\varphi}_{V}$ factors through $\tilde{\varphi}$ and has target $V({\mathbb{R}})$, (2.7) immediately yields the following trivial upper bound
+Since $\widetilde{\varphi}_{V}$ factors through $\widetilde{\varphi}$ and has target $V({\mathbb{R}})$, (2.6) immediately yields the following trivial upper bound
 (2.8) $\nu_{\mathrm{Betti},s}(T_{s}S)\leq\min\left\{\dim\varphi(S),\frac{1}{2}\dim V\right\}\quad\text{ for all }s\in S({\mathbb{C}}).$
-3. The Betti rank and Zariski closedness of the rank-strata
+We can say more.
+By Griffiths’ transversality, we must have $\nu_{\mathrm{Betti},s}(T_{s}S)<\frac{1}{2}\dim V$ if ${\mathscr{J}}({\mathbb{V}}^{\prime}_{{\mathbb{Z}}})\rightarrow S$ is not an abelian scheme (or equivalently, the Hodge type of ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}\rightarrow S$ is not $(-1,0)+(0,-1)$); see above (D.3).
+3. The Betti rank and Zariski closedness of the Betti strata
 The goal of this section is to prove one of the main theorems of this paper.
@@
 […]
+$\nu_{\mathrm{Betti},s}\colon T_{s}S\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$
+at each $s\in S({\mathbb{C}})$, and the Betti rank (1.5)
 $r(\nu)=\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S).$
@@
 Theorem 3.1.
-The Betti rank is given by the following formula:
+The following formula gives the Betti rank:
 (3.1) $r(\nu)=\min_{N}\left\{\dim\varphi_{/N}(S)+\frac{1}{2}\dim_{{\mathbb{Q}}}(V\cap N)\right\},$
@@
 Theorem 3.2.
-For each $t\geq 0$, the set $S_{{\mathscr{F}}}(t):=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ is Zariski closed in $S$.
-In particular, $r(\nu)=\dim S-\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}$.
+For each $t\geq 0$, the set $S^{\mathrm{Betti}}(t):=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ is Zariski closed in $S$.
+In particular, $r(\nu)=\dim S-\min\{t\geq 0:S^{\mathrm{Betti}}(t)=S\}$.
 The proofs of Theorem 3.1 and Theorem 3.2 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the case when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is polarizable.
-A key ingredient for our proof is the mixed Ax-Schanuel theorem, which is used multiple times.
+A key ingredient for our proof is the mixed Ax–Schanuel theorem, which is used multiple times.
 In [Gao20a], the version for universal abelian varieties [Gao20b, Thm.
 1.1] was used.
-In the current paper, we need the version for admissible VMHS independently proved by Chiu [Chi21] and Gao–Klingler [GK24].
-We also invoke [BBT23] on the algebraicity of $\varphi(S)$, whose proof builds up on o-minimal GAGA, for two reasons.
-First, it eases the notation for the proof.
-Second and more importantly, [BBT23] is necessary to prove the full version of Theorem 1.6.
-Still, we point out that [BBT23] is not needed in the case which we care the most about in this paper, i.e.
-$r(\nu)=\dim S$ if and only if the RHS of (1.6) equals $\dim S$ because we can easily reduce the proof to the case where every fiber of $\varphi$ has dimension 00.
+In the current paper, we need the version for admissible VMHS independently proved by Chiu [Chi24] and Gao–Klingler [GK24].
+We also invoke [BBT23] on the algebraicity of $\varphi(S)$, whose proof builds up on o-minimal GAGA, eases the notation for the proof.
 Another crucial input for the proof is the o-minimal structure associated with the period map [BBKT24], which we will recall in §$\S$2.3. This allows us to apply (o-minimal) definable Chow.
@@
 By (the proof of) Lemma 2.1, more precisely the second equality of (2.1), we have
-(3.3) $S_{{\mathscr{F}}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime}_{{\mathscr{F}}}(t-r)\right).$
+(3.3) $S^{\mathrm{Betti}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime}_{{\mathscr{F}}}(t-r)\right).$
 This equality allows us to replace $S$ by $\varphi(S)$ to study the Betti rank and the Betti strata.
@@
   1. (i)
-    Let $\tilde{Y}\subseteq{\mathscr{D}}$ be a complex analytic irreducible subset.
-The weakly special closure of $\tilde{Y}$, denoted by $\tilde{Y}^{\mathrm{ws}}$, is the smallest weak Mumford–Tate domain in ${\mathscr{D}}$ which contains $\tilde{Y}$.
+    Let $\widetilde{Y}\subseteq{\mathscr{D}}$ be a complex analytic irreducible subset.
+The weakly special closure of $\widetilde{Y}$, denoted by $\widetilde{Y}^{\mathrm{ws}}$, is the smallest weak Mumford–Tate domain in ${\mathscr{D}}$ which contains $\widetilde{Y}$.
   2. (ii)
@@
 The weakly special closure of $Y$, denoted by $Y^{\mathrm{ws}}$, is $u(\widetilde{Y}^{\mathrm{ws}})$ for one (hence any) complex analytic irreducible component $\widetilde{Y}$ of $u^{-1}(Y)$.
-The following Ax–Schanuel theorem for VMHS was independently proved by Chiu [Chi21] and Gao–Klingler [GK24]. We refer to Definition B.2 for the algebraic structure on ${\mathscr{D}}$.
+The following Ax–Schanuel theorem for VMHS was independently proved by Chiu [Chi24] and Gao–Klingler [GK24]. We refer to Definition B.2 for the algebraic structure on ${\mathscr{D}}$.
 Theorem 3.5 (weak Ax–Schanuel for VMHS).
-Let $\tilde{Z}\subseteq u^{-1}(S)$ be a complex analytic irreducible subset. Then
-(3.4) $\dim\tilde{Z}^{\mathrm{Zar}}+\dim u(\tilde{Z})^{\mathrm{Zar}}\geq\dim\tilde{Z}^{\mathrm{ws}}+\dim\tilde{Z},$
-where $\tilde{Z}^{\mathrm{ws}}$ is the smallest weak Mumford–Tate domain which contains $\tilde{Z}$.
+Let $\widetilde{Z}\subseteq u^{-1}(S)$ be a complex analytic irreducible subset. Then
+(3.4) $\dim\widetilde{Z}^{\mathrm{Zar}}+\dim u(\widetilde{Z})^{\mathrm{Zar}}\geq\dim\widetilde{Z}^{\mathrm{ws}}+\dim\widetilde{Z},$
+where $\widetilde{Z}^{\mathrm{ws}}$ is the smallest weak Mumford–Tate domain which contains $\widetilde{Z}$.
 Proof.
-Let $Y:=u(\tilde{Z})^{\mathrm{Zar}}$.
-Let $\mathscr{Z}:=\{(z,y)\in\tilde{Z}\times Y({\mathbb{C}}):u(z)=y\}$, then $\mathscr{Z}$ is a complex analytic irreducible subset of ${\mathscr{D}}\times_{\Gamma\backslash{\mathscr{D}}}Y^{\prime}$.
-The Zariski closure of $\mathscr{Z}$ in ${\mathscr{D}}\times Y$ is contained in $\tilde{Z}^{\mathrm{Zar}}\times Y$, and $\dim\mathscr{Z}=\dim\tilde{Z}$.
+Let $Y:=u(\widetilde{Z})^{\mathrm{Zar}}$.
+Let $\mathscr{Z}:=\{(z,y)\in\widetilde{Z}\times Y({\mathbb{C}}):u(z)=y\}$, then $\mathscr{Z}$ is a complex analytic irreducible subset of ${\mathscr{D}}\times_{\Gamma\backslash{\mathscr{D}}}Y^{\prime}$.
+The Zariski closure of $\mathscr{Z}$ in ${\mathscr{D}}\times Y$ is contained in $\widetilde{Z}^{\mathrm{Zar}}\times Y$, and $\dim\mathscr{Z}=\dim\widetilde{Z}$.
 Then (3.4) is a direct consequence of the mixed Ax–Schanuel theorem [GK24, Thm.
@@
 Definition 3.6.
-An irreducible algebraic subvariety $Y$ of $S$ is called weakly optimal if the following holds true: $Y\subsetneq Y^{\prime}\subseteq S\Rightarrow\delta_{\mathrm{ws}}(Y)<\delta_{\mathrm{ws}}(Y^{\prime})$, for any $Y^{\prime}\subseteq S$ irreducible.
+An irreducible algebraic subvariety $Y$ of $S$ is called weakly optimal if the following holds: $Y\subsetneq Y^{\prime}\subseteq S\Rightarrow\delta_{\mathrm{ws}}(Y)<\delta_{\mathrm{ws}}(Y^{\prime})$, for any $Y^{\prime}\subseteq S$ irreducible.
 3.3. Applications of Ax–Schanuel
@@
 Proposition 3.7.
-For each $t\geq 0$, $S_{{\mathscr{F}}}(t)$ is contained in the union of weakly optimal subvarieties $Y\subseteq S$ satisfying
+For each $t\geq 0$, $S^{\mathrm{Betti}}(t)$ is contained in the union of weakly optimal subvarieties $Y\subseteq S$ satisfying
 (3.6) $\dim Y\geq\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})+t.$
@@
   1. (i)
-    $S_{{\mathscr{F}}}(t)$ is covered by the union of irreducible subvarieties $Y\subseteq S$ satisfying (3.6) (without requiring $Y$ to be weakly optimal);
+    $S^{\mathrm{Betti}}(t)$ is covered by the union of irreducible subvarieties $Y\subseteq S$ satisfying (3.6) (without requiring $Y$ to be weakly optimal);
   2. (ii)
     If $Y\subseteq S$ is an irreducible subvariety satisfying (3.6) and is maximal for this property with respect to inclusions, then $Y$ is weakly optimal.
-Let us prove (i). By Lemma 2.1, more precisely the first equality of (2.1), $S_{{\mathscr{F}}}(t)$ is covered by irreducible subvarieties $Y\subseteq S$ such that
-$Y:=\overline{u(\{a\}\times\tilde{C})}^{\mathrm{Zar}},$
-for some complex analytic irreducible $\tilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\tilde{C}=t$ and some $a\in V({\mathbb{R}})$.
-Apply mixed Ax–Schanuel in this context (Theorem 3.5 to $\{a\}\times\tilde{C}$). Then we get
-$\dim\overline{\{a\}\times\tilde{C}}^{\mathrm{Zar}}+\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}+t.$
-By Lemma B.4, $\overline{\{a\}\times\tilde{C}}^{\mathrm{Zar}}=\{a\}\times\overline{\tilde{C}}^{\mathrm{Zar}}$. Hence
-$\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\overline{\tilde{C}}^{\mathrm{Zar}}+t\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\tilde{C}^{\mathrm{ws}}+t.$
-The last inequality holds because $\overline{\tilde{C}}^{\mathrm{Zar}}\subseteq\tilde{C}^{\mathrm{ws}}$. So
-$\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\tilde{C}^{\mathrm{ws}}+t.$
-Now, to prove (i), it suffices to prove $\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$ and $\dim\tilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$.
-Let us prove $u((\{a\}\times\tilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$; the upshot is $\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$.
-By definition of $Y$, we have $u(\{a\}\times\tilde{C})\subseteq Y$.
-Hence $u((\{a\}\times\tilde{C})^{\mathrm{ws}})\subseteq Y^{\mathrm{ws}}$.
-On the other hand, $u((\{a\}\times\tilde{C})^{\mathrm{ws}})$ is closed algebraic by Proposition 3.3, so $Y\subseteq u((\{a\}\times\tilde{C})^{\mathrm{ws}})$.
-So $Y^{\mathrm{ws}}\subseteq u((\{a\}\times\tilde{C})^{\mathrm{ws}})$.
-Now we have established $u((\{a\}\times\tilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$.
-Similarly, we have $\dim\tilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$. Hence, we are done for (i).
+Let us prove (i). By Lemma 2.1, more precisely the first equality of (2.1), $S^{\mathrm{Betti}}(t)$ is covered by irreducible subvarieties $Y\subseteq S$ such that
+$Y:=\overline{u(\{a\}\times\widetilde{C})}^{\mathrm{Zar}},$
+for some complex analytic irreducible $\widetilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\widetilde{C}=t$ and some $a\in V({\mathbb{R}})$.
+Apply mixed Ax–Schanuel in this context (Theorem 3.5 to $\{a\}\times\widetilde{C}$). Then we get
+$\dim\overline{\{a\}\times\widetilde{C}}^{\mathrm{Zar}}+\dim Y\geq\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}+t.$
+By Lemma B.4, $\overline{\{a\}\times\widetilde{C}}^{\mathrm{Zar}}=\{a\}\times\overline{\widetilde{C}}^{\mathrm{Zar}}$. Hence
+$\dim Y\geq\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}-\dim\overline{\widetilde{C}}^{\mathrm{Zar}}+t\geq\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}-\dim\widetilde{C}^{\mathrm{ws}}+t.$
+The last inequality holds because $\overline{\widetilde{C}}^{\mathrm{Zar}}\subseteq\widetilde{C}^{\mathrm{ws}}$. So
+$\dim Y\geq\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}-\dim\widetilde{C}^{\mathrm{ws}}+t.$
+Now, to prove (i), it suffices to prove $\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$ and $\dim\widetilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$.
+Let us prove $u((\{a\}\times\widetilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$; the upshot is $\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$.
+By definition of $Y$, we have $u(\{a\}\times\widetilde{C})\subseteq Y$.
+Hence $u((\{a\}\times\widetilde{C})^{\mathrm{ws}})\subseteq Y^{\mathrm{ws}}$.
+On the other hand, $u((\{a\}\times\widetilde{C})^{\mathrm{ws}})$ is closed algebraic by Proposition 3.3, so $Y\subseteq u((\{a\}\times\widetilde{C})^{\mathrm{ws}})$.
+So $Y^{\mathrm{ws}}\subseteq u((\{a\}\times\widetilde{C})^{\mathrm{ws}})$.
+Now we have established $u((\{a\}\times\widetilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$.
+Similarly, we have $\dim\widetilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$. Hence, we are done for (i).
 For (ii), let $Y\subseteq Y^{\prime}\subseteq X$. Assume $\delta_{\mathrm{ws}}(Y)\geq\delta_{\mathrm{ws}}(Y^{\prime})$, i.e.
@@
 Proposition 3.8.
-There exist finitely many pairs $({\mathscr{D}}^{\prime}_{1},N_{1}),\ldots,({\mathscr{D}}^{\prime}_{k},N_{k})$, with each ${\mathscr{D}}^{\prime}_{j}$ a Mumford–Tate domain contained in ${\mathscr{D}}$ and $N_{j}$ a normal subgroup of $\mathrm{MT}({\mathscr{D}}^{\prime}_{j})$, such that the following holds true.
+There exist finitely many pairs $({\mathscr{D}}^{\prime}_{1},N_{1}),\ldots,({\mathscr{D}}^{\prime}_{k},N_{k})$, with each ${\mathscr{D}}^{\prime}_{j}$ a Mumford–Tate domain contained in ${\mathscr{D}}$ and $N_{j}$ a normal subgroup of $\mathrm{MT}({\mathscr{D}}^{\prime}_{j})$, such that the following holds.
 For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}$ is the image of an $N_{j}({\mathbb{R}})^{+}$-orbit contained in ${\mathscr{D}}^{\prime}_{j}$ under $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$ for some $j\in\{1,\ldots,k\}$.
-The same statement (with $Y^{\mathrm{ws}}$ replaced by $\varphi(Y)^{\mathrm{ws}}$) still holds true without replacing $S$ by $\varphi(S)$.
+The same statement (with $Y^{\mathrm{ws}}$ replaced by $\varphi(Y)^{\mathrm{ws}}$) still holds without replacing $S$ by $\varphi(S)$.
 Denote by $\Gamma_{j}=\Gamma\cap\mathrm{MT}({\mathscr{D}}^{\prime}_{j})({\mathbb{Q}})$ and $\Gamma_{j,/N_{j}}=\Gamma_{j}/(\Gamma_{j}\cap N_{j}({\mathbb{Q}}))$.
@@
 For each $t\geq 0$, we have
-$S_{{\mathscr{F}}}(t)\subseteq\bigcup\nolimits_{j=1}^{k}E_{j}(t).$
+$S^{\mathrm{Betti}}(t)\subseteq\bigcup\nolimits_{j=1}^{k}E_{j}(t).$
 Proof.
 Let $t\geq 0$.
-By Proposition 3.7, $S_{{\mathscr{F}}}(t)$ is covered by weakly optimal $Y\subseteq S$ such that $\dim Y\geq\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})+t$.
+By Proposition 3.7, $S^{\mathrm{Betti}}(t)$ is covered by weakly optimal $Y\subseteq S$ such that $\dim Y\geq\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})+t$.
 Then by Proposition 3.8, $Y^{\mathrm{ws}}$ is a fiber $[p_{N_{j}}]$ for some $j\in\{1,\ldots,k\}$, and hence $\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})=\frac{1}{2}\dim(V\cap N_{j})$.
@@
 Let us prove $\leq$”.
-By (2.7) we have $r(\nu)=\max_{\tilde{s}\in\tilde{S}}(\mathrm{d}\tilde{\varphi}_{V})_{\tilde{s}}$.
+By (2.6) we have $r(\nu)=\max_{\widetilde{s}\in\widetilde{S}}(\mathrm{d}\widetilde{\varphi}_{V})_{\widetilde{s}}$.
 Hence $r(\nu)\leq\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)$ for any normal subgroup $N\lhd\mathbf{G}$.
@@
 Let $t=\dim S-r(\nu)$.
-Then $S_{{\mathscr{F}}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$.
+Then $S^{\mathrm{Betti}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$.
 By (3.3) and Proposition 3.9 (which should be applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
-$S_{{\mathscr{F}}}(t)\subseteq S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
+$S^{\mathrm{Betti}}(t)\subseteq S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
 Each $S_{\geq r}$ is Zariski closed in $S$, and each $E_{j}(t-r)$ is Zariski closed in $S^{\prime}=\varphi(S)$.
@@
 Taking the Zariski closure of both sides, we then have $S$ equal to a member on the right-hand side.
-If $S=S_{\geq t}$, then $\geq$” holds true already for $N=\{1\}$.
+If $S=S_{\geq t}$, then $\geq$” holds already for $N=\{1\}$.
 Assume $S=S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right)$ for some $10\leq r\leq t-$ 1 and some $j$.
@@
 Lemma 3.10.
-For each $t\geq 0$ and each $j\in\{1,\ldots,k\}$, we have $E_{j}(t)\subseteq S_{{\mathscr{F}}}(t)$.
+For each $t\geq 0$ and each $j\in\{1,\ldots,k\}$, we have $E_{j}(t)\subseteq S^{\mathrm{Betti}}(t)$.
 Proof.
@@
 Notice that this $v\in V({\mathbb{R}})$ is fixed.
-For each $s\in E_{j}(t)$, by definition there exist an irreducible $\tilde{Y}\subseteq u^{-1}(S)\cap{\mathscr{D}}^{\prime}_{j}$ such that $s\in u(\tilde{Y})$, $\dim\tilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, and that $\tilde{Y}$ is contained in a fiber of the quotient ${\mathscr{D}}^{\prime}_{j}\rightarrow{\mathscr{D}}^{\prime}_{j}/N_{j}$.
-The last condition implies that $\tilde{Y}\subseteq N_{j}({\mathbb{R}})^{+}x$ for some $x\in{\mathscr{D}}^{\prime}_{j}$.
-Hence by the discussion above, $\tilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$ for a fixed $v\in V({\mathbb{R}})$.
-Now that $\dim_{{\mathbb{C}}}\tilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, the following property holds true: For each $(a,x_{0})\in\tilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$, there exists a complex analytic subset $\tilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\tilde{C}\geq t$ such that $\{a\}\times\tilde{C}\subseteq\tilde{Y}$.
-Hence $s\in S_{{\mathscr{F}}}(t)$ by Lemma 2.1 (more precisely, the first equality in (2.1)).
+For each $s\in E_{j}(t)$, by definition there exist an irreducible $\widetilde{Y}\subseteq u^{-1}(S)\cap{\mathscr{D}}^{\prime}_{j}$ such that $s\in u(\widetilde{Y})$, $\dim\widetilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, and that $\widetilde{Y}$ is contained in a fiber of the quotient ${\mathscr{D}}^{\prime}_{j}\rightarrow{\mathscr{D}}^{\prime}_{j}/N_{j}$.
+The last condition implies that $\widetilde{Y}\subseteq N_{j}({\mathbb{R}})^{+}x$ for some $x\in{\mathscr{D}}^{\prime}_{j}$.
+Hence by the discussion above, $\widetilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$ for a fixed $v\in V({\mathbb{R}})$.
+Now that $\dim_{{\mathbb{C}}}\widetilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, the following property holds: For each $(a,x_{0})\in\widetilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$, there exists a complex analytic subset $\widetilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\widetilde{C}\geq t$ such that $\{a\}\times\widetilde{C}\subseteq\widetilde{Y}$.
+Hence $s\in S^{\mathrm{Betti}}(t)$ by Lemma 2.1 (more precisely, the first equality in (2.1)).
 Now, the conclusion of the lemma holds as $s$ runs over $E_{j}(t)$.
@@
 By (3.3), Proposition 3.9 and Lemma 3.10 (both applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
-(3.8) $S_{{\mathscr{F}}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
-So $S_{{\mathscr{F}}}(t)$ is Zariski closed in $S$ because each member in the union on the right-hand side is.
-The “In particular” part is easy to check once we have established the Zariski closedness of $S_{{\mathscr{F}}}(t)$.
+(3.8) $S^{\mathrm{Betti}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
+So $S^{\mathrm{Betti}}(t)$ is Zariski closed in $S$ because each member in the union on the right-hand side is.
+The “In particular” part is easy to check once we have established the Zariski closedness of $S^{\mathrm{Betti}}(t)$.
 ∎
@@
 The goal of this section is to prove Proposition 3.8, the finiteness result regarding weakly optimal subvarieties for admissible VMHS of weight $-1$ and 00.
-Before moving on to the proof, let us take a step back and look at this proposition from a historical point of view.
+Before moving on to our proof, let us take a step back and look at this proposition from a historical point of view.
 In studying the André–Oort conjecture, Ullmo [Ull14, Thm.
@@
 Weakly optimal subvarieties are called monodromically atypical maximal, and Proposition 3.8 is called Geometric Zilber–Pink as in [BKU24].
-Their proof does not use o-minimality, and gives some effective results.
+Their proof does not use o-minimality and gives some effective results.
 Retain the notation in (2.3).
@@
 Definition 4.1.
-A complex analytic irreducible subset $\tilde{Y}$ of $u^{-1}(S)$ is called Zariski optimal if the following holds true: $\tilde{Y}\subsetneq\tilde{Y}^{\prime}\subseteq u^{-1}(S)\Rightarrow\delta_{\mathrm{Zar}}(\tilde{Y})<\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$, with $\tilde{Y}^{\prime}$ complex analytic irreducible.
+A complex analytic irreducible subset $\widetilde{Y}$ of $u^{-1}(S)$ is called Zariski optimal if the following holds: $\widetilde{Y}\subsetneq\widetilde{Y}^{\prime}\subseteq u^{-1}(S)\Rightarrow\delta_{\mathrm{Zar}}(\widetilde{Y})<\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})$, with $\widetilde{Y}^{\prime}$ complex analytic irreducible.
 The following theorem is, in fact, an equivalent statement to the weak mixed Ax–Schanuel theorem (Theorem 3.5). In this paper, we only need one deduction.
 Theorem 4.2.
-Assume $\tilde{Y}\subseteq u^{-1}(S)$ is Zariski optimal.
-Then $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ and $\tilde{Y}$ is a complex analytic irreducible component of $\tilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$.
+Assume $\widetilde{Y}\subseteq u^{-1}(S)$ is Zariski optimal.
+Then $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$ and $\widetilde{Y}$ is a complex analytic irreducible component of $\widetilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$.
 Proof.
-Let $\tilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal.
-Then $\tilde{Y}$ is a complex analytic irreducible component of $\tilde{Y}^{\mathrm{Zar}}\cap u^{-1}(S)$.
-Hence it suffices to prove $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$.
-Let $\tilde{Y}^{\prime}$ be a complex analytic irreducible component of $\tilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$ which contains $\tilde{Y}$.
-Then $u(\tilde{Y}^{\prime})$ is closed algebraic in $S$ because $u(\tilde{Y}^{\mathrm{ws}})\cap S$ is a closed algebraic subset of $S$ by Proposition 3.3.
-Assume $\tilde{Y}\not=\tilde{Y}^{\prime}$. Then $\delta_{\mathrm{Zar}}(\tilde{Y})<\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$ by the Zariski optimality of $\tilde{Y}$. So
-$\dim\tilde{Y}^{\mathrm{Zar}}-\dim\tilde{Y}<\dim\tilde{Y}^{\prime\mathrm{Zar}}-\dim\tilde{Y}^{\prime}\leq\dim\tilde{Y}^{\mathrm{ws}}-\dim\tilde{Y}^{\prime}.$
-Now that $u(\tilde{Y})\subseteq u(\tilde{Y}^{\prime})$, we have $u(\tilde{Y})^{\mathrm{Zar}}\subseteq u(\tilde{Y}^{\prime})$ because the right hand side is closed algebraic in $S$. So
-$\dim u(\tilde{Y})^{\mathrm{Zar}}\leq\dim u(\tilde{Y}^{\prime})=\dim\tilde{Y}^{\prime}.$
+Let $\widetilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal.
+Then $\widetilde{Y}$ is a complex analytic irreducible component of $\widetilde{Y}^{\mathrm{Zar}}\cap u^{-1}(S)$.
+Hence it suffices to prove $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$.
+Let $\widetilde{Y}^{\prime}$ be a complex analytic irreducible component of $\widetilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$ which contains $\widetilde{Y}$.
+Then $u(\widetilde{Y}^{\prime})$ is closed algebraic in $S$ because $u(\widetilde{Y}^{\mathrm{ws}})\cap S$ is a closed algebraic subset of $S$ by Proposition 3.3.
+Assume $\widetilde{Y}\not=\widetilde{Y}^{\prime}$. Then $\delta_{\mathrm{Zar}}(\widetilde{Y})<\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})$ by the Zariski optimality of $\widetilde{Y}$. So
+$\dim\widetilde{Y}^{\mathrm{Zar}}-\dim\widetilde{Y}<\dim\widetilde{Y}^{\prime\mathrm{Zar}}-\dim\widetilde{Y}^{\prime}\leq\dim\widetilde{Y}^{\mathrm{ws}}-\dim\widetilde{Y}^{\prime}.$
+Now that $u(\widetilde{Y})\subseteq u(\widetilde{Y}^{\prime})$, we have $u(\widetilde{Y})^{\mathrm{Zar}}\subseteq u(\widetilde{Y}^{\prime})$ because the right hand side is closed algebraic in $S$. So
+$\dim u(\widetilde{Y})^{\mathrm{Zar}}\leq\dim u(\widetilde{Y}^{\prime})=\dim\widetilde{Y}^{\prime}.$
 These two inequalities together yield.
-$\dim\tilde{Y}^{\mathrm{Zar}}+\dim u(\tilde{Y})^{\mathrm{Zar}}<\dim\tilde{Y}^{\mathrm{ws}}+\dim\tilde{Y}.$
-This contradicts the weak mixed Ax–Schanuel theorem (Theorem 3.5) for $\tilde{Y}$.
-Hence we must have $\tilde{Y}=\tilde{Y}^{\prime}$. In particular, $u(\tilde{Y})$ is algebraic. So $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$. ∎
-For $Y\subseteq S$ irreducible, let $\tilde{Y}$ be a complex analytic irreducible component of $u^{-1}(Y)$.
+$\dim\widetilde{Y}^{\mathrm{Zar}}+\dim u(\widetilde{Y})^{\mathrm{Zar}}<\dim\widetilde{Y}^{\mathrm{ws}}+\dim\widetilde{Y}.$
+This contradicts the weak mixed Ax–Schanuel theorem (Theorem 3.5) for $\widetilde{Y}$.
+Hence we must have $\widetilde{Y}=\widetilde{Y}^{\prime}$. In particular, $u(\widetilde{Y})$ is algebraic. So $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$. ∎
+For $Y\subseteq S$ irreducible, let $\widetilde{Y}$ be a complex analytic irreducible component of $u^{-1}(Y)$.
 Corollary 4.3.
-If $Y\subseteq X$ is weakly optimal, then $\tilde{Y}$ is Zariski optimal in $u^{-1}(S)$.
+If $Y\subseteq X$ is weakly optimal, then $\widetilde{Y}$ is Zariski optimal in $u^{-1}(S)$.
 Proof.
-Let $\tilde{Y}^{\prime}\supseteq\tilde{Y}$ be a complex analytic irreducible subset of $u^{-1}(S)$ such that $\delta_{\mathrm{Zar}}(\tilde{Y})\geq\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$.
-We may and do assume that $\tilde{Y}^{\prime}$ is Zariski optimal.
-Set $Y^{\prime\prime}:=u(\tilde{Y}^{\prime})^{\mathrm{Zar}}$.
-Then $Y^{\prime\prime}\subseteq u(\tilde{Y}^{\prime,\mathrm{ws}})\cap S$ because $u(\tilde{Y}^{\prime,\mathrm{ws}})\cap S$ is closed and algebraic.
+Let $\widetilde{Y}^{\prime}\supseteq\widetilde{Y}$ be a complex analytic irreducible subset of $u^{-1}(S)$ such that $\delta_{\mathrm{Zar}}(\widetilde{Y})\geq\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})$.
+We may and do assume that $\widetilde{Y}^{\prime}$ is Zariski optimal.
+Set $Y^{\prime\prime}:=u(\widetilde{Y}^{\prime})^{\mathrm{Zar}}$.
+Then $Y^{\prime\prime}\subseteq u(\widetilde{Y}^{\prime,\mathrm{ws}})\cap S$ because $u(\widetilde{Y}^{\prime,\mathrm{ws}})\cap S$ is closed and algebraic.
 Thus
-$\delta_{\mathrm{ws}}(Y^{\prime\prime})=\dim Y^{\prime\prime,\mathrm{ws}}-\dim Y^{\prime\prime}\leq\dim\tilde{Y}^{\prime,\mathrm{ws}}-\dim\tilde{Y}^{\prime}.$
-Since $\tilde{Y}^{\prime}$ is Zariski optimal, we have $\tilde{Y}^{\prime,\mathrm{Zar}}=\tilde{Y}^{\prime,\mathrm{ws}}$ by Theorem 4.2. So, the inequality above further implies
-$\delta_{\mathrm{ws}}(Y^{\prime\prime})\leq\dim\tilde{Y}^{\prime,\mathrm{Zar}}-\dim\tilde{Y}^{\prime}=\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})\leq\delta_{\mathrm{Zar}}(\tilde{Y})\leq\dim\tilde{Y}^{\mathrm{ws}}-\dim\tilde{Y}=\delta_{\mathrm{ws}}(Y).$
-So $Y^{\prime\prime}=Y$ because $Y$ is weakly optimal. Thus $\tilde{Y}=\tilde{Y}^{\prime}$ is Zariski optimal. ∎
+$\delta_{\mathrm{ws}}(Y^{\prime\prime})=\dim Y^{\prime\prime,\mathrm{ws}}-\dim Y^{\prime\prime}\leq\dim\widetilde{Y}^{\prime,\mathrm{ws}}-\dim\widetilde{Y}^{\prime}.$
+Since $\widetilde{Y}^{\prime}$ is Zariski optimal, we have $\widetilde{Y}^{\prime,\mathrm{Zar}}=\widetilde{Y}^{\prime,\mathrm{ws}}$ by Theorem 4.2. So, the inequality above further implies
+$\delta_{\mathrm{ws}}(Y^{\prime\prime})\leq\dim\widetilde{Y}^{\prime,\mathrm{Zar}}-\dim\widetilde{Y}^{\prime}=\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})\leq\delta_{\mathrm{Zar}}(\widetilde{Y})\leq\dim\widetilde{Y}^{\mathrm{ws}}-\dim\widetilde{Y}=\delta_{\mathrm{ws}}(Y).$
+So $Y^{\prime\prime}=Y$ because $Y$ is weakly optimal. Thus $\widetilde{Y}=\widetilde{Y}^{\prime}$ is Zariski optimal. ∎
 4.2. A parametrization of Zariski optimal subsets
@@
 Fix a Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$.
-Then for the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), the action of $\mathbf{G}({\mathbb{R}})$ on ${\mathscr{D}}$ is given by $(v,g_{0})\cdot(x_{V},x_{0})=(v+g_{0}x_{V},g_{0}x_{0})$.
+Then for the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), the action of $\mathbf{G}({\mathbb{R}})$ on ${\mathscr{D}}$ is given by
+$(v,g_{0})\cdot(x_{V},x_{0})=(v+g_{0}x_{V},g_{0}x_{0}).$
 […]
@@
     the set of orbits $\{H({\mathbb{R}})x:(x,H)\in{\mathscr{N}}^{0}\}$;
-    $\{\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}:\tilde{Y}\subseteq u^{-1}(S)\text{ Zariski optimal, with }\tilde{Y}\cap\mathfrak{F}\not=\emptyset\}$.
+    $\{\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}:\widetilde{Y}\subseteq u^{-1}(S)\text{ Zariski optimal, with }\widetilde{Y}\cap\mathfrak{F}\not=\emptyset\}$.
 Moreover, $H({\mathbb{R}})x$ is a weak Mumford–Tate domain for each $(x,H)\in{\mathscr{N}}^{0}$.
@@
 Take $(x,H)\in{\mathscr{N}}^{0}$.
-We wish to prove that $H({\mathbb{R}})x$ equals $\tilde{Y}^{\mathrm{Zar}}$ for some Zariski optimal $\tilde{Y}\subseteq u^{-1}(S)$ which passes through $x\in\mathfrak{F}$.
-Let $\tilde{Y}^{\prime}$ be a complex analytic irreducible component of $u^{-1}(S)\cap H({\mathbb{R}})x$ which passes through $x$ with $\dim\tilde{Y}^{\prime}=d(x,H)$.
-Take $\tilde{Y}\supseteq\tilde{Y}^{\prime}$ with $\tilde{Y}\subseteq u^{-1}(S)$ complex analytic irreducible and $\delta_{\mathrm{Zar}}(\tilde{Y})\leq\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$.
-We may and do assume that $\tilde{Y}$ is Zariski optimal.
-Then $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain by Theorem 4.2, and hence $\tilde{Y}^{\mathrm{Zar}}=H^{\prime}({\mathbb{R}})x$ for some $(x,H^{\prime})\in{\mathscr{N}}$.
-Thus $\delta_{\mathrm{Zar}}(\tilde{Y})=\delta(x,H^{\prime})$.
-Thus $\tilde{Y}^{\prime}$ is contained in some irreducible component of $H({\mathbb{R}})x\cap H^{\prime}({\mathbb{R}})x$, which by Lemma 4.4.(ii) is $H^{\prime\prime}({\mathbb{R}})x$ for some $(x,H^{\prime\prime})\in{\mathscr{N}}$.
+We wish to prove that $H({\mathbb{R}})x$ equals $\widetilde{Y}^{\mathrm{Zar}}$ for some Zariski optimal $\widetilde{Y}\subseteq u^{-1}(S)$ which passes through $x\in\mathfrak{F}$.
+Let $\widetilde{Y}^{\prime}$ be a complex analytic irreducible component of $u^{-1}(S)\cap H({\mathbb{R}})x$ which passes through $x$ with $\dim\widetilde{Y}^{\prime}=d(x,H)$.
+Take $\widetilde{Y}\supseteq\widetilde{Y}^{\prime}$ with $\widetilde{Y}\subseteq u^{-1}(S)$ complex analytic irreducible and $\delta_{\mathrm{Zar}}(\widetilde{Y})\leq\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})$.
+We may and do assume that $\widetilde{Y}$ is Zariski optimal.
+Then $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain by Theorem 4.2, and hence $\widetilde{Y}^{\mathrm{Zar}}=H^{\prime}({\mathbb{R}})x$ for some $(x,H^{\prime})\in{\mathscr{N}}$.
+Thus $\delta_{\mathrm{Zar}}(\widetilde{Y})=\delta(x,H^{\prime})$.
+Thus $\widetilde{Y}^{\prime}$ is contained in some irreducible component of $H({\mathbb{R}})x\cap H^{\prime}({\mathbb{R}})x$, which by Lemma 4.4.(ii) is $H^{\prime\prime}({\mathbb{R}})x$ for some $(x,H^{\prime\prime})\in{\mathscr{N}}$.
 Hence $d(x,H)=d(x,H^{\prime\prime})$, so by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime\prime}({\mathbb{R}})x$.
@@
 But $\delta(x,H)=\delta(x,H^{\prime})$.
-So by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime}({\mathbb{R}})x$, and hence $H({\mathbb{R}})x=\tilde{Y}^{\mathrm{Zar}}$.
+So by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime}({\mathbb{R}})x$, and hence $H({\mathbb{R}})x=\widetilde{Y}^{\mathrm{Zar}}$.
 Now we have proved that the first set is contained in the second set and the “Moreover” part of the proposition.
-Conversely let $\tilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal with $\tilde{Y}\cap\mathfrak{F}\not=\emptyset$.
-By Theorem 4.2, $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain, and hence equals $H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}$.
-We may furthermore choose this $x$ to be a smooth point of $H({\mathbb{R}})x\cap u^{-1}(S)\cap\mathfrak{F}$; then $\dim\tilde{Y}=d(x,H)$.
+Conversely let $\widetilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal with $\widetilde{Y}\cap\mathfrak{F}\not=\emptyset$.
+By Theorem 4.2, $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain, and hence equals $H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}$.
+We may furthermore choose this $x$ to be a smooth point of $H({\mathbb{R}})x\cap u^{-1}(S)\cap\mathfrak{F}$; then $\dim\widetilde{Y}=d(x,H)$.
 We wish to prove $(x,H)\in{\mathscr{N}}^{0}$. Let us check the two properties which define ${\mathscr{N}}^{0}$ by contradiction.
 Assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x$ and $d(x,H)=d(x,H^{\prime})$.
-Then $\dim\tilde{Y}=d(x,H^{\prime})$ and hence $\tilde{Y}\subseteq H^{\prime}({\mathbb{R}})x$ by analytic continuation.
-But then $\tilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$ by Lemma 4.4.(i).
-So $\tilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x=\tilde{Y}^{\mathrm{Zar}}$, which is impossible.
+Then $\dim\widetilde{Y}=d(x,H^{\prime})$ and hence $\widetilde{Y}\subseteq H^{\prime}({\mathbb{R}})x$ by analytic continuation.
+But then $\widetilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$ by Lemma 4.4.(i).
+So $\widetilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x=\widetilde{Y}^{\mathrm{Zar}}$, which is impossible.
 So for any $(x,H^{\prime})\in{\mathscr{N}}$, we have $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x\Rightarrow d(x,H)>d(x,H^{\prime})$.
 Next, assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x$ and $\delta(x,H)\geq\delta(x,H^{\prime})$.
-Take a complex analytic irreducible component $\tilde{Y}^{\prime}$ of $u^{-1}(S)\bigcap H^{\prime}({\mathbb{R}})x$ such that $\dim\tilde{Y}^{\prime}=d(x,H^{\prime})$.
-Then $\tilde{Y}\subseteq\tilde{Y}^{\prime}$ by analytic continuation.
-By Lemma 4.4.(i), we have $\tilde{Y}^{\prime,\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$.
-So $\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})\leq\delta(x,H^{\prime})\leq\delta(x,H)=\delta_{\mathrm{Zar}}(\tilde{Y})$.
-Hence $\tilde{Y}^{\prime}=\tilde{Y}$ because $\tilde{Y}$ is Zariski optimal.
+Take a complex analytic irreducible component $\widetilde{Y}^{\prime}$ of $u^{-1}(S)\bigcap H^{\prime}({\mathbb{R}})x$ such that $\dim\widetilde{Y}^{\prime}=d(x,H^{\prime})$.
+Then $\widetilde{Y}\subseteq\widetilde{Y}^{\prime}$ by analytic continuation.
+By Lemma 4.4.(i), we have $\widetilde{Y}^{\prime,\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$.
+So $\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})\leq\delta(x,H^{\prime})\leq\delta(x,H)=\delta_{\mathrm{Zar}}(\widetilde{Y})$.
+Hence $\widetilde{Y}^{\prime}=\widetilde{Y}$ because $\widetilde{Y}$ is Zariski optimal.
 So $\delta(x,H^{\prime})=\delta(x,H)$ and $d(x,H^{\prime})=d(x,H)$, and so
@@
 Let $Y$ be a weakly optimal subvariety of $S$.
-Take a complex analytic irreducible component $\tilde{Y}$ of $u^{-1}(Y)$ such that $\tilde{Y}\cap\mathfrak{F}\not=\emptyset$.
-Then $\tilde{Y}$ is Zariski optimal in $u^{-1}(S)$ by Corollary 4.3.
-Then by Proposition 4.5, we have $\tilde{Y}^{\mathrm{ws}}=H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}^{0}$.
-Hence by the finiteness result Proposition 4.6, there exist finitely many ${\mathbb{Q}}$-groups $N_{1},\ldots,N_{k}<\mathbf{G}$ satisfying the following property: For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}=u(\tilde{Y}^{\mathrm{ws}})$ equals $u(N_{j}({\mathbb{R}})^{+}x)$ for some $j\in\{1,\ldots,k\}$ and some $x\in{\mathscr{D}}$ with $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})$.
+Take a complex analytic irreducible component $\widetilde{Y}$ of $u^{-1}(Y)$ such that $\widetilde{Y}\cap\mathfrak{F}\not=\emptyset$.
+Then $\widetilde{Y}$ is Zariski optimal in $u^{-1}(S)$ by Corollary 4.3.
+Then by Proposition 4.5, we have $\widetilde{Y}^{\mathrm{ws}}=H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}^{0}$.
+Hence by the finiteness result Proposition 4.6, there exist finitely many ${\mathbb{Q}}$-groups $N_{1},\ldots,N_{k}<\mathbf{G}$ satisfying the following property: For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}=u(\widetilde{Y}^{\mathrm{ws}})$ equals $u(N_{j}({\mathbb{R}})^{+}x)$ for some $j\in\{1,\ldots,k\}$ and some $x\in{\mathscr{D}}$ with $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})$.
 Then ${\mathscr{D}}^{\prime}_{j}:=\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})^{+}x$ is a Mumford–Tate domain and is independent of the choice of such $x$.
@@
 ∎
-5. Application to non-degeneracy in case of irreducible VHSs
-In this section, we give two applications of Theorem 3.1: when the VHS is irreducible and when it has a simple algebraic monodromy group.
+5. Application to non-degeneracy in some geometric cases
+In this section, we give two applications of Theorem 3.1: when the VHS is irreducible or when it has a simple algebraic monodromy group.
 Both cases apply to the Gross-Shoen and the Ceresa normal functions.
@@
 Recall the period map $\varphi=\varphi_{\nu}\colon S\rightarrow\Gamma\backslash{\mathscr{D}}$, the ${\mathbb{Q}}$-group $\mathbf{G}$ and its unipotent radical $V$ which is a vector group from (2.3).
-We emphasize that $V$ is in general not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$ and its geometric meaning is given by Remark 2.2: $\frac{1}{2}\dim V$ is the relative dimension of ${\mathscr{J}}_{\nu}$, the smallest intermediate Jacobian of sub-VHS of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ translated by a torsion multisection which contains $\nu(S)$.
+We emphasize that $V$ is, in general, not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$ and its geometric meaning is given by Remark 2.2: $\frac{1}{2}\dim V$ is the relative dimension of ${\mathscr{J}}_{\nu}$, the smallest intermediate Jacobian of sub-VHS ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ of ${\mathbb{V}}_{{\mathbb{Z}}}\rightarrow S$ translated by a torsion multisection which contains $\nu(S)$.
 Notice that the trivial upper bound (2.8) yields the following necessary condition for $r(\nu)=\dim S$:
@@
 In particular, if futhermore $\dim\varphi(S)=\dim S$ and $\dim{\mathbb{V}}_{{\mathbb{Q}},s}\geq 2\dim S$, then we have $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$.
+We can say more in the situation of Theorem 5.1.
+In fact in this case ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}={\mathbb{V}}_{{\mathbb{Z}}}$.
+By the comment below (2.8), the equality $r(\nu)=\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}$ can be attained only if ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is an abelian scheme.
+Hence $\dim{\mathbb{V}}_{{\mathbb{Q}},s}>2\dim\varphi(S)$ in the situation of Theorem 5.1 if ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is not an abelian scheme.
 Proof.
-Set $\mathbf{G}_{0}=\mathbf{G}/V$. Then $V$ is a $\mathbf{G}_{0}$-submodule of ${\mathbb{V}}_{{\mathbb{Q}},s}$ for one (and hence all) $s\in S({\mathbb{C}})$.
-By (i), ${\mathbb{V}}_{{\mathbb{Q}},s}$ is irreducible as a $\mathbf{G}_{0}$-module. By (ii), $V\not=\{0\}$. Hence $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
+Set $\mathbf{G}_{0}=\mathbf{G}/V$.
+Then $V$ is a $\mathbf{G}_{0}$-submodule of ${\mathbb{V}}_{{\mathbb{Q}},s}$ for one (and hence all) $s\in S({\mathbb{C}})$, and $\mathbf{G}_{0}$ is by definition a subgroup of ${\mathrm{GL}}({\mathbb{V}}_{{\mathbb{Q}},s})$.
+By (i), ${\mathbb{V}}_{{\mathbb{Q}},s}$ is irreducible as a $\mathbf{G}_{0}$-module.
+By (ii), $V\not=\{0\}$.
+Hence $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
+Thus we have $\mathbf{G}_{0}\subset{\mathrm{GL}}(V)$.
 Let $N$ be a normal subgroup of $\mathbf{G}$.
@@
 Hence by (i), either $V\cap N=\{0\}$ or $V\cap N=V$.
-Assume $V\cap N=\{0\}$.
-Then $N\lhd\mathbf{G}_{0}=\mathbf{G}/V$ is reductive and it acts trivially on $V=V/(V\cap N)$.
-Hence, $N({\mathbb{R}})^{+}x_{0}$ is a point for any $x_{0}\in{\mathscr{D}}_{0}$, and therefore $N$ is contained in the center of $\mathbf{G}_{0}$.
-Using again the fact that $N$ acts trivially on $V$, we see that $N$ is contained in the center of $\mathbf{G}$.
-So $\dim\varphi_{/N}(S)=\dim\varphi(S)$ in this case.
-On the other hand, it is clearly true that $\min_{N,~{}V\cap N=V}\left\{\dim\varphi_{/N}(S)\right\}+\frac{1}{2}\dim V$ is attained at $N=\mathbf{G}$ and hence equals $\frac{1}{2}\dim V$.
-Now we are done.
-∎
+In the first case, we have the image $N_{0}$ of $N$ acting trivially on $V$. Thus $N_{0}=0$. This shows that $N$ is trivial. So $r(\nu)=\dim\varphi(S)$.
+In the second case, $V\cap N=V$. It is clearly true that
+$\min_{N,~{}V\cap N=V}\left\{\dim\varphi_{/N}(S)\right\}+\frac{1}{2}\dim V$
+is attained at $N=\mathbf{G}$ and hence equals $\frac{1}{2}\dim V$. ∎
 Theorem 5.2.
@@
 $r(\nu)=\min\left\{\dim\varphi(S),\frac{1}{2}\dim V\right\}.$
+An important case where assumption (i) is satisfied is when the Mumford–Tate group of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is quasi-simple; this follows from Deligne [Del87, §$\S$7.5].
+And similar to the comment below Theorem 5.1, we must have $\dim V>2\dim\varphi(S)$ in the situation of Theorem 5.2 if ${\mathscr{J}}({\mathbb{V}}^{\prime}_{{\mathbb{Z}}})\rightarrow S$ is not an abelian scheme.
 Proof.
-Set $\mathbf{G}_{0}=\mathbf{G}/V$. By Deligne, $H\lhd\mathbf{G}^{\mathrm{der}}_{0}$.
+Set $\mathbf{G}_{0}=\mathbf{G}/V$. By Deligne [Del87, §$\S$7.5], $H\lhd\mathbf{G}^{\mathrm{der}}_{0}$.
 Let $N$ be a normal subgroup of $\mathbf{G}$.
-We may and do assume $N\lhd\mathbf{G}^{\mathrm{der}}$.
-The reductive part $\mathbf{G}_{N}:=N/(V\cap N)$ is a normal subgroup of $\mathbf{G}$.
-Now that $\mathbf{G}_{N}\cap H$ is a normal subgroup of $H$, we have either $\mathbf{G}_{N}\cap H=\{1\}$ or $H<\mathbf{G}_{N}$ by (i).
-Assume $H<\mathbf{G}_{N}$.
+The reductive part $N_{0}:=N/(V\cap N)$ is a normal subgroup of $\mathbf{G}_{0}$.
+Now that $N_{0}\cap H$ is a normal subgroup of $H$, by (i), we have that either $N_{0}\cap H$ is finite or $H<N_{0}$.
+Assume $H<N_{0}$.
 Since $\mathbf{G}_{0}$ is reductive, we can decompose $V=(V\cap N)\bigoplus(V\cap N)^{\perp}$ as $\mathbf{G}_{0}$-modules.
 Now $(V\cap N)^{\perp}$ gives rise to a sub-VHS of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$.
-But $H$ acts trivially on $(V\cap N)^{\perp}$ because $\mathbf{G}_{N}$ acts trivially on $V/(V\cap N)$.
+But $H$ acts trivially on $(V\cap N)^{\perp}$ because $N_{0}$ acts trivially on $V/(V\cap N)$.
 So $(V\cap N)^{\perp}=\{0\}$ by (ii).
 Hence $V\cap N=V$, and hence $V<N$.
-Thus $\varphi_{/N}(S)$ is a point.
-So $\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)=\frac{1}{2}\dim V$.
-If $\mathbf{G}_{N}\cap H=\{1\}$, then $\dim\varphi_{/N}(S)=\dim\varphi(S)$. Now we are done. ∎
-Now, we apply these two theorems to the Gross-Shoen and the Ceresa normal functions.
+So in this case we have
+$\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)\geq\frac{1}{2}\dim V.$
+The equality obtains when $N=\mathbf{G}$.
+Assume that $N_{0}\cap H$ is finite.
+To handle this case, we again use the universal period map of the VHS $\widetilde{\varphi}_{0}\colon\widetilde{S}\rightarrow{\mathscr{D}}_{0}$ associated with the VHS ${\mathbb{V}}_{{\mathbb{Z}}}\rightarrow S$ from (2.3), where ${\mathscr{D}}_{0}$ is a $\mathbf{G}_{0}({\mathbb{R}})^{+}$-orbit and the natural projection $p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}$ is induced by the quotient $\mathbf{G}\rightarrow\mathbf{G}_{0}$.
+Moreover, we have $\widetilde{\varphi}_{0}=p\circ\widetilde{\varphi}$.
+By logarithmic Ax [GK24, Thm.
+7.2], we have $\widetilde{\varphi}_{0}(\widetilde{S})\subseteq H({\mathbb{R}})^{+}x_{0}$ for some $x_{0}\in{\mathscr{D}}_{0}$.
+So $\widetilde{\varphi}(\widetilde{S})\subseteq p^{-1}(\widetilde{\varphi}_{0}(\widetilde{S}))\subseteq p^{-1}(H({\mathbb{R}})^{+}x_{0})$.
+Notice that $p^{-1}(H({\mathbb{R}})^{+}x_{0})=(V\rtimes H)({\mathbb{R}})^{+}x$ for some $x\in{\mathscr{D}}$.
+So $\widetilde{\varphi}(\widetilde{S})\subseteq(V\rtimes H)({\mathbb{R}})^{+}x$, and hence $\varphi(S)\subseteq u((V\rtimes H)({\mathbb{R}})^{+}x)$ for $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$.
+[…]
+$p_{N}\colon\mathbf{G}\xrightarrow{p_{V\cap N}}\mathbf{G}/(V\cap N)\xrightarrow{p_{N_{0}}}\mathbf{G}/N,$
+and this induces
+$[p_{N}]\colon\Gamma\backslash{\mathscr{D}}\xrightarrow{[p_{V\cap N}]}\Gamma_{/V\cap N}\backslash({\mathscr{D}}/(V\cap N))\xrightarrow{[p_{N_{0}}]}\Gamma_{/N}\backslash({\mathscr{D}}/N).$
+It is clearly true that $\dim\varphi(S)=\dim([p_{V\cap N}]\circ\varphi)(S)$.
+On the other hand, $\dim([p_{V\cap N}]\circ\varphi)(S)=\dim([p_{N_{0}}]\circ[p_{V\cap N}]\circ\varphi)(S)$ by the conclusion of the previous paragraph.
+So $\dim\varphi_{/N}(S)=\dim\varphi(S)$ in this case.
+Now, we are done by combining the two cases above. ∎
+Now, we apply these two theorems to the Gross–Schoen and the Ceresa normal functions.
 Let $\nu_{{\mathrm{GS}}}$ (resp.
@@
 We shall apply Theorem 5.1 and let us check the assumptions.
-Assumption (i) holds by Proposition C.10.
-(iii).
-Assumption (ii) holds by Proposition C.10.
-(ii).
+Assumption (i) holds by part (iii) of Proposition C.10.
+Assumption (ii) holds by part (ii) of Proposition C.10.
 Moreover the VHS ${\mathbb{V}}_{{\mathbb{Z}}}$ on $S$ has maximal moduli, so $\dim\varphi_{{\mathrm{GS}}}({\mathscr{M}}_{g})=\dim\varphi_{{\mathrm{Ce}}}({\mathscr{M}}_{g})=\dim{\mathscr{M}}_{g}=3g-3$ where $\varphi_{{\mathrm{GS}}}$ and $\varphi_{{\mathrm{Ce}}}$ are the period maps.
@@
 Denote for simplicity by $S={\mathscr{M}}_{g}$. Let $\nu$ be either $\nu_{{\mathrm{GS}}}$ or $\nu_{{\mathrm{Ce}}}$.
-Set $U:=S\setminus S_{{\mathscr{F}}}(1)$, which is Zariski open by Theorem 3.2.
-By Corollary 5.3 and the “In particular” part of Theorem 3.2, we have that $\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}=0$. Thus $S_{{\mathscr{F}}}(1)\not=S$. Now, part (i) follows from Proposition C.7.
-For (ii), notice that by definition $S_{{\mathscr{F}}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion.
+Set $U:=S\setminus S^{\mathrm{Betti}}(1)$, which is Zariski open by Theorem 3.2.
+By Corollary 5.3 and the “In particular” part of Theorem 3.2, we have that $\min\{t\geq 0:S^{\mathrm{Betti}}(t)=S\}=0$. Thus $S^{\mathrm{Betti}}(1)\not=S$. Now, part (i) follows from Proposition C.7.
+For (ii), notice that by definition $S^{\mathrm{Betti}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion.
 Thus, the set $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is discrete, and hence at most countable.
 ∎
-6. Northcott property for Gross–Schoen and Ceresa cycles
+6. Beilinson–Bloch height of Gross–Schoen and Ceresa cycles
 In this section, we turn to the arithmetic applications.
@@
 (6.2) ¯$\overline{{\mathscr{L}}}:=\langle\overline{{\mathscr{Q}}},\overline{{\mathscr{Q}}},\overline{{\mathscr{Q}}}\rangle\in\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}.$
-The following theorem is a reformulation of the second-named author’s [Zha10, Thm. 2.3.5].
+The line bundle $\overline{{\mathscr{L}}}$ has been constructed [Yua21, §$\S$3.3.4] in slightly different step.
+The following theorem is a reformulation of the second-named author’s [Zha10, Thm.
+2.3.5].
+Recall the notation ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{s})$ with $\xi_{s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ such that $(2g-2)\xi_{s}$ being the canonical divisor class on ${\mathscr{C}}_{s}$.
 Theorem 6.1.
@@
 Proof.
-In some neighborhood $U$ of $s$ in ${\mathscr{M}}_{g}$, $\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})=\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is an element of $Z^{2}({\mathscr{C}}_{U}^{3})$, i.e.
+As notation, for any subset $U\subseteq{\mathscr{M}}_{g}$, set $\xi_{U}\in\mathrm{Pic}^{1}({\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}U/U)$ to be such that $(2g-2)\xi_{U}$ is the canonical divisor class of ${\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}U/U$.
+In some neighborhood $U$ of $s$ in ${\mathscr{M}}_{g}$, $\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})=\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$ is an element of $Z^{2}({\mathscr{C}}_{U}^{3})$, i.e.
 it is a $2$-cocycle of ${\mathscr{C}}_{U}^{3}$.
@@
 Up to shrinking $U$ we can take rational sections $\ell_{1},\ell_{2},\ell_{3}$ of ${\mathscr{Q}}$ over ${\mathscr{C}}_{U}^{2}$ with the following property: For their divisors $t_{1},t_{2},t_{3}\in\mathrm{Div}({\mathscr{C}}_{U}^{2})$, we have
-$|t_{1}|\cap|t_{2}|\cap|t_{3}|=\emptyset,\qquad|\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})|\cap|(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})|=\emptyset,$
+$|t_{1}|\cap|t_{2}|\cap|t_{3}|=\emptyset,\qquad|\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})|\cap|(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})|=\emptyset,$
 where $t_{1}\otimes t_{2}\otimes t_{3}$ is seen as a correspondence of ${\mathscr{C}}_{U}^{3}$.
@@
 The right hand side is precisely $\frac{1}{[{\mathbb{Q}}(s):{\mathbb{Q}}]}\widehat{\deg}\overline{{\mathscr{L}}}|_{s}=h_{\overline{{\mathscr{L}}}}(s)$.
-The left hand side is $\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}$ because, by (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$.
+The left hand side is $\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}$ because, by (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$.
 Hence, we are done.
@@
 Let $\nu_{{\mathrm{GS}}}$ be the Gross–Schoen normal function ${\mathscr{M}}_{g}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defined in (C.6) (with ${\mathbb{V}}_{{\mathbb{Z}}}$ from Proposition C.10).
-By (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ Hence, they define the same normal functions $\nu_{{\mathrm{GS}}}$ on $U$.
+By (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$ Hence, they define the same normal function $\nu_{{\mathrm{GS}}}$ on $U$.
 Set
@@
 By [YZ21, Thm.
-5.3.5], lower bounds of the height function $h_{\overline{{\mathscr{L}}}}$ correspond to bigness properties of $\overline{{\mathscr{L}}}$.
-In this paper, we prove the bigness of the generic fiber of $\overline{{\mathscr{L}}}$ and deduce the desired height comparison from it.
-There is a natural base change of adelic line bundles $\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}\rightarrow\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Q}})_{{\mathbb{Q}}}$.
-Denote by $\tilde{{\mathscr{L}}}$ the image of $\overline{{\mathscr{L}}}$ under this base change map; $\tilde{{\mathscr{L}}}$ is called the generic fiber of $\overline{{\mathscr{L}}}$.
-We start by stating the following formula for $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})$.
-Notice that it does not follow from arithmetic Hilbert–Samuel because we have not proved the nefness of $\tilde{{\mathscr{L}}}$.
-Proposition 6.3.
-We have
-$\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})^{3g-3}.$
-The proof of this proposition shall be postponed to §$\S$6.4.
+5.3.5] (which we reformulate as Theorem E.9), lower bounds of the height function $h_{\overline{{\mathscr{L}}}}$ correspond to bigness properties of $\overline{{\mathscr{L}}}$.
+In this paper, we prove the bigness of the generic fiber $\widetilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$ and deduce the desired height comparison from it.
+The definition of $\widetilde{{\mathscr{L}}}$ is recalled in §$\S$§E.5.
+Roughly speaking, there is a natural base change of adelic line bundles $\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}\rightarrow\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Q}})_{{\mathbb{Q}}}$, and $\widetilde{{\mathscr{L}}}$ is the image of $\overline{{\mathscr{L}}}$ under this base change.
+The arithmetic volume of $\widetilde{{\mathscr{L}}}$ is defined in (E.11).
+We will prove the following formula for $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}})$.
+Notice that it does not follow from arithmetic Hilbert–Samuel because we have yet to prove the nefness of $\widetilde{{\mathscr{L}}}$.
 Theorem 6.3.
-Assume $g\geq 3$. Then the adelic line bundle $\tilde{{\mathscr{L}}}$ is big, i.e. $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})>0$.
+Let $S\subseteq{\mathscr{M}}_{g}$ be an irreducible subvariety, and let ~$\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S}$ be the image of $\widetilde{{\mathscr{L}}}$ under the natural base change $\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Q}})_{{\mathbb{Q}}}\rightarrow\widehat{\mathrm{Pic}}(S/{\mathbb{C}})_{{\mathbb{Q}}}$.
+Then we have
+(6.4) $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S})=\int_{S({\mathbb{C}})}c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})^{\wedge\dim S}=\int_{S({\mathbb{C}})}\beta_{{\mathrm{GS}}}^{\wedge\dim S}.$
+The proof of this theorem shall be postponed to §$\S$6.4.
+An interesting result of this volume computation is as follows.
+For each subvariety $S$ of ${\mathscr{M}}_{g}$, consider the Betti strata (1.2) with $\nu=\nu_{{\mathrm{GS}}}$.
+Corollary 6.4.
+Let $S$ be a subvariety of ${\mathscr{M}}_{g}$ defined over a subfield $k$ of ${\mathbb{C}}$.
+Then $S^{\mathrm{Betti}}(1)$ is also defined over $k$.
+In particular ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is defined over ${\mathbb{Q}}$.
 Proof.
-By Proposition 6.3 and Proposition 6.2, we have $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}$.
-Now $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\geq 0$ since $\beta_{{\mathrm{GS}}}$ is semi-positive.
-By Corollary 5.3, $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\not\equiv 0$.
-Hence $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})>0$.
+Let $S^{\prime}$ be an irreducible component of $S^{\mathrm{Betti}}(1)$.
+Then $S^{\prime}=S^{\prime}_{{\mathscr{F}}}(1)$, and hence $\beta_{{\mathrm{GS}}}^{\wedge\dim S^{\prime}}\equiv 0$ on $S^{\prime}$; see (1.6) and Corollary D.7.
+So $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S^{\prime}})=0$ by Theorem 6.3.
+By Definition E.11, for any automorphism $\sigma\in{\mathrm{Aut}}({\mathbb{C}}/k)$ we have
+$\widehat{\mathrm{vol}}(\widetilde{\mathscr{L}}_{{\mathbb{C}}}\otimes_{\sigma}{\mathbb{C}}|_{S^{\prime}\otimes_{\sigma}{\mathbb{C}}})=\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S^{\prime}})=0.$
+Let $S^{\prime\sigma}$ denote $S^{\prime}\otimes_{\sigma}{\mathbb{C}}$.
+Then $S^{\prime\sigma}$ is also a closed subvariety of $S$.
+As $\widetilde{\mathscr{L}}$ is defined over ${\mathbb{Q}}$, Proposition 6.3 implies
+$\int_{S^{\prime\sigma}({\mathbb{C}})}c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})^{\wedge\dim S^{\prime}}=0.$
+Thus $\beta_{{\mathrm{GS}}}^{\wedge\dim S^{\prime}}\equiv 0$ on $S^{\prime\sigma}$ by Proposition 6.2, and therefore
+$S^{\prime\sigma}({\mathbb{C}})=S^{\prime\sigma}_{\mathscr{F}}(1)\subseteq S^{\mathrm{Betti}}(1).$
+This shows that $S^{\mathrm{Betti}}(1)\otimes_{\sigma}{\mathbb{C}}=S^{\mathrm{Betti}}(1)$. Thus $S^{\mathrm{Betti}}(1)$ is defined over $k$. ∎
+Theorem 6.5.
+Assume $g\geq 3$. Then the adelic line bundle $\widetilde{{\mathscr{L}}}$ is big, i.e. $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}})>0$.
+Proof.
+By the flatness of extension ${\mathbb{Q}}\subseteq{\mathbb{C}}$, we have $\widehat{\mathrm{vol}}(\widetilde{\mathscr{L}})=\widehat{\mathrm{vol}}(\widetilde{\mathscr{L}}_{\mathbb{C}})$.
+So $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}$ by Theorem 6.3.
+Now $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\geq 0$ since $\beta_{{\mathrm{GS}}}$ is semi-positive, and $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\not\equiv 0$ by Corollary 5.3.
+So $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}})>0$.
 ∎
+Now, we are ready to prove our main theorem.
 Proof of Theorem 1.1.
-The claim for $\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}$ is a direct consequence of Theorem 6.4 and [YZ21, Thm.
-5.3.5.(iii)].
+It suffices to prove the claim for $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}$ because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple.
+Next by [Zha10, Cor.
+2.5.2], it suffices to prove for $\xi=\xi_{s}$.
+We start by showing there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$ such that
+(6.5) $\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{{\mathrm{Fal}}}(s)-c$
+for all $s\in U(\overline{\mathbb{Q}})$.
+The conclusion follows immediately from Theorem 6.5 and [YZ21, Thm.
+5.3.5.(iii)] (which we reformulated as Theorem E.9).
 Notice that one can either take the adelic line bundle on ${\mathscr{M}}_{g}$, which defines the Faltings height (it exists by [YZ21, §$\S$2.6.2]), or one can take any ample line bundle on a suitable compactification of ${\mathscr{M}}_{g}$ and then use the comparison of the logarithmic Weil height and the Faltings height.
-The claim for $\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}$ follows because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple.
-∎
-6.4. Proof of Proposition 6.3
-Denote for simplicity by ${\mathscr{S}}={\mathscr{M}}_{g}$.
-By the flatness of extension ${\mathbb{Q}}\subseteq{\mathbb{C}}$, we have $\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{\mathbb{Q}})=\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{\mathbb{C}})$.
-By [YZ21, Theorem 5.2.1], it suffices to construct a sequence of model Hermitian line bundles $({\mathscr{S}}_{i},\overline{\mathscr{L}}_{i})$ with limit $({\mathscr{S}},\overline{\mathscr{L}})$ so that
-$\lim_{i{\longrightarrow}\infty}\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{i})=\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})^{3g-3}.$
-Let $\overline{\mathscr{N}}$ be an ample Hermitian line bundle on some integral model ${\mathscr{S}}$ of ${\mathscr{S}}$.
-Following proof of [YZ21, Theorem 6.1.1], we have a sequence of model line bundles $({\mathscr{J}}_{i},\overline{\mathscr{P}}^{\Delta}_{i},\ell_{i})$ with the limit $({\mathscr{J}},\overline{\mathscr{P}}^{\Delta})$, so that $\overline{\mathscr{P}}^{\Delta}_{i}+4^{-i}\pi^{*}\overline{\mathscr{N}}$ is nef.
-For any $n\in{\mathbb{N}}$, we have an action of ${\mathscr{J}}[n]$ on ${\mathscr{J}}$ by translating torsion points:
-$m_{n}\colon{\mathscr{J}}[n]\times{\mathscr{J}}\overset{(m,p)}{{\longrightarrow}}{\mathscr{J}}\times{\mathscr{J}},\qquad(t,x)\mapsto(x+t,x)$
-By the Theorem of square, ${\mathrm{N}}_{p}m_{n}^{*}{\mathscr{P}}^{\Delta}=({\mathscr{P}}^{\Delta})^{n^{2g}}$. Thus, we obtain an adelic metrized line bundle
-$\overline{\mathscr{P}}^{\Delta}_{i,n}:=n^{-2g}{\mathrm{N}}_{p}m_{n}^{*}\overline{\mathscr{P}}^{\Delta}_{i}.$
-This bundle is again realized on some model ${\mathscr{J}}_{i,n}$ of ${\mathscr{J}}$ with connection morphism $\ell_{i,n}:{\mathscr{P}}^{\Delta}{\longrightarrow}{\mathscr{P}}^{\Delta}_{i,n,{\mathscr{J}}}$.
+We can do better to get an explicit $U$.
+Suppose $U_{1},U_{2}$ are open subsets of ${\mathscr{M}}_{g}$ over ${\mathbb{Q}}$ such that (6.5) holds true for all $s\in U_{1}(\overline{\mathbb{Q}})$ and $s\in U_{2}(\overline{\mathbb{Q}})$, then (6.5) holds true for all $s\in(U_{1}\cup U_{2})(\overline{\mathbb{Q}})$.
+Therefore, there exists a maximal Zariski open dense subset $U^{\prime}$ of ${\mathscr{M}}_{g}$ on which (6.5) holds.
+Let $S:={\mathscr{M}}_{g}\setminus U^{\prime}$.
+Then $S$ is proper Zariski closed in ${\mathscr{M}}_{g}$, and the height inequality (6.5) cannot hold for any non-empty open subvariety $V$ of $S$.
+By Theorem E.9, $\widetilde{{\mathscr{L}}}$ cannot be big on any irreducible component $Z$ of $S$, and hence $c_{1}(\overline{\mathscr{L}}|_{Z({\mathbb{C}})})^{\wedge\dim Z}\equiv 0$ by Proposition 6.3.
+Thus $Z_{{\mathbb{C}}}=Z_{{\mathscr{F}}}(1)\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+So $S_{{\mathbb{C}}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+So $S\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ by Corollary 6.4.
+Thus we can take ${\mathscr{M}}_{g}^{\mathrm{amp}}$ to be ${\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, which is Zariski open dense and is defined over ${\mathbb{Q}}$, in Theorem 1.1.∎
+6.4. Proof of Theorem 6.3
+The second equality is immediate by Proposition 6.2.
+We will prove the first equality.
+We start to work on ${\mathscr{M}}_{g}$.
+To ease notation, denote by ${\mathscr{M}}={\mathscr{M}}_{g}$ and ${\mathscr{J}}={\mathscr{J}}_{g}$.
+By [YZ21, Theorem 5.2.1], it suffices to construct a sequence of model (Hermitian) line bundles $({\mathscr{M}}_{i},\overline{\mathscr{L}}_{i})$ which converges to $({\mathscr{M}},\overline{\mathscr{L}})$ (see the beginning of §$\S$§E.1 for notation) such that
+(6.6) $\lim_{i\rightarrow\infty}\mathrm{vol}({\mathscr{L}}_{i,{\mathbb{C}}}|_{S_{i}})=\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})^{\wedge\dim S},$
+where $S_{i}$ is the Zariski closure of $S\subseteq{\mathscr{M}}\subseteq{\mathscr{M}}_{i,{\mathbb{C}}}$.
+We start the construction. First, the proof of [YZ21, Thm. 6.1.1] yields
+  • a sequence of projective integral models $\pi_{i}\colon{\mathscr{J}}_{i}\rightarrow{\mathscr{M}}$ of ${\mathscr{J}}\rightarrow{\mathscr{M}}$,
+  • an ample Hermitian line bundle $\overline{\mathscr{N}}$ on ${\mathscr{M}}$,
+  • a sequence of model line bundles $({\mathscr{J}}_{i},\overline{\mathscr{P}}^{\Delta}_{i})$ converging to $({\mathscr{J}},\overline{\mathscr{P}}^{\Delta})$,
+such that $\overline{\mathscr{P}}^{\Delta}_{i}+4^{-i}\pi_{i}^{*}\overline{\mathscr{N}}$ is nef.
+Denote by $\ell_{i}$ the connection morphism ${\mathscr{P}}^{\Delta}\xrightarrow{\sim}{\mathscr{P}}^{\Delta}_{i,{\mathscr{J}}}:={\mathscr{P}}^{\Delta}_{i}\times_{{\mathscr{J}}_{i}}{\mathscr{J}}$.
+For each $n\in{\mathbb{N}}$, we have an action of ${\mathscr{J}}[n]$ on ${\mathscr{J}}$ by translating torsion points:
+$m_{n}\colon{\mathscr{J}}[n]\times_{S}{\mathscr{J}}{\longrightarrow}{\mathscr{J}},\qquad(t,x)\mapsto x+t.$
+By Theorem of the square, the norm line bundle ${\mathrm{N}}_{p}m_{n}^{*}{\mathscr{P}}^{\Delta}$ is $({\mathscr{P}}^{\Delta})^{\otimes n^{2g}}$, where $p$ is the natural projection ${\mathscr{J}}[n]\times_{S}{\mathscr{J}}\rightarrow{\mathscr{J}}$.
+Thus, we obtain an adelic line bundle
+$\overline{\mathscr{P}}^{\Delta}_{i,n}:=n^{-2g}{\mathrm{N}}_{p}m_{n}^{*}\overline{\mathscr{P}}^{\Delta}_{i}$
+which is realized on some model ${\mathscr{J}}_{i,n}$ of ${\mathscr{J}}$ with connection morphism $\ell_{i,n}:{\mathscr{P}}^{\Delta}\xrightarrow{\sim}{\mathscr{P}}^{\Delta}_{i,n,{\mathscr{J}}}$.
 Moreover ${\mathrm{div}}(\ell_{i,n})$ is in fact bounded by ${\mathrm{div}}(\ell_{i})$.
@@
 It follows that these forms converge to $c_{1}(\overline{\mathscr{P}}^{\Delta})$ uniformly in any compact subset of ${\mathscr{J}}({\mathbb{C}})$.
-These bundles also induce a double sequence of model line bundles $({\mathscr{S}}_{i},\overline{\mathscr{L}}_{i,n})$ of $({\mathscr{S}},\overline{\mathscr{L}})$ so that they convergent to $({\mathscr{S}},\overline{\mathscr{L}})$ as $i{\longrightarrow}\infty$, and that the metric $c_{1}(\overline{\mathscr{L}}_{i,n,{\mathbb{C}}})$ uniformly convergent to $c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})$ as $n{\longrightarrow}\infty$.
-More precisely, we let $\Omega_{i}$ be an increasing sequence of relatively compact open subsets of ${\mathscr{S}}({\mathbb{C}})$ so that ${\mathscr{S}}({\mathbb{C}})=\cup\Omega_{i}$ and $\epsilon_{i}$ be a decreasing sequence of positive numbers convergent to 00 so that on $\Omega_{i}$,
-$c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})\leq\epsilon_{i}^{-1}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
-Then for each $i$, we choose $n_{i}$ so that
+These bundles also induce a double sequence of model line bundles $({\mathscr{M}}_{i},\overline{\mathscr{L}}_{i,n})$ of $({\mathscr{M}},\overline{\mathscr{L}})$ so that they convergent to $({\mathscr{M}},\overline{\mathscr{L}})$ as $i\rightarrow\infty$, and that the metric $c_{1}(\overline{\mathscr{L}}_{i,n,{\mathbb{C}}})$ convergent uniformly to $c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})$ as $n\rightarrow\infty$.
+More precisely, we let $\Omega_{i}$ be an increasing sequence of relatively compact open subsets of ${\mathscr{M}}({\mathbb{C}})$ with ${\mathscr{M}}({\mathbb{C}})=\cup\Omega_{i}$ and $\epsilon_{i}$ be a decreasing sequence of positive numbers convergent to 00 such that
+(6.7) $c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})\leq\epsilon_{i}^{-1}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\quad\text{ on }\Omega_{i}.$
+Then for each $i$, choose $n_{i}$ such that
 (6.8) $-\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\leq c_{1}(\overline{\mathscr{L}}_{i,n_{i},{\mathbb{C}}})-c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})\leq\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})$
-as hermitian forms on the tangent bundle on $\Omega_{i}$.
-We simply write ¯$\overline{\mathscr{L}}_{i}=\overline{\mathscr{L}}_{i,n_{i}}$.
+as Hermitian forms on the tangent bundle of $\Omega_{i}$.
+For each $i\geq 1$, set ¯$\overline{\mathscr{L}}_{i}:=\overline{\mathscr{L}}_{i,n_{i}}$.
+We will show that the sequence $({\mathscr{M}}_{i},\overline{\mathscr{L}}_{i})$ is the desired sequence of model line bundles such that (6.6) holds.
 Using the reference curvature $c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})$, we may talk about eigenvalues of $c_{1}(\overline{\mathscr{L}}_{{\mathbb{C}}})$ and $c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})$.
-Now we apply Demailly’s Morse inequality [Dem91] for the bundle ¯$\overline{\mathscr{L}}_{i}$ on ${\mathscr{S}}_{i}({\mathbb{C}})$.
-For each $q\in{\mathbb{N}}$, let ${\mathscr{S}}_{i,q}$ denote the subset of ${\mathscr{S}}_{i}({\mathbb{C}})$ of points where $c_{1}(\overline{\mathscr{L}}_{i})$ has $q$-negative eigenvalues and $n-q$ positive eigenvalues.
-Then by [Dem91, (1.3), (1.5)], we have the following estimate as $k{\longrightarrow}\infty$:
-$h^{q}(k{\mathscr{L}}_{i})\leq\frac{k^{d}}{d!}\left|\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}\right|+o(k^{d}),\qquad d:=3g-3$
-$\sum_{q}(-1)^{q}h^{q}(k{\mathscr{L}}_{i})=\frac{k^{d}}{d!}\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}+o(k^{d}).$
+Recall that $S_{i}$ is the closure of $S\subseteq{\mathscr{M}}\subseteq{\mathscr{M}}_{i,{\mathbb{C}}}$ for each $i$.
+From now on, to ease notation we use ${\mathscr{L}}_{i,{\mathbb{C}}}$ to denote ${\mathscr{L}}_{i,{\mathbb{C}}}|_{S_{i}}$.
+Let us apply Demailly’s Morse inequality [Dem91] to the bundle ¯$\overline{\mathscr{L}}_{i,{\mathbb{C}}}$ on $S_{i}({\mathbb{C}})$.
+For each $q\in{\mathbb{N}}$, let $S_{i,q}$ denote the subset of $S_{i}({\mathbb{C}})$ of points where $c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})$ has $q$ negative eigenvalues and $n-q$ positive eigenvalues.
+Then by [Dem91, (1.3), (1.5)], we have the following estimates as $k\rightarrow\infty$, with $d:=\dim S$
+$h^{q}(k{\mathscr{L}}_{i,{\mathbb{C}}})\leq\frac{k^{d}}{d!}\left|\int_{S_{i,q}}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}\right|+o(k^{d}),$
+$\sum_{q}(-1)^{q}h^{q}(k{\mathscr{L}}_{i,{\mathbb{C}}})=\frac{k^{d}}{d!}\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}+o(k^{d}).$
+It follows, combined with the definition of $\mathrm{vol}({\mathscr{L}}_{i,{\mathbb{C}}})$, that
+$\left|\mathrm{vol}({\mathscr{L}}_{i,{\mathbb{C}}})-\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}\right|\leq\sum_{q>0}\left|\int_{S_{i,q}}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}\right|.$
+By [YZ21, Thm. 5.4.4],
+$\lim_{i\rightarrow\infty}\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}=\int_{S({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{{\mathbb{C}}})^{\wedge d}.$
+Thus to prove (6.6), it remains to prove
+(6.9) $\lim_{i\rightarrow\infty}\int_{S_{i,q}}c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}=0\quad\text{ for each }q>0.$
+We will prove (6.9) on $\Omega_{i}\cap S_{i,q}$ and on its complement $S_{i,q}\setminus\Omega_{i}$ respectively.
+On $\Omega_{i}\cap S_{i,q}$, by (6.7) and (6.8) we have
+$-\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\leq c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})\leq(\epsilon_{i}^{-1}+\epsilon_{i}^{d})c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
+Thus on $\Omega_{i}\cap S_{i,q}$, $c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})$ has all eigenvalue $\leq\epsilon_{i}^{-1}+\epsilon_{i}^{d}$ and one negative eigenvalue with absolute value bounded by $\leq\epsilon_{i}^{d}$.
+It follows that $|c_{1}({\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}|$ is bounded by
+$\epsilon_{i}^{d}(\epsilon_{i}^{-1}+\epsilon_{i}^{d})^{d-1}c_{1}(\overline{\mathscr{N}}_{{\mathbb{C}}})^{\wedge d}=\epsilon_{i}(1+\epsilon_{i}^{d+1})^{d-1}c_{1}(\overline{\mathscr{N}}_{{\mathbb{C}}})^{d}.$
 It follows that
-$\left|\widehat{\mathrm{vol}}({\mathscr{L}}_{i})-\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}\right|\leq\sum_{q>0}\left|\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}\right|.$
-By [YZ21, Theorem 5.4.4],
-$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}=\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}})^{d}.$
-Thus, it remains to the following estimate for each $q>0$:
-$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}=0.$
-We want to prove this on $\Omega_{i}\cap{\mathscr{S}}_{i,q}$ and on its complement ${\mathscr{S}}_{i,q}\setminus\Omega_{i}$ respectively.
-By (6.4), we have
-$-\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\leq c_{1}(\overline{\mathscr{L}}_{i})\leq(\epsilon^{-1}+\epsilon^{d})c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
-Thus on $\Omega_{i}\cap{\mathscr{S}}_{i,q}$, $c_{1}(\overline{\mathscr{L}}_{i})$ has all eigenvalue $\leq\epsilon^{-1}+\epsilon^{d}$ and one negative eigenvalue with absolute value bounded by $\leq\epsilon^{d}$.
-It follows that $|c_{1}({\mathscr{L}}_{i})^{d}|$ is bounded by
-$\epsilon^{d}(\epsilon^{-1}+\epsilon^{d})^{d-1}c_{1}(\overline{\mathscr{N}})^{d}=\epsilon_{i}(1+\epsilon_{i}^{d+1})^{d-1}c_{1}(\overline{\mathscr{N}})^{d}.$
-It follows that
-$\int_{{\mathscr{S}}_{i,q}\cap\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{d}|=O(\epsilon_{i}).$
-It remains to treat the integral over ${\mathscr{S}}_{i,q}\setminus\Omega_{i}$. Using decomposition
+(6.10) $\int_{S_{i,q}\cap\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{\wedge d}|=O(\epsilon_{i}).$
+On $S_{i,q}\setminus\Omega_{i}$, using
 ${\mathscr{P}}=\frac{1}{2}(m^{*}{\mathscr{P}}^{\Delta}-p_{1}^{*}{\mathscr{P}}^{\Delta}-p_{2}^{*}{\mathscr{P}}^{\Delta}),$
@@
 Then we have
-$\int_{{\mathscr{S}}_{i,q}\setminus\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{d}|\leq\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}.$
-Now let $\Omega_{i}^{\prime}\subset\Omega_{i}$ be another increasing sequence of relatively compact open subsets so that $\cup\Omega_{i}^{\prime}={\mathscr{S}}({\mathbb{C}})$.
-Then we can construct an increasing sequence of continuous functions $f_{i}$ so that $f_{i}(x)=1$ on ${\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}$ and $f_{i}=0$ on $\Omega_{i}^{\prime}$.
+(6.11) $\int_{S_{i,q}\setminus\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}|\leq\int_{S({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}.$
+Now let $\Omega_{i}^{\prime}\subseteq\Omega_{i}$ be another increasing sequence of relatively compact open subsets so that $\cup\Omega_{i}^{\prime}=S({\mathbb{C}})$.
+Then we can construct an increasing sequence of continuous functions $f_{i}$ so that $f_{i}\equiv 1$ on $S({\mathbb{C}})\setminus\Omega_{i}$ and $f_{i}\equiv 0$ on $\Omega_{i}^{\prime}$.
 Then for any $i\geq j$ we have
-$\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{i}\cdot c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}.$
+$\int_{S({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}\leq\int_{S({\mathbb{C}})}f_{i}\cdot c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}\leq\int_{S({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}.$
 Fix $j$ and take $i{\longrightarrow}\infty$, we get
-$\limsup_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{j}^{\prime}}c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{d}.$
-Let $j{\longrightarrow}\infty$. We get
-$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}=0.$
-We are done. ∎
+$\limsup_{i\rightarrow\infty}\int_{S({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}\leq\int_{S({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{\wedge d}\leq\int_{S({\mathbb{C}})\setminus\Omega_{j}^{\prime}}c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{\wedge d}.$
+Letting $j\rightarrow\infty$, we get
+(6.12) $\lim_{i\rightarrow\infty}\int_{S({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}=0.$
+Now (6.9) follows immediately from (6.10), (6.11) and (6.12). We are done. ∎
 Appendix A Variation of mixed Hodge structures
@@
 $00{\longrightarrow}M_{-1}{\longrightarrow}M{\longrightarrow}M_{0}{\longrightarrow}$ 0 .
-It is clear that $h({\mathbb{S}})$ stabilizes this exact sequence and induced identity on $N$.
+It is clear that $h({\mathbb{S}})$ stabilizes this exact sequence and induced identity on $M_{0}$.
 Let $\mathbf{G}_{0}$ denote the Mumford–Tate group of $M_{-1}$.
 Then, we have an exact sequence of reductive groups:
-$0{\longrightarrow}V{\longrightarrow}\mathbf{G}{\longrightarrow}\mathbf{G}_{0}{\longrightarrow}0,$
-where $V$ is a vector group included into ${\mathrm{Hom}}(M_{0},M_{-1})$. If $M_{-1}$ is polarized, then $\mathbf{G}_{0}$ is reductive by Deligne. Thus, $V$ is the unipotent radical of $\mathbf{G}$.
-The base change to ${\mathbb{R}}$ has a natural splitting $M_{\mathbb{R}}=M_{-1,{\mathbb{R}}}\oplus M_{0,{\mathbb{R}}}$ of Hodge structures given by the inverse of the isomorphism $M_{\mathbb{R}}\cap F^{0}M_{\mathbb{C}}{\overset{\sim}{{\longrightarrow}}}M_{0,{\mathbb{R}}}$.
+$0{\longrightarrow}\mathbf{G}_{-1}{\longrightarrow}\mathbf{G}{\longrightarrow}\mathbf{G}_{0}{\longrightarrow}0,$
+where $\mathbf{G}_{-1}$ is a vector group included into ${\mathrm{Hom}}(M_{0},M_{-1})$.
+If $M_{-1}$ is polarized, then $\mathbf{G}_{0}$ is reductive by Deligne.
+Thus, $\mathbf{G}_{-1}$ is the unipotent radical of $\mathbf{G}$.
+The base change to ${\mathbb{R}}$ has a natural splitting $M_{\mathbb{R}}=V_{{\mathbb{R}}}\oplus M_{0,{\mathbb{R}}}$ of Hodge structures given by the inverse of the isomorphism $M_{\mathbb{R}}\cap F^{0}M_{\mathbb{C}}{\overset{\sim}{{\longrightarrow}}}M_{0,{\mathbb{R}}}$.
 This induces splittings
-$\mathbf{G}_{\mathbb{R}}=V_{\mathbb{R}}\rtimes\mathbf{G}_{0,{\mathbb{R}}}.$
+$\mathbf{G}_{\mathbb{R}}=\mathbf{G}_{-1,{\mathbb{R}}}\rtimes\mathbf{G}_{0,{\mathbb{R}}}.$
 Moreover $h({\mathbb{S}})$ is included into $G_{0,{\mathbb{R}}}$. Thus, we have proved the following:
@@
   2. (ii)
-    The unipotent radical $V$ of $\mathbf{G}$ is a vector group.
+    The unipotent radical of $\mathbf{G}$ is a vector group.
 A.3. Variation of mixed Hodge structures and admissibility
@@
   1. (1)
-    Any VMHS arising from geometry is admissible [EZ86] (see also [BZ14]).
+    Any VMHS arising from geometry is admissible ([SZ85] and [Kas86]).
   2. (2)
@@
 B.1. Classifying space
-Let $M_{-1}$ be a finite dimensional ${\mathbb{Q}}$-vector space. Let $M=M_{-1}\oplus{\mathbb{Q}}$.
+Let $V$ be a finite-dimensional ${\mathbb{Q}}$-vector space. Let $E=V\oplus{\mathbb{Q}}$.
 B.1.1. Pure Hodge structures
-Consider the polarized Hodge data on $M_{-1}$: a non-degenerate skew pairing $Q_{-1}\colon M_{-1}\otimes M_{-1}\rightarrow{\mathbb{Q}}(1)$, and a partition $\{h^{p,q}_{M_{-1}}\}_{p,q\in{\mathbb{Z}}}$ of $\dim M_{-1,{\mathbb{C}}}$ into non-negative integers with $p+q=-1$ such that $h^{p,q}_{M_{-1}}=h^{q,p}_{M_{-1}}$.
-Then there exists a classifying space ${\mathscr{M}}_{0}$ parametrizing ${\mathbb{Q}}$-Hodge structures on $M_{-1}$ of weight $-1$ with a polarization by $Q_{-1}$ such that the $(p,q)$-constituent of $M_{-1,{\mathbb{C}}}$ has complex dimension $h^{p,q}$.
-Moreover, the ${\mathbb{Q}}$-group $\mathbf{G}_{0}^{{\mathscr{M}}}:=\mathrm{Aut}(M_{-1},Q_{-1})$, the associated real Lie group $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ acts transitively on ${\mathscr{M}}_{0}$, i.e.
+Consider the polarized Hodge data on $V$: a non-degenerate skew pairing $Q_{-1}\colon V\otimes V\rightarrow{\mathbb{Q}}(1)$, and a partition $\{h^{p,q}_{V}\}_{p,q\in{\mathbb{Z}}}$ of $\dim V_{{\mathbb{C}}}$ into non-negative integers with $p+q=-1$ such that $h^{p,q}_{V}=h^{q,p}_{V}$.
+Then there exists a classifying space ${\mathscr{M}}_{0}$ parametrizing ${\mathbb{Q}}$-Hodge structures on $V$ of weight $-1$ with a polarization by $Q_{-1}$ such that the $(p,q)$-constituent of $V_{{\mathbb{C}}}$ has complex dimension $h^{p,q}$.
+Moreover, the ${\mathbb{Q}}$-group $\mathbf{G}_{0}^{{\mathscr{M}}}:=\mathrm{Aut}(V,Q_{-1})$, the associated real Lie group $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ acts transitively on ${\mathscr{M}}_{0}$, i.e.
 ${\mathscr{M}}_{0}=\mathbf{G}_{0}({\mathbb{R}})^{+}x_{0}$
@@
 For each $x_{0}\in{\mathscr{M}}_{0}$, we have a Hodge decomposition and a Hodge filtration $F^{\bullet}_{x_{0}}$
-(B.1) $M_{-1,{\mathbb{C}}}=\bigoplus\nolimits_{p+q=-1}(M_{-1,x_{0}})^{p,q},\qquad F^{p}_{x_{0}}M_{-1,{\mathbb{C}}}=\bigoplus\nolimits_{p^{\prime}\geq p}(M_{-1,x_{0}})^{p^{\prime},q^{\prime}}$
-with $(M_{-1,x_{0}})^{q,p}=\overline{(M_{-1,x_{0}})^{p,q}}$. The inclusion ${\mathscr{M}}_{0}\subseteq{\mathscr{M}}_{0}^{\vee}$ is given by $x_{0}\mapsto F^{\bullet}_{x_{0}}$.
-For the Deligne torus ${\mathbb{S}}=\mathrm{Res}_{{\mathbb{C}}/{\mathbb{R}}}{\mathbb{G}}_{\mathrm{m},{\mathbb{C}}}$, the bi-grading decomposition above defines a morphism $h_{x_{0}}\colon{\mathbb{S}}\rightarrow\mathrm{GL}(M_{-1,{\mathbb{R}}})$, with $(M_{-1,x_{0}})^{p,q}$ the eigenspace of the character $z\mapsto z^{-p}\bar{z}^{-q}$ of ${\mathbb{S}}$.
+(B.1) $V_{{\mathbb{C}}}=\bigoplus\nolimits_{p+q=-1}(V_{x_{0}})^{p,q},\qquad F^{p}_{x_{0}}V_{{\mathbb{C}}}=\bigoplus\nolimits_{p^{\prime}\geq p}(V_{x_{0}})^{p^{\prime},q^{\prime}}$
+with $(V_{x_{0}})^{q,p}=\overline{(V_{x_{0}})^{p,q}}$. The inclusion ${\mathscr{M}}_{0}\subseteq{\mathscr{M}}_{0}^{\vee}$ is given by $x_{0}\mapsto F^{\bullet}_{x_{0}}$.
+For the Deligne torus ${\mathbb{S}}=\mathrm{Res}_{{\mathbb{C}}/{\mathbb{R}}}{\mathbb{G}}_{\mathrm{m},{\mathbb{C}}}$, the bi-grading decomposition above defines a morphism $h_{x_{0}}\colon{\mathbb{S}}\rightarrow\mathrm{GL}(V_{{\mathbb{R}}})$, with $(V_{x_{0}})^{p,q}$ the eigenspace of the character $z\mapsto z^{-p}\bar{z}^{-q}$ of ${\mathbb{S}}$.
 It is known that $h_{x_{0}}({\mathbb{S}})<\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})$ for all $x_{0}\in{\mathscr{M}}_{0}$.
@@
 So we will view ${\mathscr{M}}_{0}$ as a subset of ${\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}})$.
-We close this subsubsection with the following remark on the Mumford–Tate group $\mathrm{MT}_{x_{0}}$ of the pure Hodge structure on $M_{-1}$ determined by $x_{0}$.
+We close this subsubsection with the following remark on the Mumford–Tate group $\mathrm{MT}_{x_{0}}$ of the pure Hodge structure on $V$ determined by $x_{0}$.
 We have that $\mathrm{MT}_{x_{0}}$ is a subgroup of $\mathbf{G}_{0}^{{\mathscr{M}}}$ for all $x_{0}\in{\mathscr{M}}_{0}$ and equals $\mathbf{G}_{0}^{{\mathscr{M}}}$ for some $x_{0}\in{\mathscr{M}}_{0}$.
+Moreover, ${\mathrm{MT}}_{x_{0}}$ is a reductive group for each $x_{0}\in{\mathscr{M}}_{0}$.
 B.1.2. Weight $-1$ and 00
 Next, we turn to mixed Hodge structures of weight of $-1$ and 00.
-Fix the following data on $M=M_{-1}\oplus{\mathbb{Q}}$: the weight filtration $W_{\bullet}:=(0=W_{-2}M\subseteq W_{-1}M=M_{-1}\subseteq W_{0}M=M)$; the partition $\{h^{p,q}\}_{p,q\in{\mathbb{Z}}}$ of $\dim M_{{\mathbb{C}}}$ into non-negative integers, with $h^{p,q}=h^{p,q}_{M_{-1}}$ for $p+q=-1$ and $h^{0,0}=1$ and $h^{p,q}=0$ otherwise.
-There exists the classifying space ${\mathscr{M}}$ parametrizing ${\mathbb{Q}}$-mixed Hodge structures $(M,W_{\bullet},F^{\bullet})$ of weight $-1$ and 00 such that:
+Fix the following data on $E=V\oplus{\mathbb{Q}}$: the weight filtration
+$W_{\bullet}:=(0=W_{-2}E\subseteq W_{-1}E=V\subseteq W_{0}E=E);$
+the partition $\{h^{p,q}\}_{p,q\in{\mathbb{Z}}}$ of $\dim E_{{\mathbb{C}}}$ into non-negative integers, with $h^{p,q}=h^{p,q}_{V}$ for $p+q=-1$ and $h^{0,0}=1$ and $h^{p,q}=0$ otherwise.
+There exists the classifying space ${\mathscr{M}}$ parametrizing ${\mathbb{Q}}$-mixed Hodge structures $(E,W_{\bullet},F^{\bullet})$ of weight $-1$ and 00 such that:
   1. (a)
-    the $(p,q)$-constituent $\mathrm{Gr}^{p}_{F}\mathrm{Gr}^{W}_{p+q}M_{\mathbb{C}}$ has complex dimension $h^{p,q}$;
+    the $(p,q)$-constituent $\mathrm{Gr}^{p}_{F}\mathrm{Gr}^{W}_{p+q}E_{\mathbb{C}}$ has complex dimension $h^{p,q}$;
   2. (b)
-    $\mathrm{Gr}^{W}_{-1}M=M_{-1}$ is polarized by $Q_{-1}$.
+    $\mathrm{Gr}^{W}_{-1}E=V$ is polarized by $Q_{-1}$.
 See for example [Pea00, below (3.7) to the Remark below Lem.
 3.9].
-Notice that $\mathrm{Gr}^{W}_{0}M={\mathbb{Q}}$ is polarized by $Q_{0}\colon{\mathbb{Q}}\otimes{\mathbb{Q}}\rightarrow{\mathbb{Q}}$ is $a\otimes b\mapsto ab$.
-So $M$ is graded-polarized.
+Notice that $\mathrm{Gr}^{W}_{0}E={\mathbb{Q}}$ is polarized by $Q_{0}\colon{\mathbb{Q}}\otimes{\mathbb{Q}}\rightarrow{\mathbb{Q}}$ is $a\otimes b\mapsto ab$.
+So $E$ is graded-polarized.
 In our case, we need a better understanding of the structure of ${\mathscr{M}}$ than [Pea00].
-The map $x\mapsto F_{x}^{\bullet}$ realizes ${\mathscr{M}}$ as a semi-algebraic open subset of a suitable flag variety ${\mathscr{M}}^{\vee}$, which is easily seen to be an orbit under $\mathbf{G}({\mathbb{C}})$ for the ${\mathbb{Q}}$-group
-(B.3) $\mathbf{G}^{{\mathscr{M}}}:=M_{-1}\rtimes\mathrm{Aut}(M_{-1},Q_{-1})=M_{-1}\rtimes\mathbf{G}_{0}^{{\mathscr{M}}}.$
+The map $x\mapsto F_{x}^{\bullet}$ realizes ${\mathscr{M}}$ as a semi-algebraic open subset of a suitable flag variety ${\mathscr{M}}^{\vee}$, which is easily seen to be an orbit under $\mathbf{G}^{{\mathscr{M}}}({\mathbb{C}})$ for the ${\mathbb{Q}}$-group
+(B.3) $\mathbf{G}^{{\mathscr{M}}}:={\mathrm{Aut}}(E,Q_{-1},W)=V\rtimes\mathrm{Aut}(V,Q_{-1})=V\rtimes\mathbf{G}_{0}^{{\mathscr{M}}}.$
 […]
@@
 (B.5) ${\mathscr{M}}\longrightarrow{\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{{\mathbb{R}}}),\quad x\mapsto h_{x}.$
-As in the pure case, for the Mumford–Tate group $\mathrm{MT}_{x}$ for the mixed Hodge structure on $M$ determined by $x$, we have that $\mathrm{MT}_{x}<\mathbf{G}^{{\mathscr{M}}}$ for all $x\in{\mathscr{M}}$, and $\mathrm{MT}_{x}=\mathbf{G}^{{\mathscr{M}}}$ for some $x\in{\mathscr{M}}$.
+As in the pure case, for the Mumford–Tate group $\mathrm{MT}_{x}$ for the mixed Hodge structure on $E$ determined by $x$, we have that $\mathrm{MT}_{x}<\mathbf{G}^{{\mathscr{M}}}$ for all $x\in{\mathscr{M}}$, and $\mathrm{MT}_{x}=\mathbf{G}^{{\mathscr{M}}}$ for some $x\in{\mathscr{M}}$.
+Unlike the pure case, $\mathrm{MT}_{x}$ is not a reductive group in general.
 B.1.3.
-For the quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}=\mathbf{G}^{{\mathscr{M}}}/M_{-1}$, consider the following surjective $\mathbf{G}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map
-(B.6) $p\colon{\mathscr{M}}\rightarrow{\mathscr{M}}_{0},\quad g\cdot x\mapsto p(g)\cdot\bar{h}_{x}$
-with $\bar{h}_{x}=p\circ h_{x}\colon{\mathbb{S}}\rightarrow\mathbf{G}^{{\mathscr{M}}}_{{\mathbb{R}}}\rightarrow\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}}$.
-This map sends the mixed Hodge structure $(M,W_{\bullet}M,F^{\bullet}M_{{\mathbb{C}}})$ to the associated pure Hodge structure on $M_{-1}=\mathrm{Gr}_{W}^{-1}M$.
+For each $x\in{\mathscr{M}}$, we have an induced Hodge structure $p(x)$ on $V$ and a unique vector $q(x)\in V_{\mathbb{R}}$ such that $(q(x),1)\in F^{0}_{x}E_{\mathbb{C}}$.
+Thus, we have a Hodge decomposition
+$E_{\mathbb{C}}=\bigoplus_{p+q=-1,0}E_{x}^{p,q},\qquad E_{x}^{p,q}:=F_{x}^{p}E_{\mathbb{C}}\cap\overline{F_{x}^{q}E_{\mathbb{C}}}.$
+The sum over $p+q=-1$ is the Hodge decomposition on $V$ and $E_{x}^{0,0}={\mathbb{C}}(q(x),1)$. This decomposition induces a bijection
+(B.6) $(q,p)\colon{\mathscr{M}}{\overset{\sim}{{\longrightarrow}}}V_{\mathbb{R}}\times{\mathscr{M}}_{0},\qquad x\mapsto(q(x),p(x))$
+where $p(x)$ is the induced Hodge structure on $V=W_{-1}E$.
+For the inverse map, for each $(v,x_{0})\in V_{\mathbb{R}}\times{\mathscr{M}}_{0}$ we take $x\in{\mathscr{M}}$ be the unique Hodge structure $h$ on $E$ extending $x_{0}$ on $V_{\mathbb{C}}$ so that $F^{0}_{x}E_{\mathbb{C}}=F^{0}_{x_{0}}V_{\mathbb{C}}+{\mathbb{C}}(v,1)$.
+The real semi-algebraic structure on ${\mathscr{M}}$ is given by the product in the above isomorphism.
+The complex structure on ${\mathscr{M}}$ is given by making ${\mathscr{M}}$ into a vector bundle over ${\mathscr{M}}_{0}$ with fibers $V_{\mathbb{R}}{\overset{\sim}{{\longrightarrow}}}V_{\mathbb{C}}/F^{0}_{x}V_{\mathbb{C}}$.
+Notice that for each $a\in V({\mathbb{R}})$, the submanifold $(q,p)^{-1}(\{a\}\times{\mathscr{M}}_{0})$ is a semi-algebraic and complex submanifold of ${\mathscr{M}}$.
+By abuse of notation, we also denote the group quotient still by $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}=\mathbf{G}^{{\mathscr{M}}}/V$.
+For each $x\in{\mathscr{M}}$, we have a surjective morphism ${\mathrm{MT}}_{x}{\longrightarrow}{\mathrm{MT}}_{p(x)}$ with kernel $V_{x}:=V\cap{\mathrm{MT}}_{x}$.
+Thus $V_{x}$ is the unipotent radical of ${\mathrm{MT}}_{x}$.
+This surjection has a section $s\colon{\mathrm{MT}}_{p(x)}{\longrightarrow}{\mathrm{MT}}_{x}$.
+Via (B.3), this section can be written as $s(g)=(\sigma(g),g)$ with $\sigma$ a cocycle in $Z^{1}({\mathrm{MT}}_{x},V)$.
+But $H^{1}({\mathrm{MT}}_{p(x)},V)=0$ since ${\mathrm{MT}}_{p(x)}$ is reductive.
+Thus $\sigma$ is a coboundary, i.e.
+$\sigma(g)=v_{0}-gv_{0}$ for some $v_{0}\in V$ unique up to addition by invariant vectors of $V$ under action by ${\mathrm{MT}}_{p(x)}$.
+Thus, we get a precise Levi decomposition
+(B.7) ${\mathrm{MT}}_{x}=V_{x}\rtimes{\mathrm{MT}}_{x_{0}}={\mathrm{Ad}}(v_{0})(V_{x}\rtimes{\mathrm{MT}}_{p(x)}),$
+where $V_{x}=V\cap{\mathrm{MT}}_{x}$, $x_{0}=(v_{0},p(x))$ is a point of ${\mathscr{M}}$ via the isomorphism (B.6).
+With this isomorphism, we see that $x=(v_{1}+v_{0},p(x))$ with $v_{1}\in V_{\mathbb{R}}$ which has Zariski closure $V_{x}$ over ${\mathbb{Q}}$.
 We will study the fibered structure given by $p$ more carefully. We will do this in the finer setting of Mumford–Tate domains.
@@
 Now, let us take a closer look at the semi-algebraic structure and a complex structure on ${\mathscr{D}}$.
-The unipotent radical $V$ of $\mathbf{G}=\mathrm{MT}({\mathscr{D}})$ equals $M_{-1}\cap\mathbf{G}$ by reason of weight.
-Let $\mathbf{G}_{0}:=\mathbf{G}/V$ be the reductive part.
+The unipotent radical of $\mathbf{G}=\mathrm{MT}({\mathscr{D}})$ equals $\mathbf{G}_{-1}:=V\cap\mathbf{G}$ by reason of weight.
+Let $\mathbf{G}_{0}:=\mathbf{G}/\mathbf{G}_{-1}$ be the reductive part.
 Set ${\mathscr{D}}_{0}:=p({\mathscr{D}})\subseteq{\mathscr{M}}_{0}$ for the map $p$ defined in (B.6).
@@
 By abuse of notation, we also use $p$ to denote the natural projections
-(B.7) $p\colon\mathbf{G}\rightarrow\mathbf{G}_{0}=\mathbf{G}/V\quad\text{ and }\quad p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}.$
-Fix a Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$. Identify $\mathbf{G}_{0}$ with $\{0\}\times\mathbf{G}_{0}$.
-Recall that each $x_{0}\in{\mathscr{M}}_{0}$ endows $M_{-1}$ with a Hodge structure of weight $-1$.
-For $x_{0}\in{\mathscr{D}}_{0}$, $V$ is a sub-Hodge structure because $V$ is a $\mathbf{G}_{0}$-submodule of $M_{-1}$ and that $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$.
-Finally, consider the constant bundle $V({\mathbb{C}})\times{\mathscr{D}}_{0}\rightarrow{\mathscr{D}}_{0}$.
-Define the holomorphic subbundle $F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})$ to be such that the fiber over each $x_{0}\in{\mathscr{D}}_{0}$ is $F^{0}_{x_{0}}V_{{\mathbb{C}}}$.
+(B.8) $p\colon\mathbf{G}\rightarrow\mathbf{G}_{0}=\mathbf{G}/\mathbf{G}_{-1}\quad\text{ and }\quad p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}.$
+Fix a Levi decomposition $\mathbf{G}=\mathbf{G}_{-1}\rtimes\mathbf{G}_{0}$, such as one given by B.7. Identify $\mathbf{G}_{0}$ with $\{0\}\times\mathbf{G}_{0}$.
+Recall that each $x_{0}\in{\mathscr{M}}_{0}$ endows $V$ with a Hodge structure of weight $-1$.
+For $x_{0}\in{\mathscr{D}}_{0}$, $\mathbf{G}_{-1}$ is a sub-Hodge structure because $\mathbf{G}_{-1}$ is a $\mathbf{G}_{0}$-submodule of $V$ and that $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$.
+Finally, consider the constant bundle $\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0}\rightarrow{\mathscr{D}}_{0}$.
+Define the holomorphic subbundle $F^{0}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})$ to be such that the fiber over each $x_{0}\in{\mathscr{D}}_{0}$ is $F^{0}_{x_{0}}\mathbf{G}_{-1,{\mathbb{C}}}$.
 Proposition B.3.
@@
     The semi-algebraic structure on ${\mathscr{D}}\subseteq{\mathrm{Hom}}({\mathbb{S}},\mathbf{G}_{{\mathbb{R}}})$ is given by
-    $V({\mathbb{R}})\times{\mathscr{D}}_{0}\xrightarrow{\sim}{\mathscr{D}},\quad(v,x_{0})\mapsto\mathrm{Int}(v)\circ h_{x_{0}}.$
+    $\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\xrightarrow{\sim}{\mathscr{D}},\quad(v,x_{0})\mapsto\mathrm{Int}(v)\circ h_{x_{0}}.$
   2. (ii)
-    The complex structure on ${\mathscr{D}}$ is given by ${\mathscr{D}}=(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})$.
+    The complex structure on ${\mathscr{D}}$ is given by ${\mathscr{D}}{\overset{\sim}{{\longrightarrow}}}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})$.
   3. (iii)
     These two structures are related by the natural bijection
-    (B.8) $V({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq V({\mathbb{C}})\times{\mathscr{D}}_{0}\longrightarrow(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0}).$
+    (B.9) $\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0}\longrightarrow(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0}).$
 Proof.
 Fix $x\in{\mathscr{D}}$ and let $h_{x}\colon{\mathbb{S}}\rightarrow\mathbf{G}_{{\mathbb{R}}}$ be the corresponding homomorphism.
-The Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ induces a bijection $\mathbf{G}({\mathbb{R}})^{+}x\cong V({\mathbb{R}})\times{\mathscr{D}}_{0}$, which is semi-algebraic.
+The Levi decomposition $\mathbf{G}=\mathbf{G}_{-1}\rtimes\mathbf{G}_{0}$ induces a bijection $\mathbf{G}({\mathbb{R}})^{+}x\simeq\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}$, which is semi-algebraic.
 This establishes (i).
-Set ${\mathscr{X}}:=\mathbf{G}({\mathbb{R}})^{+}V({\mathbb{C}})\cdot h_{x}\subseteq{\mathrm{Hom}}({\mathbb{S}}_{{\mathbb{C}}},\mathbf{G}_{{\mathbb{C}}})$, where the action is given by conjugation.
-The quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}$ induces a natural surjective map ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and by [Pin89, 1.8(a)] each fiber of this map is a $V({\mathbb{C}})$-torsor.
-The Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ induces a global section of ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and hence an isomorphism ${\mathscr{X}}\cong V({\mathbb{C}})\times{\mathscr{D}}_{0}$ over ${\mathscr{D}}_{0}$.
+Set ${\mathscr{X}}:=\mathbf{G}({\mathbb{R}})^{+}\mathbf{G}_{-1}({\mathbb{C}})\cdot h_{x}\subseteq{\mathrm{Hom}}({\mathbb{S}}_{{\mathbb{C}}},\mathbf{G}_{{\mathbb{C}}})$, where the action is given by conjugation.
+The quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}$ induces a natural surjective map ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and by [Pin89, 1.8(a)] each fiber of this map is a $\mathbf{G}_{-1}({\mathbb{C}})$-torsor.
+The Levi decomposition $\mathbf{G}=\mathbf{G}_{-1}\rtimes\mathbf{G}_{0}$ induces a global section of ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and hence an isomorphism ${\mathscr{X}}\simeq\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0}$ over ${\mathscr{D}}_{0}$.
 Consider the following surjective equivariant map
 (B.10) $\varphi\colon{\mathscr{X}}\rightarrow{\mathscr{D}},\qquad gh_{x}g^{-1}\mapsto g\cdot x.$
-By [Pin89, 1.8(b)], for each $x\in{\mathscr{D}}$, the fiber $\varphi^{-1}(x)$ is a principle homogeneous space under $F^{0}_{x_{0}}V_{{\mathbb{C}}}$ with $x_{0}=p(x)$.
-Hence $\varphi$ gives a bijection $(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})\rightarrow{\mathscr{D}}$ over ${\mathscr{D}}_{0}$.
+By [Pin89, 1.8(b)], for each $x\in{\mathscr{D}}$, the fiber $\varphi^{-1}(x)$ is a principle homogeneous space under $F^{0}_{x_{0}}\mathbf{G}_{-1,{\mathbb{C}}}$ with $x_{0}=p(x)$.
+Hence $\varphi$ gives a bijection $(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})\rightarrow{\mathscr{D}}$ over ${\mathscr{D}}_{0}$.
 Moreover, the complex structure of ${\mathscr{D}}$ is precisely given by this bijection; see [GK24, Proof of Prop.
@@
 Corollary B.4.
-Let $\tilde{Z}_{0}\subseteq{\mathscr{D}}_{0}$ be an irreducible algebraic subset.
-Then for any $a\in V({\mathbb{R}})$, the subset $\{a\}\times\tilde{Z}_{0}\subseteq V({\mathbb{R}})\times{\mathscr{D}}_{0}\cong{\mathscr{D}}$ is irreducible algebraic.
+Let $\widetilde{Z}_{0}\subseteq{\mathscr{D}}_{0}$ be an irreducible algebraic subset.
+Then for any $a\in\mathbf{G}_{-1}({\mathbb{R}})$, the subset $\{a\}\times\widetilde{Z}_{0}\subseteq\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\simeq{\mathscr{D}}$ is irreducible algebraic.
 Proof.
-$\{a\}\times\tilde{Z}_{0}$ is clearly semi-algebraic in ${\mathscr{D}}$. In view of (B.8), $\{a\}\times\tilde{Z}_{0}$ is also complex analytic in ${\mathscr{D}}$. Hence, we are done. ∎
+$\{a\}\times\widetilde{Z}_{0}$ is clearly semi-algebraic in ${\mathscr{D}}$. In view of (B.9), $\{a\}\times\widetilde{Z}_{0}$ is also complex analytic in ${\mathscr{D}}$. Hence, we are done. ∎
 Lemma B.5.
 Let ${\mathscr{D}}^{\prime}$ be a sub-Mumford–Tate domain of ${\mathscr{D}}$.
-Then under the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), we have ${\mathscr{D}}^{\prime}=(V^{\prime}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime})$ for some $v_{0}\in V({\mathbb{Q}})$.
+Then under the identification ${\mathscr{D}}=\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), we have ${\mathscr{D}}^{\prime}=(\mathbf{G}_{-1}^{\prime}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime})$ for some $v_{0}\in\mathbf{G}_{-1}({\mathbb{Q}})$.
 Proof.
-Denote by $\mathbf{G}^{\prime}=\mathrm{MT}({\mathscr{D}}^{\prime})$, $V^{\prime}:=V\cap\mathbf{G}^{\prime}$, and $\mathbf{G}^{\prime}_{0}:=\mathbf{G}^{\prime}/V^{\prime}$.
-Then, since ${\mathscr{D}}^{\prime}$ is a sub-Mumford–Tate domain of ${\mathscr{D}}$ and by reason of weight, $V^{\prime}$ is the unipotent radical of $\mathbf{G}^{\prime}$.
-Set $\mathbf{G}^{\prime\prime}:=V\mathbf{G}^{\prime}$.
-Then $\mathbf{G}^{\prime\prime}$ is a subgroup of $\mathbf{G}$, with unipotent radical $V$ and reductive part $\mathbf{G}^{\prime}_{0}$.
-We can construct two Levi decompositions of $\mathbf{G}^{\prime\prime}$ as follows.
-First, $\mathbf{G}^{\prime}_{0}=\mathbf{G}^{\prime}/V^{\prime}=\mathbf{G}^{\prime}/(V\cap\mathbf{G}^{\prime})<\mathbf{G}/V=\mathbf{G}_{0}$.
-Under the Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ fixed above Proposition B.3, we have $\{0\}\times\mathbf{G}^{\prime}_{0}<\mathbf{G}^{\prime\prime}$.
-The composite $\{0\}\times\mathbf{G}^{\prime}_{0}<\mathbf{G}^{\prime\prime}\rightarrow\mathbf{G}^{\prime\prime}/V=\mathbf{G}^{\prime}_{0}$ is the natural isomorphism (the projection to the second factor).
-Thus, we get a Levi decomposition $\mathbf{G}^{\prime\prime}=V\rtimes\mathbf{G}^{\prime}_{0}$, which is compatible with our fixed Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$.
-Next, fix a Levi decomposition $\mathbf{G}^{\prime}=V^{\prime}\rtimes^{\prime}\mathbf{G}^{\prime}_{0}$.
-It is uniquely determined by a section of $\mathbf{G}^{\prime}\rightarrow\mathbf{G}^{\prime}_{0}$, i.e.
-an injective morphism $s^{\prime}\colon\mathbf{G}^{\prime}_{0}\rightarrow\mathbf{G}^{\prime}$ such that $s^{\prime}\circ p_{\mathbf{G}^{\prime}}=1_{\mathbf{G}^{\prime}_{0}}$.
-Now abuse of notation denote by $s^{\prime}$ the composite $\mathbf{G}^{\prime}_{0}\xrightarrow{s^{\prime}}\mathbf{G}^{\prime}<\mathbf{G}^{\prime\prime}$.
-Then $s^{\prime}$ defines a Levi decomposition $\mathbf{G}^{\prime\prime}=V\rtimes^{\prime}\mathbf{G}_{0}^{\prime}$.
-By the general theory of algebraic groups, any two Levi decompositions of $\mathbf{G}^{\prime\prime}$ differ from the conjugation by some $v_{0}\in V({\mathbb{Q}})$.
-Now, the conclusion follows from Proposition B.3.(i).∎
+It suffices to apply (B.7) to both ${\mathrm{MT}}({\mathscr{D}})$ and ${\mathrm{MT}}({\mathscr{D}}^{\prime})$. ∎
 B.3. Quotient by a normal subgroup and weak Mumford–Tate domains
@@
 It is clearly true that $p_{N}$ is semi-algebraic.
-More precisely, let $M_{-1}/N$ be the maximal quotient of $M_{-1}$ on which $N$ acts trivially, and let $M/N=M_{-1}/N\oplus{\mathbb{Q}}$.
+More precisely, let $V/N$ be the maximal quotient of $V$ on which $N$ acts trivially, and let $M/N=V/N\oplus{\mathbb{Q}}$.
 For each $x\in{\mathscr{D}}$, the composition
 $\bar{h}_{x}:{\mathbb{S}}\overset{h_{x}}{{\longrightarrow}}{\mathrm{GL}}(M_{{\mathbb{R}}}){\longrightarrow}{\mathrm{GL}}((M/N)_{{\mathbb{R}}})$
-defines some Hodge structure on $M/N$ with weight $-1$ on $M_{-1}/N$ and weight 00 on the quotient ${\mathbb{Q}}$.
+defines some Hodge structure on $M/N$ with weight $-1$ on $V/N$ and weight 00 on the quotient ${\mathbb{Q}}$.
 The Hodge numbers on $M/N$ defined by $h_{x}$ do not depend on the choice of $x$.
@@
 Assume $\Gamma$ is an arithmetic subgroup of $\mathbf{G}({\mathbb{Q}})$.
-Then the quotient $\Gamma\backslash{\mathscr{D}}$ is an orbifold.
+Then, the quotient $\Gamma\backslash{\mathscr{D}}$ is an orbifold.
 Denote by $\Gamma_{/N}$ the image of $\Gamma$ under the quotient $\mathbf{G}\rightarrow\mathbf{G}/N$.
@@
 $[p_{N}]\colon\Gamma\backslash{\mathscr{D}}\rightarrow\Gamma_{/N}\backslash({\mathscr{D}}/N).$
-We will pay special attention to the fibers of $p_{N}$ (and $[p_{N}]$). More generally, we define:
+We will pay special attention to the fibers of $p_{N}$ (and $[p_{N}]$). More generally, we define
 Definition B.6.
 A subset ${\mathscr{D}}_{N}$ of ${\mathscr{D}}$ is called a weak Mumford–Tate domain if there exist $x\in{\mathscr{D}}$ and a normal subgroup $N$ of $\mathrm{MT}(x)$ such that ${\mathscr{D}}_{N}=N({\mathbb{R}})^{+}x$.
-In this definition, if $x$ is taken to be a Hodge generic point, i.e. $\mathrm{MT}(x)=\mathbf{G}$, then the weakly Mumford–Tate domain thus obtained is a fiber of $p_{N}$.
+In this definition, if $x$ is taken to be a Hodge generic point, i.e. $\mathrm{MT}(x)=\mathbf{G}$, then the weak Mumford–Tate domain thus obtained is a fiber of $p_{N}$.
 Appendix C Intermediate Jacobians and normal functions
@@
 For each $n\geq 1$, the relative intermediate Jacobian of the (polarizable) VHS $R^{2n-1}f_{*}{\mathbb{Z}}(n)_{X}$ is called the $n$-th intermediate Jacobian of $X\rightarrow S$ and is denoted by ${\mathscr{J}}^{n}(X/S)$.
-If $S$ is a point and $X=X$, then we simply write ${\mathscr{J}}^{n}(X)$.
+If $S$ is a point, then we simply write ${\mathscr{J}}^{n}(X)$.
 For each $s\in S$ we have $H_{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{Z}})={\mathbb{V}}_{{\mathbb{Z}},s}$ as ${\mathbb{Z}}$-modules.
@@
 The dual of $H^{1,0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\subseteq H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$ gives $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})^{\vee}\rightarrow\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$.
-Thus the de Rham–Betti comparison $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\cong H^{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{C}})$ gives rise to a Hodge filtration ${\mathscr{F}}_{{\mathscr{J}}}^{\bullet}$ on ${\mathscr{V}}$ defined by ${\mathscr{F}}_{{\mathscr{J}}}^{1}=0$, ${\mathscr{F}}_{{\mathscr{J}}}^{-1}={\mathscr{V}}$ and
+Thus the de Rham–Betti comparison $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\simeq H^{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{C}})$ gives rise to a Hodge filtration ${\mathscr{F}}_{{\mathscr{J}}}^{\bullet}$ on ${\mathscr{V}}$ defined by ${\mathscr{F}}_{{\mathscr{J}}}^{1}=0$, ${\mathscr{F}}_{{\mathscr{J}}}^{-1}={\mathscr{V}}$ and
 $({\mathscr{F}}_{{\mathscr{J}}}^{0})_{s}=\ker({\mathscr{V}}_{s}=H_{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{C}})\rightarrow\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}))\quad\text{for all }s\in S.$
@@
 Proof.
-We computed in the proof of Lemma C.2 that ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{V}}_{s}/({\mathscr{F}}^{n}_{s}+{\mathbb{V}}_{{\mathbb{Z}},s})$.
+We computed in the proof of Lemma C.2 that ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathbb{V}}_{{\mathbb{Z}},s}\backslash{\mathscr{V}}_{s}/{\mathscr{F}}^{n}_{s}$.
 Hence $\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})={\mathscr{V}}_{s}/{\mathscr{F}}^{n}_{s}$.
@@
 Write $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ for the natural projection and $d$ for the relative dimension.
-The holomorphic vector bundle ${\mathscr{V}}$ is endowed with a connection $\nabla$ whose sheaf of the horizontal sections is ${\mathbb{V}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}}{\mathbb{C}}$.
 For each open subset $\Delta\subseteq S$, We have the following exact sequence.
@@
 Notice that $({\mathbb{R}}^{2d}\times\Delta)\cap{\mathscr{F}}_{{\mathscr{J}}}^{0}{\mathscr{V}}_{\Delta}$ is 00 on each fiber by weight reasons.
-Hence ${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta\rightarrow\mathrm{Lie}(\pi^{-1}(\Delta)/\Delta)$ is a real analytic diffeomorphism, and over each $s\in\Delta$ it becomes a group homomorphism.
+Hence
+${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta\rightarrow\mathrm{Lie}(\pi^{-1}(\Delta)/\Delta)$
+is a real analytic diffeomorphism, and over each $s\in\Delta$ it becomes a group homomorphism.
 Moreover, the image of $\{r\}\times\Delta$ is easily seen to be complex analytic for any $r\in{\mathbb{R}}^{2d}$.
@@
 Each leaf is holomorphic to the discussion above.
-In fact, the Betti foliation is the unique foliation on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ which is everywhere transverse to the fibers of $\pi$ and whose set of leaves contains all torsion multisections.
-For each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, the Betti foliation induces a decomposition $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$, and $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ induces a natural isomorphism $T_{x}{\mathscr{F}}_{\mathrm{Betti}}\cong T_{\pi(x)}S$.
+In fact, the Betti foliation is the unique foliation on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, which is everywhere transverse to the fibers of $\pi$ and whose set of leaves contains all torsion multisections.
+For each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, the Betti foliation induces a decomposition
+$T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)},$
+and $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ induces a natural isomorphism $T_{x}{\mathscr{F}}_{\mathrm{Betti}}\simeq T_{\pi(x)}S$.
 The translation on the torus ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$ yields a canonical isomorphism $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})=T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})$.
@@
 Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers of dimension $d$.
-Let $Z$ be a family of homologically trivial $n$-cocycle in $X/S$, i.e.
+Let $Z$ be a family of homologically trivial $n$-cocycles in $X/S$, i.e.
 a formal sum of integral subschemes of $X$ which are flat and dominant over $S$ such that each fiber $Z_{s}$ is a homologically trivial $n$-cocycle of $X_{s}$.
@@
 The holomorphic section
-$\nu_{Z}\colon S\rightarrow{\mathscr{J}}^{n}(X/S),\qquad s\mapsto{\mathrm{AJ}}(Z_{s})$
-It is an admissible normal function.
+(C.5) $\nu_{Z}\colon S\rightarrow{\mathscr{J}}^{n}(X/S),\qquad s\mapsto{\mathrm{AJ}}(Z_{s})$
+is an admissible normal function.
 The following proposition is a simple application of the Abel–Jacobi map and the Betti foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on intermediate Jacobians introduced in §$\S$§C.2.
+Recall the definition $S^{\mathrm{amp}}:=S\setminus S^{\mathrm{Betti}}(1)$.
 Proposition C.7.
-Set $S^{\circ}:=S\setminus S_{{\mathscr{F}}}(1)$.
-Assume $X/S$ and $Z$ are defined over $\overline{\mathbb{Q}}$. Then $[Z_{s}]\in\mathrm{CH}^{n}(X_{s})$ is non-torsion for any $s\in S^{\circ}({\mathbb{C}})\setminus S^{\circ}(\overline{\mathbb{Q}})$.
+Assume $X/S$ and $Z$ are defined over $\overline{\mathbb{Q}}$.
+Then $[Z_{s}]\in\mathrm{Ch}^{n}(X_{s})$ is non-torsion for any $s\in S^{\mathrm{amp}}({\mathbb{C}})\setminus S^{\mathrm{amp}}(\overline{\mathbb{Q}})$.
 Proof.
@@
 So $\nu_{Z}(\bar{s})\subseteq{\mathscr{F}}_{\mathrm{Betti}}$.
-If $s\not\in S(\overline{\mathbb{Q}})$, then $\dim\bar{s}\geq 1$, and hence $\bar{s}\subseteq S_{{\mathscr{F}}}(1)$ by definition.
+If $s\not\in S(\overline{\mathbb{Q}})$, then $\dim\bar{s}\geq 1$, and hence $\bar{s}\subseteq S^{\mathrm{Betti}}(1)$ by definition.
 Now we are done.
@@
 C.6. Normal functions associated with Gross–Schoen cycles and Ceresa cycles
-Let $C$ be an irreducible smooth projective curve defined over a field $k$ of genus $g\geq 3$. Let $\xi\in C(k)$. We have:
-    For each subset $T\subseteq\{1,2,3\}$, the modified diagonal $\Delta_{T}(C)=\{(x_{1},x_{2},x_{3}):x_{i}=\xi\text{ for }i\not\in T,~{}x_{j}=x_{j^{\prime}}\text{ for all }j,j^{\prime}\in T\}\in\mathrm{CH}^{2}(C^{3})$.
-    The (classical) Abel–Jacobi map $i_{\xi}\colon C\rightarrow\mathrm{Jac}(C)$ sending $x\mapsto[x-\xi]$.
-Both $\Delta_{T}(C)$ and $i_{\xi}$ can be extended to any $\xi=\sum_{i}n_{i}e_{i}\in\mathrm{Div}^{1}(C)(k)=\mathrm{Pic}^{1}(C)(k)$.
-This is classical and direct for $i_{\xi}$, and for $\Delta_{T}(C)$ one can define $\Delta_{123}(C)=\{(c,c,c):c\in C\}$, $\Delta_{12}(C)=\sum n_{i}\{(c,c,e_{i}):c\in C\}$, $\Delta_{1}(C)=\sum_{i,j}n_{i}n_{j}\{(c,e_{i},e_{j}):c\in C\}$, etc.
-For any $\xi\in\mathrm{Pic}^{1}(C)$, define the Gross–Schoen cycle based at $\xi$ (resp. the Ceresa cycle based at $\xi$) to be:
-  • (Gross–Schoen cycle based at $\xi$) $\Delta_{{\mathrm{GS}},\xi}(C):=\Delta_{123}(C)-\Delta_{12}(C)-\Delta_{13}(C)-\Delta_{23}(C)+\Delta_{1}(C)+\Delta_{2}(C)+\Delta_{3}(C)\in\mathrm{CH}^{2}(C^{3})$.
-  • (Ceresa cycle based at $\xi$) ${\mathrm{Ce}}_{\xi}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in\mathrm{CH}^{g-1}(\mathrm{Jac}(C))$.
-${\mathrm{Ce}}_{\xi}(C)$ is clearly homologically trivial because $[-1]$ acts trivially on even-degree cohomology groups.
-It is not hard to check that $\Delta_{{\mathrm{GS}},\xi}(C)$ is also homologically trivial; this follows for example from [GS95, Prop.
-3.1] and because the map $\mathrm{Pic}^{1}(C)\rightarrow\mathrm{CH}^{2}(C^{3})$, $\xi\mapsto\Delta_{{\mathrm{GS}},\xi}(C)$ is a group homomorphism.
-Let $\omega_{C}$ be the canonical divisor on $C$.
+Let $C$ be an irreducible smooth projective curve defined over a field $k$ of genus $g\geq 3$. Let $e\in C(k)$. We have:
+    For each subset $T\subseteq\{1,2,3\}$, the modified diagonal $\Delta_{T}(C)=\{(x_{1},x_{2},x_{3}):x_{i}=e\text{ for }i\not\in T,~{}x_{j}=x_{j^{\prime}}\text{ for all }j,j^{\prime}\in T\}\in\mathrm{Ch}^{2}(C^{3})$.
+    The (classical) Abel–Jacobi map $i_{e}\colon C\rightarrow\mathrm{Jac}(C)$ sending $x\mapsto[x-e]$.
+Both $\Delta_{T}(C)$ and $i_{e}$ can be extended to any $e=\sum_{i}n_{i}e_{i}\in\mathrm{Div}^{1}(C)(k)=\mathrm{Pic}^{1}(C)(k)$.
+This is classical and direct for $i_{e}$, and for $\Delta_{T}(C)$ one can define $\Delta_{123}(C)=\{(c,c,c):c\in C\}$, $\Delta_{12}(C)=\sum n_{i}\{(c,c,e_{i}):c\in C\}$, $\Delta_{1}(C)=\sum_{i,j}n_{i}n_{j}\{(c,e_{i},e_{j}):c\in C\}$, etc.
+For any $e\in\mathrm{Pic}^{1}(C)$, define the Gross–Schoen cycle based at $e$ (resp. the Ceresa cycle based at $e$) to be:
+  • (Gross–Schoen cycle based at $e$) $\Delta_{{\mathrm{GS}},e}(C):=\Delta_{123}(C)-\Delta_{12}(C)-\Delta_{13}(C)-\Delta_{23}(C)+\Delta_{1}(C)+\Delta_{2}(C)+\Delta_{3}(C)\in\mathrm{Ch}^{2}(C^{3})$.
+  • (Ceresa cycle based at $e$) ${\mathrm{Ce}}_{e}(C):=i_{e}(C)-[-1]_{*}i_{e}(C)\in\mathrm{Ch}^{g-1}(\mathrm{Jac}(C))$.
+${\mathrm{Ce}}_{e}(C)$ is clearly homologically trivial because $[-1]$ acts trivially on even-degree cohomology groups.
+It is not hard to check that $\Delta_{{\mathrm{GS}},e}(C)$ is also homologically trivial; this follows for example from [GS95, Prop.
+3.1] and because the map $\mathrm{Pic}^{1}(C)\rightarrow\mathrm{Ch}^{2}(C^{3})$, $e\mapsto\Delta_{{\mathrm{GS}},e}(C)$ is a group homomorphism.
+Let $\omega_{C}$ be the canonical divisor on $C$ and $[\omega_{C}]$ its divisor class.
 Definition C.8.
@@
   • (Ceresa cycle) ${\mathrm{Ce}}(C):={\mathrm{Ce}}_{\xi}(C)\in\mathrm{Ch}^{g-1}(\mathrm{Jac}(C))_{\hom}$.
-There are finitely many $\xi$’s with $(2g-2)\xi=\omega_{C}$ and each two differ from a $(2g-2)$-torsion.
+There are finitely many $\xi$’s with $(2g-2)\xi=[\omega_{C}]$ and each two differ from a $(2g-2)$-torsion.
 Hence, the Gross–Schoen cycle and the Ceresa cycle of $C$ are well-defined at $(2g-2)$-torsions.
-We will associate normal functions with the Schoen cycles and Ceresa cycles.
+We will associate normal functions with the Gross–Schoen cycles and Ceresa cycles.
 First we have ${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))\in{\mathscr{J}}^{2}(C^{3})={\mathscr{J}}(H^{3}(C^{3},{\mathbb{Z}})(2))$ and ${\mathrm{AJ}}({\mathrm{Ce}}(C))\in{\mathscr{J}}^{g-1}(\mathrm{Jac}(C))={\mathscr{J}}(H^{2g-3}(\mathrm{Jac}(C),{\mathbb{Z}})(g-1))$.
@@
 Next, $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$, as a subspace of $H^{3}(C^{3},{\mathbb{Z}})$, has a basis consisting of $\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3}$, with $\alpha_{j}$ the pullback of an element in $H^{1}(C,{\mathbb{Z}})$ under the $j$-th projection $C^{3}\rightarrow C$.
-The symmetric group $S_{3}$ acts naturally on $C^{3}$, and this induces an action of $S_{3}$ on $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$ with $\sigma(\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3})=\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$ for each $\sigma\in S_{3}$.
+The symmetric group $S_{3}$ acts naturally on $C^{3}$, and this induces an action of $S_{3}$ on $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$ with
+$\sigma(\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3})=\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$
+for each $\sigma\in S_{3}$.
 Then $(H^{1}(C,{\mathbb{Z}})^{\otimes 3})^{S_{3}}=\bigwedge^{3}H^{1}(C,{\mathbb{Z}})$ for this action, with each member having a basis consisting of $\sum_{\sigma\in S_{3}}\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$.
@@
 $(\iota_{3})_{*}\Delta_{{\mathrm{GS}}}(C)-3{\mathrm{Ce}}(C)=\sum\nolimits_{j\geq 2}a_{j}C_{j}$
-For appropriate numbers $a_{j}$.
-The multiplication $[n]\colon\mathrm{Jac}(C)\rightarrow\mathrm{Jac}(C)$ induces $\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})\rightarrow\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})$, $x\mapsto n^{3}x$.
+For appropriate numbers $a_{j}$. The multiplication $[n]\colon\mathrm{Jac}(C)\rightarrow\mathrm{Jac}(C)$ induces
+$\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})\rightarrow\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}}),\quad x\mapsto n^{3}x.$
 Hence $n^{2+j}{\mathrm{AJ}}(C_{j})=n^{3}{\mathrm{AJ}}(C_{j})$ for all $n\in{\mathbb{Z}}$, and so ${\mathrm{AJ}}(C_{j})=0$ for all $j\geq 2$.
@@
 We have a VHS $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}\rightarrow{\mathscr{M}}_{g}$.
-The fiberwise polarization $q_{H}\colon H^{1}(C,{\mathbb{Z}})\otimes H^{1}(C,{\mathbb{Z}})\xrightarrow{\cup}H^{2}(C,{\mathbb{Z}})\cong{\mathbb{Z}}(-1)$, with $\cup$ the cup product (the dual of the intersection pairing) on $H^{1}(C,{\mathbb{Z}})$, induces a polarization on $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}$.
+The fiberwise polarization
+$q_{H}\colon H^{1}(C,{\mathbb{Z}})\otimes H^{1}(C,{\mathbb{Z}})\xrightarrow{\cup}H^{2}(C,{\mathbb{Z}})\simeq{\mathbb{Z}}(-1),$
+with $\cup$ the cup product (the dual of the intersection pairing) on $H^{1}(C,{\mathbb{Z}})$, induces a polarization on $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}$.
 Now we have two normal functions
@@
 where $\eta_{W_{s}}$ is a Green’s current for $W_{s}$. This pairing is known to be symmetric.
-The two families of homologically trivial cocycles give rise to two admissible normal functions.
+The two families of homologically trivial cocycles give rise to two admissible normal functions
 $\nu_{Z}\colon S\rightarrow{\mathscr{J}}^{p}(X/S)\quad\text{ and }\quad\nu_{W}\colon S\rightarrow{\mathscr{J}}^{q}(X/S).$
 A particularly important case is when $p=q=\frac{d+1}{2}$ and that $Z$ and $W$ are rationally equivalent.
-In this case, we obtain the height pairings $\langle Z_{s},Z_{s}\rangle_{\infty}$, or simply the height of $Z_{s}$.
+In this case, we obtain the height pairings $\langle Z_{s},Z_{s}\rangle_{\infty}$, or simply the archimedean local height of $Z_{s}$.
 More precisely, let $Z$ be a family of homologically trivial $p$-cocycle in $X/S$ with $p=\frac{d+1}{2}$, and let $\eta$ be the generic point of $S$.
@@
 Notice that in this case $\nu_{Z}=\nu_{W}$.
+We can extend this definition to $p\leq\frac{d+1}{2}$ as follows.
+Fix a relatively ample line bundle $L$ on $X/S$.
+Then intersection induces a family of maps $\cdot L_{s}^{d+1-2p}\colon{\mathrm{Ch}}^{p}(X_{s})_{\hom}\rightarrow{\mathrm{Ch}}^{q}(X_{s})_{\hom}$ for each $s\in S({\mathbb{C}})$.
+Therefore we can define $\langle Z_{s},Z_{s}\rangle_{\infty}$ to be $\langle Z_{s},L_{s}^{d+1-2p}(Z_{s})\rangle_{\infty}$; here one may also need to apply Moving Lemma.
+In some cases, we can also extend this definition to $p>\frac{d+1}{2}$.
+In this case $q=d+1-p\leq\frac{d+1}{2}$, and we have $\cdot L_{s}^{d+1-2q}\colon{\mathrm{Ch}}^{q}(X_{s})_{\hom}\rightarrow{\mathrm{Ch}}^{p}(X_{s})_{\hom}$.
+Now if $L_{s}^{d+1-2q}$ is an isomorphism, then we can define $\langle Z_{s},Z_{s}\rangle_{\infty}$ to be $\langle Z_{s},(L_{s}^{d+1-2q})^{-1}(Z_{s})\rangle_{\infty}$.
+This happens, for example, when $X_{s}$ is an abelian variety.
 Hain [Hai90, §$\S$3.3] related the local height pairing at archimedean places (D.1) to the metrized Poincaré bundle. We will review this in the next subsection.
@@
   1. (i)
-    Over each $s\in S({\mathbb{C}})$, we have ${\mathscr{P}}|_{\{0\}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}\cong{\mathscr{O}}_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}$,
+    Over each $s\in S({\mathbb{C}})$, we have ${\mathscr{P}}|_{\{0\}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}\simeq{\mathscr{O}}_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}$,
   2. (ii)
@@
   3. (iii)
-    $\epsilon^{*}{\mathscr{P}}\cong{\mathscr{O}}_{S}$ for the zero section $\epsilon$ of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}\rightarrow S$.
-Moreover, ${\mathscr{P}}$ can be endowed a canonical Hermitian metric $\|\cdot\|_{\mathrm{can}}$ uniquely determined by the following properties: (i) the curvature of $\overline{{\mathscr{P}}}$ is translation invariant on each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}$; (ii) $\epsilon^{*}\overline{{\mathscr{P}}}\cong({\mathscr{O}}_{S},\|\cdot\|_{\mathrm{triv}})$ for the trivial metric on ${\mathscr{O}}_{S}$.
+    $\epsilon^{*}{\mathscr{P}}\simeq{\mathscr{O}}_{S}$ for the zero section $\epsilon$ of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}\rightarrow S$.
+Moreover, ${\mathscr{P}}$ can be endowed a canonical Hermitian metric $\|\cdot\|_{\mathrm{can}}$ uniquely determined by the following properties: (i) the curvature of $\overline{{\mathscr{P}}}$ is translation invariant on each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}$; (ii) $\epsilon^{*}\overline{{\mathscr{P}}}\simeq({\mathscr{O}}_{S},\|\cdot\|_{\mathrm{triv}})$ for the trivial metric on ${\mathscr{O}}_{S}$.
 Definition D.1.
@@
 $\Delta_{{\mathscr{Q}}}=(1,i_{{\mathscr{Q}}})\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}.$
-This is the case, for example, for Poincaré duality, where the polarization is given by the cup product.
+This is the case, for example, for Poincaré duality, where the cup product gives the polarization.
 Definition D.3.
@@
 ∎
+Appendix E Adelic line bundles over quasi-projective varieties
+The goal of this appendix is to review some basic notions and facts about adelic line bundles over quasi-projective varieties developed by the second-named author and Yuan [YZ21], generalizing [Zha95].
+See also [Yua21, §$\S$2.1] for another quick summary.
+Let $K$ be a number field and let ${\mathscr{O}}_{K}$ be its ring of integers.
+Through the whole appendix, $X$ is a flat and quasi-projective variety defined over $K$.
+Via the natural inclusions ${\mathbb{Z}}\subseteq{\mathbb{Q}}\subseteq K$, we can see $X$ as a scheme over ${\mathrm{Spec}}{\mathbb{Z}}$.
+By an arithmetic variety, we mean an integral scheme that is flat, separated, and of finite type over ${\mathrm{Spec}}{\mathbb{Z}}$.
+E.1. Adelic line bundles on $X$ as limits of model Hermitian line bundles
+An adelic line bundle $\overline{L}$ on $X$ is, roughly speaking, the limit of a sequence of model (Hermitian) line bundles $({\mathscr{X}}_{i},\overline{\mathscr{L}}_{i})$, where ${\mathscr{X}}_{i}$ is a projective arithmetic variety such that $X$ is Zariski open in ${\mathscr{X}}_{i,{\mathbb{Q}}}$.
+We say that the sequence $({\mathscr{X}}_{i},\overline{\mathscr{L}}_{i})$ converges to $(X,\overline{L})$.
+For this limit/convergence to make sense, we will construct the group of isomorphism classes of adelic line bundles on $X$, denoted by $\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})$, in the following steps.
+  1. (i)
+    Consider all quasi-projective models ${\mathscr{U}}$ of $X$, i.e.
+${\mathscr{U}}$ is an integral scheme which is quasi-projective over ${\mathrm{Spec}}{\mathscr{O}}_{K}$, such that $X$ is open in the generic fiber of ${\mathscr{U}}\rightarrow{\mathrm{Spec}}{\mathscr{O}}_{K}$.
+These quasi-projective models of $X$ form an inverse system.
+  2. (ii)
+    Define for each quasi-projective model ${\mathscr{U}}$ of $X$ the groups of isomorphism classes adelic line bundles $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})$ and of adelic divisors $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})$ on ${\mathscr{U}}$.
+They are defined as suitable completions of the groups of isomorphism classes of model adelic line bundles $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$ and of model adelic divisors $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$.
+  3. (iii)
+[…]
+Steps (i) and (iii) are formal.
+We will explain step (ii) from two different points of view: in §$\S$§E.3 by constructing $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})$ and $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})$ via a limit process, and in §$\S$§E.4 by explaining how to see adelic line bundles on ${\mathscr{U}}$ as metrized line bundles on the (Berkovich) analytification ${\mathscr{U}}^{\mathrm{an}}$ of ${\mathscr{U}}$ and to see adelic divisors on ${\mathscr{U}}$ as arithmetic divisors on ${\mathscr{U}}^{\mathrm{an}}$.
+Before doing so, we need to recall some basic definitions of Hermitian line bundles and arithmetic divisors on projective arithmetic varieties in §$\S$§E.2.
+E.2. Hermitian line bundles and arithmetic divisors
+Let ${\mathscr{X}}$ be a projective arithmetic variety.
+[…]
+$\widehat{\mathrm{Pic}}({\mathscr{X}}):=\{\text{Hermitian line bundles on }{\mathscr{X}}\}/\simeq.$
+Then $\widehat{\mathrm{Pic}}({\mathscr{X}})$ has a natural group structure, with the identity element being $({\mathscr{O}}_{X},\|\cdot\|_{\mathrm{triv}})$ and the group law being the tensor product.
+An arithmetic divisor $\overline{D}$ on ${\mathscr{X}}$ is a pair $(D,g)$ where $D$ is a Cartier divisor on ${\mathscr{X}}$ and $g$ is a continuous Green’s function of $|D|({\mathbb{C}})$ on ${\mathscr{X}}({\mathbb{C}})$ which is invariant under the complex conjugation.
+Here a Green’s function is a continuous function $g\colon{\mathscr{X}}({\mathbb{C}})\setminus|D|({\mathbb{C}})\rightarrow{\mathbb{R}}$ such that for any meromorphic function $f$ on an open subset $U$ of ${\mathscr{X}}$ with $\mathrm{div}(f)=D|_{U}$, the function $g+\log|f|$ extends to a continuous function on $U$.
+Next, a principle arithmetic divisor is defined to be an arithmetic divisor of the form $(\mathrm{div}(f),-\log|f|)$ for some non-zero rational function $f$ on ${\mathscr{X}}$.
+Denote by
+$\widehat{{\mathrm{Div}}}({\mathscr{X}}):=\{\text{the group of arithmetic divisors on }{\mathscr{X}}\}$
+and $\widehat{\mathrm{Prin}}({\mathscr{X}}):=\{\text{the group of principal arithmetic divisors on }{\mathscr{X}}\}$.
+There exists a canonical isomorphism
+(E.1) $\widehat{{\mathrm{Div}}}({\mathscr{X}})/\widehat{\mathrm{Prin}}({\mathscr{X}})\xrightarrow{\sim}\widehat{\mathrm{Pic}}({\mathscr{X}}),$
+[…]
+[…]
+An arithmetic divisor $\overline{D}=(D,g)\in\widehat{{\mathrm{Div}}}({\mathscr{X}})$ is called effective if $D\geq 0$ and $g\geq 0$ on ${\mathscr{X}}({\mathbb{C}})\setminus|D|({\mathbb{C}})$.
+It is called strictly effective if furthermore $g>0$ on ${\mathscr{X}}({\mathbb{C}})\setminus D({\mathbb{C}})$.
+E.3. Adelic line bundles on ${\mathscr{U}}$ via limit process
+Let ${\mathscr{U}}$ be a quasi-projective arithmetic variety.
+In this subsection, we construct the group of isomorphism classes of adelic line bundles $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})$ via a limit process on the projective models of ${\mathscr{U}}$, where a projective model ${\mathscr{X}}$ of ${\mathscr{U}}$ is a projective arithmetic variety ${\mathscr{X}}$ which contains ${\mathscr{U}}$ as an open subscheme.
+A model adelic line bundle on ${\mathscr{U}}$ is a pair $({\mathscr{X}},\overline{{\mathscr{L}}})$ where ${\mathscr{X}}$ is a projective model of ${\mathscr{U}}$ and ¯$\overline{{\mathscr{L}}}\in\widehat{\mathrm{Pic}}({\mathscr{X}})_{{\mathbb{Q}}}$ such that ${\mathscr{L}}|_{{\mathscr{U}}}\in\mathrm{Pic}({\mathscr{U}})$.
+The group of isomorphism classes of model adelic line bundles on ${\mathscr{U}}$ is denoted by $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$.
+An adelic line bundle on ${\mathscr{U}}$ is a pair $({\mathscr{L}},({\mathscr{X}}_{i},\overline{{\mathscr{L}}}_{i},\ell_{i})_{i\geq 1})$ where ${\mathscr{L}}\in\mathrm{Pic}({\mathscr{U}})$, $({\mathscr{X}}_{i},\overline{{\mathscr{L}}}_{i})$ is a model adelic line bundle for each $i\geq 1$ and $\ell_{i}\colon{\mathscr{L}}\rightarrow{\mathscr{L}}_{i}|_{{\mathscr{U}}}$ is an isomorphism in $\mathrm{Pic}({\mathscr{U}})_{{\mathbb{Q}}}$ for each $i\geq 1$, such that the sequence $({\mathscr{X}}_{i},\overline{{\mathscr{L}}}_{i},\ell_{i})_{i\geq 1}$ satisfies a suitable Cauchy condition.
+It is more convenient to define this Cauchy condition on model adelic divisors.
+So now let us construct the group of adelic divisors $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})$ in the following steps.
+  1. (a)
+    Consider all projective models ${\mathscr{X}}$ of ${\mathscr{U}}$. These projective models of ${\mathscr{U}}$ form an inverse system.
+  2. (b)
+    For each projective model ${\mathscr{X}}$ of ${\mathscr{U}}$, define $\widehat{{\mathrm{Div}}}({\mathscr{X}},{\mathscr{U}})$ to be the fiber product $\widehat{{\mathrm{Div}}}({\mathscr{X}})_{{\mathbb{Q}}}\times_{{\mathrm{Div}}({\mathscr{X}})_{{\mathbb{Q}}}}{\mathrm{Div}}({\mathscr{U}})$, where the fiber product is taken with respect to the forgetful map $\widehat{{\mathrm{Div}}}({\mathscr{X}})_{{\mathbb{Q}}}\rightarrow{\mathrm{Div}}({\mathscr{X}})_{{\mathbb{Q}}}$ and the natural map ${\mathrm{Div}}({\mathscr{U}})\rightarrow{\mathrm{Div}}({\mathscr{X}})\rightarrow{\mathrm{Div}}({\mathscr{X}})_{{\mathbb{Q}}}$.
+Roughly speaking, $\widehat{{\mathrm{Div}}}({\mathscr{X}},{\mathscr{U}})$ consists of ${\mathbb{Q}}$-arithmetic divisors $\overline{D}=(D,g)$ on ${\mathscr{X}}$ (with $D\in{\mathrm{Div}}({\mathscr{X}})_{{\mathbb{Q}}}$) such that $D|_{{\mathscr{U}}}$ lies in in ${\mathrm{Div}}({\mathscr{U}})$.
+  3. (c)
+[…]
+  4. (d)
+    Finally, define $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})$ to be the equivalence classes of Cauchy sequences in $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$ with respect to the boundary topology.
+We need to explain the boundary topology on $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$ in the last step.
+First, notice that there is a partial order $\leq$ on $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$: Let $\overline{D}_{1},\overline{D}_{2}\in\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$, then $\overline{D}_{1},\overline{D}_{2}\in\widehat{{\mathrm{Div}}}({\mathscr{X}},{\mathscr{U}})$ for some projective model ${\mathscr{X}}$, and we say that $\overline{D}_{1}\leq\overline{D}_{2}$ if the images of $\overline{D}_{2}-\overline{D}_{1}\in\widehat{{\mathrm{Div}}}({\mathscr{X}},{\mathscr{U}})$ in $\widehat{{\mathrm{Div}}}({\mathscr{X}})_{{\mathbb{Q}}}$ and in ${\mathrm{Div}}({\mathscr{U}})$ are both effective.
+Next, fix a pair $({\mathscr{X}}_{0},\overline{{\mathscr{E}}}_{0})$ with ${\mathscr{X}}_{0}$ a projective model of ${\mathscr{U}}$ and ¯$\overline{{\mathscr{E}}}_{0}$ a strictly effective arithmetic divisor on ${\mathscr{X}}_{0}$ supporting at ${\mathscr{X}}_{0}\setminus{\mathscr{U}}$.
+It defines the following boundary norm
+¯$\|\cdot\|_{\overline{{\mathscr{E}}}_{0}}\colon\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}\rightarrow[0,\infty]$
+by sending $\overline{D}\mapsto\inf\{\epsilon\in{\mathbb{Q}}_{>0}:-\epsilon\overline{{\mathscr{E}}}_{0}\leq\overline{D}\leq\epsilon\overline{{\mathscr{E}}}_{0}\}$.
+This boundary norm induces a topology on $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$, which we call the boundary topology.
+We remark that the boundary topology does not depend on the choice of the pair $({\mathscr{X}}_{0},\overline{{\mathscr{E}}}_{0})$.
+[…]
+E.4. Adelic line bundles on ${\mathscr{U}}$ as metrized line bundles
+Let ${\mathscr{U}}$ be a quasi-projective arithmetic variety.
+A more intrinsic way to understand adelic line bundles is that they are still line bundles with a suitable metric, but now on the (Berkovich) analytification ${\mathscr{U}}^{\mathrm{an}}$ of ${\mathscr{U}}$.
+We start by constructing ${\mathscr{U}}^{\mathrm{an}}$.
+If ${\mathscr{U}}={\mathrm{Spec}}A$, then ${\mathscr{U}}^{\mathrm{an}}$ is defined to be ${\mathscr{M}}(A)$ of multiplicative semi-norms ||$\cdot$|| ⋅ | on $A$, endowed with the weakest topology such that ${\mathscr{M}}(A)\rightarrow{\mathbb{R}}$, $|\cdot|\mapsto|f|$ is continuous for all $f\in A$.
+Denote by $|\cdot|_{x}$ the semi-norm on $A$ corresponding to $x\in{\mathscr{M}}(A)$.
+The contraction is the map
+$\kappa\colon({\mathrm{Spec}}A)^{\mathrm{an}}\rightarrow{\mathrm{Spec}}A$
+defined by sending each $x\in{\mathscr{M}}(A)$ to the prime ideal $\bar{x}:=\ker(|\cdot|_{x})$.
+For each $x\in({\mathrm{Spec}}A)^{\mathrm{an}}$, the residue field $H_{x}$ of $x$ is defined to be the completion of $\mathrm{Frac}(A/\ker(|\cdot|_{x}))$.
+Now $|\cdot|_{x}$ factors through
+$|\cdot|_{x}\colon A\rightarrow H_{x}\xrightarrow{|\cdot|}{\mathbb{R}}.$
+In general ${\mathscr{U}}$ is covered by an affine open cover $\{{\mathrm{Spec}}A_{i}\}_{i}$, and ${\mathscr{U}}^{\mathrm{an}}$ is defined to be the union of ${\mathscr{M}}(A_{i})$, glued in the canonical way, such that each ${\mathscr{M}}(A_{i})$ is an open subspace of ${\mathscr{U}}^{\mathrm{an}}$.
+Then we can define the contraction
+$\kappa\colon{\mathscr{U}}^{\mathrm{an}}\rightarrow{\mathscr{U}},\qquad x\mapsto\bar{x}$
+and the residue field $H_{x}$ of each $x\in{\mathscr{U}}^{\mathrm{an}}$ since the definitions are local.
+We again have a semi-norm $|\cdot|\colon H_{x}\rightarrow{\mathbb{R}}$ induced by $|\cdot|_{x}$ for each $x\in{\mathscr{U}}^{\mathrm{an}}$.
+Notice that in the constructions above, we did not use the flatness of ${\mathscr{U}}\rightarrow{\mathrm{Spec}}{\mathbb{Z}}$.
+For example, the construction works with any quasi-projective variety $X$ defined over a number field $K$, viewed as a scheme over ${\mathrm{Spec}}{\mathbb{Z}}$ via ${\mathbb{Z}}\subseteq{\mathbb{Q}}\subseteq K$.
+Example E.1.
+Let us look at $({\mathrm{Spec}}{\mathbb{Z}})^{\mathrm{an}}$. It is the union of the closed-line segments
+$[0,1]_{\infty}:=\{|\cdot|_{\infty}^{t}:0\leq t\leq 1\}$
+and the closed-line segments
+$[0,\infty]_{p}:=\{|\cdot|_{p}^{t}:0\leq t\leq\infty\}$
+for all finite prime numbers $p>0$, by identifying the endpoints $|\cdot|_{\infty}^{0}$ and $|\cdot|_{p}^{0}$ for all $p$ with the trivial norm $|\cdot|_{0}$ on ${\mathbb{Z}}$.
+In particular, $({\mathrm{Spec}}{\mathbb{Z}})^{\mathrm{an}}$ is compact and path-connected.
+The contraction map leaves $|\cdot|_{p}^{\infty}$ stable and sends all other points to $|\cdot|_{0}$.
+For each finite prime $p$, the residue field of $|\cdot|_{p}^{t}$ is ${\mathbb{Q}}_{p}$ when $t\in(0,\infty)$ and is ${\mathbb{F}}_{p}$ when $t=\infty$.
+The residue field of $|\cdot|_{\infty}^{t}$ is ${\mathbb{R}}$; this includes the case of $|\cdot|_{0}$.
+And $({\mathrm{Spec}}{\mathbb{Q}})^{\mathrm{an}}$ is $({\mathrm{Spec}}{\mathbb{Z}})^{\mathrm{an}}$ with $|\cdot|_{p}^{\infty}$ removed for all $p>0$.
+In general, the structural morphism ${\mathscr{U}}\rightarrow{\mathrm{Spec}}{\mathbb{Z}}$ gives a structural map ${\mathscr{U}}^{\mathrm{an}}\rightarrow({\mathrm{Spec}}{\mathbb{Z}})^{\mathrm{an}}$.
+This divides ${\mathscr{U}}^{\mathrm{an}}$ into
+(E.2) ${\mathscr{U}}^{\mathrm{an}}={\mathscr{U}}^{\mathrm{an}}[f]\coprod{\mathscr{U}}^{\mathrm{an}}[\infty],$
+where ${\mathscr{U}}^{\mathrm{an}}[f]$ is the inverse image of $(0,\infty]_{p}$ and ${\mathscr{U}}^{\mathrm{an}}[\infty]$ is the inverse image of $[0,1]_{\infty}$.
+Denote by ${\mathscr{U}}^{\mathrm{an}}_{\mathrm{triv}}$ the fiber over the point $|\cdot|_{0}\in({\mathrm{Spec}}{\mathbb{Z}})^{\mathrm{an}}$ corresponding to the trivial norm on ${\mathbb{Z}}$.
+Now we are ready to define metrized line bundles and arithmetic divisors on ${\mathscr{U}}^{\mathrm{an}}$.
+[…]
+Similarly, an arithmetic divisor on ${\mathscr{U}}^{\mathrm{an}}$ is a pair $\overline{D}=(D,g)$ where $D$ is a Cartier divisor on ${\mathscr{U}}$ and $g$ is a continuous Green’s function of $|D|^{\mathrm{an}}$ on ${\mathscr{U}}^{\mathrm{an}}$, i.e.
+a continuous function $g\colon{\mathscr{U}}^{\mathrm{an}}\setminus|D|^{\mathrm{an}}\rightarrow{\mathbb{R}}$ such that for any meromorphic function $f$ on an open subset ${\mathscr{V}}$ of ${\mathscr{U}}$ with $\mathrm{div}(f)=D|_{{\mathscr{V}}}$, the function $g+\log|f|$ extends to a continuous function on ${\mathscr{V}}^{\mathrm{an}}$.
+The group of arithmetic divisors on ${\mathscr{U}}^{\mathrm{an}}$ is denoted by $\widehat{{\mathrm{Div}}}({\mathscr{U}}^{\mathrm{an}})$.
+We also have the group of principal arithmetic divisors on ${\mathscr{U}}^{\mathrm{an}}$, denoted by $\widehat{\mathrm{Prin}}({\mathscr{U}}^{\mathrm{an}})$, consisting of arithmetic divisors of the form $\widehat{{\mathrm{div}}}(f):=({\mathrm{div}}(f),-\log|f|)$ for some non-zero rational function $f$ on ${\mathscr{U}}$.
+As in the case of Hermitian line bundles and arithmetic divisors on projective arithmetic varieties, we also have a canonical isomorphism
+(E.3) $\widehat{{\mathrm{Div}}}({\mathscr{U}}^{\mathrm{an}})/\widehat{\mathrm{Prin}}({\mathscr{U}}^{\mathrm{an}})\xrightarrow{\sim}\widehat{\mathrm{Pic}}({\mathscr{U}}^{\mathrm{an}})$
+with a similar construction as (E.1).
+By [YZ21, Prop. 3.4.1], there exists a canonical injective map
+(E.4) $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}}){\longrightarrow}\widehat{\mathrm{Pic}}({\mathscr{U}}^{\mathrm{an}}).$
+Moreover, any ¯$\overline{{\mathscr{L}}}\in\widehat{\mathrm{Pic}}({\mathscr{U}}^{\mathrm{an}})$ in the image of this map is norm-equivariant, i.e.
+for any rational section $s$ of ${\mathscr{L}}$ on ${\mathscr{U}}$ and any points $x,x^{\prime}\in{\mathscr{U}}^{\mathrm{an}}\setminus|{\mathrm{div}}(s)|^{\mathrm{an}}$ satisfying $|\cdot|_{x}=|\cdot|_{x^{\prime}}^{t}$ (with $0\leq t<\infty$) locally on ${\mathscr{O}}_{{\mathscr{U}}}$, we have $\|s\|_{x}=\|s\|_{x^{\prime}}^{t}$.
+Similarly there exists a canonical injective map $\widehat{{\mathrm{Div}}}({\mathscr{U}}/{\mathbb{Z}}){\longrightarrow}\widehat{{\mathrm{Div}}}({\mathscr{U}}^{\mathrm{an}})$ mapping $\widehat{\mathrm{Prin}}({\mathscr{U}}/{\mathbb{Z}})$ into $\widehat{\mathrm{Prin}}({\mathscr{U}}^{\mathrm{an}})$.
+In this point of view, the completion process to get $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})$ from $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$ can be taken as follows.
+Fix a pair $({\mathscr{X}}_{0},\overline{{\mathscr{E}}}_{0})$ with ${\mathscr{X}}_{0}$ a projective model of ${\mathscr{U}}$ and ¯$\overline{{\mathscr{E}}}_{0}=({\mathscr{E}}_{0},g_{0})$ an arithmetic divisor on ${\mathscr{X}}_{0}$ with $|{\mathscr{E}}_{0}|={\mathscr{X}}_{0}\setminus{\mathscr{U}}$ and $g_{0}>0$.
+Then ¯$\overline{{\mathscr{E}}}_{0}$ induces an arithmetic divisor $({\mathscr{E}}_{0},\widetilde{g}_{0})$ on ${\mathscr{X}}_{0}^{\mathrm{an}}$ which restricts to a continuous function ${\mathscr{U}}^{\mathrm{an}}\rightarrow{\mathbb{R}}_{\geq 0}$; see [YZ21, Lem.
+3.3.2] for more details[2][2][2]On ${\mathscr{X}}_{0}^{\mathrm{an}}[f]$ it is defined using the local meromorphic function on ${\mathscr{X}}_{0}$.
+On ${\mathscr{X}}_{0}^{\mathrm{an}}[\infty]$, it is defined using the Green’s function $g_{0}$ and then extended using norm-equivariance..
+Thus one can define the boundary topology on $C({\mathscr{U}}^{\mathrm{an}})$, the space of continuous functions on ${\mathscr{U}}^{\mathrm{an}}$, by the extended norm
+$\|f\|_{\overline{{\mathscr{E}}}_{0}}:=\sup_{x\in{\mathscr{U}}^{\mathrm{an}},~{}\widetilde{g}_{0}(x)>0}\frac{|f(x)|}{\widetilde{g}_{0}(x)}.$
+Now, we say that a sequence {¯$\{\overline{{\mathscr{L}}}_{i}=({\mathscr{L}}_{i},\|\cdot\|_{i})\}_{i\geq 1}$ in $\widehat{\mathrm{Pic}}({\mathscr{U}}^{\mathrm{an}})$ converges to ¯$\overline{{\mathscr{L}}}:=({\mathscr{L}},\|\cdot\|)\in\widehat{\mathrm{Pic}}({\mathscr{U}}^{\mathrm{an}})$ if there exist isomorphisms $\tau_{i}\colon{\mathscr{L}}\rightarrow{\mathscr{L}}_{i}$ such that the sequence $\{-\log\frac{\tau_{i}^{*}\|\cdot\|_{i}}{\|\cdot\|}\}_{i\geq 1}$ converges to 00 in $C({\mathscr{U}}^{\mathrm{an}})$ under the boundary topology above.
+Thus an adelic line bundle ¯$\overline{{\mathscr{L}}}\in\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})$ is a metrized line bundle in $\widehat{\mathrm{Pic}}({\mathscr{U}}^{\mathrm{an}})$ which is isomorphic to the limit of a sequence in $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{mod}}$.
+E.5. The generic fiber of an adelic line bundle
+For each adelic line bundle ¯$\overline{{\mathscr{L}}}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})$, we give two equivalent constructions of its generic fiber (or its geometric part) $\widetilde{{\mathscr{L}}}$.
+The first way is to construct $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})$.
+The construction is similar to $\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})$, but much easier because there are no arithmetic inputs now.
+First define $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\mathrm{mod}}$ to be the group of isomorphism classes of pairs $(Y,L)$ where $Y$ is a projective model of $X$ (i.e.
+$Y$ is a projective variety defined over ${\mathbb{Q}}$ which contains $X$ as an open subset) and $L\in\mathrm{Pic}(Y)_{{\mathbb{Q}}}$ such that $L|_{X}\in\mathrm{Pic}(X)$.
+Next define $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})$ as the group of isomorphism classes of pairs $(L,(Y_{i},L_{i},\ell_{i})_{i\geq 1})$ where $L\in\mathrm{Pic}(X)$, $(Y_{i},L_{i})\in\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\mathrm{mod}}$ and $\ell_{i}\colon L\rightarrow L_{i}|_{X}$ is an isomorphism in $\mathrm{Pic}(X)_{{\mathbb{Q}}}$ such that the sequence $\{{\mathrm{div}}(\ell_{i}\ell_{1}^{-1})\}_{i\geq 1}$ satisfies the Cauchy condition defined using the boundary topology on $\mathrm{Div}(X/{\mathbb{Q}})_{\mathrm{mod}}$.
+Here, the boundary topology is defined using any fixed projective model $Y_{0}$ of $X$ such that $Y_{0}\setminus X$ is a divisor, which defines a boundary norm on $\mathrm{Div}(X/{\mathbb{Q}})_{\mathrm{mod}}$ by sending $\widetilde{D}\mapsto\inf\{\epsilon\in{\mathbb{Q}}_{>0}:-\epsilon(Y_{0}\setminus X)\leq\widetilde{D}\leq\epsilon(Y_{0}\setminus X)\}$.
+Notice that there is a natural map
+(E.5) $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})\rightarrow\mathrm{Pic}(X),\quad(L,(Y_{i},L_{i},\ell_{i})_{i\geq 1})\mapsto L.$
+Remark E.2.
+Let $k$ be any field (notably a field finitely generated over ${\mathbb{Q}}$), and let $X^{\prime}$ be a variety defined over $k$.
+The construction above defines $\widehat{\mathrm{Pic}}(X^{\prime}/k)$ by replacing ${\mathbb{Q}}$ by $k$ and $X$ by $X^{\prime}$.
+Now for any projective model ${\mathscr{X}}$ of $X$ over ${\mathbb{Z}}$, the generic fiber ${\mathscr{X}}_{{\mathbb{Q}}}$ is a projective model of $X$, and hence we have a natural map $\widehat{\mathrm{Pic}}({\mathscr{X}})\rightarrow\mathrm{Pic}({\mathscr{X}}_{{\mathbb{Q}}})$.
+This induces the desired map
+(E.6) $\widehat{\mathrm{Pic}}(X/{\mathbb{Z}}){\longrightarrow}\widehat{\mathrm{Pic}}(X/{\mathbb{Q}}),\qquad\overline{{\mathscr{L}}}\mapsto\widetilde{{\mathscr{L}}},$
+with $\widetilde{{\mathscr{L}}}$ called the generic fiber of $\overline{{\mathscr{L}}}$.
+The second way is to see $\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})$ as metrized line bundles on the (Berkovich) analytification $X^{\mathrm{an}}$ of $X$; see the comment above Example E.1.
+Then the structural morphism $X\rightarrow{\mathrm{Spec}}K$ gives a structural map $X^{\mathrm{an}}\rightarrow({\mathrm{Spec}}K)^{\mathrm{an}}$.
+Denote by $X^{\mathrm{an}}_{\mathrm{triv}}$ the fiber over the point $v_{0}\in({\mathrm{Spec}}K)^{\mathrm{an}}$ which corresponds to the trivial absolute value $|\cdot|_{0}$ on $K$, i.e.
+$|0|_{0}=0$ and $|a|_{0}=1$ for all $a\in K^{*}$.
+We have the pullback
+(E.7) $\widehat{\mathrm{Pic}}(X^{\mathrm{an}}){\longrightarrow}\widehat{\mathrm{Pic}}(X^{\mathrm{an}}_{\mathrm{triv}})$
+of the groups of isomorphism classes of metrized line bundles.
+The canonical injective map (E.4) induces a canonical injective map $\widehat{\mathrm{Pic}}(X/{\mathbb{Z}}){\longrightarrow}\widehat{\mathrm{Pic}}(X^{\mathrm{an}})$.
+Similarly there is a canonical injective map $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}}){\longrightarrow}\widehat{\mathrm{Pic}}(X^{\mathrm{an}}_{\mathrm{triv}})$.
+Thus, we obtain a diagram
+(E.8) ${\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{\mathrm{Pic}}(X^{\mathrm{an}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{\mathrm{Pic}}(X^{\mathrm{an}}_{\mathrm{triv}}).}$
+The following proposition says that the generic fiber of an adelic line bundle is obtained by pulling back the corresponding metrized line bundle to the fiber over the trivial norm.
+Proposition E.3.
+The diagram above commutes.
+E.6. Nefness and integrability
+Definition E.4.
+Let ${\mathscr{U}}$ be a quasi-projective arithmetic variety, and let ¯$\overline{{\mathscr{L}}}\in\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})$.
+  • $\overline{{\mathscr{L}}}$ is called strongly nef if it is the limit of a sequence of model adelic line bundles $({\mathscr{X}}_{i},\overline{{\mathscr{L}}}_{i})$ with ¯$\overline{{\mathscr{L}}}_{i}$ nef as a Hermitian ${\mathbb{Q}}$-line bundle on ${\mathscr{X}}_{i}$;
+  • $\overline{{\mathscr{L}}}$ is called nef if there exists a strongly nef adelic line bundle $\overline{{\mathscr{M}}}$ on ${\mathscr{U}}$ such that $a\overline{{\mathscr{L}}}+\overline{{\mathscr{M}}}$ is strongly nef for all $a\in{\mathbb{Z}}_{>0}$;
+  • $\overline{{\mathscr{L}}}$ is called integrable if it is isomorphic to ¯$\overline{{\mathscr{L}}}_{1}-\overline{{\mathscr{L}}}_{2}$ for two strongly nef adelic line bundles ¯$\overline{{\mathscr{L}}}_{1}$ and ¯$\overline{{\mathscr{L}}}_{2}$ on ${\mathscr{U}}$.
+With this definition, we can write the following subgroups of $\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})$:
+$\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{snef}},\quad\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{nef}},\quad\text{and}\quad\widehat{\mathrm{Pic}}({\mathscr{U}}/{\mathbb{Z}})_{\mathrm{int}}$
+of isomorphism classes of strongly nef, nef, and integrable adelic line bundles on ${\mathscr{U}}$.
+[…]
+[…]
+[…]
+[…]
+The adelic line bundles on $X$ in these subgroups are said to be strongly nef, nef, integrable respectively.
+We can also define similar notions for the generic fiber of an adelic line bundle on $X$.
+Recall the group $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})$ defined at the beginning of §$\S$§E.5.
+A $\widetilde{L}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})$ is called strongly nef if it is the limit of $\{(Y_{i},L_{i})\}_{i\geq 1}$ in $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\mathrm{mod}}$ with each $L_{i}$ a nef ${\mathbb{Q}}$-line bundle on $Y_{i}$.
+And $\widetilde{L}$ is called nef if there exists a strongly nef $\widetilde{M}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})$ such that $a\widetilde{L}+\widetilde{M}$ is strongly nef for all $a\in{\mathbb{Z}}_{>0}$, and $\widetilde{L}$ is called integrable if it is isomorphic to the difference of two strongly nef elements in $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})$.
+We also use $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\mathrm{snef}}$, $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\mathrm{nef}}$, and $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\mathrm{int}}$ to denote the relevant groups.
+Then, the operation of taking the generic fiber (E.6) yields
+$\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})_{\textbf{P}}{\longrightarrow}\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\textbf{P}},\qquad\overline{L}\mapsto\widetilde{L}$
+with P being snef, nef, or int.
+E.7. Height defined by adelic line bundles
+Extending Gillet–Soulé’s arithmetic intersection theory, one has an arithmetic intersection pairing
+(E.9) $\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})^{\dim X+1}_{\mathrm{int}}{\longrightarrow}{\mathbb{R}},\qquad(\overline{L}_{1},\ldots,\overline{L}_{\dim X+1})\mapsto\langle\overline{L}_{1},\ldots,\overline{L}_{\dim X+1}\rangle,$
+such that $\langle\overline{L}_{1},\ldots,\overline{L}_{\dim X+1}\rangle\geq 0$ if $\overline{L}_{i}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})_{\mathrm{nef}}$ for all $i\in\{1,\ldots,\dim X+1\}$.
+When $\dim X=0$, then this intersection is called the arithmetic degree of the adelic line bundle and is denoted by
+$\widehat{\deg}\colon\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})_{\mathrm{int}}{\longrightarrow}{\mathbb{R}}.$
+Definition E.5.
+Let $\overline{L}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})_{\mathrm{int}}$. The height function defined by $\overline{L}$ is
+$h_{\overline{L}}\colon X(\overline{\mathbb{Q}}){\longrightarrow}{\mathbb{R}},\qquad x\mapsto\frac{\widehat{\deg}(\overline{L}|_{x})}{[K(x):K]}.$
+As a remark, we point out that $\overline{L}$ needs not be integrable to define the height function.
+Denote by $(L,\|\cdot\|)\in\widehat{\mathrm{Pic}}(X^{\mathrm{an}})$ the image of $\overline{L}$ under the canonical injective map $\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})\rightarrow\widehat{\mathrm{Pic}}(X^{\mathrm{an}})$.
+Then the formula for the height function is:
+(E.10) $h_{\overline{L}}(x)=-\frac{1}{[K(x):K]}\sum_{p\in M_{{\mathbb{Q}}}}\sum_{z\in\mathrm{Gal}(\overline{\mathbb{Q}}/K)x\times_{{\mathbb{Q}}}{\mathbb{Q}}_{p}}\log\|s(z^{\mathrm{an}})\|^{\deg_{{\mathbb{Q}}_{p}}z}$
+for any non-zero rational section $s$ of $L$ on $X$ with $x\not\in|{\mathrm{div}}(s)|$; here $p$ can be infinity.
+E.8. Volume, bigness, and height inequality
+Let $\overline{L}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})$.
+Write $(L,\|\cdot\|)\in\widehat{\mathrm{Pic}}(X^{\mathrm{an}})$ for the image of $\overline{L}$ under the canonical injective map $\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})\rightarrow\widehat{\mathrm{Pic}}(X^{\mathrm{an}})$.
+To get lower bounds on $h_{\overline{L}}$, it is important to find “small” sections, as suggested by (E.10). Define the $\sup$-norm of a section $s\in H^{0}(X,L)$ to be
+$\|s\|_{\sup}:=\sup_{x\in X^{\mathrm{an}}}\|s(x)\|.$
+Define the arithmetic $H^{0}$ to be
+$\widehat{H}^{0}(X,\overline{L}):=\{s\in H^{0}(X,L):\|s\|_{\sup}\leq 1\}.$
+Now that $H^{0}(X,L)$ is a lattice in $H^{0}(X,L)_{{\mathbb{R}}}$ and $\widehat{H}^{0}(X,\overline{L})$ is contained in the unit ball, we have that $\widehat{H}^{0}(X,\overline{L})$ is a finite set.
+Define
+$\widehat{h}^{0}(X,\overline{L}):=\log\#\widehat{H}^{0}(X,\overline{L}).$
+Theorem-Definition E.6 ([YZ21, Thm. 5.2.1.(1)]).
+The following limit exists
+$\widehat{{\mathrm{vol}}}(X,\overline{L}):=\lim_{m\rightarrow\infty}\frac{(\dim X+1)!}{m^{\dim X+1}}\widehat{h}^{0}(X,m\overline{L}),$
+and is defined to be the (arithmetic) volume of $\overline{L}$.
+Moreover, part (2) of [YZ21, Thm.
+5.2.1] asserts that $\widehat{{\mathrm{vol}}}(X,\overline{L})$ can be computed as the limit of the arithmetic volumes of any sequence of model (Hermitian) line bundles converging to $\overline{L}$.
+Definition E.7.
+An adelic line bundle $\overline{L}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})$ is said to be big if $\widehat{{\mathrm{vol}}}(X,\overline{L})>0$.
+Theorem E.8 ([YZ21, Thm. 5.2.2 and 5.2.9]).
+We have:
+    (Arithmetic Hilbert–Samuel) If $\overline{L}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})$ is nef, then $\widehat{{\mathrm{vol}}}(X,\overline{L})=\overline{L}^{\cdot(\dim X+1)}$.
+    (Arithmetic Siu) Let $\overline{L},\overline{M}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})_{\mathrm{nef}}$. Then
+    $\widehat{{\mathrm{vol}}}(X,\overline{L}-\overline{M})\geq\overline{L}^{\cdot(\dim X+1)}-(\dim X+1)\overline{L}^{\cdot\dim X}\cdot\overline{M}.$
+    (continuity) Let $\overline{L},\overline{M}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})$.
+Then $\lim_{t\rightarrow 0}\widehat{{\mathrm{vol}}}(\overline{L}+t\overline{M})=\widehat{{\mathrm{vol}}}(\overline{L})$.
+The discussion above applies to $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\mathrm{int}}$. For each $\widetilde{L}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\mathrm{int}}$, define
+$\widehat{h}^{0}(X,\widetilde{L}):=\dim\{s\in H^{0}(X,L):\sup_{x\in X^{\mathrm{an}}_{\mathrm{triv}}}\|s(x)\|\leq 1\}.$
+The limit
+(E.11) $\widehat{{\mathrm{vol}}}(X,\widetilde{L}):=\lim_{m\rightarrow\infty}\frac{(\dim X)!}{m^{\dim X}}\widehat{h}^{0}(X,m\widetilde{L})$
+exists and is defined to be the volume of $\widetilde{L}$.
+Then $\widetilde{L}$ is said to be big if $\widehat{{\mathrm{vol}}}(X,\widetilde{L})>0$.
+We also have arithmetic Hilbert–Samuel and Siu for elements in $\widehat{\mathrm{Pic}}(X/{\mathbb{Q}})_{\mathrm{nef}}$, with $\dim X$ replaced by $\dim X-1$ in the formulae.
+With these preparations, we are ready to state the height of inequality.
+Theorem E.9 ([YZ21, Thm. 5.3.5]).
+Let $\pi\colon X\rightarrow S$ be a morphism of quasi-projective varieties defined over a number field $K$.
+Let $\overline{L}\in\widehat{\mathrm{Pic}}(X/{\mathbb{Z}})_{\mathrm{int}}$ and $\overline{M}\in\widehat{\mathrm{Pic}}(M/{\mathbb{Z}})_{\mathrm{int}}$.
+If the generic fiber $\widetilde{L}$ of $\overline{L}$ is big, then there exists a Zariski open dense subset $U$ of $X$ and positive constants $\epsilon$ and $c$, such that
+$h_{\overline{L}}(x)\geq\epsilon h_{\overline{M}}(\pi(x))-c,\qquad\text{for all}x\in U(\overline{\mathbb{Q}}).$
+If, moreover, $\overline{L}$ is big, then one can remove the constant $c$ from the inequality above.
+When $\overline{L}$ is big, Arithmetic Siu and continuity together imply $\widehat{{\mathrm{vol}}}(\overline{L}-\epsilon\pi^{*}\overline{M})>0$ by definition of bigness.
+Thus, the conclusion follows immediately from (E.10).
+The general case is reduced to the case where $\overline{L}$ is big by [YZ21, Lem.
+5.2.10], which asserts that the bigness of $\widetilde{L}$ is not far from the bigness of the whole $\overline{L}$.

### v3

@@
   • the Ceresa cycle based at $\xi$ defined by ${\mathrm{Ce}}_{\xi}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in{\mathrm{Ch}}_{1}({\mathrm{Jac}}(C))$ defined by embedding $i_{\xi}\colon C{\longrightarrow}{\mathrm{Jac}}(C)$ via $\xi$ ([Cer83]).
-Let ${\mathscr{M}}_{g}$ be the moduli space of smooth projective curves of genus $g$, let ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$ be the universal curve, and let $\pi\colon\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g}){\longrightarrow}{\mathscr{M}}_{g}$ be the relative Picard group of degree $1$ on ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$.
+Let ${\mathscr{M}}_{g}$ be the coarse moduli space of smooth projective curves of genus $g$.
+Let ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$ be the universal curve, and let $\pi\colon\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g}){\longrightarrow}{\mathscr{M}}_{g}$ be the relative Picard group of degree $1$ on ${\mathscr{C}}_{g}/{\mathscr{M}}_{g}$, both as Deligne–Mumford stacks.
 For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathscr{C}}_{s}$ the curve parametrized by $s$.
@@
 for $s=\pi(\xi)\in{\mathscr{M}}_{g}({\mathbb{C}})$.
-For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, a particularly important element $\xi_{s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ is such that the divisor class of $(2g-2)\xi_{s}$ is the canonical divisor class on ${\mathscr{C}}_{s}$; it is well-defined up to $(2g-2)$-torsions.
+For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, a particularly important element $\xi_{\mathrm{can},s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ is such that the divisor class of $(2g-2)\xi_{\mathrm{can},s}$ is the canonical divisor class on ${\mathscr{C}}_{s}$; it is well-defined up to $(2g-2)$-torsions.
 Here are our main results about the Beilinson–Bloch heights for these cycles unconditionally defined by Gross–Schoen [GS95] and Künnemann [Kün01] respectively.
 Theorem 1.1.
-For each $g\geq 3$, there exist a Zariski open dense subset ${\mathscr{M}}_{g}^{\mathrm{amp}}$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$, such that for any $s\in{\mathscr{M}}_{g}^{\mathrm{amp}}(\overline{\mathbb{Q}})$ and any $\xi\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$, we have
-(1.1) $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{s}))-c,$
-$\langle{{\mathrm{Ce}}(\xi),{\mathrm{Ce}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{s}))-c,$
+For each $g\geq 3$, there exist a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$, such that for any $s\in U(\overline{\mathbb{Q}})$ and any $\xi\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$, we have
+(1.1) $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{\mathrm{can},s}))-c,$
+$\langle{{\mathrm{Ce}}(\xi),{\mathrm{Ce}}(\xi)}\rangle_{\mathrm{BB}}\geq\epsilon(h_{\mathrm{Fal}}(s)+h_{{\mathrm{NT}}}(\xi-\xi_{\mathrm{can},s}))-c,$
 where $h_{\mathrm{Fal}}$ is the Faltings height on ${\mathscr{M}}_{g}(\overline{\mathbb{Q}})$ and $h_{\mathrm{NT}}$ is the canonical Néron–Tate height on ${\mathrm{Jac}}({\mathscr{C}}_{s})(\overline{\mathbb{Q}})$.
-The proof will be executed in §$\S$6.3, in which we give a construction of ${\mathscr{M}}_{g}^{\mathrm{amp}}$.
-Indeed, we will take ${\mathscr{M}}_{g}^{\mathrm{amp}}$ to be ${\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, where ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is the Betti degeneracy locus (1.3) with the normal function $\nu$ being the Gross–Schoen normal function (C.6).
+If (1.1) holds over Zariski open dense subsets $U_{1}$ and $U_{2}$ of ${\mathscr{M}}_{g}$, then it holds also over $U_{1}\cup U_{2}$. Thus, we can propose the following definition:
+  1. (i)
+    the ample locus of the Gross–Schoen and Ceresa cycles, denoted by ${\mathscr{M}}_{g}^{\mathrm{amp}}$, is the maximal open subset of ${\mathscr{M}}_{g}$ over ${\mathbb{Q}}$ so that (1.1) holds;
+  2. (ii)
+    ${\mathscr{M}}_{g}^{\mathrm{slim}}:={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{amp}}$ the slim locus of the Gross–Schoen and Ceresa cycles.
+Theorem 1.1 is then equivalent to: ${\mathscr{M}}_{g}^{\mathrm{amp}}$ is non-empty if $g\geq 3$.
+The proof will be executed in §$\S$6.3, in which we give an explicit construction of ${\mathscr{M}}_{g}^{\mathrm{amp}}$ (or equivalently its complement ${\mathscr{M}}_{g}^{\mathrm{slim}}$).
+Indeed, we will show that ${\mathscr{M}}_{g}^{\mathrm{slim}}$ is precisely the Betti degeneracy locus ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ defined in (1.3) with the normal function $\nu$ being the Gross–Schoen normal function (C.6).
 A priori ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is only a real analytic set.
@@
 Finally, we show that it is defined over ${\mathbb{Q}}$ by volume calculation for adelic line bundles; see Corollary 6.4.
-Because of the importance of this Zariski open dense set ${\mathscr{M}}_{g}^{\mathrm{amp}}$ on which (1.1) holds, we will make the following definition: We call
-  1. (i)
-    ${\mathscr{M}}_{g}^{\mathrm{amp}}$ the ample locus of the Gross–Schoen and Ceresa cycles;
-  2. (ii)
-    ${\mathscr{M}}_{g}^{\mathrm{slim}}:={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{amp}}$ the slim locus of the Gross–Schoen and Ceresa cycles.
-By the explanation above, it is the ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
-It certainly contains the hyperelliptic locus and also the locus of Picard curves by Laga–Shnidman [LS24, Thm.
+We point out that ${\mathscr{M}}_{g}^{\mathrm{slim}}$ contains the hyperelliptic locus and also the locus of Picard curves by Laga–Shnidman [LS24, Thm.
 C].
-We expect ${\mathscr{M}}_{g}^{\mathrm{amp}}$ is the largest open subset of ${\mathscr{M}}_{g}$ so that the height inequality (1.1) in Theorem 1.1 holds; see Conjecture 1.12.
 Theorem 1.1 yields the following Northcott property immediately.
@@
 In the course of the proof, we also establish the following geometric result on the torsion locus.
-For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s}):={\mathrm{Ce}}(\xi_{s})$.
+For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{\mathrm{can},s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s}):={\mathrm{Ce}}(\xi_{\mathrm{can},s})$.
 Theorem 1.3.
@@
 Moreover, $S^{\mathrm{amp}}$ is the largest open subvariety of $S$ on which $s\mapsto\langle{Z_{s},Z_{s}}\rangle_{\mathrm{BB}}$ is a dominant height.
+In practice, in addition to this conjecture one also needs to check whether $S^{\mathrm{amp}}$ is empty, by using (1.6) and Theorem 1.6.
 The second part of Conjecture 1.12 is equivalent to: For any immersion $S\subseteq{\mathbb{P}}^{N}$ defined over $\overline{\mathbb{Q}}$, there exists a Zariski dense sequence $s_{i}\in S^{\mathrm{slim}}(\bar{K})$ such that
 $\lim_{i}h_{\mathrm{Weil}}(s_{i})=\infty,\qquad\lim_{i}\frac{\langle{Z_{s_{i}},Z_{s_{i}}}\rangle_{\mathrm{BB}}}{h_{\mathrm{Weil}}(s_{i})}=0.$
-The proof of Theorem 1.1 executed in §$\S$6.3 in fact proves the first part of Conjecture 1.12 in the case of Gross–Schoen and Ceresa cycles for any subvariety of ${\mathscr{M}}_{g}$.
-A key ingredient is the volume identity (6.4).
+The proof of Theorem 1.1 executed in §$\S$6.3 in fact proves the following:
+Theorem 1.13.
+Conjecture 1.12 holds in the case of Gross–Schoen and Ceresa cycles for any subvariety of ${\mathscr{M}}_{g}$.
+A key ingredient to prove this theorem is the volume identity (6.4).
 The same argument can be used to show that the following conjectured volume identity (1.10) implies the full Conjecture 1.12 with the help of [YZ21, Thm.
-6.3.6].
-To check whether $S^{\mathrm{amp}}$ is non-empty, one can use (1.6) and Theorem 1.6.
+5.3.8].
 In [Blo89], Bloch constructed a biextension line bundle $L$ on $S$ defined over $K$.
@@
 5.2.2] (which we recall as Theorem E.8.(i)), one consequence of this conjecture is the following volume identity for the generic fiber $\widetilde{L}$ of $\overline{L}$ (see (E.6) for the definition): For any archimedean place $v$ of $K$ and any subvariety $S^{\prime}$ of $S_{v,{\mathbb{C}}}$, we have
-(1.10) $\widehat{\mathrm{vol}}(\widetilde{L}|_{S^{\prime}})=\int_{S^{\prime}({\mathbb{C}})}c_{1}(\bar{L}_{v})^{\dim S}.$
+(1.10) $\widehat{\mathrm{vol}}(\widetilde{L}|_{S^{\prime}})=\int_{S^{\prime}({\mathbb{C}})}c_{1}(\bar{L}_{v})^{\wedge\dim S^{\prime}}.$
 For the Gross–Schoen and Ceresa cycles for any subvariety of ${\mathscr{M}}_{g}$, we prove the existence of the extension in §$\S$6.1 and the volume identity as Theorem 6.3, but not the nefness of the adelic line bundle in question.
@@
 2.5.5] the desired $\overline{{\mathscr{L}}}$ can be obtained from a suitable Deligne pairing.
-This $\overline{{\mathscr{L}}}$ has been constructed by Yuan [Yua21, §$\S$3.3.4] is slightly different steps.
-The required bigness property is the bigness of the generic fiber $\widetilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$.
+This $\overline{{\mathscr{L}}}$ has been constructed by Yuan [Yua21, §$\S$3.3.4] in slightly different steps, where he also asked the question of nefness and bigness of $\overline{{\mathscr{L}}}$.
+In our paper, the required bigness property is the bigness of the generic fiber $\widetilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$.
 We furthermore prove the volume identity as Theorem 6.3, which implies that the bigness of $\widetilde{{\mathscr{L}}}$ will follow from the non-vanishing of $c_{1}(\overline{{\mathscr{L}}})^{\wedge\dim S}$ with $c_{1}(\overline{{\mathscr{L}}})$ the curvature form.
@@
 Acknowledgements
-We would like to thank Ari Shnidman for asking us the question about the Northcott property for Ceresa cycles, Richard Hain for numerous inspiring discussions and for sharing his notes, Xinyi Yuan for his help during the preparation of this project, Dick Gross for suggesting we look at $\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ instead of just ${\mathscr{M}}_{g}$, and Peter Sarnak for asking the question about the positivity of the Beilinson–Bloch heights.
+We would like to thank Ari Shnidman for asking us the question about the Northcott property for Ceresa cycles, Richard Hain for numerous inspiring discussions and for sharing his notes, Xinyi Yuan for his help during the preparation of this project and comments on a previous version of the paper, Dick Gross for suggesting we look at $\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ instead of just ${\mathscr{M}}_{g}$, and Peter Sarnak for asking the question about the positivity of the Beilinson–Bloch heights.
 Part of this work was completed when ZG visited SZ at Princeton University in January, February, and May 2024.
@@
 Set ${\mathscr{J}}^{\prime}:=S^{\prime}\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})$; it is an intermediate Jacobian over $S^{\prime}$.
-Then the inclusion $S^{\prime}\subseteq\Gamma\backslash{\mathscr{D}}$ yields a section $\nu^{\prime}$ of ${\mathscr{J}}^{\prime}\rightarrow S^{\prime}$, and thus we can define $S^{\prime}_{{\mathscr{F}}}(t)$ for each $t\geq 0$ with respect to $\nu^{\prime}$.
+Then the inclusion $S^{\prime}\subseteq\Gamma\backslash{\mathscr{D}}$ yields a section $\nu^{\prime}$ of ${\mathscr{J}}^{\prime}\rightarrow S^{\prime}$, and thus we can define $S^{\prime\mathrm{Betti}}(t)$ for each $t\geq 0$ with respect to $\nu^{\prime}$.
 We have the following commutative diagram, with $\varphi_{{\mathscr{J}}}$ induced by $\varphi$, such that $\varphi_{{\mathscr{J}}}\circ\nu=\nu^{\prime}\circ\varphi$:
@@
 By (the proof of) Lemma 2.1, more precisely the second equality of (2.1), we have
-(3.3) $S^{\mathrm{Betti}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime}_{{\mathscr{F}}}(t-r)\right).$
+(3.3) $S^{\mathrm{Betti}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime\mathrm{Betti}}(t-r)\right).$
 This equality allows us to replace $S$ by $\varphi(S)$ to study the Betti rank and the Betti strata.
@@
 Then $S^{\mathrm{Betti}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$.
-By (3.3) and Proposition 3.9 (which should be applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
+By (3.3) and Proposition 3.9 (which should be applied to $S^{\prime\mathrm{Betti}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
 $S^{\mathrm{Betti}}(t)\subseteq S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
@@
 […]
-By (3.3), Proposition 3.9 and Lemma 3.10 (both applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
+By (3.3), Proposition 3.9 and Lemma 3.10 (both applied to $S^{\prime\mathrm{Betti}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
 (3.8) $S^{\mathrm{Betti}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
@@
 2.3.5].
-Recall the notation ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{s})$ with $\xi_{s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ such that $(2g-2)\xi_{s}$ being the canonical divisor class on ${\mathscr{C}}_{s}$.
+Recall the notation ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{\mathrm{can},s})$ with $\xi_{\mathrm{can},s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ such that $(2g-2)\xi_{\mathrm{can},s}$ being the canonical divisor class on ${\mathscr{C}}_{s}$.
 Theorem 6.1.
@@
 Let $S^{\prime}$ be an irreducible component of $S^{\mathrm{Betti}}(1)$.
-Then $S^{\prime}=S^{\prime}_{{\mathscr{F}}}(1)$, and hence $\beta_{{\mathrm{GS}}}^{\wedge\dim S^{\prime}}\equiv 0$ on $S^{\prime}$; see (1.6) and Corollary D.7.
+Then $S^{\prime}=S^{\prime\mathrm{Betti}}(1)$, and hence $\beta_{{\mathrm{GS}}}^{\wedge\dim S^{\prime}}\equiv 0$ on $S^{\prime}$; see (1.6) and Corollary D.7.
 So $\widehat{\mathrm{vol}}(\widetilde{{\mathscr{L}}}_{{\mathbb{C}}}|_{S^{\prime}})=0$ by Theorem 6.3.
@@
 Thus $\beta_{{\mathrm{GS}}}^{\wedge\dim S^{\prime}}\equiv 0$ on $S^{\prime\sigma}$ by Proposition 6.2, and therefore
-$S^{\prime\sigma}({\mathbb{C}})=S^{\prime\sigma}_{\mathscr{F}}(1)\subseteq S^{\mathrm{Betti}}(1).$
+$S^{\prime\sigma}({\mathbb{C}})=(S^{\prime\sigma})^{\mathrm{Betti}}(1)\subseteq S^{\mathrm{Betti}}(1).$
 This shows that $S^{\mathrm{Betti}}(1)\otimes_{\sigma}{\mathbb{C}}=S^{\mathrm{Betti}}(1)$. Thus $S^{\mathrm{Betti}}(1)$ is defined over $k$. ∎
@@
 Now, we are ready to prove our main theorem.
-Proof of Theorem 1.1.
-It suffices to prove the claim for $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}$ because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple.
+Proof of Theorem 1.1 and of ${\mathscr{M}}_{g}^{\mathrm{amp}}={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+We start by proving Theorem 1.1.
+It suffices to prove the height comparison (1.1) for $\langle{{\mathrm{GS}}(\xi),{\mathrm{GS}}(\xi)}\rangle_{\mathrm{BB}}$ because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple.
 Next by [Zha10, Cor.
-2.5.2], it suffices to prove for $\xi=\xi_{s}$.
-We start by showing there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$ such that
+2.5.2], it suffices to prove (1.1) for $\xi=\xi_{\mathrm{can},s}$, i.e.
+there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$ such that
 (6.5) $\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{{\mathrm{Fal}}}(s)-c$
@@
 5.3.5.(iii)] (which we reformulated as Theorem E.9).
-Notice that one can either take the adelic line bundle on ${\mathscr{M}}_{g}$, which defines the Faltings height (it exists by [YZ21, §$\S$2.6.2]), or one can take any ample line bundle on a suitable compactification of ${\mathscr{M}}_{g}$ and then use the comparison of the logarithmic Weil height and the Faltings height.
-We can do better to get an explicit $U$.
-Suppose $U_{1},U_{2}$ are open subsets of ${\mathscr{M}}_{g}$ over ${\mathbb{Q}}$ such that (6.5) holds true for all $s\in U_{1}(\overline{\mathbb{Q}})$ and $s\in U_{2}(\overline{\mathbb{Q}})$, then (6.5) holds true for all $s\in(U_{1}\cup U_{2})(\overline{\mathbb{Q}})$.
-Therefore, there exists a maximal Zariski open dense subset $U^{\prime}$ of ${\mathscr{M}}_{g}$ on which (6.5) holds.
-Let $S:={\mathscr{M}}_{g}\setminus U^{\prime}$.
-Then $S$ is proper Zariski closed in ${\mathscr{M}}_{g}$, and the height inequality (6.5) cannot hold for any non-empty open subvariety $V$ of $S$.
-By Theorem E.9, $\widetilde{{\mathscr{L}}}$ cannot be big on any irreducible component $Z$ of $S$, and hence $c_{1}(\overline{\mathscr{L}}|_{Z({\mathbb{C}})})^{\wedge\dim Z}\equiv 0$ by Proposition 6.3.
-Thus $Z_{{\mathbb{C}}}=Z_{{\mathscr{F}}}(1)\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
-So $S_{{\mathbb{C}}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
-So $S\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ by Corollary 6.4.
-Thus we can take ${\mathscr{M}}_{g}^{\mathrm{amp}}$ to be ${\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, which is Zariski open dense and is defined over ${\mathbb{Q}}$, in Theorem 1.1.∎
+Notice that one can either take the adelic line bundle $\overline{{\mathscr{M}}}$ on ${\mathscr{M}}_{g}$ which defines the Faltings height (it exists by [YZ21, §$\S$2.6.2]), or one can take any ample line bundle on a suitable compactification of ${\mathscr{M}}_{g}$ and then use the comparison of the logarithmic Weil height and the Faltings height.
+Now we prove ${\mathscr{M}}_{g}^{\mathrm{amp}}={\mathscr{M}}_{g}\setminus{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, or equivalently ${\mathscr{M}}_{g}^{\mathrm{slim}}={\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+Notice that ${\mathscr{M}}_{g}^{\mathrm{slim}}$ is by definition defined over ${\mathbb{Q}}$ and ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ is defined over ${\mathbb{Q}}$ by Corollary 6.4.
+We start by showing ${\mathscr{M}}_{g}^{\mathrm{slim}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+By definition of ${\mathscr{M}}_{g}^{\mathrm{slim}}$, the height inequality (6.5) cannot hold for any non-empty open subvariety $V$ of ${\mathscr{M}}_{g}^{\mathrm{slim}}$.
+By Theorem E.9, $\widetilde{{\mathscr{L}}}$ cannot be big on any irreducible component $S$ of ${\mathscr{M}}_{g,\overline{\mathbb{Q}}}^{\mathrm{slim}}$, and hence $c_{1}(\overline{\mathscr{L}}|_{S({\mathbb{C}})})^{\wedge\dim S}\equiv 0$ by Proposition 6.3.
+Thus $S=S^{\mathrm{Betti}}(1)\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+So ${\mathscr{M}}_{g,\overline{\mathbb{Q}}}^{\mathrm{slim}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+So ${\mathscr{M}}_{g}^{\mathrm{slim}}\subseteq{\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$ by Corollary 6.4.
+It remains to prove that $S:={\mathscr{M}}_{g}^{\mathrm{Betti}}(1)\setminus{\mathscr{M}}_{g}^{\mathrm{slim}}={\mathscr{M}}_{g}^{\mathrm{Betti}}(1)\cap{\mathscr{M}}_{g}^{\mathrm{amp}}$ is empty.
+By definition of ${\mathscr{M}}_{g}^{\mathrm{amp}}$, the inequality (6.5) holds for all $s\in S(\overline{\mathbb{Q}})$ and $S$ is Zariski open in ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+If $S\not=\emptyset$, then ~$\widetilde{{\mathscr{L}}}|_{Z}$ is big by [YZ21, Thm.
+5.3.8] for some irreducible component $Z$ of ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$, and hence $c_{1}(\overline{{\mathscr{L}}}|_{Z({\mathbb{C}})})^{\dim Z}\not\equiv 0$.
+But this contradicts the construction of ${\mathscr{M}}_{g}^{\mathrm{Betti}}(1)$.
+So $S=\emptyset$, and we are done.
+∎
 6.4. Proof of Theorem 6.3
@@
 With these preparations, we are ready to state the height of inequality.
-Theorem E.9 ([YZ21, Thm. 5.3.5]).
+Theorem E.9 ([YZ21, Thm. 5.3.7]).
 Let $\pi\colon X\rightarrow S$ be a morphism of quasi-projective varieties defined over a number field $K$.